### v1

# Understanding Practical Membership Privacy of Deep Learning

Abstract
We apply a state-of-the-art membership inference attack (MIA) to systematically test the practical privacy vulnerability of fine-tuning large image classification models. We focus on understanding the properties of data sets and samples that make them vulnerable to membership inference. In terms of data set properties, we find a strong power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate. For an individual sample, large gradients at the end of training are strongly correlated with MIA vulnerability.
1 Introduction
Machine learning models are prone to memorising their training data, which makes them vulnerable to privacy attacks such as membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and reconstruction attacks (e.g. Balle et al., 2022; Nasr et al., 2023). Differential privacy (DP; Dwork et al., 2006b) provides protection against these attacks, but strong formal protection often comes at the cost of significant loss of model utility.
Finding the correct balance between making models resistant to attacks while maintaining a high utility is important for many applications. In health, for example, many European countries and soon also the EU within the European Health Data Space have requirements that models trained on health data that are made publicly available must be anonymous, i.e. they must not contain information that can be linked to an identifiable individual. On the other hand, loss of utility of the model may compromise the health benefits that might be gained from it.
In this paper, our aim is to apply a state-of-the-art MIA systematically to help understand practical privacy risks training deep-learning-based classifiers without DP protections. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep learning models to MIA. Since the earliest MIA methods, there has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017). Similarly, there is evidence that models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023). Based on extensive experimental data over many data sets with varying sizes, we uncover a power law describing the vulnerability remarkably well.
Previously Tobaben et al. (2023) and Yu et al. (2023) have presented limited studies of how data set properties affect the MIA vulnerability. Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots (i.e. the number of examples available per class). Yu et al. (2023) studied how the MIA vulnerability of a class in an image classifier trained from scratch depends on the average individual privacy parameters of that class. Our work significantly expands on both of these by more systematic study and modelling of data set properties and by providing a more detailed look at individual samples, not just classes.
List of contributions
In this work we perform an extensive empirical study on the MIA vulnerability of deep learning models. We will focus on transfer learning setting on computer vision tasks, where a large pre-trained neural network is fine-tuned on a sensitive data set, and the MIA is performed on the final classifier. We find that the MIA vulnerability has a strong correlation with certain data set properties (Section 5) and make the following contributions:
  1. 1.
    Power law: We uncover a power law between the number of examples per class and the vulnerability to MIA (tpr at fixed low fpr) based on extensive experimental data over many data sets with varying sizes (See Figure 1(a)).
  2. 2.
    Regression model: We utilize the observations to train a regression model to predict MIA vulnerability (tpr at fixed low fpr) based on examples per class ($S$) and number of classes ($C$) and show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor (See Figure 3).
Additionally, in Section 6 we extend the analysis to individual samples and discover a correlation between vulnerability of individual samples and the gradient norms during training. We further study the correlation between sample’s vulnerability and the similarity to the other samples in the feature space.
2 Background
Notation for the properties of the training data set $\mathcal{\mathcal{D}}$:
  • $C$ for the number of classes
  • $S$ for shots (examples per class)
  • $\mathcal{|\mathcal{D}|}$ for training data set size ($\mathcal{|\mathcal{D}|}=CS$)
We denote the number of MIA shadow models with $M$.
Membership inference attacks (MIAs) aim to infer whether a particular sample was part of the training set of the targeted model Shokri et al. (2017). Thus, they can be used to determine lower bounds on the privacy leakage of models to complement the theoretical upper bounds obtained through differential privacy.
Likelihood Ratio attack (LiRA; Carlini et al., 2022) While many different MIAs have been proposed Hu et al. (2022), in this work we consider the SOTA: the Likelihood Ratio Attack (LiRA). LiRA assumes an attacker that has black-box access to the attacked model, knows the training data distribution, the training set size, the model architecture, hyperparameters and training algorithm. Based on this information, the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack but for which the attacker knows the training data set.
LiRA exploits the observation that the value of the loss function used to train a model is often lower for examples that have been part of the training set than for examples that have not been. For a target sample $x$, LiRA trains the shadow models: 1 with $x$ as a part of the training set ($x\in\mathcal{D}$) 2 and without $x$ in the training set ($x\notin\mathcal{D}$) . After training the shadow models, $x$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed, one for the losses of $x\in\mathcal{D}$ shadow models and one for the $x\notin\mathcal{D}$. Finally the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $x\in\mathcal{D}$ or $x\notin\mathcal{D}$. Carlini et al. (2022) proposed an optimization for performing LiRA for multiple models and points without training a computationally infeasible amount of shadow models. We utilize this optimization in our experiments.
In Appendix A we provide background on Differential privacy (DP; Dwork et al., 2006b) and DP stochastic gradient descent (DP-SGD) that we use for training DP models.
3 Methods
In this Section, we introduce the preliminaries for evaluating MIA vulnerability.
3.1 Measuring MIA vulnerability
Using the likelihood-ratios from LiRA Carlini et al. (2022) as a score, we can build a binary classifier to predict whether sample belongs to the training data or not. The operational characteristics of such classifier can be used to measure the success of the MIA. More specifically, throughout the rest of the paper, we will use the true positive rate (tpr) at a specific false positive rate (fpr) as a measure for the vulnerability. Identifying even a small number of samples with high confidence is considered harmful (Carlini et al., 2022) and thus we focus on the regions of small fpr.
The tpr and fpr can be also connected to DP. Kairouz et al. (2015) have shown in Equation 1 that any classifier distinguishing the training samples based on the results of a DP algorithm has an upper bound for the tpr that depends on the privacy parameters $(\epsilon,\delta)$. Since MIA is trying to build exactly these types of classifiers, we can use the upper bounds by Kairouz et al. (2015) to validate the privacy claims, and also to better understand the gap between the theoretical privacy guarantees and the realistic attacks.
Theorem 1 (Kairouz et al. (2015)).
A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP if and only if for all adjacent $\mathcal{D}\sim\mathcal{D}^{\prime}$
$\mathrm{TPR}\leq\min\{e^{\epsilon}\mathrm{FPR}+\delta,1-e^{-\epsilon}(1-\delta-\mathrm{FPR})\}\ .$ (1)
3.2 Measuring the uncertainty for TPR
The tpr values from the LiRA based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples. Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimator. Therefore, when we report the tpr values for single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals Clopper & Pearson (1934). Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
$\begin{split}&B(\alpha/2;\textsc{tp},\textsc{p}-\textsc{tp}+1)<\textsc{tpr}\\ &\textsc{tpr}<B(1-\alpha/2;\textsc{tp}+1,\textsc{p}-\textsc{tp}),\end{split}$ (2)
where $B(q;a,b)$ is the $q$th-quantile of $\text{Beta}(a,b)$ distribution.
4 Experiment overview
Throughout the experiments, we utilize LiRA to understand the vulnerability of the trained models.
We fine-tune pre-trained models on sensitive downstream data sets and assess the vulnerability using LiRA. We use the following setup for all experiments:
  • Pre-trained models. BiT-M-R50x1 (R-50; Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021) with 85.8M parameters, both pre-trained on the ImageNet-21K data set (Russakovsky et al., 2015)
  • Parameterization. Due to computational cost we only consider fine-tuning subsets of the model parameters. We consider the following configurations: 1 Head: training a linear last layer classifier on top of the pre-trained model 2 FiLM: Head and additionally training parameter-efficient FiLM (Perez et al., 2018) adapters scattered throughout the network (see Section B.1)
  • Fine-tuning data sets. We utilize different fine-tuning data sets (see Section B.3 for details) with different $S$, $C$ and $|\mathcal{D}|$ as specified in the experiments.
  • Hyperparameter optimization (HPO). We utilize the library Optuna Akiba et al. (2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations. We optimize the batch size, learning rate and number of epochs. Additionally for DP-SGD we optimize the clipping norm. Details in Section B.2.
  • DP-SGD. We are using Opacus Yousefpour et al. (2021) and compute the noise multiplier based on the target ($\epsilon,\delta$).
  • DP Accounting. We use numerical privacy accounting for accurate evaluation of the cumulative privacy loss of the optimisation (Koskela et al., 2020, 2021) using the PRV accountant (Gopi et al., 2021).
  • LiRA. We base our experiments on the publically available source code of Carlini et al. (2022). We utilize $M=256$ shadow models unless specified otherwise.
We illustrate the effectiveness of LiRA and the effect of different levels of DP on the MIA vulnerability (tpr at $\textsc{fpr}=0.001$) in Figure 1. It shows the MIA vulnerability as a function of privacy budget (additional versions for different fprs in Figure A.1) and two upper bounds on the tpr at fpr emerging from the theoretical guarantees of DP-SGD.
Figure 1: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of privacy budget ($\epsilon$ at $\delta=10^{-5}$) when attacking different backbones fine-tuned on CIFAR10 and CIFAR100 at $S\in\{25,50\}$. The solid line displays the median and the error bars the minimum of the lower bounds and the maximum of the upper bounds for the Clopper-Pearson CIs over 15 seeds. The upper bound is the theoretical DP bound based on Kairouz et al. (2015). We present the bound computed with $\delta=10^{-5}$ as well as the optimal bound based on multiple $\delta$.
We compute the upper bounds using the method of Kairouz et al. (2015) in Equation 1. Simply using $\delta=10^{-5}$ and the associated $\epsilon$ gives a relatively weak bound, that can be sharpened considerably by finding an optimal $\delta$ and corresponding $\epsilon(\delta)$, which is made possible by the particular mechanism deployed, that gives the tightest possible bound. For the attacked few-shot models the bounds are quite tight for the medium privacy budget of $\epsilon=1$ but very loose at low privacy budgets of $\epsilon=8$ and thus meaningless for low privacy levels.
5 Predicting MIA vulnerability for data sets
In this Section, we investigate how different properties of data sets affect the MIA vulnerability in the non-DP setting. Based on our observations, we propose a method to predict the vulnerability to MIA using these properties of the data set.
We base our experiments on a subset of the few-shot benchmark VTAB  (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by practitioners. For more detailed descriptions of the data sets, see Table A2. In this Section, we will focus on a fine-tuned last layer classifier (Head) trained on top of a ViT-B, pre-trained on ImageNet-21k Russakovsky et al. (2015). The results for using R-50 as a backbone can be found in Section C.1.
5.1 Correlations with data set properties
Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 3.1). We make the following observations:
  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in Figures A.2, A3 and A4 in the Appendix.
  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 1(b) with further evidence in Figures A.3, A5 and A6 in the Appendix. However, the trend w.r.t. $C$ is not as clear as with $S$.
(a) as function of $S$ (shots)
(b) as function of $C$ (classes)
Figure 2: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of dataset properities when attacking a ViT-B Head fine-tuned without DP on different data sets. The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over multiple seeds (6 for Figure 1(a) and 12 for Figure 1(b))
5.2 Model to predict data set vulnerability
Based on the trends seen in Figures 1(b) and 1(a), we fit a linear regression model using the logarithmic transform of $C,S$ to predict the tpr (also logarithmically transformed) with statsmodels Seabold & Perktold (2010). The general form of the model can be found in Equation 3 where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
$\log_{10}(\textsc{tpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (3)
We utilize MIA results of ViT-B (see Table A3) as the training data and use R-50 (see Table A4) as the test data to investigate if the vulnerability prediction model generalizes to a completely different backbone.
The parameters of the prediction model fitted to the training data (note that the parameters are specific to $\textsc{fpr}=0.001$) are $\beta_{S}=-0.492,\beta_{C}=0.357$ and $\beta_{0}=-0.486$. Based on the $R^{2}$ score ($R^{2}=0.917$), our model fits the data extremely well. Furthermore, Figure 3 shows that the model is robust to a change of feature extractor, as it is able to predict the tpr for R-50 despite being trained on ViT-B (test $R^{2}=0.740$). See Section C.1.3 for further supporting results at $\textsc{fpr}\in\{0.1,0.01\}$.
Figure 3: Predicted MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of $S$ (shots) using a model based on Equation 3 fitted Table A3 (ViT-B). The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100).
6 MIA vulnerability for data samples
This Section extends our investigation from the data set level properties studied in Section 5 to individual data sample level properties. Previously Yu et al. (2023) have studied correlations between classes and privacy risk before, but in comparison to their study, we look at data samples independent from classes.
We will focus on a fine-tuned last layer classifier (Head) trained on top of a ViT-B, pre-trained on ImageNet-21k Russakovsky et al. (2015). We investigate two data sample properties that could potentially influence the vulnerability (tpr at fpr) of data points:
  • Per-sample gradients: Per-sample gradients form the basis for DP individual accounting Feldman & Zrnic (2021); Koskela et al. (2023). In individual accounting a larger sum of gradient norms results in a higher individual privacy loss. Our hypothesis is that this holds also for MIA vulnerability in non-DP settings.
  • Vector space distances: The pre-trained feature extractors project images into a vector space in which one can compute distances between individual data samples. Our hypothesis is that data samples that are further away from other data samples are easier to distinguish.
We study individual samples, but aggregate the results over multiple data samples by binning the samples based on their data sample properties to reduce the noise. For each data sample we compute its property and assign the sample to a bin based on its percentile in relation to other samples. We then compute the tpr at fpr for each bin. We finally report for each data property bin the percentile of the tpr at fpr.
Figure 4: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of data sample properties (binned among a seed). The results have been obtained over 6 seeds with $S=256$ (Patch Camelyon) and $S=64$ (CIFAR10).
Figure 4 shows the observed relationships between the MIA vulnerability and the average individual gradient norms and the vector space distance. We find that:
  • Per-sample gradients: The average gradient norms are correlated with the vulnerability in all our experiments, but the strength of the correlation depends how many steps are taken into account. Generally, we observe that the correlation is stronger when we only consider the gradients at the end of the training (either the last 20 or 5 steps). This is different from the individual accounting methods that rely on the whole gradient trace for DP accounting.
  • Vector space distances: The distance of data samples to their nearest neighbor correlates with the vulnerability for CIFAR10, but not for Patch Camelyon. We leave a more comprehensive study of this correlation and its limitations for future work.
7 Discussion
As DP reduces the utility of models, it is important to understand also when it is necessary. Our results on MIA vulnerability as a function of data set properties display a strong power law dependence on the number of examples per class. This leads to an unfortunate dilemma: DP seems most important in the small data regime where it likely reduces the model utility the most.
Limitations
Due to computational constraints, we only report results on transfer learning with fine-tuning only a part of the model.
Acknowledgements
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. The authors wish to thank the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources. We thank John F. Bronskill for helpful discussions regarding few-shot learning.
Appendix A Background
Differential privacy (DP; Dwork et al., 2006b) is the gold standard for formalizing privacy guarantees. $(\epsilon,\delta)$-DP (Dwork et al., 2006a) has a privacy budget consisting of $\epsilon\geq 0$ and $\delta\in[0,1]$, where smaller values correspond to a stronger privacy guarantee. We refer to Dwork & Roth (2014) for a comprehensive intro to DP.
DP deep learning utilizes DP-SGD (Rajkumar & Agarwal, 2012; Song et al., 2013; Abadi et al., 2016), which is the DP adaptation of stochastic gradient descent. DP-SGD clips per-sample gradients to a pre-determined norm to limit the influence of each sample and adds Gaussian noise the sum of the clipped gradients.
Clipping the gradients and adding Gaussian noise in DP-SGD typically comes with a cost in accuracy. This decrease in accuracy can be alleviated through transfer-learning. Publically available pre-trained models are fine-tuned with DP-SGD on sensitive downstream data sets. Both in NLP (Li et al., 2022; Yu et al., 2022) and computer vision (Kurakin et al., 2022; De et al., 2022; Mehta et al., 2022; Cattan et al., 2022; Tobaben et al., 2023) transfer-learning significantly closes the gap between non-private and private training. Tramèr et al. (2022) critically discuss this line of work and the implication of relying on pre-trained models.
Appendix B Training details
B.1 Parameterization
We utilize pre-trained feature extractors BiT-M-R50x1 (R-50) (Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B) (Dosovitskiy et al., 2021) with 85.8M parameters, both pretrained on the ImageNet-21K data set (Russakovsky et al., 2015). We download the feature extractor checkpoints from the respective repositories.
Following Tobaben et al. (2023) that show the favorable trade-off of parameter-efficient fine-tuning between computational cost, utility and privacy even for small data sets, we only consider fine-tuning subsets of all feature extractor parameters. We consider the following configurations:
  • Head: We train a linear layer on top of the pre-trained feature extractor.
  • FiLM: In addition to the linear layer from Head, we fine-tune parameter-efficient FiLM (Perez et al., 2018) adapters scattered throughout the network. While a diverse set of adapters has been proposed, we utilize FiLM as it has been shown to be competitive in prior work (Shysheya et al., 2023; Tobaben et al., 2023).
B.2 Hyperparameter tuning
Our hyperparameter tuning is heavily inspired by the comprehensive few-shot experiments by Tobaben et al. (2023). We utilize their hyperparameter tuning protocol as it has been proven to yield SOTA results for (DP) few-shot models.
[…]
Table A1: Hyperparameter ranges used for the Bayesian optimization with Optuna.
lower bound upper bound
batch size 10 $|\mathcal{D}|$
clipping norm 0.2 10
epochs 1 200
learning rate 1e-7 1e-2
noise multiplier Based on target $\epsilon$
B.3 Datasets
Table A2 shows the used data sets in the paper. We base our experiments on a subset of the the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by a practitioner. Additionally, we add CIFAR10 that is not part of the original VTAB benchmark.
Table A2: Used datasets in the paper, their minimum and maximum shots $S$ and maximum number of classes $C$ and their test accuracy when fine-tuning a non-DP ViT-B Head. The test accuracy for EuroSAT and Resics45 is computed on the part of the training split that is not used for training the particular model due to both datasets missing an official test split. Note that LiRA requires $2S$ for training the shadow models and thus $S$ is smaller than when only performing fine-tuning.
dataset (max.) min. max. test accuracy test accuracy
$C$ $S$ $S$ (min. $S$) (max. $S$)
Patch Camelyon (Veeling et al., 2018) 2 256 65536 82.8% 85.6%
CIFAR10 (Krizhevsky, 2009) 10 8 2048 92.7% 97.7%
EuroSAT (Helber et al., 2019) 10 8 512 80.2% 96.7%
Pets (Parkhi et al., 2012) 37 8 32 82.3% 90.7%
Resics45 (Cheng et al., 2017) 45 32 256 83.5% 91.6%
CIFAR100 (Krizhevsky, 2009) 100 16 128 82.2% 87.6%
Appendix C Additional results
In this section, we provide tabular results for our experiments and additional figures that did not fit into the main paper.
(a) tpr at $\textsc{fpr}=0.1$
(b) tpr at $\textsc{fpr}=0.01$
(c) tpr at $\textsc{fpr}=0.001$
Figure A.1: MIA vulnerability as a function of privacy budget ($\epsilon$ at $\delta=10^{-5}$) when attacking different backbones fine-tuned on CIFAR10 and CIFAR100 at $S\in\{25,50\}$. The solid line displays the median and the error bars the minimum of the lower bounds and the maximum of the upper bounds for the Clopper-Pearson CIs over 15 seeds. The DP bounds are the theoretical DP bounds based on Kairouz et al. (2015) where the naive DP bound is computed with $\delta=10^{-5}$ and the tight DP bound over an interval of deltas and their respective $\epsilon$.
C.1 Additional results for Section 5
This section contains additional results for Section 5.
C.1.1 Vulnerability as a function of shots
This section displays additional results to Figure 1(a) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A3 and A4.
(a) ViT-B Head tpr at $\textsc{fpr}=0.1$
(b) R-50 Head tpr at $\textsc{fpr}=0.1$
(c) ViT-B Head tpr at $\textsc{fpr}=0.01$
(d) R-50 Head tpr at $\textsc{fpr}=0.01$
(e) ViT-B Head tpr at $\textsc{fpr}=0.001$
(f) R-50 Head tpr at $\textsc{fpr}=0.001$
Figure A.2: MIA vulnerability as a function of shots (examples per class) when attacking a pre-trained ViT-B and R-50 Head trained without DP on different downstream data sets. The errorbars display the min and max Clopper-Pearson CIs over 6 seeds and the solid line the median.
Table A3: Median MIA vulnerability over six seeds as a function of $S$ (shots) when attacking a Head trained without DP on-top of a ViT-B. The ViT-B is pre-trained on ImageNet-21k.
dataset classes ($C$) shots ($S$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 2 256 0.266 0.086 0.032 512 0.223 0.059 0.018 1024 0.191 0.050 0.015 2048 0.164 0.037 0.009 4096 0.144 0.028 0.007 8192 0.128 0.021 0.005 16384 0.118 0.017 0.003 32768 0.109 0.014 0.002 65536 0.105 0.012 0.002 CIFAR10 (Krizhevsky, 2009) 10 8 0.910 0.660 0.460 16 0.717 0.367 0.201 32 0.619 0.306 0.137 64 0.345 0.132 0.067 128 0.322 0.151 0.082 256 0.227 0.096 0.054 512 0.190 0.068 0.032 1024 0.168 0.056 0.025 2048 0.148 0.039 0.013 EuroSAT (Helber et al., 2019) 10 8 0.921 0.609 0.408 16 0.738 0.420 0.234 32 0.475 0.222 0.113 64 0.400 0.159 0.074 128 0.331 0.155 0.084 256 0.259 0.104 0.049 512 0.213 0.080 0.037 Pets (Parkhi et al., 2012) 37 8 0.648 0.343 0.160 16 0.745 0.439 0.259 32 0.599 0.311 0.150 Resics45 (Cheng et al., 2017) 45 32 0.672 0.425 0.267 64 0.531 0.295 0.168 128 0.419 0.212 0.115 256 0.323 0.146 0.072 CIFAR100 (Krizhevsky, 2009) 100 16 0.814 0.508 0.324 32 0.683 0.445 0.290 64 0.538 0.302 0.193 128 0.433 0.208 0.114
Table A4: Median MIA vulnerability over six seeds as a function of $S$ (shots) when attacking a Head trained without DP on-top of a R-50. The R-50 is pre-trained on ImageNet-21k.
dataset classes ($C$) shots ($S$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 2 256 0.272 0.076 0.022 512 0.195 0.045 0.011 1024 0.201 0.048 0.011 2048 0.178 0.041 0.010 4096 0.163 0.033 0.008 8192 0.143 0.026 0.006 16384 0.124 0.019 0.004 32768 0.118 0.016 0.003 65536 0.106 0.012 0.002 CIFAR10 (Krizhevsky, 2009) 10 8 0.911 0.574 0.324 16 0.844 0.526 0.312 32 0.617 0.334 0.183 64 0.444 0.208 0.106 128 0.334 0.159 0.084 256 0.313 0.154 0.086 512 0.251 0.103 0.051 1024 0.214 0.082 0.038 EuroSAT (Helber et al., 2019) 10 8 0.846 0.517 0.275 16 0.699 0.408 0.250 32 0.490 0.236 0.121 64 0.410 0.198 0.105 128 0.332 0.151 0.075 256 0.269 0.111 0.056 512 0.208 0.077 0.036 Pets (Parkhi et al., 2012) 37 8 0.937 0.631 0.366 16 0.745 0.427 0.227 32 0.588 0.321 0.173 Resics45 (Cheng et al., 2017) 45 32 0.671 0.405 0.235 64 0.534 0.289 0.155 128 0.445 0.231 0.121 256 0.367 0.177 0.088 CIFAR100 (Krizhevsky, 2009) 100 16 0.897 0.638 0.429 32 0.763 0.549 0.384 64 0.634 0.414 0.269
C.1.2 Vulnerability as a function of the number of classes
This section displays additional results to Figure 1(b) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.3, A5 and A6.
(a) ViT-B Head tpr at $\textsc{fpr}=0.1$
(b) R-50 Head tpr at $\textsc{fpr}=0.1$
(c) ViT-B Head tpr at $\textsc{fpr}=0.01$
(d) R-50 Head tpr at $\textsc{fpr}=0.01$
(e) ViT-B Head tpr at $\textsc{fpr}=0.001$
(f) R-50 Head tpr at $\textsc{fpr}=0.001$
Figure A.3: MIA vulnerability as a function of $C$ (classes) when attacking a ViT-B and R-50 Head fine-tuned without DP on different data sets where the classes are randomly sub-sampled and $S=32$. The solid line displays the median and the errorbars the min and max clopper-pearson CIs over 12 seeds.
Table A5: Median MIA vulnerability over 12 seeds as a function of $C$ (classes) when attacking a Head trained without DP on-top of a ViT-B. The Vit-B is pre-trained on ImageNet-21k.
dataset shots ($S$) classes ($C$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 32 2 0.467 0.192 0.080 CIFAR10 (Krizhevsky, 2009) 32 2 0.494 0.167 0.071 4 0.527 0.217 0.115 8 0.574 0.262 0.123 EuroSAT (Helber et al., 2019) 32 2 0.306 0.100 0.039 4 0.298 0.111 0.047 8 0.468 0.211 0.103 Pets (Parkhi et al., 2012) 32 2 0.232 0.045 0.007 4 0.324 0.092 0.035 8 0.296 0.094 0.035 16 0.406 0.158 0.069 32 0.553 0.269 0.136 Resics45 (Cheng et al., 2017) 32 2 0.272 0.084 0.043 4 0.322 0.119 0.056 8 0.496 0.253 0.148 16 0.456 0.204 0.108 32 0.580 0.332 0.195 CIFAR100 (Krizhevsky, 2009) 32 2 0.334 0.088 0.035 4 0.445 0.150 0.061 8 0.491 0.223 0.121 16 0.525 0.256 0.118 32 0.553 0.276 0.153 64 0.612 0.350 0.211
Table A6: Median MIA vulnerability over 12 seeds as a function of $C$ (classes) when attacking a Head trained without DP on-top of a R-50. The R-50 is pre-trained on ImageNet-21k.
dataset shots ($S$) classes ($C$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 32 2 0.452 0.151 0.041 CIFAR10 (Krizhevsky, 2009) 32 2 0.404 0.146 0.060 4 0.560 0.266 0.123 8 0.591 0.318 0.187 EuroSAT (Helber et al., 2019) 32 2 0.309 0.111 0.050 4 0.356 0.144 0.064 8 0.480 0.233 0.123 Pets (Parkhi et al., 2012) 32 2 0.249 0.068 0.029 4 0.326 0.115 0.056 8 0.419 0.173 0.075 16 0.493 0.245 0.127 32 0.559 0.294 0.166 Resics45 (Cheng et al., 2017) 32 2 0.310 0.103 0.059 4 0.415 0.170 0.083 8 0.510 0.236 0.119 16 0.585 0.311 0.174 32 0.644 0.382 0.218 CIFAR100 (Krizhevsky, 2009) 32 2 0.356 0.132 0.054 4 0.423 0.176 0.087 8 0.545 0.288 0.163 16 0.580 0.338 0.196 32 0.648 0.402 0.244 64 0.711 0.476 0.320
C.1.3 Predicting data set vulnerability as function of $S$ and $C$
Table A7: Results for fitting Equation 3 with statsmodels Seabold & Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.924 0.898 -0.237 0.008 -30.994 $0$0.252 -0.222 0.01 0.937 0.839 -0.380 0.012 -31.818 $0$0.403 -0.356 0.001 0.917 0.740 -0.492 0.019 -26.143 $0$0.530 -0.455 $\beta_{C}$ (for $C$) 0.1 0.924 0.898 0.095 0.014 6.992 0.000 0.068 0.121 0.01 0.937 0.839 0.226 0.021 10.680 0.000 0.184 0.268 0.001 0.917 0.740 0.357 0.033 10.677 0.000 0.291 0.423 $\beta_{0}$ (intercept) 0.1 0.924 0.898 -0.008 0.029 -0.274 0.785 -0.064 0.049 0.01 0.937 0.839 -0.224 0.045 -5.006 $0$0.312 -0.136 0.001 0.917 0.740 -0.486 0.071 -6.879 $0$0.626 -0.347
(a) tpr at $\textsc{fpr}=0.1$
(b) tpr at $\textsc{fpr}=0.01$
(c) tpr at $\textsc{fpr}=0.001$
Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation 3 fitted Table A3 (ViT-B). The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds. Note that the dots for $C=10$ are for EuroSAT.

### v2

@@
-Understanding Practical Membership Privacy of Deep Learning
+On the Impact of Dataset Properties on Membership Privacy of Deep Learning
 Abstract
@@
 In terms of data set properties, we find a strong power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate.
-For an individual sample, large gradients at the end of training are strongly correlated with MIA vulnerability.
+We train a linear model to predict true positive rate based on data set properties and observe good fit for MIA vulnerability on unseen data.
+To analyse the phenomenon theoretically, we reproduce the result on a simplified model of membership inference that behaves similarly to our experimental data.
+We prove that in this model, the logarithm of the difference of true and false positive rates depends linearly on the logarithm of the number of examples per class.
+For an individual sample, the gradient norm is predictive of its vulnerability.
 1 Introduction
@@
 Balle et al., 2022; Nasr et al., 2023).
-Differential privacy (DP; Dwork et al., 2006b) provides protection against these attacks, but strong formal protection often comes at the cost of significant loss of model utility.
+Differential privacy (DP; Dwork et al., 2006) provides protection against these attacks, but strong formal protection often comes at the cost of significant loss of model utility.
 Finding the correct balance between making models resistant to attacks while maintaining a high utility is important for many applications.
@@
 they must not contain information that can be linked to an identifiable individual.
-On the other hand, loss of utility of the model may compromise the health benefits that might be gained from it.
-In this paper, our aim is to apply a state-of-the-art MIA systematically to help understand practical privacy risks training deep-learning-based classifiers without DP protections.
+On the other hand, loss of utility of the model due to privacy constraints may compromise the health benefits that might be gained from it.
+In this paper, our aim is to apply a state-of-the-art MIA systematically, to help understand practical privacy risks when training deep-learning-based classifiers without DP protections.
 Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep learning models to MIA.
@@
 Based on extensive experimental data over many data sets with varying sizes, we uncover a power law describing the vulnerability remarkably well.
-Previously Tobaben et al.
+Previously, Tobaben et al.
 (2023) and Yu et al.
@@
 Our work significantly expands on both of these by more systematic study and modelling of data set properties and by providing a more detailed look at individual samples, not just classes.
-List of contributions
-In this work we perform an extensive empirical study on the MIA vulnerability of deep learning models.
-We will focus on transfer learning setting on computer vision tasks, where a large pre-trained neural network is fine-tuned on a sensitive data set, and the MIA is performed on the final classifier.
-We find that the MIA vulnerability has a strong correlation with certain data set properties (Section 5) and make the following contributions:
+List of contributions In this work we perform an extensive empirical study on the MIA vulnerability of deep learning models.
+We focus on a transfer learning setting for image classification tasks, where a large pre-trained neural network is fine-tuned on a sensitive data set, and the MIA is performed on the final classifier.
+We find that the MIA vulnerability has a strong correlation with certain data set properties (Section 3).
+In Section 4 we analyse the reasons for these findings more thoroughly and we extend the analysis to individual samples in Section 5.
+The primary contributions of this work are:
   1. 1.
-    Power law: We uncover a power law between the number of examples per class and the vulnerability to MIA (tpr at fixed low fpr) based on extensive experimental data over many data sets with varying sizes (See Figure 1(a)).
+    Few-shot MIA: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) of target models trained using many different data sets with varying sizes and uncover a power law between the number of examples per class and the vulnerability to MIA (see Figure 1).
   2. 2.
-    Regression model: We utilize the observations to train a regression model to predict MIA vulnerability (tpr at fixed low fpr) based on examples per class ($S$) and number of classes ($C$) and show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor (See Figure 3).
-Additionally, in Section 6 we extend the analysis to individual samples and discover a correlation between vulnerability of individual samples and the gradient norms during training.
-We further study the correlation between sample’s vulnerability and the similarity to the other samples in the feature space.
-2 Background
-Notation for the properties of the training data set $\mathcal{\mathcal{D}}$:
-  • $C$ for the number of classes
-  • $S$ for shots (examples per class)
-  • $\mathcal{|\mathcal{D}|}$ for training data set size ($\mathcal{|\mathcal{D}|}=CS$)
+    Regression model: We utilise our empirical observations to train a regression model to predict MIA vulnerability ($\log(\textsc{tpr}-\textsc{fpr})$ at fixed low fpr) based on examples per class ($\log S$) and number of classes ($\log C$).
+We show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Figure 3).
+  3. 3.
+    Theoretical analysis using a simplified model of MIA: To analyse the phenomenon theoretically, we reproduce the result on a simplified model of membership inference that behaves similarly to our experimental data.
+In this model we can formally prove that $\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log S$ (see Section 4).
+  4. 4.
+    Individual sample MIA: We extend the study to individual samples and find: 1 The gradient norm is predictive of vulnerability of an individual sample.
+2 Gradient norms at different stages of learning have different impact, but how depends on the data set (see Section 5).
+2 Background and Methods
+Notation for the properties of the training data set $\mathcal{\mathcal{D}}$: 1 $C$for the number of classes 2 $S$for shots (examples per class) 3 $\mathcal{|\mathcal{D}|}$for training data set size ($\mathcal{|\mathcal{D}|}=CS$).
 We denote the number of MIA shadow models with $M$.
@@
 Based on this information, the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack but for which the attacker knows the training data set.
-LiRA exploits the observation that the value of the loss function used to train a model is often lower for examples that have been part of the training set than for examples that have not been.
+LiRA exploits the observation that the loss function value used to train a model is often lower for examples that were part of the training set compared to those that were not.
 For a target sample $x$, LiRA trains the shadow models: 1 with $x$ as a part of the training set ($x\in\mathcal{D}$) 2 and without $x$ in the training set ($x\notin\mathcal{D}$) .
-After training the shadow models, $x$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed, one for the losses of $x\in\mathcal{D}$ shadow models and one for the $x\notin\mathcal{D}$.
-Finally the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $x\in\mathcal{D}$ or $x\notin\mathcal{D}$.
+After training the shadow models, $x$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $x\in\mathcal{D}$ shadow models, and one for the $x\notin\mathcal{D}$.
+Finally, the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $x\in\mathcal{D}$ or $x\notin\mathcal{D}$.
 Carlini et al.
-(2022) proposed an optimization for performing LiRA for multiple models and points without training a computationally infeasible amount of shadow models.
-We utilize this optimization in our experiments.
-In Appendix A we provide background on Differential privacy (DP; Dwork et al., 2006b) and DP stochastic gradient descent (DP-SGD) that we use for training DP models.
-3 Methods
-In this Section, we introduce the preliminaries for evaluating MIA vulnerability.
-3.1 Measuring MIA vulnerability
-Using the likelihood-ratios from LiRA Carlini et al.
-(2022) as a score, we can build a binary classifier to predict whether sample belongs to the training data or not.
+(2022) proposed an optimisation for performing LiRA for multiple models and points without training a computationally infeasible number of shadow models.
+We utilise this optimisation.
+Measuring MIA vulnerability Using the likelihood-ratios from LiRA as a score, we can build a binary classifier to predict whether a sample belongs to the training data or not.
 The operational characteristics of such classifier can be used to measure the success of the MIA.
@@
 Identifying even a small number of samples with high confidence is considered harmful (Carlini et al., 2022) and thus we focus on the regions of small fpr.
-The tpr and fpr can be also connected to DP.
-Kairouz et al.
-(2015) have shown in Equation 1 that any classifier distinguishing the training samples based on the results of a DP algorithm has an upper bound for the tpr that depends on the privacy parameters $(\epsilon,\delta)$.
-Since MIA is trying to build exactly these types of classifiers, we can use the upper bounds by Kairouz et al.
-(2015) to validate the privacy claims, and also to better understand the gap between the theoretical privacy guarantees and the realistic attacks.
-Theorem 1 (Kairouz et al. (2015)).
-A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP if and only if for all adjacent $\mathcal{D}\sim\mathcal{D}^{\prime}$
-$\mathrm{TPR}\leq\min\{e^{\epsilon}\mathrm{FPR}+\delta,1-e^{-\epsilon}(1-\delta-\mathrm{FPR})\}\ .$ (1)
-3.2 Measuring the uncertainty for TPR
-The tpr values from the LiRA based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples.
-Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimator.
-Therefore, when we report the tpr values for single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals Clopper & Pearson (1934).
+Measuring the uncertainty for tpr The tpr values from the LiRA-based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples.
+Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimators.
+Therefore, when we report the tpr values for a single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals Clopper and Pearson (1934).
 Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
@@
 where $B(q;a,b)$ is the $q$th-quantile of $\text{Beta}(a,b)$ distribution.
-4 Experiment overview
-Throughout the experiments, we utilize LiRA to understand the vulnerability of the trained models.
-We fine-tune pre-trained models on sensitive downstream data sets and assess the vulnerability using LiRA. We use the following setup for all experiments:
-  • Pre-trained models.
-BiT-M-R50x1 (R-50; Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021) with 85.8M parameters, both pre-trained on the ImageNet-21K data set (Russakovsky et al., 2015)
-  • Parameterization.
-Due to computational cost we only consider fine-tuning subsets of the model parameters.
-We consider the following configurations: 1 Head: training a linear last layer classifier on top of the pre-trained model 2 FiLM: Head and additionally training parameter-efficient FiLM (Perez et al., 2018) adapters scattered throughout the network (see Section B.1)
-  • Fine-tuning data sets. We utilize different fine-tuning data sets (see Section B.3 for details) with different $S$, $C$ and $|\mathcal{D}|$ as specified in the experiments.
-  • Hyperparameter optimization (HPO).
-We utilize the library Optuna Akiba et al.
-(2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations.
-We optimize the batch size, learning rate and number of epochs.
-Additionally for DP-SGD we optimize the clipping norm.
-Details in Section B.2.
-  • DP-SGD. We are using Opacus Yousefpour et al. (2021) and compute the noise multiplier based on the target ($\epsilon,\delta$).
-  • DP Accounting.
-We use numerical privacy accounting for accurate evaluation of the cumulative privacy loss of the optimisation (Koskela et al., 2020, 2021) using the PRV accountant (Gopi et al., 2021).
-  • LiRA. We base our experiments on the publically available source code of Carlini et al. (2022). We utilize $M=256$ shadow models unless specified otherwise.
-We illustrate the effectiveness of LiRA and the effect of different levels of DP on the MIA vulnerability (tpr at $\textsc{fpr}=0.001$) in Figure 1.
-It shows the MIA vulnerability as a function of privacy budget (additional versions for different fprs in Figure A.1) and two upper bounds on the tpr at fpr emerging from the theoretical guarantees of DP-SGD.
-Figure 1: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of privacy budget ($\epsilon$ at $\delta=10^{-5}$) when attacking different backbones fine-tuned on CIFAR10 and CIFAR100 at $S\in\{25,50\}$.
-The solid line displays the median and the error bars the minimum of the lower bounds and the maximum of the upper bounds for the Clopper-Pearson CIs over 15 seeds.
-The upper bound is the theoretical DP bound based on Kairouz et al.
+3 MIA Vulnerability and Data Set Properties
+In this section, we investigate how different properties of data sets affect the MIA vulnerability.
+Based on our observations, we propose a method to predict the vulnerability to MIA using these properties.
+3.1 Experimental Setup
+We focus on a image classification setting where we fine-tune pre-trained models on sensitive downstream data sets and assess the MIA vulnerability using LiRA with $M=256$ shadow models.
+We base our experiments on a subset of the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a test classification accuracy $>80\%$ (see Table A2).
+We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k Russakovsky et al.
 (2015).
-We present the bound computed with $\delta=10^{-5}$ as well as the optimal bound based on multiple $\delta$.
-We compute the upper bounds using the method of Kairouz et al.
-(2015) in Equation 1.
-Simply using $\delta=10^{-5}$ and the associated $\epsilon$ gives a relatively weak bound, that can be sharpened considerably by finding an optimal $\delta$ and corresponding $\epsilon(\delta)$, which is made possible by the particular mechanism deployed, that gives the tightest possible bound.
-For the attacked few-shot models the bounds are quite tight for the medium privacy budget of $\epsilon=1$ but very loose at low privacy budgets of $\epsilon=8$ and thus meaningless for low privacy levels.
-5 Predicting MIA vulnerability for data sets
-In this Section, we investigate how different properties of data sets affect the MIA vulnerability in the non-DP setting.
-Based on our observations, we propose a method to predict the vulnerability to MIA using these properties of the data set.
-We base our experiments on a subset of the few-shot benchmark VTAB  (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by practitioners.
-For more detailed descriptions of the data sets, see Table A2.
-In this Section, we will focus on a fine-tuned last layer classifier (Head) trained on top of a ViT-B, pre-trained on ImageNet-21k Russakovsky et al.
-(2015).
-The results for using R-50 as a backbone can be found in Section C.1.
-5.1 Correlations with data set properties
-Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 3.1). We make the following observations:
-  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in Figures A.2, A3 and A4 in the Appendix.
+The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section B.1.
+We optimise the hyperparameters (batch size, learning rate and number of epochs) using the library Optuna Akiba et al.
+(2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section A.2).
+We provide the the code for reproducing the experiments in the supplementary material.
+3.2 Experimental Results
+Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 2). We make the following observations:
+  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in the Appendix (Figures A.1, A3 and A4).
   • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 1(b) with further evidence in Figures A.2, A5 and A6 in the Appendix.
@@
 (a) as function of $S$ (shots)
-(b) as function of $C$ (classes)
+(b) as function of $C$ (classes) at $S=32$
 Figure 1: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of data set properities when attacking a ViT-B Head fine-tuned without DP on different data sets.
-The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over multiple seeds (6 for Figure 1(a) and 12 for Figure 1(b))
+The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over multiple seeds (six for Figure 1(a) and 12 for Figure 1(b))
 3.3 Model to Predict Data Set Vulnerability
-Based on the trends seen in Figures 1(b) and 1(a), we fit a linear regression model using the logarithmic transform of $C,S$ to predict the tpr (also logarithmically transformed) with statsmodels Seabold & Perktold (2010).
-The general form of the model can be found in Equation 3 where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
-$\log_{10}(\textsc{tpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (3)
-We utilize MIA results of ViT-B (see Table A3) as the training data and use R-50 (see Table A4) as the test data to investigate if the vulnerability prediction model generalizes to a completely different backbone.
-The parameters of the prediction model fitted to the training data (note that the parameters are specific to $\textsc{fpr}=0.001$) are $\beta_{S}=-0.492,\beta_{C}=0.357$ and $\beta_{0}=-0.486$.
-Based on the $R^{2}$ score ($R^{2}=0.917$), our model fits the data extremely well.
-Furthermore, Figure 3 shows that the model is robust to a change of feature extractor, as it is able to predict the tpr for R-50 despite being trained on ViT-B (test $R^{2}=0.740$).
-See Section C.1.3 for further supporting results at $\textsc{fpr}\in\{0.1,0.01\}$.
-Figure 3: Predicted MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of $S$ (shots) using a model based on Equation 3 fitted Table A3 (ViT-B).
-The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100).
+[…]
+$\log_{10}(\textsc{tpr}-\textsc{fpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (2)
+[…]
+We utisee MIA results of ViT-B (see Table A3) as the training data.
+Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well.
+We provide further evidence for other fpr in Figure A.3 and Table A8 in the Appendix.
+Figure 2 shows the parameters of the prediction model fitted to the training data.
+Figure 2: Coefficient values for different fpr when fitting a regression model based on Equation 2 fitted on data from ViT-B (Head) (Table A3).
+The error bars display the 95% confidence intervals based on Student’s t-distribution.
+Prediction quality on other MIA target models
+We analyse how the regression model trained on the ViT-B (Head) data generalizes to other target models. The main points are:
+  • R-50 (Head): Figure 3(a) shows that the regression model is robust to a change of the feature extractor, as it is able to predict the tpr for R-50 (Head) (test $R^{2}=0.790$).
+  • R-50 (FiLM): Figure 3(b) shows that the prediction quality is good for R-50 (FiLM) models.
+These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section A.1).
+Tobaben et al.
+(2023) demonstrated that FiLM layers are a competitive alternative to training all parameters.
+We supplement the MIA results of Tobaben et al.
+(2023) with own FiLM training runs.
+Refer to Table A7 in the Appendix.
+  • From-Scratch-Training: Carlini et al.
+(2022) provide limited results on from-scratch-training.
+To the best of our knowledge these are the only published LiRA results on image classification models.
+Figure 3(b) displays that our prediction model underestimates the vulnerability of the from-scratch trained target models.
+We have identified two potential explanations for this 1 In from-scratch-training all weights of the model need to be trained from the sensitive data and thus potentially from-scratch-training could be more vulnerable than fine-tuning.
+2 The strongest attack in Carlini et al.
+(2022) uses data augmentations to improve the performance.
+We are not using this optimization.
+[…]
+(b) Evaluating prediction performance for target models 1 left: fine-tuned with FiLM (data partially from Tobaben et al.
+(2023), see Table A7) 2 right: trained from scratch (data from Carlini et al.
+(2022))
+Figure 3: Performance of the regression model based on Equation 2 fitted on data from Table A3.
+4 Understanding the MIA Prediction Model
+In this section, we focus on understanding why the previously observed trends exist. Our goal is to formally prove a similar relationship as in the predictive model of Section 3.3.
+4.1 A Simplified Model of Membership Inference
+In this section, we construct a simplified model of membership inference that streamlines the data generation and shadow model training.
+We observe that the trends of the vulnerability of the simplified model is a good approximation of the trends previously observed with the image classification example.
+Instead of training linear classifiers to classify images, we sample vectors on a high dimensional unit sphere and classify based on inner product with estimated class mean.
+This model is easier to analyse theoretically, and allows us to simulate arbitrary combinations of $S$ and $C$ without relying on the existence of a specific image classification downstream data set as especially high $C$ image classification downstream data sets that are computationally feasible to fine-tune on have small $S$.
+We generate the data and form the classifiers (which are our target models) as follows:
+  1. 1.
+    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in C:m_{i}\perp m_{j}\lor i=j$).
+  2. 2.
+    We sample $2S$ vectors $x_{c}$ for each class.
+We assume that they are Gaussian distributed around the the true class mean $x_{c}\sim\mathcal{N}(m_{c},s^{2}I)$ where the $s^{2}$ is the in-class variance.
+  3. 3.
+    For each "target model" we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $r_{c}$.
+  4. 4.
+    The computed mean is used to classify sample $(x,y)$ by computing the inner product $\langle x,r_{c}\rangle$ as a metric of similarity.
+The attacker has to infer which vectors have been used for training the classifier.
+Instead of utilising the logits (like in the image classification example), the attacker can use the inner products of a point with the cluster means.
+In Figure 4 we compare the attacker’s performance as a function of $C$ and $S$ to the the previous image classification example.
+This simplified model resembles the linear (Head) classifier.
+In the linear classifier, we find a matrix $A$ and biases $b$, to optimize the cross-entropy between the labels and logits $Az+b$, where $z$ denotes the feature space representation of the data.
+In the simplified model, the rows of $A$ are replaced by the cluster means and we do not include the bias term in the classification.
+(a) as a function of $S$ (shots)
+(b) as a function of $C$ (classes)
+[…]
+4.2 Theoretical Analysis
+Similarly to the setting in Section 4.1, we focus on a classification model that aims to classify the sample $(x,y)$ based on the inner product $\langle x,r_{c}\rangle$, where $r_{c}$ is the mean of the training data points $x_{m}$ with $y_{m}=c$.
+However, we only focus on the scores of the target class, i.e.
+$c=y$, and therefore the score used of LiRA simplifies to $\langle x,r_{y}\rangle$.
+The fpr for LiRA is given as
+$\textsc{fpr}=\Pr\left(\frac{\mathcal{N}(t_{x};\mu_{\textrm{in}},\sigma_{\textrm{in}}^{2})}{\mathcal{N}(t_{x};\mu_{\textrm{out}},\sigma_{\textrm{out}}^{2})}>\exp(\tau)\mid x\notin\mathcal{D}_{\textrm{target}}\right),$ (3)
+where $t_{x}=\langle x,r_{y}^{\textrm{(target)}}\rangle$ is the score for the target model, $\tau$ is the decision threshold for the classifier, and $\mu_{\textrm{in}},\mu_{\textrm{out}},\sigma_{\textrm{in}}$ and $\sigma_{\textrm{out}}$ are the LiRA statistics computed from the shadow models.
+We assume that $\sigma_{\textrm{in}}=\sigma_{\textrm{out}}=\sigma$, which allows to write Equation 3 as
+fpr $=\Pr\left(\log\frac{\mathcal{N}(t_{x};\mu_{\textrm{in}},\sigma^{2})}{\mathcal{N}(t_{x};\mu_{\textrm{out}},\sigma^{2})}>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right)$ (4)
+$=\Pr\left(-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right).$ (5)
+From now on, we will denote the $-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]$ with $\lambda_{\textsc{fpr}}$ when $x\notin\mathcal{D}_{\textrm{target}}$ and with $\lambda_{\textsc{tpr}}$ otherwise.
+Furthermore we will drop the condition $x\in\mathcal{D}_{\textrm{target}}$ or $x\notin\mathcal{D}_{\textrm{target}}$ from the fpr and tpr expressions
+fpr $:=\Pr(\lambda_{\textsc{fpr}}>\tau)$ (6)
+tpr $:=\Pr(\lambda_{\textsc{tpr}}>\tau).$ (7)
+Lemma 1.
+In the limit of infinite shadow models and with large enough $S$, the LiRA test statistics $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ obtain the following forms
+$\lambda_{\textsc{fpr}}$ $=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right)$ (8)
+$\lambda_{\textsc{tpr}}$ $\approx-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle-\frac{2}{S}||x||^{2}+\frac{1}{S}\langle x,x-m_{y}\rangle\right),$ (9)
+where $z\sim\mathcal{N}(0,I_{d})$.
+Proof.
+We defer the proof to Appendix C. ∎
+The factors independent of $S$ in for both $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ correspond to the $-(\mu_{\textrm{in}}-\mu_{\textrm{out}})/(2\sigma^{2})$.
+Figure A.5 illustrates how this factor behaves as a function of $S$ in our experiments with the simple model.
+As predicted by the theory, the factor (i.e.
+the difference in means divided by the variance) remains fairly constant across different values of $S$.
+Using Lemma 1 we can derive a rate for the decision threshold $\tau$:
+Corollary 2.
+In order to maintain constant fpr, we need to adjust $\tau$ as $O(1/\sqrt{S})$.
+According to Corollary 2, the $\tau:=\tau(S)$ evolves as $O(1/\sqrt{S})$, and we will approximate it as $\tau(S)=\gamma_{1}\frac{1}{\sqrt{S}}$, with constant $\gamma_{1}$ depending on the target fpr.
+Theorem 3.
+$\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log(S)$.
+Proof.
+From Lemma 1, we see that $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ differ only in a single term which evolves as $O(1/S)$.
+Thus for sufficiently large $S$ we can consider this additional r.v.
+as a small constant shift to the distribution of $\lambda_{\textsc{fpr}}$.
+Let us denote this variable now with $q$.
+When $S$ grows, $q$ becomes small, and we have
+[…]
+$\approx p_{\lambda_{\textsc{fpr}}}(\tau)q.$ (11)
+Taking the log yields
+$\log(\textsc{tpr}-\textsc{fpr})$ $\approx\log\left(p_{\lambda_{\textsc{fpr}}}\left(\tau\right)\right)+\log q$ (12)
+$=\log\left(p_{\lambda_{\textsc{fpr}}}\left(\tau\right)\right)-\log S+\gamma_{2}.$ (13)
+From Lemma 1, we can see that a scaled r.v.
+$\hat{\lambda}_{\textsc{fpr}}:=\sqrt{S}\lambda_{\textsc{fpr}}$ becomes almost independent of $S$ for large enough $S$.
+Hence using the change of variables formula, we have that $p_{\lambda_{\textsc{fpr}}}(t)=p_{\hat{\lambda}_{\textsc{fpr}}}(\sqrt{S}t)\sqrt{S}$.
+Now, evaluating $p_{\lambda_{\textsc{fpr}}}(t)$ at $t=\tau(S)$, we get $p_{\lambda_{\textsc{fpr}}}(t)=p_{\hat{\lambda}_{\textsc{fpr}}}(\gamma_{1})\sqrt{S}$, where the $p_{\hat{\lambda}_{\textsc{fpr}}}(\gamma_{1})$ is independent of $S$.
+Therefore, ignoring additive constants,
+$\log(\textsc{tpr}-\textsc{fpr})$ $\approx\log\sqrt{S}-\log S=\frac{1}{2}\log S-\log S=-\frac{1}{2}\log S.$ (14)
+This shows that $\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log S$. ∎
+Compared to the model in Section 3.3, our theoretical model predicts $\beta_{S}=-1/2$.
+Our empirical estimates match this for high fpr.
+For lower FPR, $\beta_{S}$ decreases, which indicates there may be finer details of $p_{\lambda_{\textsc{fpr}}}$ we are missing in this analysis.
 5 MIA Vulnerability For Data Samples
-This Section extends our investigation from the data set level properties studied in Section 5 to individual data sample level properties.
+This section extends our investigation from the data-set-level properties (Section 3) to individual data-sample-level properties.
 Previously Yu et al.
 (2023) have studied correlations between classes and privacy risk before, but in comparison to their study, we look at data samples independent from classes.
-We will focus on a fine-tuned last layer classifier (Head) trained on top of a ViT-B, pre-trained on ImageNet-21k Russakovsky et al.
-(2015).
-We investigate two data sample properties that could potentially influence the vulnerability (tpr at fpr) of data points:
-  • Per-sample gradients: Per-sample gradients form the basis for DP individual accounting Feldman & Zrnic (2021); Koskela et al.
-(2023).
-In individual accounting a larger sum of gradient norms results in a higher individual privacy loss.
-Our hypothesis is that this holds also for MIA vulnerability in non-DP settings.
-  • Vector space distances: The pre-trained feature extractors project images into a vector space in which one can compute distances between individual data samples.
+We investigate two properties that could potentially influence the vulnerability of data points:
+  • Per-sample gradients: Per-sample gradients form the basis for DP individual accounting (Feldman and Zrnic, 2021; Koskela et al., 2023).
+Under individual Gaussian DP (GDP, Dong et al., 2022) accounting (Koskela et al., 2023), the cumulative $\mu$-GDP privacy loss of individual $i$ with individual gradient norms $g_{i,t}$ is $\mu_{\text{i,tot}}\propto\sqrt{\sum_{t}g_{i,t}^{2}}$.
+Our hypothesis is that this similarly holds also for MIA vulnerability in non-DP settings.
+  • Feature space distances: The pre-trained feature extractors project images into a feature space in which one can compute distances between individual data samples.
 Our hypothesis is that data samples that are further away from other data samples are easier to distinguish.
-We study individual samples, but aggregate the results over multiple data samples by binning the samples based on their data sample properties to reduce the noise.
+5.1 Experimental Setup
+Our model training is done as in Section 3.1.
+The target models are fine-tuned pre-trained ViT-B (Head) and R-50 (Head) models.
+Below we describe how we measure the sample properties and individual tpr at fpr.
+Recording of gradient norms During the training of the target models we record the per-sample gradient norm for every sample at every step (independent of being part of the current batch) efficiently using Opacus (Yousefpour et al., 2021).
+The recording of the gradient norms is identical to the DP individual accounting, with the exception that we record the gradient norms of out-samples as well.
+In our analysis, we are utilising the square root of the sum of squared gradient norm per sample over deciles of steps (i.e.
+the sum of squared individual gradient norms over the first 10% of steps, second 10% of steps, …, last 10% of steps).
+$G_{i}^{(\alpha)}=\sqrt{\sum_{t=\lfloor T(\alpha-0.1)\rfloor}^{\lfloor T\alpha\rfloor}g_{i,t}^{2}},\;\alpha=(0.1,0.2,\ldots,1.0)$ (15)
+Computation of distance in feature space All samples are projected into a feature space using the pre-trained model.
+We measure the Euclidean distance of all examples (in and out) to all in-samples in the feature space and report the distance to the nearest in-sample.
+tpr at fpr for individual samples We study individual samples, but aggregate the results over multiple data samples by binning the samples based on their data sample properties to reduce the noise.
 For each data sample we compute its property and assign the sample to a bin based on its percentile in relation to other samples.
@@
 We finally report for each data property bin the percentile of the tpr at fpr.
-Figure 4: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of data sample properties (binned among a seed).
-The results have been obtained over 6 seeds with $S=256$ (Patch Camelyon) and $S=64$ (CIFAR10).
-Figure 4 shows the observed relationships between the MIA vulnerability and the average individual gradient norms and the vector space distance. We find that:
-  • Per-sample gradients: The average gradient norms are correlated with the vulnerability in all our experiments, but the strength of the correlation depends how many steps are taken into account.
-Generally, we observe that the correlation is stronger when we only consider the gradients at the end of the training (either the last 20 or 5 steps).
-This is different from the individual accounting methods that rely on the whole gradient trace for DP accounting.
-  • Vector space distances: The distance of data samples to their nearest neighbor correlates with the vulnerability for CIFAR10, but not for Patch Camelyon.
-We leave a more comprehensive study of this correlation and its limitations for future work.
+Throughout our experiments we noticed that the amount of data points needs to be sufficiently large, in order to allow for an appropriate estimation of the tpr at fpr.
+Thus, we use a sufficient amount of shots ($S$) so that the size of in- and out-set combined is at least $10002|\mathcal{D}|=$ 1000 (see Table A10).
+5.2 Individual Gradient Norms Are Predictive of Individual MIA Vulnerability
+Figure 5 shows the observed relationships between the MIA vulnerability and the average individual gradient norms at $\textsc{fpr}=0.001$.
+When plotting the data for all deciles of steps at once, we find that the average gradient norms are correlated with the vulnerability in all our experiments, but the strength of the correlation depends on the data set.
+The $R^{2}$ (coefficient of determination) compares the observed relationship to a scenario where individual gradient norms would be a perfect predictor (corresponding to a scenario where all points are on the diagonal line).
+See more results in Table A10 and Figure A.6 in the Appendix.
+Figure 5: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed).
+Data obtained over six seeds.
+The $R^{2}$ score compares the observed relationship to a scenario where individual gradient norms would be a perfect predictor.
+Which parts of the training are more predictive?
+In Figure 6, we examine whether the gradient norms at different stages of training are a better predictor of vulnerability by computing the ratio between the $R^{2}$ per iteration decile and $R^{2}$ over the whole training.
+For most data sets the ratio is constantly at a value of roughly 1, which suggests that gradient norms over even short segments can be as predictive as the norms of the whole trace.
+Furthermore, there are deviations that indicate that for some data sets, gradient norms during the early iterations can be relatively more predictive than the norms of the entire trace.
+This is interesting, as the DP bound would directly depend on the norms of the entire trace.
+Figure 6: Ratio between $R^{2}$ per iteration decile and $R^{2}$ over the whole training.
+A ratio $>1$ means that the decile is as more predictive than whole trace and a ratio $<1$ means that it is less predicitive.
+5.3 Distance to Neighbors in Feature Space of Pre-trained Model
+Figure 7 shows the observed relationships between the MIA vulnerability and distance to the nearest neighbor in the feature space.
+The distance of data samples to their nearest neighbor correlates with the vulnerability all analysed data sets but Patch Camelyon and Resics45.
+The observation holds over all fpr and for R-50 (Head) as well, refer to Figures A.9 and A.10 in the Appendix.
+Figure 7: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
 6 Discussion
-As DP reduces the utility of models, it is important to understand also when it is necessary.
-Our results on MIA vulnerability as a function of data set properties display a strong power law dependence on the number of examples per class.
-This leads to an unfortunate dilemma: DP seems most important in the small data regime where it likely reduces the model utility the most.
+Under the GDPR and similar legal regimes, machine learning (ML) models trained on personal data that memorise the data are personal data and need to be carefully protected.
+Our work analyses in which cases trained models would most likely be personal data and in which cases they might not be.
+This will help in evaluating the risk of different kinds of models, favouring less risky models when possible and paying extra attention to more risky cases.
+As the best means of protecting privacy, DP, reduces the utility of models, it is important to understand when it is necessary.
+Our results highlight that models are most vulnerable to MIA when the number of examples per class is low.
+This leads to an unfortunate dilemma: DP seems most important in the regime where it likely reduces the model utility the most.
+Our model singles out the number of examples per class $S$ as the key predictor of model vulnerability.
+While our results are based on balanced data sets, it seems likely that minority classes and outliers would be more vulnerable as predicted by our model given their smaller $S$.
+Studying the vulnerability in unbalanced data and in the presence of outliers is an important area of future work.
 Limitations
-Due to computational constraints, we only report results on transfer learning with fine-tuning only a part of the model.
+Our work measures the vulnerability through running one particular MIA.
+There is no guarantee that models that are not vulnerable to this attack might not be vulnerable to some stronger attack.
+Formal bounds on attack susceptibility would require something like DP.
+Due to computational constraints, our predictive model is based only on transfer learning with fine-tuning only a part of the model.
+Limited comparison of our model predictions to previous results indicates it may underestimate their vulnerability.
+Broader Impact
+Our work systematically studies factors influencing privacy risk of trained ML models.
+This has significant positive impact on users training ML models on personal data by allowing them to understand and limit the risks.
+Reproducibility
+We provide the code for reproducing the experiments in the supplementary material and will publish the code in an open repository after acceptance of the paper.
+All used pre-trained models and data sets are publicly available.
 Acknowledgements
@@
 Bronskill for helpful discussions regarding few-shot learning.
-Appendix A Background
-Differential privacy (DP; Dwork et al., 2006b) is the gold standard for formalizing privacy guarantees.
-$(\epsilon,\delta)$-DP (Dwork et al., 2006a) has a privacy budget consisting of $\epsilon\geq 0$ and $\delta\in[0,1]$, where smaller values correspond to a stronger privacy guarantee.
-We refer to Dwork & Roth (2014) for a comprehensive intro to DP.
-DP deep learning utilizes DP-SGD (Rajkumar & Agarwal, 2012; Song et al., 2013; Abadi et al., 2016), which is the DP adaptation of stochastic gradient descent.
-DP-SGD clips per-sample gradients to a pre-determined norm to limit the influence of each sample and adds Gaussian noise the sum of the clipped gradients.
-Clipping the gradients and adding Gaussian noise in DP-SGD typically comes with a cost in accuracy.
-This decrease in accuracy can be alleviated through transfer-learning.
-Publically available pre-trained models are fine-tuned with DP-SGD on sensitive downstream data sets.
-Both in NLP (Li et al., 2022; Yu et al., 2022) and computer vision (Kurakin et al., 2022; De et al., 2022; Mehta et al., 2022; Cattan et al., 2022; Tobaben et al., 2023) transfer-learning significantly closes the gap between non-private and private training.
-Tramèr et al.
-(2022) critically discuss this line of work and the implication of relying on pre-trained models.
-Appendix B Training details
-B.1 Parameterization
-We utilize pre-trained feature extractors BiT-M-R50x1 (R-50) (Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B) (Dosovitskiy et al., 2021) with 85.8M parameters, both pretrained on the ImageNet-21K data set (Russakovsky et al., 2015).
+Appendix A Training details
+A.1 Parameterization
+We utilise pre-trained feature extractors BiT-M-R50x1 (R-50) (Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B) (Dosovitskiy et al., 2021) with 85.8M parameters, both pretrained on the ImageNet-21K data set (Russakovsky et al., 2015).
 We download the feature extractor checkpoints from the respective repositories.
@@
   • FiLM: In addition to the linear layer from Head, we fine-tune parameter-efficient FiLM (Perez et al., 2018) adapters scattered throughout the network.
-While a diverse set of adapters has been proposed, we utilize FiLM as it has been shown to be competitive in prior work (Shysheya et al., 2023; Tobaben et al., 2023).
-B.2 Hyperparameter tuning
-Our hyperparameter tuning is heavily inspired by the comprehensive few-shot experiments by Tobaben et al.
-(2023).
-We utilize their hyperparameter tuning protocol as it has been proven to yield SOTA results for (DP) few-shot models.
+While a diverse set of adapters has been proposed, we utilise FiLM as it has been shown to be competitive in prior work (Shysheya et al., 2023; Tobaben et al., 2023).
+A.1.1 Licenses and Access
+The licenses and means to access the model checkpoints can be found below.
+A.2 Hyperparameter tuning
 […]
@@
 learning rate 1e-7 1e-2
-noise multiplier Based on target $\epsilon$
-B.3 Datasets
-Table A2 shows the used data sets in the paper.
+A.3 Data sets
+Table A2 shows the data sets used in the paper.
 We base our experiments on a subset of the the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by a practitioner.
-Additionally, we add CIFAR10 that is not part of the original VTAB benchmark.
+Additionally, we add CIFAR10 which is not part of the original VTAB benchmark.
 Table A2: Used datasets in the paper, their minimum and maximum shots $S$ and maximum number of classes $C$ and their test accuracy when fine-tuning a non-DP ViT-B Head.
@@
 CIFAR100 (Krizhevsky, 2009) 100 16 128 82.2% 87.6%
-Appendix C Additional results
+A.3.1 Licenses and Access
+The licenses and means to access the data sets can be found below. We downloaded all data sets from TensorFlow datasets https://www.tensorflow.org/datasets but Resics45 which required manual download.
+A.4 Compute Resources
+All experiments but the R-50 (FiLM) experiments are run on CPU with 8 cores and 16 GB of host memory.
+The training time depends on the model (ViT is cheaper than R-50), number of shots $S$ and the number of classes $C$ but ranges for the training of one model from some minutes to an hour.
+This assumes that the images are passed once through the pre-trained backbone and then cached as feature vectors.
+The provided code implements this optimization.
+In Section 5 we record the per-sample gradients during the training.
+This makes the training much more expensive despite the use of efficient per-sample gradient computation using Opacus Yousefpour et al.
+(2021).
+The training of the considered models does still not exceed one hour per model as we only consider settings with relatively low $S$ for this part of the experiments.
+The R-50 (FiLM) experiments are significantly more expensive and utilise a NVIDIA V100 with 40 GB VRAM, 10 CPU cores and 64 GB of host memory.
+The training of 257 shadow models then does not exceed 24h for the settings that we consider.
+We estimate that in total we spend around 7 days of V100 and some dozens of weeks of CPU core time but more exact measurements are hard to make.
+Appendix B Additional results
 In this section, we provide tabular results for our experiments and additional figures that did not fit into the main paper.
-(a) tpr at $\textsc{fpr}=0.1$
-(b) tpr at $\textsc{fpr}=0.01$
-(c) tpr at $\textsc{fpr}=0.001$
-Figure A.1: MIA vulnerability as a function of privacy budget ($\epsilon$ at $\delta=10^{-5}$) when attacking different backbones fine-tuned on CIFAR10 and CIFAR100 at $S\in\{25,50\}$.
-The solid line displays the median and the error bars the minimum of the lower bounds and the maximum of the upper bounds for the Clopper-Pearson CIs over 15 seeds.
-The DP bounds are the theoretical DP bounds based on Kairouz et al.
-(2015) where the naive DP bound is computed with $\delta=10^{-5}$ and the tight DP bound over an interval of deltas and their respective $\epsilon$.
-C.1 Additional results for Section 5
+B.1 Additional results for Section 3
 This Section contains additional results for Section 3.
-C.1.1 Vulnerability as a function of shots
+B.1.1 Vulnerability as a function of shots
 This section displays additional results to Figure 1(a) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.1, A3 and A4.
@@
 Figure A.1: MIA vulnerability as a function of shots (examples per class) when attacking a pre-trained ViT-B and R-50 Head trained without DP on different downstream data sets.
-The errorbars display the min and max Clopper-Pearson CIs over 6 seeds and the solid line the median.
+The errorbars display the minimum and maximum Clopper-Pearson CIs over six seeds and the solid line the median.
 Table A3: Median MIA vulnerability over six seeds as a function of $S$ (shots) when attacking a Head trained without DP on-top of a ViT-B. The ViT-B is pre-trained on ImageNet-21k.
@@
 dataset classes ($C$) shots ($S$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 2 256 0.272 0.076 0.022 512 0.195 0.045 0.011 1024 0.201 0.048 0.011 2048 0.178 0.041 0.010 4096 0.163 0.033 0.008 8192 0.143 0.026 0.006 16384 0.124 0.019 0.004 32768 0.118 0.016 0.003 65536 0.106 0.012 0.002 CIFAR10 (Krizhevsky, 2009) 10 8 0.911 0.574 0.324 16 0.844 0.526 0.312 32 0.617 0.334 0.183 64 0.444 0.208 0.106 128 0.334 0.159 0.084 256 0.313 0.154 0.086 512 0.251 0.103 0.051 1024 0.214 0.082 0.038 EuroSAT (Helber et al., 2019) 10 8 0.846 0.517 0.275 16 0.699 0.408 0.250 32 0.490 0.236 0.121 64 0.410 0.198 0.105 128 0.332 0.151 0.075 256 0.269 0.111 0.056 512 0.208 0.077 0.036 Pets (Parkhi et al., 2012) 37 8 0.937 0.631 0.366 16 0.745 0.427 0.227 32 0.588 0.321 0.173 Resics45 (Cheng et al., 2017) 45 32 0.671 0.405 0.235 64 0.534 0.289 0.155 128 0.445 0.231 0.121 256 0.367 0.177 0.088 CIFAR100 (Krizhevsky, 2009) 100 16 0.897 0.638 0.429 32 0.763 0.549 0.384 64 0.634 0.414 0.269
-C.1.2 Vulnerability as a function of the number of classes
+B.1.2 Vulnerability as a function of the number of classes
 This section displays additional results to Figure 1(b) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A5 and A6.
@@
 dataset shots ($S$) classes ($C$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 32 2 0.452 0.151 0.041 CIFAR10 (Krizhevsky, 2009) 32 2 0.404 0.146 0.060 4 0.560 0.266 0.123 8 0.591 0.318 0.187 EuroSAT (Helber et al., 2019) 32 2 0.309 0.111 0.050 4 0.356 0.144 0.064 8 0.480 0.233 0.123 Pets (Parkhi et al., 2012) 32 2 0.249 0.068 0.029 4 0.326 0.115 0.056 8 0.419 0.173 0.075 16 0.493 0.245 0.127 32 0.559 0.294 0.166 Resics45 (Cheng et al., 2017) 32 2 0.310 0.103 0.059 4 0.415 0.170 0.083 8 0.510 0.236 0.119 16 0.585 0.311 0.174 32 0.644 0.382 0.218 CIFAR100 (Krizhevsky, 2009) 32 2 0.356 0.132 0.054 4 0.423 0.176 0.087 8 0.545 0.288 0.163 16 0.580 0.338 0.196 32 0.648 0.402 0.244 64 0.711 0.476 0.320
-C.1.3 Predicting data set vulnerability as function of $S$ and $C$
-Table A7: Results for fitting Equation 3 with statsmodels Seabold & Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$.
+B.1.3 Data for FiLM and from scratch training
+Table A7: MIA vulnerability data used in Figure 3(b).
+Note that the data from Carlini et al.
+(2022) is only partially tabular, thus we estimated the tpr at fpr from the plots in the Appendix of their paper.
+model dataset classes shots source tpr@ tpr@ tpr@ ($C$) ($S$) fpr=0.1 fpr=0.01 fpr=0.001 R-50 FiLM CIFAR10 10 50 This work 0.482 0.275 0.165 (Krizhevsky, 2009) CIFAR100 100 10 Tobaben et al.
+(2023) 0.933 0.788 0.525 (Krizhevsky, 2009) 25 Tobaben et al.
+(2023) 0.766 0.576 0.449 50 Tobaben et al.
+(2023) 0.586 0.388 0.227 100 Tobaben et al.
+(2023) 0.448 0.202 0.077 EuroSAT 10 8 This work 0.791 0.388 0.144 (Helber et al., 2019) Patch Camelyon 2 256 This work 0.379 0.164 0.076 (Veeling et al., 2018) Pets 37 8 This work 0.956 0.665 0.378 (Parkhi et al., 2012) Resics45 45 32 This work 0.632 0.379 0.217 (Cheng et al., 2017) from scratch CIFAR10 10 2500 Carlini et al.
+(2022) 0.300 0.110 0.084 (Krizhevsky, 2009) (wide ResNet) CIFAR100 100 250 Carlini et al.
+(2022) 0.700 0.400 0.276 (Krizhevsky, 2009)
+B.1.4 Predicting data set vulnerability as function of $S$ and $C$
+This section provides additional results for the model based on Equation 2
+Table A8: Results for fitting Equation A1 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$.
 We utilize an ordinary least squares.
@@
 $[0.025$ coeff.
-$0.975]$ $\beta_{S}$ (for $S$) 0.1 0.924 0.898 -0.237 0.008 -30.994 $0$0.252 -0.222 0.01 0.937 0.839 -0.380 0.012 -31.818 $0$0.403 -0.356 0.001 0.917 0.740 -0.492 0.019 -26.143 $0$0.530 -0.455 $\beta_{C}$ (for $C$) 0.1 0.924 0.898 0.095 0.014 6.992 0.000 0.068 0.121 0.01 0.937 0.839 0.226 0.021 10.680 0.000 0.184 0.268 0.001 0.917 0.740 0.357 0.033 10.677 0.000 0.291 0.423 $\beta_{0}$ (intercept) 0.1 0.924 0.898 -0.008 0.029 -0.274 0.785 -0.064 0.049 0.01 0.937 0.839 -0.224 0.045 -5.006 $0$0.312 -0.136 0.001 0.917 0.740 -0.486 0.071 -6.879 $0$0.626 -0.347
+$0.975]$ $\beta_{S}$ (for $S$) 0.1 0.952 0.907 -0.506 0.011 -44.936 $0$0.529 -0.484 0.01 0.946 0.854 -0.555 0.014 -39.788 $0$0.582 -0.527 0.001 0.930 0.790 -0.627 0.019 -32.722 $0$0.664 -0.589 0.0001 0.852 0.618 -0.741 0.035 -21.467 $0$0.809 -0.673 0.00001 0.837 0.404 -0.836 0.045 -18.690 $0$0.924 -0.748 $\beta_{C}$ (for $C$) 0.1 0.952 0.907 0.090 0.021 4.231 0.000 0.048 0.131 0.01 0.946 0.854 0.182 0.026 6.960 0.000 0.131 0.234 0.001 0.930 0.790 0.300 0.036 8.335 0.000 0.229 0.371 0.0001 0.852 0.618 0.363 0.065 5.616 0.000 0.236 0.491 0.00001 0.837 0.404 0.569 0.085 6.655 0.000 0.400 0.737 $\beta_{0}$ (intercept) 0.1 0.952 0.907 0.314 0.045 6.953 0.000 0.225 0.402 0.01 0.946 0.854 0.083 0.056 1.491 0.137 -0.027 0.193 0.001 0.930 0.790 -0.173 0.077 -2.261 0.025 -0.324 -0.022 0.0001 0.852 0.618 -0.303 0.138 -2.202 0.029 -0.575 -0.032 0.00001 0.837 0.404 -0.615 0.180 -3.414 0.001 -0.970 -0.260
+Figure A.3 shows the performance for all considered fpr.
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure A.3: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation 2 fitted Table A3 (ViT-B).
+The triangles show the median tpr for the train set (ViT-B; Table A3) and circle the test set (R-50; Table A4) over six seeds.
+Note that the triangles and dots for $C=10$ are for EuroSAT.
+B.2 Simpler variant of the prediction model
+[…]
+The general form this variant can be found in Equation A1, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
+$\log_{10}(\textsc{tpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (A1)
+Table A9 provides tabular results on the performance of the variant.
+Table A9: Results for fitting Equation A1 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$.
+We utilize an ordinary least squares.
+The test $R^{2}$ assesses the fit to the data of R-50 Head.
+coeff.
+fpr $R^{2}$ test $R^{2}$ coeff.
+value std.
+error $t$ $p>|z|$ coeff.
+$[0.025$ coeff.
+$0.975]$ $\beta_{S}$ (for $S$) 0.1 0.908 0.764 -0.248 0.008 -30.976 $0$0.264 -0.233 0.01 0.940 0.761 -0.416 0.011 -36.706 $0$0.438 -0.393 0.001 0.931 0.782 -0.553 0.017 -32.507 $0$0.586 -0.519 0.0001 0.865 0.628 -0.697 0.031 -22.274 $0$0.758 -0.635 0.00001 0.862 0.400 -0.802 0.040 -20.311 $0$0.880 -0.725 $\beta_{C}$ (for $C$) 0.1 0.908 0.764 0.060 0.015 3.955 0.000 0.030 0.089 0.01 0.940 0.761 0.169 0.021 7.941 0.000 0.127 0.211 0.001 0.931 0.782 0.297 0.032 9.303 0.000 0.234 0.360 0.0001 0.865 0.628 0.371 0.059 6.328 0.000 0.255 0.486 0.00001 0.862 0.400 0.580 0.076 7.679 0.000 0.431 0.729 $\beta_{0}$ (intercept) 0.1 0.908 0.764 0.029 0.032 0.913 0.362 -0.034 0.093 0.01 0.940 0.761 -0.118 0.045 -2.613 0.010 -0.208 -0.029 0.001 0.931 0.782 -0.295 0.068 -4.345 $0$0.429 -0.161 0.0001 0.865 0.628 -0.387 0.125 -3.104 0.002 -0.633 -0.141 0.00001 0.862 0.400 -0.683 0.159 -4.288 $0$0.996 -0.369
+Figure A.4 plots the performance of the variant similar to Figure 3(a) in the main text.
 (a) tpr at $\textsc{fpr}=0.1$
@@
 (c) tpr at $\textsc{fpr}=0.001$
-Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation 3 fitted Table A3 (ViT-B).
-The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds.
-Note that the dots for $C=10$ are for EuroSAT.
+(d) tpr at $\textsc{fpr}=0.0001$
+(e) tpr at $\textsc{fpr}=0.00001$
+Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation A1 fitted Table A3 (ViT-B).
+The triangles show the median tpr for the train set (ViT-B; Table A3) and circle the test set (R-50; Table A4) over six seeds.
+Note that the triangles and dots for $C=10$ are for EuroSAT.
+B.3 Additional results for Section 4
+This section provides additional results for Section 4.
+B.4 Rates of In- and Out-distribution Parameters as Function of data set Properties
+LiRA assumes the attacker can build Gaussian distributions of losses for each data point $x$ through the training of shadow models (see Section 2). We denote the distributions as following:
+  • In-distribution: Distribution of the losses of shadow models using $x$ for training ($x\in\mathcal{D}$)
+  • Out-distribution: Distribution of the losses of shadow models not using $x$ for training ($x\notin\mathcal{D}$)
+In Figure A.5 we visualise the changes of the two distributions as functions of $S$ and $C$.
+(a) Distribution means
+(b) Distribution variances
+(c) Distribution standard deviations
+(d) Difference of Means
+(e) Difference of Means / std
+(f) Difference of Means / var
+Figure A.5: Means of observed statistics of the in- and out-distributions for the simplified model with in-class $\sigma=0.1$. The statistics are obtained for $C=2$.
+B.5 Additional results for Section 5
+This section provides additional results for Section 5.
+Figures A.6 and A.7 provide additional data for all considered fpr and for both ViT-B (Head) and R-50 (Head).
+Table A10 the $R^{2}$ scores for ViT-B (Head) and the experiment settings for ViT-B (Head) and R-50 (Head).
+Figure A.8 the relative $R^{2}$ scores for R-50 (Head).
+B.5.1 MIA vulnerability as a function of individual gradient norms
+Figure A.6: ViT-B (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed).
+The results have been obtained over six seeds.
+The number in the bottom right is the $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
+Figure A.7: R-50 (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed).
+The results have been obtained over six seeds.
+The number in the bottom right is the $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
+Table A10: ViT-B (Head): $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
+dataset shots ($S$) classes ($C$) data set size (|$\mathcal{D}$|) $R^{2}$ at fpr=0.1 $R^{2}$ at fpr=0.01 $R^{2}$ at fpr=0.001 Patch Camelyon (Veeling et al., 2018) 256 2 512 0.98 0.97 0.81 CIFAR10 (Krizhevsky, 2009) 64 10 640 0.98 0.98 0.92 EuroSAT (Helber et al., 2019) 64 10 640 0.98 0.97 0.92 Pets (Parkhi et al., 2012) 32 37 1184 0.95 0.93 0.89 Resics45 (Cheng et al., 2017) 32 45 1440 0.95 0.97 0.97 CIFAR100 (Krizhevsky, 2009) 16 100 1600 0.89 0.89 0.86
+Figure A.8: R-50 (Head): Ratio between $R^{2}$ per iteration decile and $R^{2}$ over the whole training.
+A ratio $>1$ means that the decile is as more predictive than whole trace and a ratio $<1$ means that it is less predictive.
+B.5.2 MIA vulnerability as a function of feature space distance
+This section provides additional data regarding the individual MIA prediction using feature space distances.
+Figure A.9: ViT-B (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
+Figure A.10: R-50 (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
+Appendix C Missing proofs of Section 4.2
+See 1
+Proof.
+Recall
+fpr $=\Pr\left(\underbrace{-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]}_{\lambda_{\textsc{fpr}}}>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right)$ (A2)
+tpr $=\Pr\left(\underbrace{-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]}_{\lambda_{\textsc{tpr}}}>\tau\mid x\in\mathcal{D}_{\textrm{target}}\right).$ (A3)
+We denote the class means for the $m$th in/out shadow data sets with $r_{y,m}^{(\textrm{in})}$ and $r_{y,m}^{(\textrm{out})}$
+$r_{y,m}^{(\textrm{in})}$ $=\frac{1}{S}\left(\sum_{i=1}^{S-1}x_{i,m,y}+x\right),\;r_{y,m}^{(\textrm{out})}=\frac{1}{S}\left(\sum_{i=1}^{S}x_{i,m,y}\right),$ (A4)
+where $x_{i,m,y}$ denotes the $i$th element of the $m$th shadow data set with label $y$.
+The corresponding scores $s_{y,m}^{(\textrm{in})}=\langle x,r_{y,m}^{(\textrm{in})}\rangle$ and $s_{y,m}^{(\textrm{out})}=\langle x,r_{y,m}^{(\textrm{out})}\rangle$ which can be written as
+$s_{y,m}^{(\textrm{in})}$ $=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S-1}x_{i,m,y}\rangle+\langle x,x\rangle\right)=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S}x_{i,m,y}\rangle+\langle x,x-x_{S,m,y}\rangle\right)$ (A5)
+$s_{y,m}^{(\textrm{out})}$ $=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S}x_{i,m,y}\rangle\right)$ (A6)
+Now, assuming we have arbitrarily many shadow models, we can write the LiRA statistics as
+$\mu_{\textrm{in}}=\operatorname{\mathbb{E}}_{D_{m}}[s_{y,m}^{(\textrm{in})}],\;\mu_{\textrm{out}}=\operatorname{\mathbb{E}}_{D_{m}}[s_{y,m}^{(\textrm{out})}]=\langle x,m_{y}\rangle$ (A7)
+$\Rightarrow$ $\mu_{\textrm{in}}-\mu_{\textrm{out}}=\frac{1}{S}\langle x,x-\mathbb{E}_{D_{m}}[x_{S,m}]\rangle=\frac{1}{S}\langle x,x-m_{y}\rangle,$ (A8)
+where $m_{y}$ is the true center of the $y$th cluster. For $\sigma$ we have
+$\sigma^{2}$ $=\operatorname{Var}(s_{y,m}^{(\textrm{out})})$ (A9)
+$=\frac{1}{S^{2}}\sum_{i=1}^{S}\operatorname{Var}(\langle x,x_{i,m}\rangle)$ (A10)
+$=\frac{1}{S}\operatorname{Var}(\langle x,x_{m}^{(j)}\rangle)$ (A11)
+$=\frac{1}{S}\operatorname{Var}(x_{m}^{(j)})||x||^{2}=\frac{1}{S}s^{2}||x||^{2},$ (A12)
+where $x_{m}^{(j)}$ denotes the $j$th dimension of the vector $x_{m}$ and $s$ is the std of each dimension of $x_{m}$. Now we can write the $\lambda_{\textsc{fpr}}$ as
+$\lambda_{\textsc{fpr}}=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(-2t_{x}+2\langle x,m_{y}\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A13)
+In the fpr case, we have that $x\notin\mathcal{D}_{\textrm{target}}$, therefore for the $t_{x}$ we have
+$t_{x}=\langle x,r_{y}^{\textrm{(target)}}\rangle\sim\langle x,m_{y}\rangle+\frac{s}{\sqrt{S}}\langle x,z\rangle,$ (A14)
+where $z\sim N(0,I_{d})$. Substituting this into $\lambda_{\textsc{fpr}}$ we have
+$\lambda_{\textsc{fpr}}=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A15)
+We see that $\lambda_{\textsc{fpr}}$ evolves as $O(1/\sqrt{S})$.
+Next, we will focus on the $\lambda_{\textsc{tpr}}$. The expression remains otherwise the same as for $\lambda_{\textsc{fpr}}$, but since $x\in\mathcal{D}_{\textrm{target}}$ the $t_{x}$ changes to
+$t_{x}\sim\langle x,m_{y}\rangle+\frac{1}{S}\langle x,x\rangle+s\frac{1}{\sqrt{S}}\sqrt{1-\frac{1}{S}}\langle x,z\rangle.$ (A16)
+Hence, for large enough $S$ we have
+$\lambda_{\textsc{tpr}}\approx-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle-\frac{2}{S}||x||^{2}+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A17)

### v3

@@
-On the Impact of Dataset Properties on Membership Privacy of Deep Learning
+Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning
 Abstract
-We apply a state-of-the-art membership inference attack (MIA) to systematically test the practical privacy vulnerability of fine-tuning large image classification models.
-We focus on understanding the properties of data sets and samples that make them vulnerable to membership inference.
-In terms of data set properties, we find a strong power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate.
-We train a linear model to predict true positive rate based on data set properties and observe good fit for MIA vulnerability on unseen data.
-To analyse the phenomenon theoretically, we reproduce the result on a simplified model of membership inference that behaves similarly to our experimental data.
-We prove that in this model, the logarithm of the difference of true and false positive rates depends linearly on the logarithm of the number of examples per class.
-For an individual sample, the gradient norm is predictive of its vulnerability.
+We analyse the relationship between privacy vulnerability and dataset properties, such as examples per class and number of classes, when applying two state-of-the-art membership inference attacks (MIAs) to fine-tuned neural networks.
+We derive per-example MIA vulnerability in terms of score distributions and statistics computed from shadow models.
+We introduce a simplified model of membership inference and prove that in this model, the logarithm of the difference of true and false positive rates depends linearly on the logarithm of the number of examples per class.
+We complement the theoretical analysis with empirical analysis by systematically testing the practical privacy vulnerability of fine-tuning large image classification models and obtain the previously derived power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate.
+Finally, we fit a parametric model of the previously derived form to predict true positive rate based on dataset properties and observe good fit for MIA vulnerability on unseen fine-tuning scenarios.
 1 Introduction
@@
 On the other hand, loss of utility of the model due to privacy constraints may compromise the health benefits that might be gained from it.
-In this paper, our aim is to apply a state-of-the-art MIA systematically, to help understand practical privacy risks when training deep-learning-based classifiers without DP protections.
-Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep learning models to MIA.
+In this paper, our aim is to theoretically understand and apply two state-of-the-art MIAs systematically, to help understand practical privacy risks when fine-tuning deep-learning-based classifiers without DP protections.
+We focus on transfer learning using fine-tuning because this is increasingly used for all practical applications of deep learning.
+Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep transfer learning models to MIA.
 Since the earliest MIA methods, there has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017).
 Similarly, there is evidence that models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023).
-Based on extensive experimental data over many data sets with varying sizes, we uncover a power law describing the vulnerability remarkably well.
+Feldman & Zhang (2020) showed that neural networks trained from scratch can memorise a large fraction of their training data, while the memorisation is greatly reduced for fine-tuning.
+In this paper, we theoretically study the relationship between examples per class, which we denote as shots ($S$), and MIA vulnerability (true positive rate tpr at fixed false positive rate fpr) for a simplified model of fine-tuning and derive a power-law relationship in the form
+$\log(\textsc{tpr}-\textsc{fpr})=-\beta_{S}\log(S)-\beta_{0}.$ (1)
+We complement the theoretical analysis with extensive experiments over many datasets with varying sizes in the transfer learning setting for image classification tasks and observe the same power law.
 Previously, Tobaben et al.
@@
 (2023) studied how the MIA vulnerability of a class in an image classifier trained from scratch depends on the average individual privacy parameters of that class.
-Our work significantly expands on both of these by more systematic study and modelling of data set properties and by providing a more detailed look at individual samples, not just classes.
-List of contributions In this work we perform an extensive empirical study on the MIA vulnerability of deep learning models.
+Our work significantly expands on both of these by a more systematic study and modelling of dataset properties.
+List of contributions We analyze the MIA vulnerability of deep transfer learning using two state-of-the-art MIAs LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024).
+Both MIAs are based on the assumption that the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack.
+We first analyze both MIAs theoretically by introducing a simplified model of membership inference (Section 3).
+We support our theoretical findings by an extensive empirical study on the MIA vulnerability of deep learning models.
 We focus on a transfer learning setting for image classification tasks, where a large pre-trained neural network is fine-tuned on a sensitive dataset, and the MIA is performed on the final classifier.
 We find that the MIA vulnerability has a strong correlation with certain dataset properties (Section 4).
-In Section 4 we analyse the reasons for these findings more thoroughly and we extend the analysis to individual samples in Section 5.
-The primary contributions of this work are:
   1. 1.
-    Few-shot MIA: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) of target models trained using many different data sets with varying sizes and uncover a power law between the number of examples per class and the vulnerability to MIA (see Figure 1).
+    Per-example vulnerability: We derive per-example LiRA and RMIA vulnerability (tpr at fixed fpr) in terms of MIA attack score distributions and statistics computed from shadow models (see Section 3.3).
   2. 2.
-    Regression model: We utilise our empirical observations to train a regression model to predict MIA vulnerability ($\log(\textsc{tpr}-\textsc{fpr})$ at fixed low fpr) based on examples per class ($\log S$) and number of classes ($\log C$).
+    Power-law in simplified model of MI: We formulate a simplified model of membership inference to quantitatively relate dataset properties and MIA vulnerability.
+For this model, we prove a power-law relationship between the per-example LiRA and RMIA vulnerability and the number of examples per class.
+We then extend the the power-law relationship to the average-case LiRA and RMIA vulnerability (See Section 3.4).
+  3. 3.
+    Few-shot MIA: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power law between the number of examples per class and the vulnerability to MIA (see Figure 1).
+  4. 4.
+    Regression model: We utilise our empirical observations to fit a regression model to predict MIA vulnerability ($\log(\textsc{tpr}-\textsc{fpr})$ at fixed low fpr) based on examples per class ($\log S$) and number of classes ($\log C$), which follows the functional form of the theoretically derived power-law.
 We show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Figure 4).
-  3. 3.
-    Theoretical analysis using a simplified model of MIA: To analyse the phenomenon theoretically, we reproduce the result on a simplified model of membership inference that behaves similarly to our experimental data.
-In this model we can formally prove that $\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log S$ (see Section 4).
-  4. 4.
-    Individual sample MIA: We extend the study to individual samples and find: 1 The gradient norm is predictive of vulnerability of an individual sample.
-2 Gradient norms at different stages of learning have different impact, but how depends on the data set (see Section 5).
-2 Background and Methods
+2 Background
 Notation for the properties of the training dataset $\mathcal{\mathcal{D}}$: 1 $C$for the number of classes 2 $S$for shots (examples per class) 3 $\mathcal{|\mathcal{D}|}$for training dataset size ($\mathcal{|\mathcal{D}|}=CS$).
 We denote the number of MIA shadow models with $M$.
-Membership inference attacks (MIAs) aim to infer whether a particular sample was part of the training set of the targeted model Shokri et al.
-(2017).
+Membership inference attacks (MIAs) aim to infer whether a particular sample was part of the training set of the targeted model (Shokri et al., 2017).
 Thus, they can be used to determine lower bounds on the privacy leakage of models to complement the theoretical upper bounds obtained through differential privacy.
-Likelihood Ratio attack (LiRA; Carlini et al., 2022) While many different MIAs have been proposed Hu et al.
-(2022), in this work we consider the SOTA: the Likelihood Ratio Attack (LiRA).
-LiRA assumes an attacker that has black-box access to the attacked model, knows the training data distribution, the training set size, the model architecture, hyperparameters and training algorithm.
+Likelihood Ratio attack (LiRA; Carlini et al., 2022) While many different MIAs have been proposed (Hu et al., 2022), in this work we consider the Likelihood Ratio Attack (LiRA).
+LiRA is a strong attack that assumes an attacker that has black-box access to the attacked model, knows the training data distribution, the training set size, the model architecture, hyperparameters and training algorithm.
 Based on this information, the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack but for which the attacker knows the training dataset.
 LiRA exploits the observation that the loss function value used to train a model is often lower for examples that were part of the training set compared to those that were not.
-For a target sample $x$, LiRA trains the shadow models: 1 with $x$ as a part of the training set ($x\in\mathcal{D}$) 2 and without $x$ in the training set ($x\notin\mathcal{D}$) .
-After training the shadow models, $x$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $x\in\mathcal{D}$ shadow models, and one for the $x\notin\mathcal{D}$.
-Finally, the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $x\in\mathcal{D}$ or $x\notin\mathcal{D}$.
+For a target sample $(x,y_{x})$, LiRA trains the shadow models: 1 with $(x,y_{x})$ as a part of the training set ($(x,y_{x})\in\mathcal{D}$) and 2 without $x$ in the training set ($(x,y_{x})\notin\mathcal{D}$) .
+After training the shadow models, $(x,y_{x})$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $(x,y_{x})\in\mathcal{D}$ shadow models, and one for the $(x,y_{x})\notin\mathcal{D}$.
+Finally, the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $(x,y_{x})\in\mathcal{D}$ or $(x,y_{x})\notin\mathcal{D}$.
+We use an optimization by Carlini et al.
+(2022) for performing LiRA for multiple models and points without training a computationally infeasible number of shadow models.
+It relies on sampling the shadow datasets in a way that each sample is in expectation half of the time included in the training dataset of a shadow model and half of the time not.
+At attack time each model will be attacked once using all other models as shadow models.
+Robust Membership Inference Attack (RMIA; Zarifzadeh et al., 2024) Recently Zarifzadeh et al.
+(2024) proposed a new MIA algorithm called RMIA, which aims to improve performance when the number of shadow models is limited.
+Similar to LiRA, RMIA is based on shadow model training and computing the attack statistics based on a likelihood ratio.
+The main difference to LiRA is that RMIA does not compute the likelihood ratio based on aggregated IN/OUT statistics, but instead compares the target data point against random samples $(z,y_{z})$ from the target data distribution.
+After computing the likelihood ratios over multiple $(z,y_{z})$ values, the MIA score is estimated as a proportion of the ratios exceeding a preset bound.
+This approach makes RMIA a more effective attack when the number of shadow models is low.
+Measuring MIA vulnerability Using the chosen MIA score of our attack, we can build a binary classifier to predict whether a sample belongs to the training data or not.
+The accuracy profile of such classifier can be used to measure the success of the MIA.
+More specifically, throughout the rest of the paper, we will use the true positive rate (tpr) at a specific false positive rate (fpr) as a measure for the vulnerability.
+Identifying even a small number of examples with high confidence is considered harmful (Carlini et al., 2022) and thus we focus on the regions of small fpr.
+3 Theoretical analysis
+In this section, we seek to theoretically understand the impact of the dataset properties on the MIA vulnerability.
+It is known that different data points exhibit different levels of MIA vulnerability depending on the underlying distribution (e.g.
+Aerni et al., 2024; Leemann et al., 2024).
+Therefore, we start with analysing per-example vulnerabilities for LiRA and RMIA.
+In order to quantitatively relate dataset properties to these vulnerabilities, a simplified model is formulated.
+Within this model, we prove a power-law between the per-example vulnerability and the number $S$ of examples per class.
+Finally, the per-example power-law is analytically extended to average-case MIA vulnerability, for which we provide empirical evidence in Section 4.
+3.1 Preliminaries
+First, let us restate the MIA score from LiRA as defined by Carlini et al.
+(2022).
+Denoting the logit of a target model $\mathcal{M}$ applied on a target data point $(x,y_{x})$ as $\ell(\mathcal{M}(x),y_{x})$, the LiRA computes the MIA score as the likelihood ratio
+$\mathrm{LR}(x)=\frac{p(\ell(\mathcal{M}(x),y_{x})\mid\mathbb{Q}_{\text{in}}(x,y_{x}))}{p(\ell(\mathcal{M}(x),y_{x})\mid\mathbb{Q}_{\text{out}}(x,y_{x}))},$ (2)
+where the $\mathbb{Q}_{\text{in}/\text{out}}$ denote the hypotheses that $(x,y_{x})$ was or was not in the training set of $\mathcal{M}$.
 Carlini et al.
-(2022) proposed an optimisation for performing LiRA for multiple models and points without training a computationally infeasible number of shadow models.
-We utilise this optimisation.
-Measuring MIA vulnerability Using the likelihood-ratios from LiRA as a score, we can build a binary classifier to predict whether a sample belongs to the training data or not.
-The operational characteristics of such classifier can be used to measure the success of the MIA.
-More specifically, throughout the rest of the paper, we will use the true positive rate (tpr) at a specific false positive rate (fpr) as a measure for the vulnerability.
-Identifying even a small number of samples with high confidence is considered harmful (Carlini et al., 2022) and thus we focus on the regions of small fpr.
+(2022) approximate the IN/OUT hypotheses as normal distributions.
+Denoting $t_{x}=\ell(\mathcal{M}(x),y_{x})$, the score becomes
+$\mathrm{LR}(x)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}^{2}(x))}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}^{2}(x))},$ (3)
+where the $\hat{\mu}_{\text{in/out}}(x)$ and $\hat{\sigma}_{\text{in/out}}(x)$ are the means and standard deviations for the IN/OUT shadow model losses for $(x,y_{x})$.
+Larger values of $\mathrm{LR}(x)$ suggest that $(x,y_{x})$ is more likely in the training set and vice versa.
+Now, to build a classifier from this score, the LiRA tests if $\mathrm{LR}(x)>\beta$ for some threshold $\beta$.
+Next, let us restate how RMIA (Zarifzadeh et al., 2024) builds the MIA score.
+RMIA augments the likelihood-ratio with a sample from the target data distribution to calibrate how likely you would obtain the target model if $(x,y_{x})$ is replaced with another sample $(z,y_{z})$.
+Denoting the target model parameters with $\theta$, RMIA computes
+$\mathrm{LR}(x,z)=\frac{p(\theta\mid x,y_{x})}{p(\theta\mid z,y_{z})},$ (4)
+and the corresponding MIA score is given as
+$\text{Score}_{\text{RMIA}}(x)=\Pr_{(z,y_{z})\sim\mathbb{D}}(\mathrm{LR}(x,z)>\gamma),$ (5)
+where $\mathbb{D}$ denotes the training data distribution.
+Similar to LiRA, the classifier is built by checking if $\text{Score}_{\text{RMIA}}(x)>\beta$.
+In the following, we will use the direct computation of likelihood-ratio as described in Equation 11 of Zarifzadeh et al.
+(2024) which approximates $\mathrm{LR}(x,z)$ using normal distributions.
+3.2 Computing the tpr for LiRA and RMIA
+Using the LiRA formulation of Equation 3, the tpr for the target point $(x,y_{x})$ for LiRA is defined as
+$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}(x)^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})}\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\right),$ (6)
+where $\beta$ is a threshold that defines a rejection region of the likelihood ratio test, $\hat{\mu}_{\mathrm{in}}(x),\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{in}}(x)$ and $\hat{\sigma}_{\mathrm{out}}(x)$ are LiRA statistics computed from shadow models, and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Appendix A for derivation).
+For theoretical analysis of RMIA, we focus on the direct approach that is an approximation of the efficient Bayesian approach, as Zarifzadeh et al.
+(2024) empirically demonstrates that these approaches exhibit similar performances.
+Let $\hat{\mu}_{a,b}$ and $\hat{\sigma}_{a,b}$ denote, respectively, the mean and standard deviation of $t_{b}$ estimated from shadow models, where $a$ denotes which of $(x,y_{x})$ and $(z,y_{z})$ is in the training set.
+By Equation 11 in (Zarifzadeh et al., 2024), the per-example performance for RMIA is given as
+$\textsc{tpr}_{\mathrm{RMIA}}(x)=\\ \Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\Pr_{(z,y_{z})\sim{\mathbb{D}}}(\mathrm{LR}(x,z)\geq\gamma)\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\notin\mathcal{D}_{\mathrm{target}}\right)$ (7)
+$\mathrm{LR}(x,z)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{x,x},\hat{\sigma}_{x,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{x,z},\hat{\sigma}_{x,z}^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{z,x},\hat{\sigma}_{z,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{z,z},\hat{\sigma}_{z,z}^{2})},$ (8)
+where $t_{z}$ is the score on $z$ similar to $t_{x}$ and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Appendix A for derivation).
+We define the average-case tprs for LiRA and RMIA by taking the expectation over the data distribution:
+$\overline{\textsc{tpr}}_{\mathrm{LiRA}}=\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}[\textsc{tpr}_{\mathrm{LiRA}}(x)]$ (9)
+$\overline{\textsc{tpr}}_{\mathrm{RMIA}}=\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}[\textsc{tpr}_{\mathrm{LiRA}}(x)]$ (10)
+3.3 Per-example MIA vulnerability
+Although LiRA models $t_{x}$ by a normal distribution, we consider a more general case where the true distribution of $t_{x}$ is of the location-scale family. That is,
+$t_{x}=\begin{cases}\mu_{\mathrm{in}}(x)+\sigma_{\mathrm{in}}(x)t&\text{ if }(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\\ \mu_{\mathrm{out}}(x)+\sigma_{\mathrm{out}}(x)t&\text{ if }(x,y_{x})\notin\mathcal{D}_{\mathrm{target}},\end{cases}$ (11)
+where $t$ has the standard location and unit scale, and $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and $\sigma_{\mathrm{in}}(x),\sigma_{\mathrm{out}}(x)$ are the locations and scales of IN/OUT distributions of $t_{x}$.
+We assume that the target and shadow datasets have a sufficient number of examples.
+This allows us to also assume that $\hat{\sigma}(x)=\hat{\sigma}_{\mathrm{in}}(x)=\hat{\sigma}_{\mathrm{out}}(x)$ and $\sigma(x)=\sigma_{\mathrm{in}}(x)=\sigma_{\mathrm{out}}(x)$, where $\hat{\sigma}(x)$ is the standard deviation of $t_{x}$ estimated from shadow models and $\sigma(x)$ is the true scale parameter of $t_{x}$.
+(See Appendix B for the validity of these assumptions).
+The following result reduces the LiRA vulnerability to the location and scale parameters of $t_{x}$.
+Lemma 1 (Per-example LiRA vulnerability).
+Suppose that the true distribution of $t_{x}$ is of location-scale family with locations $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and scale $\sigma(x)$, and that LiRA models $t_{x}$ by $\mathcal{N}(\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}(x))$ and $\mathcal{N}(\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}(x))$.
+Assume that an attacker has access to the underlying distribution ${\mathbb{D}}$.
+Then for a large enough number of examples per class and infinitely many shadow models, the LiRA vulnerability of a fixed target example is
+$\textsc{tpr}_{\mathrm{LiRA}}(x)=\begin{cases}1-F_{t}\left(F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)>\hat{\mu}_{\mathrm{out}}(x)\\ F_{t}\left(F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)<\hat{\mu}_{\mathrm{out}}(x),\end{cases}$ (12)
+where $F_{t}$ is the cdf of $t$ with the standard location and unit scale.
+Proof.
+See Section C.1. ∎
+Here we assume that an attacker trains shadow models with the true underlying distribution.
+However, in real-world settings the precise underlying distribution may not be available for an attacker.
+We relax this assumption in Appendix B so that the attacker only needs an approximated underlying distribution for the optimal LiRA as in Lemma 1.
+Next we focus on the per-example RMIA performance. As in the case of LiRA, we assume that $t_{x}$ and $t_{z}$ follow distributions of the location-scale family. We have
+$t_{x}$ =$=\begin{cases}\mu_{x,x}+\sigma_{x,x}t&\text{ if }(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{z,x}+\sigma_{z,x}t&\text{ if }(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\in\mathcal{D}_{\mathrm{target}}\end{cases}$ (13)
+$t_{z}$ =$=\begin{cases}\mu_{x,z}+\sigma_{x,z}t&\text{ if }(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{z,z}+\sigma_{z,z}t&\text{ if }(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\in\mathcal{D}_{\mathrm{target}}.\end{cases}$ (14)
+It is important to note that $\mu_{a,b}$ and $\sigma_{a,b}$ denote, respectively, a location and a scale, while previously defined $\hat{\mu}_{a,b}$ and $\hat{\sigma}_{a,b}$ are, respectively, a mean and a standard deviation.
+As for the analysis of LiRA, we assume that the target and shadow sets have a sufficient number of examples per class, and that $\sigma_{x}=\sigma_{x,x}=\sigma_{z,x}$, $\sigma_{z}=\sigma_{x,z}=\sigma_{z,z}$, $\hat{\sigma}_{x}=\hat{\sigma}_{x,x}=\hat{\sigma}_{z,x}$ and $\hat{\sigma}_{z}=\hat{\sigma}_{x,z}=\hat{\sigma}_{z,z}$, where $\sigma_{x}$ and $\sigma_{z}$ are, respectively, the true scales of $t_{x}$ and $t_{z}$, and $\hat{\sigma}_{x}$ and $\hat{\sigma}_{z}$ are, respectively, standard deviations of $t_{x}$ and $t_{z}$ estimated from shadow models (see Appendix B for the validity of these assumptions).
+Lemma 2 (Per-example RMIA vulnerability).
+Suppose that the true distributions of $t_{x}$ and $t_{z}$ are of location-scale family with locations $\mu_{x,x},\mu_{z,x},\mu_{x,z},\mu_{z,z}$ and scales $\sigma_{x},\sigma_{z}$, and that RMIA models $t_{x}$ and $t_{z}$ by normal distributions with parameters computed from shadow models.
+For a large enough number of examples per class and infinitely many shadow models, the RMIA vulnerability of a fixed target example is bounded by
+$\textsc{tpr}_{\mathrm{RMIA}}(x)\leq\begin{cases}1-F_{t}\left(F_{t}^{-1}(1-\alpha)-\frac{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[q]}{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]}\right)&\text{ if }\ \mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]>0\\ F_{t}\left(F_{t}^{-1}(\alpha)-\frac{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[q]}{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]}\right)&\text{ if }\ \mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]<0,\end{cases}$ (15)
+for some constant $\alpha\geq\textsc{fpr}_{\mathrm{RMIA}}(x)$, where
+$q$ $=\frac{(\mu_{x,x}-\mu_{z,x})(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})}{\hat{\sigma}_{x}^{2}}-\frac{(\mu_{x,z}-\mu_{z,z})(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}{\hat{\sigma}_{z}^{2}}$ (16)
+$A$ $=\frac{\sigma_{x}}{\hat{\sigma}_{x}^{2}}(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\sigma_{z}}{\hat{\sigma}_{z}^{2}}(\hat{\mu}_{x,z}-\hat{\mu}_{z,z}).$ (17)
+Proof.
+See Section C.2. ∎
+Note that here we must assume that the attacker has access to the underlying distribution for the optimal RMIA as the Equations 16 and 17 depend on the parameters computed from shadow models.
+3.4 A simplified model of membership inference
+Now we construct a simplified model of membership inference that streamlines the data generation and shadow model training.
+We sample vectors on a high dimensional unit sphere and classify them based on inner product with estimated class mean.
+This model is easier to analyse theoretically than real-world deep learning examples.
+We generate the data and form the classifiers (which are our target models) as follows:
+  1. 1.
+    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in\{1,\dots,C\}:m_{i}\perp m_{j}\lor i=j$).
+  2. 2.
+    We sample $2S$ vectors $x_{c}$ for each class.
+We assume that they are Gaussian distributed around the the true class mean $x_{c}\sim\mathcal{N}(m_{c},s^{2}I)$ where the $s^{2}$ is the in-class variance.
+  3. 3.
+    For each ”target model” we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $r_{c}$.
+  4. 4.
+    The computed mean is used to classify sample $x$ by computing the inner product $\langle x,r_{c}\rangle$ as a metric of similarity.
+The attacker has to infer which vectors have been used for training the classifier.
+Instead of utilising the logits (like in many image classification tasks), the attacker can use the inner products of a point with the cluster means.
+This simplified model resembles a linear (Head) classifier often used in transfer learning when adapting to a new dataset.
+We also focus on the linear (Head) classifier in our empirical evaluation in Section 4.
+In the linear classifier, we find a matrix $W$ and biases $b$, to optimize the cross-entropy between the labels and logits $Wv+b$, where $v$ denotes the feature space representation of the data.
+In the simplified model, the rows of $W$ are replaced by the cluster means and we do not include the bias term in the classification.
+Now, applying Lemma 1 to the simplified model yields the following result.
+Theorem 3 (Per-example LiRA power-law).
+Fix a target example $(x,y_{x})$. For the simplified model with arbitrary $C$ and infinitely many shadow models, the per-example LiRA vulnerability is given as
+$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))+\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}s||x||}\right),$ (18)
+where $m_{x}$ is the true mean of class $y_{x}$. In addition, for large $S$ we have
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}+\log\frac{|\langle x,x-m_{x}\rangle|}{||x||s\sqrt{2\pi}}.$ (19)
+Proof.
+See Section C.3. ∎
+An immediate upper bound is obtained from Theorem 3 by the Cauchy-Schwarz inequality:
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x))\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}+\log\frac{||x-m_{x}||}{s\sqrt{2\pi}}.$ (20)
+This implies that if $||x-m_{x}||$ is bounded, then the worst-case vulnerability is also bounded.
+Hence we can have the LiRA vulnerability arbitrarily close to zero in this non-DP setting by simply increasing $S$.
+Similarly, employing Lemma 2 and the simplified model, we obtain the following upper bound for RMIA performance.
+Theorem 4 (Per-example RMIA power-law).
+Fix a target example $(x,y_{x})$. For the simplified model with infinitely many shadow models, the per-example RMIA vulnerability is given as
+$\textsc{tpr}_{\mathrm{RMIA}}(x)\leq 1-F_{|t|}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\psi(x,C)}{\sqrt{S}s}\right),$ (21)
+where $\alpha\geq\textsc{fpr}_{\mathrm{RMIA}}(x)$ and
+$\psi(x,C)=\\ \frac{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[2||x-z||^{2}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[(||x-m_{x}||^{2}+||z-m_{z}||^{2})\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[2||x-z||\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[(||x-m_{x}||+||z-m_{z}||)\mid y_{z}\not=y_{x}\right]}.$ (22)
+In addition, for large $S$ we have
+$\log(\textsc{tpr}_{\mathrm{RMIA}}(x)-\textsc{fpr}_{\mathrm{RMIA}}(x))\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\alpha)^{2}+\log\frac{\psi(x,C)}{\sqrt{\pi/2}}.$ (23)
+Proof.
+See Section C.4. ∎
+As for the LiRA power-law, bounding $||x-z||,||x-m_{x}||$ and $||z-m_{z}||$ will provide a worst-case upper bound for which the power-law holds.
+Now the following corollaries extend the power-law to the average-case MIA vulnerabilities.
+We will also empirically validate these results in Section 4.
+Corollary 5 (Average-case LiRA power-law).
+For the simplified model with arbitrary $C$, sufficiently large $S$ and infinitely many shadow models, we have
+$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\langle x,x-m_{x}\rangle}{\sqrt{2\pi}||x||s}\right]\right).$ (24)
+Proof.
+See Section C.5. ∎
+Corollary 6 (Average-case RMIA power-law).
+For the simplified model with sufficiently large $S$ and infinitely many shadow models, we have
+$\log(\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}})\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\alpha)^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\psi(x,C)}{\sqrt{2\pi}}\right]\right).$ (25)
+Proof.
+See Section C.6. ∎
+4 Empirical evaluation of MIA vulnerability and dataset properties
+In this section, we investigate how different properties of datasets affect the MIA vulnerability.
+Based on our observations, we propose a method to predict the vulnerability to MIA using these properties.
+4.1 Experimental setup
+We focus on a image classification setting where we fine-tune pre-trained models on sensitive downstream datasets and assess the MIA vulnerability using LiRA and RMIA with $M=256$ shadow/reference models.
+We base our experiments on a subset of the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a test classification accuracy $>80\%$ (see Table A3).
+We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k (Russakovsky et al., 2015).
+The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section F.1.
+We optimise the hyperparameters (batch size, learning rate and number of epochs) using the library Optuna (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section E.2).
+We provide the the code for reproducing the experiments in the supplementary material.
 Measuring the uncertainty for tpr The tpr values from the LiRA-based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples.
 Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimators.
-Therefore, when we report the tpr values for a single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals Clopper and Pearson (1934).
+Therefore, when we report the tpr values for a single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals (Clopper & Pearson, 1934).
 Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
@@
 where $B(q;a,b)$ is the $q$th-quantile of $\text{Beta}(a,b)$ distribution.
-3 MIA Vulnerability and Data Set Properties
-In this section, we investigate how different properties of data sets affect the MIA vulnerability.
-Based on our observations, we propose a method to predict the vulnerability to MIA using these properties.
-3.1 Experimental Setup
-We focus on a image classification setting where we fine-tune pre-trained models on sensitive downstream data sets and assess the MIA vulnerability using LiRA with $M=256$ shadow models.
-We base our experiments on a subset of the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a test classification accuracy $>80\%$ (see Table A2).
-We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k Russakovsky et al.
-(2015).
-The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section B.1.
-We optimise the hyperparameters (batch size, learning rate and number of epochs) using the library Optuna Akiba et al.
-(2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section A.2).
-We provide the the code for reproducing the experiments in the supplementary material.
 4.2 Experimental results
-Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 2). We make the following observations:
-  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in the Appendix (Figures A.1, A3 and A4).
-  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 1(b) with further evidence in Figures A.2, A5 and A6 in the Appendix.
+Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 2) using LiRA.
+We make the following observations:
+  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in the Appendix (Figures A.2, A4 and A5).
+  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 1(b) with further evidence in Figures A.3, A6 and A7 in the Appendix.
 However, the trend w.r.t.
@@
 (b) as function of $C$ (classes) at $S=32$
-Figure 1: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of data set properities when attacking a ViT-B Head fine-tuned without DP on different data sets.
-The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over multiple seeds (six for Figure 1(a) and 12 for Figure 1(b))
+[…]
+RMIA In Figure 2 we compare the vulnerability of the models to LiRA and RMIA as a function of the number of $S$ (shots) at $\textsc{fpr}=0.1$.
+We observe the power-law for both attacks, but the RMIA is more unstable than LiRA (especially for lower fpr).
+More results for RMIA are in Figures A.6, A.7 and A.8 in the Appendix.
+[…]
 4.3 Model to predict dataset vulnerability
-[…]
+The trends seen in Figure 1 suggest the same power law relationship that we derived for the simplified model of membership inference in Section 3.
+We fit a linear regression model to predict $\log(\textsc{tpr}-\textsc{fpr})$ for each $\textsc{fpr}=10^{-k},k=1,\dots,5$ separately using the $\log C$ and $\log S$ as covariates with statsmodels (Seabold & Perktold, 2010).
+The general form of the model can be found in Equation 27, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
 $\log_{10}(\textsc{tpr}-\textsc{fpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (27)
 […]
-We utisee MIA results of ViT-B (see Table A3) as the training data.
+We utilise MIA results of ViT-B (Head) (see Table A4) as the training data.
 Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well.
@@
 Figure 3 shows the parameters of the prediction model fitted to the training data.
-Figure 2: Coefficient values for different fpr when fitting a regression model based on Equation 2 fitted on data from ViT-B (Head) (Table A3).
+For larger fpr, the coefficient $\beta_{S}$ is around $-0.5$, as our theoretical analysis predicts.
+However, the coefficient value decreases for small fpr.
+This is perhaps because the power-law in Equation (24) only holds for large $S$, and for small fpr Equation (24) significantly underestimates the vulnerability in small-$S$ regime (see Appendix D).
+Figure 3: Coefficient values for different fpr when fitting a regression model based on Equation 27 fitted on data from ViT-B (Head) with LiRA (Table A4).
 The error bars display the 95% confidence intervals based on Student’s t-distribution.
@@
   • R-50 (FiLM): Figure 4(b) shows that the prediction quality is good for R-50 (FiLM) models.
-These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section A.1).
+These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section E.1).
 Tobaben et al.
@@
 (2023), see Table A8) 2 right: trained from scratch (data from Carlini et al.
-(2022))
+(2022)) .
+The linear model is robust to changing the fine-tuning method from Head to FiLM, but from scratch training seems to be more vulnerable than predicted.
+Note that the results from Carlini et al.
+(2022) use data augmentation while we do not.
 Figure 4: Performance of the regression model based on Equation 27 fitted on data from Table A4.
-4 Understanding the MIA Prediction Model
-In this section, we focus on understanding why the previously observed trends exist. Our goal is to formally prove a similar relationship as in the predictive model of Section 3.3.
-4.1 A Simplified Model of Membership Inference
-In this section, we construct a simplified model of membership inference that streamlines the data generation and shadow model training.
-We observe that the trends of the vulnerability of the simplified model is a good approximation of the trends previously observed with the image classification example.
-Instead of training linear classifiers to classify images, we sample vectors on a high dimensional unit sphere and classify based on inner product with estimated class mean.
-This model is easier to analyse theoretically, and allows us to simulate arbitrary combinations of $S$ and $C$ without relying on the existence of a specific image classification downstream data set as especially high $C$ image classification downstream data sets that are computationally feasible to fine-tune on have small $S$.
-We generate the data and form the classifiers (which are our target models) as follows:
-  1. 1.
-    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in C:m_{i}\perp m_{j}\lor i=j$).
-  2. 2.
-    We sample $2S$ vectors $x_{c}$ for each class.
-We assume that they are Gaussian distributed around the the true class mean $x_{c}\sim\mathcal{N}(m_{c},s^{2}I)$ where the $s^{2}$ is the in-class variance.
-  3. 3.
-    For each "target model" we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $r_{c}$.
-  4. 4.
-    The computed mean is used to classify sample $(x,y)$ by computing the inner product $\langle x,r_{c}\rangle$ as a metric of similarity.
-The attacker has to infer which vectors have been used for training the classifier.
-Instead of utilising the logits (like in the image classification example), the attacker can use the inner products of a point with the cluster means.
-In Figure 4 we compare the attacker’s performance as a function of $C$ and $S$ to the the previous image classification example.
-This simplified model resembles the linear (Head) classifier.
-In the linear classifier, we find a matrix $A$ and biases $b$, to optimize the cross-entropy between the labels and logits $Az+b$, where $z$ denotes the feature space representation of the data.
-In the simplified model, the rows of $A$ are replaced by the cluster means and we do not include the bias term in the classification.
-(a) as a function of $S$ (shots)
-(b) as a function of $C$ (classes)
-[…]
-4.2 Theoretical Analysis
-Similarly to the setting in Section 4.1, we focus on a classification model that aims to classify the sample $(x,y)$ based on the inner product $\langle x,r_{c}\rangle$, where $r_{c}$ is the mean of the training data points $x_{m}$ with $y_{m}=c$.
-However, we only focus on the scores of the target class, i.e.
-$c=y$, and therefore the score used of LiRA simplifies to $\langle x,r_{y}\rangle$.
-The fpr for LiRA is given as
-$\textsc{fpr}=\Pr\left(\frac{\mathcal{N}(t_{x};\mu_{\textrm{in}},\sigma_{\textrm{in}}^{2})}{\mathcal{N}(t_{x};\mu_{\textrm{out}},\sigma_{\textrm{out}}^{2})}>\exp(\tau)\mid x\notin\mathcal{D}_{\textrm{target}}\right),$ (3)
-where $t_{x}=\langle x,r_{y}^{\textrm{(target)}}\rangle$ is the score for the target model, $\tau$ is the decision threshold for the classifier, and $\mu_{\textrm{in}},\mu_{\textrm{out}},\sigma_{\textrm{in}}$ and $\sigma_{\textrm{out}}$ are the LiRA statistics computed from the shadow models.
-We assume that $\sigma_{\textrm{in}}=\sigma_{\textrm{out}}=\sigma$, which allows to write Equation 3 as
-fpr $=\Pr\left(\log\frac{\mathcal{N}(t_{x};\mu_{\textrm{in}},\sigma^{2})}{\mathcal{N}(t_{x};\mu_{\textrm{out}},\sigma^{2})}>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right)$ (4)
-$=\Pr\left(-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right).$ (5)
-From now on, we will denote the $-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]$ with $\lambda_{\textsc{fpr}}$ when $x\notin\mathcal{D}_{\textrm{target}}$ and with $\lambda_{\textsc{tpr}}$ otherwise.
-Furthermore we will drop the condition $x\in\mathcal{D}_{\textrm{target}}$ or $x\notin\mathcal{D}_{\textrm{target}}$ from the fpr and tpr expressions
-fpr $:=\Pr(\lambda_{\textsc{fpr}}>\tau)$ (6)
-tpr $:=\Pr(\lambda_{\textsc{tpr}}>\tau).$ (7)
-Lemma 1.
-In the limit of infinite shadow models and with large enough $S$, the LiRA test statistics $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ obtain the following forms
-$\lambda_{\textsc{fpr}}$ $=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right)$ (8)
-$\lambda_{\textsc{tpr}}$ $\approx-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle-\frac{2}{S}||x||^{2}+\frac{1}{S}\langle x,x-m_{y}\rangle\right),$ (9)
-where $z\sim\mathcal{N}(0,I_{d})$.
-Proof.
-We defer the proof to Appendix C. ∎
-The factors independent of $S$ in for both $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ correspond to the $-(\mu_{\textrm{in}}-\mu_{\textrm{out}})/(2\sigma^{2})$.
-Figure A.5 illustrates how this factor behaves as a function of $S$ in our experiments with the simple model.
-As predicted by the theory, the factor (i.e.
-the difference in means divided by the variance) remains fairly constant across different values of $S$.
-Using Lemma 1 we can derive a rate for the decision threshold $\tau$:
-Corollary 2.
-In order to maintain constant fpr, we need to adjust $\tau$ as $O(1/\sqrt{S})$.
-According to Corollary 2, the $\tau:=\tau(S)$ evolves as $O(1/\sqrt{S})$, and we will approximate it as $\tau(S)=\gamma_{1}\frac{1}{\sqrt{S}}$, with constant $\gamma_{1}$ depending on the target fpr.
-Theorem 3.
-$\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log(S)$.
-Proof.
-From Lemma 1, we see that $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ differ only in a single term which evolves as $O(1/S)$.
-Thus for sufficiently large $S$ we can consider this additional r.v.
-as a small constant shift to the distribution of $\lambda_{\textsc{fpr}}$.
-Let us denote this variable now with $q$.
-When $S$ grows, $q$ becomes small, and we have
-[…]
-$\approx p_{\lambda_{\textsc{fpr}}}(\tau)q.$ (11)
-Taking the log yields
-$\log(\textsc{tpr}-\textsc{fpr})$ $\approx\log\left(p_{\lambda_{\textsc{fpr}}}\left(\tau\right)\right)+\log q$ (12)
-$=\log\left(p_{\lambda_{\textsc{fpr}}}\left(\tau\right)\right)-\log S+\gamma_{2}.$ (13)
-From Lemma 1, we can see that a scaled r.v.
-$\hat{\lambda}_{\textsc{fpr}}:=\sqrt{S}\lambda_{\textsc{fpr}}$ becomes almost independent of $S$ for large enough $S$.
-Hence using the change of variables formula, we have that $p_{\lambda_{\textsc{fpr}}}(t)=p_{\hat{\lambda}_{\textsc{fpr}}}(\sqrt{S}t)\sqrt{S}$.
-Now, evaluating $p_{\lambda_{\textsc{fpr}}}(t)$ at $t=\tau(S)$, we get $p_{\lambda_{\textsc{fpr}}}(t)=p_{\hat{\lambda}_{\textsc{fpr}}}(\gamma_{1})\sqrt{S}$, where the $p_{\hat{\lambda}_{\textsc{fpr}}}(\gamma_{1})$ is independent of $S$.
-Therefore, ignoring additive constants,
-$\log(\textsc{tpr}-\textsc{fpr})$ $\approx\log\sqrt{S}-\log S=\frac{1}{2}\log S-\log S=-\frac{1}{2}\log S.$ (14)
-This shows that $\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log S$. ∎
-Compared to the model in Section 3.3, our theoretical model predicts $\beta_{S}=-1/2$.
-Our empirical estimates match this for high fpr.
-For lower FPR, $\beta_{S}$ decreases, which indicates there may be finer details of $p_{\lambda_{\textsc{fpr}}}$ we are missing in this analysis.
-5 MIA Vulnerability For Data Samples
-This section extends our investigation from the data-set-level properties (Section 3) to individual data-sample-level properties.
-Previously Yu et al.
-(2023) have studied correlations between classes and privacy risk before, but in comparison to their study, we look at data samples independent from classes.
-We investigate two properties that could potentially influence the vulnerability of data points:
-  • Per-sample gradients: Per-sample gradients form the basis for DP individual accounting (Feldman and Zrnic, 2021; Koskela et al., 2023).
-Under individual Gaussian DP (GDP, Dong et al., 2022) accounting (Koskela et al., 2023), the cumulative $\mu$-GDP privacy loss of individual $i$ with individual gradient norms $g_{i,t}$ is $\mu_{\text{i,tot}}\propto\sqrt{\sum_{t}g_{i,t}^{2}}$.
-Our hypothesis is that this similarly holds also for MIA vulnerability in non-DP settings.
-  • Feature space distances: The pre-trained feature extractors project images into a feature space in which one can compute distances between individual data samples.
-Our hypothesis is that data samples that are further away from other data samples are easier to distinguish.
-5.1 Experimental Setup
-Our model training is done as in Section 3.1.
-The target models are fine-tuned pre-trained ViT-B (Head) and R-50 (Head) models.
-Below we describe how we measure the sample properties and individual tpr at fpr.
-Recording of gradient norms During the training of the target models we record the per-sample gradient norm for every sample at every step (independent of being part of the current batch) efficiently using Opacus (Yousefpour et al., 2021).
-The recording of the gradient norms is identical to the DP individual accounting, with the exception that we record the gradient norms of out-samples as well.
-In our analysis, we are utilising the square root of the sum of squared gradient norm per sample over deciles of steps (i.e.
-the sum of squared individual gradient norms over the first 10% of steps, second 10% of steps, …, last 10% of steps).
-$G_{i}^{(\alpha)}=\sqrt{\sum_{t=\lfloor T(\alpha-0.1)\rfloor}^{\lfloor T\alpha\rfloor}g_{i,t}^{2}},\;\alpha=(0.1,0.2,\ldots,1.0)$ (15)
-Computation of distance in feature space All samples are projected into a feature space using the pre-trained model.
-We measure the Euclidean distance of all examples (in and out) to all in-samples in the feature space and report the distance to the nearest in-sample.
-tpr at fpr for individual samples We study individual samples, but aggregate the results over multiple data samples by binning the samples based on their data sample properties to reduce the noise.
-For each data sample we compute its property and assign the sample to a bin based on its percentile in relation to other samples.
-We then compute the tpr at fpr for each bin.
-We finally report for each data property bin the percentile of the tpr at fpr.
-Throughout our experiments we noticed that the amount of data points needs to be sufficiently large, in order to allow for an appropriate estimation of the tpr at fpr.
-Thus, we use a sufficient amount of shots ($S$) so that the size of in- and out-set combined is at least $10002|\mathcal{D}|=$ 1000 (see Table A10).
-5.2 Individual Gradient Norms Are Predictive of Individual MIA Vulnerability
-Figure 5 shows the observed relationships between the MIA vulnerability and the average individual gradient norms at $\textsc{fpr}=0.001$.
-When plotting the data for all deciles of steps at once, we find that the average gradient norms are correlated with the vulnerability in all our experiments, but the strength of the correlation depends on the data set.
-The $R^{2}$ (coefficient of determination) compares the observed relationship to a scenario where individual gradient norms would be a perfect predictor (corresponding to a scenario where all points are on the diagonal line).
-See more results in Table A10 and Figure A.6 in the Appendix.
-Figure 5: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed).
-Data obtained over six seeds.
-The $R^{2}$ score compares the observed relationship to a scenario where individual gradient norms would be a perfect predictor.
-Which parts of the training are more predictive?
-In Figure 6, we examine whether the gradient norms at different stages of training are a better predictor of vulnerability by computing the ratio between the $R^{2}$ per iteration decile and $R^{2}$ over the whole training.
-For most data sets the ratio is constantly at a value of roughly 1, which suggests that gradient norms over even short segments can be as predictive as the norms of the whole trace.
-Furthermore, there are deviations that indicate that for some data sets, gradient norms during the early iterations can be relatively more predictive than the norms of the entire trace.
-This is interesting, as the DP bound would directly depend on the norms of the entire trace.
-Figure 6: Ratio between $R^{2}$ per iteration decile and $R^{2}$ over the whole training.
-A ratio $>1$ means that the decile is as more predictive than whole trace and a ratio $<1$ means that it is less predicitive.
-5.3 Distance to Neighbors in Feature Space of Pre-trained Model
-Figure 7 shows the observed relationships between the MIA vulnerability and distance to the nearest neighbor in the feature space.
-The distance of data samples to their nearest neighbor correlates with the vulnerability all analysed data sets but Patch Camelyon and Resics45.
-The observation holds over all fpr and for R-50 (Head) as well, refer to Figures A.9 and A.10 in the Appendix.
-Figure 7: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
 5 Discussion
@@
 This will help in evaluating the risk of different kinds of models, favouring less risky models when possible and paying extra attention to more risky cases.
-As the best means of protecting privacy, DP, reduces the utility of models, it is important to understand when it is necessary.
+As the best means of protecting privacy, differential privacy, reduces the utility of models, it is important to understand when it is necessary.
 Our results highlight that models are most vulnerable to MIA when the number of examples per class is low.
-This leads to an unfortunate dilemma: DP seems most important in the regime where it likely reduces the model utility the most.
+This reinforces the well-known unfortunate dilemma: differential privacy seems most important in the regime where it likely reduces the model utility the most.
+One major reason for MIA vulnerability can be memorisation of the training data.
+Feldman & Zhang (2020) experimentally test memorisation in neural network training, and find that according to their definition, a large fraction of the training data are memorised when training from scratch, while only few are when fine-tuning.
+This is aligned with our results that indicate from scratch training to be more vulnerable than fine-tuning.
 Our model singles out the number of examples per class $S$ as the key predictor of model vulnerability.
@@
 Limitations
-Our work measures the vulnerability through running one particular MIA.
-There is no guarantee that models that are not vulnerable to this attack might not be vulnerable to some stronger attack.
-Formal bounds on attack susceptibility would require something like DP.
-Due to computational constraints, our predictive model is based only on transfer learning with fine-tuning only a part of the model.
-Limited comparison of our model predictions to previous results indicates it may underestimate their vulnerability.
-Broader Impact
-Our work systematically studies factors influencing privacy risk of trained ML models.
-This has significant positive impact on users training ML models on personal data by allowing them to understand and limit the risks.
-Reproducibility
-We provide the code for reproducing the experiments in the supplementary material and will publish the code in an open repository after acceptance of the paper.
-All used pre-trained models and data sets are publicly available.
-Acknowledgements
+Our work measures the vulnerability through running two particular MIAs.
+There is no guarantee that models that are not vulnerable to this attack might not be vulnerable to some stronger attacks.
+Formal bounds on attack susceptibility would require something like differential privacy.
+Both our theoretical and empirical analysis focus on deep transfer learning using fine-tuning.
+Models trained from scratch are likely to be more vulnerable.
+Acknowledgments
 This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617).
@@
 Neither the European Union nor the granting authority can be held responsible for them.
-The authors wish to thank the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources.
-We thank John F.
+The authors wish to thank the CSC – IT Center for Science, Finland for supporting this project with computational and data storage resources.
+We thank Mikko A.
+Heikkilä and Ossi Räisä for helpful comments and suggestions and John F.
 Bronskill for helpful discussions regarding few-shot learning.
-Appendix A Training details
-A.1 Parameterization
+Appendix A Formulating LiRA and RMIA for Section 3
+Let $\mathcal{M}$ be our target model and $\ell(\mathcal{M}(x),y_{x})$ be the loss of the model on a target example $(x,y_{x})$.
+The goal of MIA is to determine whether $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$.
+This can be formulated as a hypothesis test:
+$H_{0}:(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$ (A1)
+$H_{1}:(x,y_{x})\in\mathcal{D}_{\mathrm{target}}.$ (A2)
+A.1 LiRA
+Following (Carlini et al., 2022), we formulate the Likelihood Ratio Attack (LiRA).
+LiRA exploits the difference of losses on the target model under $H_{0}$ and $H_{1}$.
+To model the IN/OUT loss distributions with few shadow models, LiRA employs a parametric modelling.
+Particularly, LiRA models $t_{x}$ by a normal distribution.
+That is, the hypothesis test formulated above can be rewritten as
+$H_{0}:t_{x}\sim\mathcal{N}(\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})$ (A3)
+$H_{1}:t_{x}\sim\mathcal{N}(\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}(x)^{2}).$ (A4)
+The likelihood ratio is now
+$\mathrm{LR}(x)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x)^{2},\hat{\sigma}_{\mathrm{in}}(x)^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})}.$ (A5)
+LiRA rejects $H_{0}$ if and only if
+$\mathrm{LR}(x)\geq\beta,$ (A6)
+concluding that $H_{1}$ is true, i.e., identifying the membership of $(x,y_{x})$. Thus, the true positive rate of this hypothesis test given as
+$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}(x)^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})}\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\right),$ (A7)
+where $\phi^{M}$ denotes the randomness in the shadow set sampling and shadow model training.
+A.2 RMIA
+By modelling $t_{z}$ by a normal distribution, Zarifzadeh et al. (2024) approximate the pairwise likelihood ratio as
+$\mathrm{LR}(x,z)=\frac{p(\theta\mid x,y_{x})}{p(\theta\mid z,y_{z})}\approx\frac{\mathcal{N}(t_{x};\hat{\mu}_{x,x},\hat{\sigma}_{x,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{x,z},\hat{\sigma}_{x,z}^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{z,x},\hat{\sigma}_{z,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{z,z},\hat{\sigma}_{z,z}^{2})},$ (A8)
+where $\hat{\mu}_{a,b}$ and $\hat{\sigma}_{a,b}$ are, respectively, the mean and standard deviation of $t_{b}$ estimated from shadow models when the training set contains $a$ but not $b$.
+Then RMIA exploits the probability of rejecting the pairwise likelihood ratio test over $(z,y_{z})\sim{\mathbb{D}}$:
+$\mathrm{Score}_{\mathrm{RMIA}}(x)=\Pr_{(z,y_{z})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,z)\geq\gamma\right).$ (A9)
+Thus, RMIA rejects $H_{0}$ if and only if
+$\Pr_{(z,y_{z})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta,$ (A10)
+identifying the membership of $x$. Hence the true positive rate of RMIA is given as
+$\textsc{tpr}_{\mathrm{RMIA}}(x)=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\Pr_{(z,y_{z})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\right),$ (A11)
+where $\phi^{M}$ denotes the randomness in the shadow set sampling and shadow model training.
+Appendix B On the assumptions in Section 3
+B.1 The assumption of shared variances
+In Section 3 we assumed that for LiRA $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}=\hat{\sigma}_{\mathrm{out}}$, and that for RMIA $\sigma_{x}=\sigma_{x,x}=\sigma_{z,x}$, $\sigma_{z}=\sigma_{x.z}=\sigma_{z,z}$, $\hat{\sigma}_{x}=\hat{\sigma}_{x,x}=\hat{\sigma}_{z,x}$ and $\hat{\sigma}_{z}=\hat{\sigma}_{x.z}=\hat{\sigma}_{z,z}$.
+Utilising the simplified model formulated in Section 3.4, we show that for large enough number $S$ of examples per class these assumptions are reasonable.
+From the proof of Theorem 3 (see Section C.3), the LiRA scores are
+$s_{y_{x}}^{\mathrm{(in)}}$ $=\langle x,\frac{1}{S}\left(\sum_{i=1}^{S-1}x_{y_{x},i}+x\right)\rangle=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle+\langle x,\frac{1}{S}(x-x_{y_{x},S})\rangle$ (A12)
+$s_{y_{x}}^{\mathrm{(out)}}$ $=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle,$ (A13)
+respectively, when $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$ and when $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$.
+Then, since for the simplified model scores follow Gaussian distributions, we have
+$\sigma_{\mathrm{in}}^{2}$ $=\hat{\sigma}_{\mathrm{in}}^{2}=\operatorname{Var}(s_{y_{x}}^{\mathrm{(in)}})=\frac{1}{S}\left(1-\frac{1}{S}\right)||x||^{2}s^{2}$ (A14)
+$\sigma_{\mathrm{out}}^{2}$ $=\hat{\sigma}_{\mathrm{out}}^{2}=\operatorname{Var}(s_{y_{x}}^{\mathrm{(out)}})=\frac{1}{S}||x||^{2}s^{2}.$ (A15)
+Thus, the differences $\sigma_{\mathrm{in}}-\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}-\hat{\sigma}_{\mathrm{out}}$ are negligible for large $S$. Similarly, we have
+$\sigma_{x,x}^{2}$ $=\hat{\sigma}_{x,x}^{2}=\operatorname{Var}(s_{y_{x}}^{(x)}(x))=\frac{1}{S}\left(1-\frac{1}{S}\right)||x||^{2}s^{2}$ (A16)
+$\sigma_{z,x}^{2}$ =$\hat{\sigma}_{z,x}^{2}=\operatorname{Var}(s_{y_{x}}^{(z)}(x))=\begin{cases}\frac{1}{S}\left(1-\frac{1}{S}\right)||x||^{2}s^{2}\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}||x||^{2}s^{2}\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A17)
+$\sigma_{x,z}^{2}$ =$\hat{\sigma}_{x,z}^{2}=\operatorname{Var}(s_{y_{z}}^{(x)}(z))=\begin{cases}\frac{1}{S}\left(1-\frac{1}{S}\right)||z||^{2}s^{2}\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}||z||^{2}s^{2}\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A18)
+$\sigma_{z,z}^{2}$ $=\hat{\sigma}_{z,z}^{2}=\operatorname{Var}(s_{y_{z}}^{(z)}(z))=\frac{1}{S}\left(1-\frac{1}{S}\right)||z||^{2}s^{2}.$ (A19)
+Therefore, the differences $\sigma_{x,x}-\sigma_{z,x}$, $\sigma_{x.z}-\sigma_{z,z}$, $\hat{\sigma}_{x,x}-\hat{\sigma}_{z,x}$ and $\hat{\sigma}_{x.z}-\hat{\sigma}_{z,z}$ are negligible for large enough $S$.
+Hence as long as the simplified model approximates classification tasks to which Lemmas 1 and 2 are applied, these assumptions are reasonably justified.
+B.2 Relaxing the assumption of Lemma 1
+In Lemma 1 we assume that an attacker has access to the true underlying distribution.
+However, this is not necessarily the case in real-world settings.
+Noting that the Equation 12 mainly relies on the true location parameters $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and scale parameter $\sigma(x)$, we may relax this assumption.
+Notice that if we completely drop this assumption so that an attacker trains shadow models with an arbitrary underlying distribution, then we may not be able to choose a desired $\textsc{fpr}_{\mathrm{LiRA}}(x)$.
+From Equation A36 in the proof of Lemma 1 we have
+$\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}=\begin{cases}F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}\end{cases}$ (A20)
+$\log\beta=\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}},\end{cases}$ (A21)
+where we abuse notations by denoting $\mu_{\mathrm{in}}$ to refer to $\mu_{\mathrm{in}}(x)$ and similarly for other parameters.
+Since we need to choose a rejection region of the likelihood ratio test such that $\beta\geq 1$, we have
+$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A22)
+Therefore, the sufficient condition about attacker’s knowledge on the underlying distribution for Lemma 1 to hold is
+$\begin{cases}\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\leq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A23)
+We summarise this discussion in the following:
+Lemma 7 (Lemma 1 with a relaxed assumption).
+Suppose that the true distribution of $t_{x}$ is of location-scale family with locations $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and scale $\sigma(x)$, and that LiRA models $t_{x}$ by $\mathcal{N}(\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}(x))$ and $\mathcal{N}(\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}(x))$.
+Assume that an attacker estimates parameters $\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}(x),\hat{\mu}_{\mathrm{out}}(x)$ and $\hat{\sigma}_{\mathrm{out}}(x)$ with an approximated underlying distribution such that
+$\begin{cases}\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}(x)+\hat{\mu}_{\mathrm{out}}(x)}{2}+\mu_{\mathrm{out}}(x)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}(x)>\hat{\mu}_{\mathrm{out}}(x)\\ \sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}(x)+\hat{\mu}_{\mathrm{out}}(x)}{2}+\mu_{\mathrm{out}}(x)\leq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}(x)<\hat{\mu}_{\mathrm{out}}(x).\end{cases}$ (A24)
+Then the LiRA vulnerability of a fixed target example is
+$\textsc{tpr}_{\mathrm{LiRA}}(x)=\begin{cases}1-F_{t}\left(F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)>\hat{\mu}_{\mathrm{out}}(x)\\ F_{t}\left(F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)<\hat{\mu}_{\mathrm{out}}(x),\end{cases}$ (A25)
+where $F_{t}$ is the cdf of $t$ with the standard location and unit scale.
+Appendix C Missing proofs of Section 3
+C.1 Proof of Lemma 1
+See 1
+Proof.
+We abuse notations by denoting $\mu_{\mathrm{in}}$ to refer to $\mu_{\mathrm{in}}(x)$ and similarly for other statistics. We have
+$\log\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}},\hat{\sigma})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}},\hat{\sigma})}\geq\log\beta$ (A26)
+$-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{\mathrm{in}}}{\hat{\sigma}}\right)^{2}+\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}}\right)^{2}\geq\log\beta$ (A27)
+$\frac{1}{2\hat{\sigma}^{2}}(2t_{x}\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{in}}^{2}-2t_{x}\hat{\mu}_{\mathrm{out}}+\hat{\mu}_{\mathrm{out}}^{2})\geq\log\beta$ (A28)
+$\frac{1}{2\hat{\sigma}^{2}}(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})(2t_{x}-\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})\geq\log\beta$ (A29)
+$\begin{cases}t_{x}\geq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ t_{x}\leq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A30)
+When $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$, $t_{x}\sim\mathcal{N}(\mu_{\mathrm{out}},\sigma)$.
+Therefore, if $\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}$, in the limit of infinitely many shadow models
+$\textsc{fpr}_{\mathrm{LiRA}}(x)$ $=\Pr_{t}\left(\mu_{\mathrm{out}}+\sigma t\geq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A31)
+$=\Pr_{t}\left(t\geq\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}\right)$ (A32)
+$=1-F_{t}\left(\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}\right),$ (A33)
+and if $\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}$, similarly,
+$\textsc{fpr}_{\mathrm{LiRA}}(x)$ $=\Pr_{t}\left(\mu_{\mathrm{out}}+\sigma t\leq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A34)
+$=F_{t}\left(\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}\right).$ (A35)
+Thus
+$\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}=\begin{cases}F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A36)
+It follows that if $\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}$,
+$\textsc{tpr}_{\mathrm{LiRA}}(x)$ $=\Pr_{t}\left(\mu_{\mathrm{in}}+\sigma t\geq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A37)
+$=\Pr_{t}\left(t\geq\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{in}}}{\sigma}\right)$ (A38)
+$=1-F_{t}\left(F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right).$ (A39)
+If $\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}$, then
+$\textsc{tpr}_{\mathrm{LiRA}}(x)$ $=\Pr_{t}\left(\mu_{\mathrm{in}}+\sigma t\leq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A40)
+$=\Pr_{t}\left(t\leq\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{in}}}{\sigma}\right)$ (A41)
+$=F_{t}\left(F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right).$ (A42)
+C.2 Proof of Lemma 2
+See 2
+Proof.
+We abuse notations by denoting probabilities and expectations over sampling $\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|}$ and $(z,y_{z})\sim{\mathbb{D}}$ by subscripts $\mathcal{D}_{\mathrm{target}}$ and $z$.
+Recall that
+$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{\mathcal{D}_{\mathrm{target}}}\left(\Pr_{z}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\right)$ (A43)
+$\textsc{fpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{\mathcal{D}_{\mathrm{target}}}\left(\Pr_{z}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta\mid(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\wedge(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\right),$ (A44)
+where
+$\mathrm{LR}(x,z)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{x,x},\hat{\sigma}_{x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{x,z},\hat{\sigma}_{z}^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{z,x},\hat{\sigma}_{x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{z,z},\hat{\sigma}_{z}^{2})}.$ (A45)
+We have
+$\lambda:=$ $\log(\mathrm{LR}(x,z))$ (A46)
+$=$ $\log\left(\frac{\frac{1}{\sqrt{2\pi}\hat{\sigma}_{x}}\exp\left(-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{x,x}}{\hat{\sigma}_{x}}\right)^{2}\right)\frac{1}{\sqrt{2\pi}\hat{\sigma}_{z}}\exp\left(-\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{x,z}}{\hat{\sigma}_{z}}\right)^{2}\right)}{\frac{1}{\sqrt{2\pi}\hat{\sigma}_{x}}\exp\left(-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{z,x}}{\hat{\sigma}_{x}}\right)^{2}\right)\frac{1}{\sqrt{2\pi}\hat{\sigma}_{z}}\exp\left(-\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{z,z}}{\hat{\sigma}_{z}}\right)^{2}\right)}\right)$ (A47)
+$=$ $-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{x,x}}{\hat{\sigma}_{x}}\right)^{2}+\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{z,x}}{\hat{\sigma}_{x}}\right)^{2}-\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{x,z}}{\hat{\sigma}_{z}}\right)^{2}+\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{z,z}}{\hat{\sigma}_{z}}\right)^{2}$ (A48)
+$=$ $\frac{1}{2\hat{\sigma}_{x}^{2}}(2t_{x}\hat{\mu}_{x,x}-\hat{\mu}_{x,x}^{2}-2t_{x}\hat{\mu}_{z,x}+\hat{\mu}_{z,x}^{2})+\frac{1}{2\hat{\sigma}_{z}^{2}}(2t_{z}\hat{\mu}_{x,z}-\hat{\mu}_{x,z}^{2}-2t_{z}\hat{\mu}_{z,z}+\hat{\mu}_{z,z}^{2})$ (A49)
+$=$ $\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2t_{x}-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2t_{z}-\hat{\mu}_{x,z}-\hat{\mu}_{z,z}).$ (A50)
+When $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$, $t_{x}=\mu_{x,x}+\sigma_{x}t$ and $t_{z}=\mu_{x,z}+\sigma_{z}t$, and thus $\lambda$ becomes
+$\lambda_{x}:=$ $\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2\mu_{x,x}+2\sigma_{x}t-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})$ (A51)
+$+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2\mu_{x,z}+2\sigma_{z}t-\hat{\mu}_{x,z}-\hat{\mu}_{z,z}).$ (A52)
+Similarly, when $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$, we have $t_{x}=\mu_{z,x}+\sigma_{x}t$ and $t_{z}=\mu_{z,z}+\sigma_{z}t$. Then $\lambda$ is now
+$\lambda_{z}:=$ $\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2\mu_{z,x}+2\sigma_{x}t-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})$ (A53)
+$+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2\mu_{z,z}+2\sigma_{z}t-\hat{\mu}_{x,z}-\hat{\mu}_{z,z})$ (A54)
+$=$ $\left(\underbrace{\frac{\sigma_{x}}{\hat{\sigma}_{x}^{2}}(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\sigma_{z}}{\hat{\sigma}_{z}^{2}}(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}_{A}\right)t$ (A55)
+$+\underbrace{\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2\mu_{z,x}-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2\mu_{z,z}-\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}_{B}$ (A56)
+$=$ $At+B.$ (A57)
+Notice that $A$ and $B$ are functions of $z$ and independent of $t$.
+Also note that taking probability over $t$ corresponds to calculating probability over $\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|}$.
+Thus, Since we can set $\gamma>1$, using Markov’s inequality and the triangle inequality, we have in the limit of infinitely many shadow models
+$\textsc{fpr}_{\mathrm{RMIA}}(x)=$ $\Pr_{t}\left(\Pr_{z}(\lambda_{z}\geq\log(\gamma))\geq\beta\right)$ (A58)
+$\leq\Pr_{t}\left(\Pr_{z}(|\lambda_{z}|\geq\log(\gamma))\geq\beta\right)$ (A59)
+$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{z}|]}{\log(\gamma)}\geq\beta\right)$ (A60)
+$=\Pr_{t}\left(\mathbb{E}_{z}[|\lambda_{z}|]\geq\beta\log(\gamma)\right)$ (A61)
+$\leq\Pr_{t}\left(\mathbb{E}_{z}[|A|]|t|+\mathbb{E}_{z}[|B|]\geq\beta\log(\gamma)\right)$ (A62)
+$=\Pr_{t}\left(|t|\geq\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}\right)$ (A63)
+Therefore, we can upper-bound $\textsc{fpr}_{\mathrm{RMIA}}(x)\leq\alpha$ by setting
+$\alpha=1-F_{|t|}\left(\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}\right).$ (A64)
+That is,
+$\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}=F_{|t|}^{-1}(1-\alpha).$ (A65)
+Now let
+$q=\lambda_{x}-\lambda_{z}=\frac{(\mu_{x,x}-\mu_{z,x})(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})}{\hat{\sigma}_{x}^{2}}+\frac{(\mu_{x,z}-\mu_{z,z})(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}{\hat{\sigma}_{z}^{2}}.$ (A66)
+Note that $q$ is also independent of $t$, thereby $\mathbb{E}_{z}[|q|]$ being a constant. By the similar argument using Markov’s inequality and the triangle inequality, we have
+$\textsc{tpr}_{\mathrm{RMIA}}(x)=$ $\Pr_{t}\left(\Pr_{z}(\lambda_{x}\geq\log(\gamma))\geq\beta\right)$ (A67)
+$\leq\Pr_{t}\left(\Pr_{z}(|\lambda_{x}|\geq\log(\gamma))\geq\beta\right)$ (A68)
+$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{x}|]}{\log(\gamma)}\geq\beta\right)$ (A69)
+$=\Pr_{t}\left(\mathbb{E}_{z}[|\lambda_{x}|]\geq\beta\log(\gamma)\right)$ (A70)
+$\leq\Pr_{t}\left(\mathbb{E}_{z}[|A|]|t|+\mathbb{E}_{z}[|B|]+\mathbb{E}_{z}[|q|]\geq\beta\log(\gamma)\right)$ (A71)
+$=\Pr_{t}\left(|t|\geq\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right)$ (A72)
+Hence we obtain
+$\textsc{tpr}_{\mathrm{RMIA}}(x)\leq 1-F_{|t|}^{-1}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right).$ (A73)
+C.3 Proof of Theorem 3
+See 3
+Proof.
+Let $\mathcal{D}_{\mathrm{target}}=\{(x_{j,1},j),...,(x_{j,S},j)\}_{j=1}^{C}$. Then the LiRA score of the target $(x,y_{x})$ is
+$s_{y_{x}}^{\mathrm{(in)}}$ $=\langle x,\frac{1}{S}\left(\sum_{i=1}^{S-1}x_{y_{x},i}+x\right)\rangle=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle+\langle x,\frac{1}{S}(x-x_{y_{x},S})\rangle$ (A74)
+$s_{y_{x}}^{\mathrm{(out)}}$ $=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle,$ (A75)
+respectively, when $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$ and when $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$. Thus, in the limit of infinitely many shadow models we obtain
+$\mu_{\mathrm{in}}-\mu_{\mathrm{out}}$ $=\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}=\mathbb{E}[s_{y_{x}}^{\mathrm{(in)}}-s_{y_{x}}^{\mathrm{(out)}}]=\frac{1}{S}\langle x,x-m_{x}\rangle$ (A76)
+$\sigma^{2}$ $=\operatorname{Var}(s_{y_{x}}^{\mathrm{(out)}})=\frac{1}{S}\operatorname{Var}(\langle x,x_{y_{x},i}\rangle)=\frac{1}{S}||x||^{2}s^{2}$ (A77)
+Noting that the LiRA score follows a normal distribution, by Lemma 1 we have
+$\textsc{tpr}_{\mathrm{LiRA}}(x)$ =$=\begin{cases}1-\Phi\left(\Phi^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}},\end{cases}$ (A78)
+=$=\begin{cases}1-\Phi\left(\Phi^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}\right)&\text{ if }\langle x,x-m_{x}\rangle>0\\ \Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}\right)&\text{ if }\langle x,x-m_{x}\rangle<0,\end{cases}$ (A79)
+By the symmetry of a normal distribution, this simplifies to
+$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))+\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}s||x||}\right),$ (A80)
+where $\langle x,x-m_{x}\rangle\not=0$ by the assumption that $\hat{\mu}_{\mathrm{in}}\not=\hat{\mu}_{\mathrm{out}}$. This completes the first half of the theorem.
+Now we have
+$\textsc{tpr}_{\mathrm{LiRA}}(x)$ $=\Pr\left(\eta\leq\gamma_{0}+\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}s||x||}\right),$ (A81)
+$\textsc{fpr}_{\mathrm{LiRA}}(x)$ $=\Pr(\eta\leq\gamma_{0}),$ (A82)
+where $\gamma_{0}$ is a tunable constant and $\eta\sim\mathcal{N}(0,1)$. Thus for large enough $S$ we have
+$\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x)$ $\approx p_{\eta}(\gamma_{0})\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}||x||s}$ (A83)
+$=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}}\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}||x||s},$ (A84)
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x))$ $\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}+\log\frac{|\langle x,x-m_{x}\rangle|}{||x||s\sqrt{2\pi}}.$ (A85)
+C.4 Proof of Theorem 4
+See 4
+Proof.
+To apply Lemma 2, we will calculate $\mathbb{E}_{z}[q]$ and $\mathbb{E}_{z}[A]$.
+Let $s_{y_{x}}^{(x)}(x)$ (resp.
+$s_{y_{x}}^{(z)}(x)$) denote the score of the target $x$ for class $y_{x}$ when the dataset contains $(x,y_{x})$ but not $(z,y_{z})$ (resp.
+when the dataset contains $(z,y_{z})$ but not $(x,y_{x})$).
+Let $s_{y_{z}}^{(x)}(z)$ and $z_{y_{z}}^{(z)}(z)$ be corresponding scores of example $z$.
+Then we have
+$s_{y_{x}}^{(x)}(x)$ $=\frac{1}{S}\langle x,\sum_{i=1}^{S}x_{i}+x-x_{S}\rangle$ (A86)
+$s_{y_{x}}^{(z)}(x)$ =$\begin{cases}\frac{1}{S}\langle x,\sum_{i=1}^{S}x_{i}+z-x_{S}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}\langle x,\sum_{i=1}^{S}x_{i}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A87)
+$s_{y_{z}}^{(x)}(z)$ =$\begin{cases}\frac{1}{S}\langle z,\sum_{i=1}^{S}z_{i}+x-z_{S}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}\langle z,\sum_{i=1}^{S}z_{i}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A88)
+$s_{y_{z}}^{(z)}(z)$ $=\frac{1}{S}\langle z,\sum_{i=1}^{S}z_{i}+z-x_{S}\rangle$ (A89)
+where $x_{i}$ are samples with label $y_{x}$ and $z_{i}$ are samples with label $y_{z}$ when $y_{x}\not=y_{z}$. Thus we obtain
+$\mu_{x,x}$ $=\langle x,m_{x}\rangle+\frac{1}{S}\langle x,x-m_{x}\rangle$ (A90)
+$\mu_{z,x}$ =$\begin{cases}\langle x,m_{x}\rangle+\frac{1}{S}\langle x,z-m_{x}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \langle x,m_{x}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A91)
+$\mu_{x,z}$ =$\begin{cases}\langle z,m_{x}\rangle+\frac{1}{S}\langle z,x-m_{x}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \langle z,m_{z}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A92)
+$\mu_{z,z}$ $=\langle z,m_{x}\rangle+\frac{1}{S}\langle z,z-m_{x}\rangle\quad$ (A93)
+$\sigma_{x}$ $=\frac{1}{\sqrt{S}}s||x||$ (A94)
+$\sigma_{z}$ $=\frac{1}{\sqrt{S}}s||z||,$ (A95)
+where $m_{z}$ is the true class mean of $y_{z}$ (see Appendix B for derivation of $\sigma_{x}$ and $\sigma_{z}$).
+Now recall that
+$q=$ $\frac{(\mu_{x,x}-\mu_{z,x})(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})}{\hat{\sigma}_{x}^{2}}+\frac{(\mu_{x,z}-\mu_{z,z})(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}{\hat{\sigma}_{z}^{2}}$ (A96)
+$A=$ $\frac{\sigma_{x}}{\hat{\sigma}_{x}^{2}}(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\sigma_{z}}{\hat{\sigma}_{z}^{2}}(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})$ (A97)
+In the limit of infinitely many shadow models, estimated statisitcs converge to true parameters, thereby yielding
+$q=$ $\left(\frac{\mu_{x,x}-\mu_{z,x}}{\sigma_{x}}\right)^{2}+\left(\frac{\mu_{x,z}-\mu_{z,z}}{\sigma_{z}}\right)^{2}$ (A98)
+$A=$ $\frac{\mu_{x,x}-\mu_{z,x}}{\sigma_{x}}+\frac{\mu_{x,z}-\mu_{z,z}}{\sigma_{z}}.$ (A99)
+Using the law of total expectation, we have
+$\mathbb{E}_{z}[|q|]=$ $\Pr_{z}(y_{z}=y_{x})\mathbb{E}_{z}[|q|\mid y_{z}=y_{x}]+\sum_{j=1,j\not=y_{x}}^{C}\Pr_{z}(y_{z}=j)\mathbb{E}_{z}[|q|\mid y_{z}=j]$ (A100)
+$=$ $\frac{1}{C}\mathbb{E}_{z}\left[\left.\left(\frac{\langle x,x-z\rangle}{\sqrt{S}s||x||}\right)^{2}+\left(\frac{\langle z,x-z\rangle}{\sqrt{S}s||z||}\right)^{2}\ \right|\ y_{z}=y_{x}\right]$ (A101)
+$+\frac{C-1}{C}\mathbb{E}_{z}\left[\left.\left(\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}\right)^{2}+\left(\frac{\langle z,z-m_{z}\rangle}{\sqrt{S}s||z||}\right)^{2}\ \right|\ y_{z}\not=y_{x}\right]$ (A102)
+$=$ $\frac{1}{CSs^{2}}\mathbb{E}_{z}\left[\left.\frac{\langle x,x-z\rangle^{2}}{||x||^{2}}+\frac{\langle z,x-z\rangle^{2}}{||z||^{2}}\ \right|\ y_{z}=y_{x}\right]$ (A103)
+$+\frac{C-1}{CSs^{2}}\mathbb{E}_{z}\left[\left.\frac{\langle x,x-m_{x}\rangle^{2}}{||x||^{2}}+\frac{\langle z,z-m_{z}\rangle^{2}}{||z||^{2}}\ \right|\ y_{z}\not=y_{x}\right],$ (A104)
+and
+$\mathbb{E}_{z}[|A|]=$ $\Pr_{z}(y_{z}=y_{x})\mathbb{E}_{z}[|A|\mid y_{z}=y_{x}]+\sum_{j=1,j\not=y_{x}}^{C}\Pr_{z}(y_{z}=j)\mathbb{E}_{z}[|A|\mid y_{z}=j]$ (A105)
+$=$ $\frac{1}{C}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-z\rangle}{\sqrt{S}s||x||}+\frac{\langle z,x-z\rangle}{\sqrt{S}s||z||}\right|\ \right|\ y_{z}=y_{x}\right]$ (A106)
+$+\frac{C-1}{C}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}+\frac{\langle z,z-m_{z}\rangle}{\sqrt{S}s||z||}\right|\ \right|\ y_{z}\not=y_{x}\right]$ (A107)
+$=$ $\frac{1}{C\sqrt{S}s}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-z\rangle}{||x||}+\frac{\langle z,x-z\rangle}{||z||}\right|\ \right|\ y_{z}=y_{x}\right]$ (A108)
+$+\frac{C-1}{C\sqrt{S}s}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-m_{x}\rangle}{||x||}+\frac{\langle z,z-m_{z}\rangle}{||z||}\right|\ \right|\ y_{z}\not=y_{x}\right].$ (A109)
+Hence we obtain by the Cauchy-Schwarz inequality
+$\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}=$ $\frac{1}{\sqrt{S}s}\cdot\frac{\mathbb{E}_{z}\left[\frac{\langle x,x-z\rangle^{2}}{||x||^{2}}+\frac{\langle z,x-z\rangle^{2}}{||z||^{2}}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[\frac{\langle x,x-m_{x}\rangle^{2}}{||x||^{2}}+\frac{\langle z,z-m_{z}\rangle^{2}}{||z||^{2}}\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{z}\left[\left|\frac{\langle x,x-z\rangle}{||x||}+\frac{\langle z,x-z\rangle}{||z||}\right|\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[\left|\frac{\langle x,x-m_{x}\rangle}{||x||}+\frac{\langle z,z-m_{z}\rangle}{||z||}\right|\mid y_{z}\not=y_{x}\right]}$ (A110)
+$\leq$ $\frac{1}{\sqrt{S}s}\cdot\frac{\mathbb{E}_{z}\left[2||x-z||^{2}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||^{2}+||z-m_{z}||^{2})\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{z}\left[2||x-z||\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||+||z-m_{z}||)\mid y_{z}\not=y_{x}\right]}.$ (A111)
+Now Lemma 2 yields
+$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $\leq 1-F_{|t|}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right)$ (A112)
+$\leq 1-F_{|t|}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\psi(x,C)}{\sqrt{S}s}\right)$ (A113)
+where $F_{|t|}$ is the cdf of the folded normal distribution and
+$\psi(x,C)=\frac{\mathbb{E}_{z}\left[2||x-z||^{2}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||^{2}+||z-m_{z}||^{2})\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{z}\left[2||x-z||\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||+||z-m_{z}||)\mid y_{z}\not=y_{x}\right]}.$ (A114)
+This completes the first half of the theorem.
+Now that from Lemma 2 we have
+$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{t}\left(\Pr_{z}(\lambda_{z}+q\geq\log\gamma)\geq\beta\right)$ (A115)
+$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{z}|]+\mathbb{E}_{z}[|q|]}{\log(\gamma)}\geq\beta\right)$ (A116)
+$=\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right)$ (A117)
+$\textsc{fpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{t}\left(\Pr_{z}(\lambda_{z}\geq\log\gamma)\geq\beta\right)$ (A118)
+$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{z}|]}{\log(\gamma)}\geq\beta\right)$ (A119)
+$=\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)\right).$ (A120)
+We claim that the bound for $\textsc{fpr}_{\mathrm{RMIA}}(x)$ (Equation A119) is tighter than that for $\textsc{tpr}_{\mathrm{RMIA}}(x)$ (Equation A116).
+Since this is not straightforward, we numerically show this in Table A1.
+Now, for large enough $S$ we obtain
+$\textsc{tpr}_{\mathrm{RMIA}}(x)-\textsc{fpr}_{\mathrm{RMIA}}(x)$ $\leq\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)-\left|\frac{\mathbb{E}_{z}[q]}{\mathbb{E}_{z}[A]}\right|\right)-\Pr_{t}(|t|\geq F_{|t|}^{-1}(1-\alpha))$ (A121)
+$\leq\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)+\frac{\psi(x,C)}{\sqrt{S}s}\right)-\Pr_{t}(|t|\leq F_{|t|}^{-1}(1-\alpha))$ (A122)
+$\approx p_{|t|}(F_{|t|}^{-1}(1-\alpha))\frac{\psi(x,C)}{\sqrt{S}}.$ (A123)
+Since $|t|$ follows the standard folded normal distribution with mean 00 and scale $1$, we have
+$p_{|t|}(F_{|t|}^{-1}(1-\alpha))=\frac{1}{\sqrt{\pi/2}}\exp\left(-\frac{1}{2}F_{|t|}^{-1}(1-\alpha)^{2}\right).$ (A124)
+Hence we obtain
+$\log(\textsc{tpr}_{\mathrm{RMIA}}(x)-\textsc{fpr}_{\mathrm{RMIA}}(x))\leq-\frac{1}{2}\log S-\frac{1}{2}F_{|t|}^{-1}(1-\alpha)^{2}+\log\frac{\psi(x,C)}{\sqrt{\pi/2}}.$ (A125)
+S FPR=0.2 FPR=0.1 FPR=0.01 FPR=0.001
+10 0.000 0.000 0.000 0.000
+100 0.000 0.000 0.000 0.000
+1000 0.000 0.000 0.041 0.068
+10000 0.329 0.353 0.365 0.368
+Table A1: The gap in the upper bound for $\textsc{tpr}_{\mathrm{RMIA}}(x)$ in Equation A116 minus the gap in the upper bound for $\textsc{fpr}_{\mathrm{RMIA}}(x)$ in Equation A119.
+5 random target examples are sampled from the simplified model distribution with $C=2,s=0.1,d=50$.
+Simulation is done with 100000 samples of test points $z$, 1000 samples of $t$ and $\gamma=1.0001$.
+Due to computational constraints, results are not stable especially for small $S$.
+C.5 Proof of Corollary 5
+See 5
+Proof.
+By theorem 3 and the law of unconscious statistician, we have for large $S$
+[…]
+[…]
+[…]
+$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\langle x,x-m_{x}\rangle}{\sqrt{2\pi}||x||s}\right].$ (A129)
+Note that here we fixed $\textsc{fpr}_{\mathrm{LiRA}}(x)=\overline{\textsc{fpr}}_{\mathrm{LiRA}}$ for all $x$. Then we obtain
+$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr})^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\langle x,x-m_{x}\rangle}{\sqrt{2\pi}||x||s}\right]\right).$ (A130)
+C.6 Proof of Corollary 6
+See 6
+Proof.
+By theorem 4 and the law of unconscious statistician, we have for large $S$
+[…]
+[…]
+[…]
+$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\alpha)^{2}}\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\psi(x,C)}{\sqrt{2\pi}}\right].$ (A134)
+Hence we obtain
+$\log(\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}})\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\alpha)^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\psi(x,C)}{\sqrt{2\pi}}\right]\right).$ (A135)
+Appendix D LiRA vulnerability for Small FPR
+In Section 3.4 we proved for the simplified model that for LiRA $\log(\textsc{tpr}-\textsc{fpr})\approx-\frac{1}{2}\log S$ ignoring additive constants when fpr is fixed (Corollary 5).
+In Section 4.3 we observed that the coefficient $\beta_{S}$ for $\log S$ is around $-0.5$ for larger fpr, aligning with the theoretical value.
+However, for smaller fpr the coefficient is smaller than $-0.5$.
+To understand this phenomenon, it is important to note that the power-law (Theorem 3 and Corollary 5) only holds for sufficiently large $S$.
+Thus, it can be explained that the difference of coefficient values $\beta_{S}$ for small and large fpr comes from the small-$S$ regime as follows.
+In the proof of Corollary 5 and Theorem 3 the only approximation that could introduce some bias is Equation A84. That is,
+$\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x)$ $=\Pr\left(\eta\leq\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))+r\right)-\Pr\left(\eta\leq\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))\right)$ (A136)
+$\approx p_{\eta}\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))\right)r,$ (A137)
+where $\eta\sim\mathcal{N}(0,1)$, $r$ is a shift that scales $O(1/\sqrt{S})$ and $p_{\eta}$ is the pdf of $\eta$.
+Figure A.1 numerically illustrates this approximation.
+It can be observed that for small $S$ the approximation does not hold, underestimating the true vulnerability.
+Particularly, this effect is remarkably larger for small $\textsc{fpr}_{\mathrm{LiRA}}(x)$.
+In other words, the simplified model overestimates the coefficient value $\beta_{S}$ for small $\textsc{fpr}_{\mathrm{LiRA}}(x)$ by underestimating the true vulnerability in the small-$S$ regime.
+(a) $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.1$
+(b) $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.00001$
+Figure A.1: True vs.
+approximated $\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x)$ for the simplified model with in-class standard deviation $s=0.1$ and dimension $d=100$.
+One target example $(x,y_{x})$ is randomly sampled from the underlying distribution.
+Numbers in the plots are percentages of the approximated vulnerabilities over the true vulnerabilities.
+Figures illustrate the approximation for different $S$ (a) when $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.1$ and (b) when $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.00001$.
+Appendix E Training details
+E.1 Parameterization
 We utilise pre-trained feature extractors BiT-M-R50x1 (R-50) (Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B) (Dosovitskiy et al., 2021) with 85.8M parameters, both pretrained on the ImageNet-21K dataset (Russakovsky et al., 2015).
@@
 While a diverse set of adapters has been proposed, we utilise FiLM as it has been shown to be competitive in prior work (Shysheya et al., 2023; Tobaben et al., 2023).
-A.1.1 Licenses and Access
+E.1.1 Licenses and access
 The licenses and means to access the model checkpoints can be found below.
-A.2 Hyperparameter tuning
+E.2 Hyperparameter tuning
 […]
@@
 learning rate 1e-7 1e-2
-A.3 Data sets
+E.3 Datasets
 Table A3 shows the datasets used in the paper.
@@
 CIFAR100 (Krizhevsky, 2009) 100 16 128 82.2% 87.6%
-A.3.1 Licenses and Access
+E.3.1 Licenses and access
 The licenses and means to access the datasets can be found below. We downloaded all datasets from TensorFlow datasets https://www.tensorflow.org/datasets but Resics45 which required manual download.
-A.4 Compute Resources
+E.4 Compute resources
 All experiments but the R-50 (FiLM) experiments are run on CPU with 8 cores and 16 GB of host memory.
@@
 The provided code implements this optimization.
-In Section 5 we record the per-sample gradients during the training.
-This makes the training much more expensive despite the use of efficient per-sample gradient computation using Opacus Yousefpour et al.
-(2021).
-The training of the considered models does still not exceed one hour per model as we only consider settings with relatively low $S$ for this part of the experiments.
 The R-50 (FiLM) experiments are significantly more expensive and utilise a NVIDIA V100 with 40 GB VRAM, 10 CPU cores and 64 GB of host memory.
@@
 We estimate that in total we spend around 7 days of V100 and some dozens of weeks of CPU core time but more exact measurements are hard to make.
-Appendix B Additional results
+Appendix F Additional results
 In this section, we provide tabular results for our experiments and additional figures that did not fit into the main paper.
-B.1 Additional results for Section 3
+F.1 Additional results for Section 4
 This Section contains additional results for Section 4.
-B.1.1 Vulnerability as a function of shots
-This section displays additional results to Figure 1(a) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.1, A3 and A4.
-(a) ViT-B Head tpr at $\textsc{fpr}=0.1$
-(b) R-50 Head tpr at $\textsc{fpr}=0.1$
-(c) ViT-B Head tpr at $\textsc{fpr}=0.01$
-(d) R-50 Head tpr at $\textsc{fpr}=0.01$
-(e) ViT-B Head tpr at $\textsc{fpr}=0.001$
-(f) R-50 Head tpr at $\textsc{fpr}=0.001$
+F.1.1 Vulnerability as a function of shots
+This section displays additional results to Figure 1(a) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A4 and A5.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Figure A.2: MIA vulnerability as a function of shots (examples per class) when attacking a pre-trained ViT-B and R-50 Head trained without DP on different downstream datasets.
@@
 dataset classes ($C$) shots ($S$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 2 256 0.272 0.076 0.022 512 0.195 0.045 0.011 1024 0.201 0.048 0.011 2048 0.178 0.041 0.010 4096 0.163 0.033 0.008 8192 0.143 0.026 0.006 16384 0.124 0.019 0.004 32768 0.118 0.016 0.003 65536 0.106 0.012 0.002 CIFAR10 (Krizhevsky, 2009) 10 8 0.911 0.574 0.324 16 0.844 0.526 0.312 32 0.617 0.334 0.183 64 0.444 0.208 0.106 128 0.334 0.159 0.084 256 0.313 0.154 0.086 512 0.251 0.103 0.051 1024 0.214 0.082 0.038 EuroSAT (Helber et al., 2019) 10 8 0.846 0.517 0.275 16 0.699 0.408 0.250 32 0.490 0.236 0.121 64 0.410 0.198 0.105 128 0.332 0.151 0.075 256 0.269 0.111 0.056 512 0.208 0.077 0.036 Pets (Parkhi et al., 2012) 37 8 0.937 0.631 0.366 16 0.745 0.427 0.227 32 0.588 0.321 0.173 Resics45 (Cheng et al., 2017) 45 32 0.671 0.405 0.235 64 0.534 0.289 0.155 128 0.445 0.231 0.121 256 0.367 0.177 0.088 CIFAR100 (Krizhevsky, 2009) 100 16 0.897 0.638 0.429 32 0.763 0.549 0.384 64 0.634 0.414 0.269
-B.1.2 Vulnerability as a function of the number of classes
-This section displays additional results to Figure 1(b) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A5 and A6.
-(a) ViT-B Head tpr at $\textsc{fpr}=0.1$
-(b) R-50 Head tpr at $\textsc{fpr}=0.1$
-(c) ViT-B Head tpr at $\textsc{fpr}=0.01$
-(d) R-50 Head tpr at $\textsc{fpr}=0.01$
-(e) ViT-B Head tpr at $\textsc{fpr}=0.001$
-(f) R-50 Head tpr at $\textsc{fpr}=0.001$
+F.1.2 Vulnerability as a function of the number of classes
+This section displays additional results to Figure 1(b) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.3, A6 and A7.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Figure A.3: MIA vulnerability as a function of $C$ (classes) when attacking a ViT-B and R-50 Head fine-tuned without DP on different datasets where the classes are randomly sub-sampled and $S=32$.
@@
 dataset shots ($S$) classes ($C$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 32 2 0.452 0.151 0.041 CIFAR10 (Krizhevsky, 2009) 32 2 0.404 0.146 0.060 4 0.560 0.266 0.123 8 0.591 0.318 0.187 EuroSAT (Helber et al., 2019) 32 2 0.309 0.111 0.050 4 0.356 0.144 0.064 8 0.480 0.233 0.123 Pets (Parkhi et al., 2012) 32 2 0.249 0.068 0.029 4 0.326 0.115 0.056 8 0.419 0.173 0.075 16 0.493 0.245 0.127 32 0.559 0.294 0.166 Resics45 (Cheng et al., 2017) 32 2 0.310 0.103 0.059 4 0.415 0.170 0.083 8 0.510 0.236 0.119 16 0.585 0.311 0.174 32 0.644 0.382 0.218 CIFAR100 (Krizhevsky, 2009) 32 2 0.356 0.132 0.054 4 0.423 0.176 0.087 8 0.545 0.288 0.163 16 0.580 0.338 0.196 32 0.648 0.402 0.244 64 0.711 0.476 0.320
-B.1.3 Data for FiLM and from scratch training
+F.1.3 Data for FiLM and from scratch training
 Table A8: MIA vulnerability data used in Figure 4(b).
@@
 (2022) 0.700 0.400 0.276 (Krizhevsky, 2009)
-B.1.4 Predicting data set vulnerability as function of $S$ and $C$
+F.1.4 Predicting dataset vulnerability as function of $S$ and $C$
 This section provides additional results for the model based on Equation 27
-Table A8: Results for fitting Equation A1 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$.
+Table A9: Results for fitting Equation A138 with statsmodels Seabold & Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$.
 We utilize an ordinary least squares.
@@
 […]
-Figure A.3: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation 2 fitted Table A3 (ViT-B).
-The triangles show the median tpr for the train set (ViT-B; Table A3) and circle the test set (R-50; Table A4) over six seeds.
-Note that the triangles and dots for $C=10$ are for EuroSAT.
-B.2 Simpler variant of the prediction model
+[…]
+F.2 Simpler variant of the prediction model
 […]
@@
 Table A10 provides tabular results on the performance of the variant.
-Table A9: Results for fitting Equation A1 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$.
+Table A10: Results for fitting Equation A138 with statsmodels Seabold & Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$.
 We utilize an ordinary least squares.
@@
 Note that the triangles and dots for $C=10$ are for EuroSAT.
-B.3 Additional results for Section 4
-This section provides additional results for Section 4.
-B.4 Rates of In- and Out-distribution Parameters as Function of data set Properties
-LiRA assumes the attacker can build Gaussian distributions of losses for each data point $x$ through the training of shadow models (see Section 2). We denote the distributions as following:
-  • In-distribution: Distribution of the losses of shadow models using $x$ for training ($x\in\mathcal{D}$)
-  • Out-distribution: Distribution of the losses of shadow models not using $x$ for training ($x\notin\mathcal{D}$)
-In Figure A.5 we visualise the changes of the two distributions as functions of $S$ and $C$.
-(a) Distribution means
-(b) Distribution variances
-(c) Distribution standard deviations
-(d) Difference of Means
-(e) Difference of Means / std
-(f) Difference of Means / var
-Figure A.5: Means of observed statistics of the in- and out-distributions for the simplified model with in-class $\sigma=0.1$. The statistics are obtained for $C=2$.
-B.5 Additional results for Section 5
-This section provides additional results for Section 5.
-Figures A.6 and A.7 provide additional data for all considered fpr and for both ViT-B (Head) and R-50 (Head).
-Table A10 the $R^{2}$ scores for ViT-B (Head) and the experiment settings for ViT-B (Head) and R-50 (Head).
-Figure A.8 the relative $R^{2}$ scores for R-50 (Head).
-B.5.1 MIA vulnerability as a function of individual gradient norms
-Figure A.6: ViT-B (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed).
-The results have been obtained over six seeds.
-The number in the bottom right is the $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
-Figure A.7: R-50 (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed).
-The results have been obtained over six seeds.
-The number in the bottom right is the $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
-Table A10: ViT-B (Head): $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
-dataset shots ($S$) classes ($C$) data set size (|$\mathcal{D}$|) $R^{2}$ at fpr=0.1 $R^{2}$ at fpr=0.01 $R^{2}$ at fpr=0.001 Patch Camelyon (Veeling et al., 2018) 256 2 512 0.98 0.97 0.81 CIFAR10 (Krizhevsky, 2009) 64 10 640 0.98 0.98 0.92 EuroSAT (Helber et al., 2019) 64 10 640 0.98 0.97 0.92 Pets (Parkhi et al., 2012) 32 37 1184 0.95 0.93 0.89 Resics45 (Cheng et al., 2017) 32 45 1440 0.95 0.97 0.97 CIFAR100 (Krizhevsky, 2009) 16 100 1600 0.89 0.89 0.86
-Figure A.8: R-50 (Head): Ratio between $R^{2}$ per iteration decile and $R^{2}$ over the whole training.
-A ratio $>1$ means that the decile is as more predictive than whole trace and a ratio $<1$ means that it is less predictive.
-B.5.2 MIA vulnerability as a function of feature space distance
-This section provides additional data regarding the individual MIA prediction using feature space distances.
-Figure A.9: ViT-B (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
-Figure A.10: R-50 (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
-Appendix C Missing proofs of Section 4.2
-See 1
-Proof.
-Recall
-fpr $=\Pr\left(\underbrace{-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]}_{\lambda_{\textsc{fpr}}}>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right)$ (A2)
-tpr $=\Pr\left(\underbrace{-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]}_{\lambda_{\textsc{tpr}}}>\tau\mid x\in\mathcal{D}_{\textrm{target}}\right).$ (A3)
-We denote the class means for the $m$th in/out shadow data sets with $r_{y,m}^{(\textrm{in})}$ and $r_{y,m}^{(\textrm{out})}$
-$r_{y,m}^{(\textrm{in})}$ $=\frac{1}{S}\left(\sum_{i=1}^{S-1}x_{i,m,y}+x\right),\;r_{y,m}^{(\textrm{out})}=\frac{1}{S}\left(\sum_{i=1}^{S}x_{i,m,y}\right),$ (A4)
-where $x_{i,m,y}$ denotes the $i$th element of the $m$th shadow data set with label $y$.
-The corresponding scores $s_{y,m}^{(\textrm{in})}=\langle x,r_{y,m}^{(\textrm{in})}\rangle$ and $s_{y,m}^{(\textrm{out})}=\langle x,r_{y,m}^{(\textrm{out})}\rangle$ which can be written as
-$s_{y,m}^{(\textrm{in})}$ $=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S-1}x_{i,m,y}\rangle+\langle x,x\rangle\right)=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S}x_{i,m,y}\rangle+\langle x,x-x_{S,m,y}\rangle\right)$ (A5)
-$s_{y,m}^{(\textrm{out})}$ $=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S}x_{i,m,y}\rangle\right)$ (A6)
-Now, assuming we have arbitrarily many shadow models, we can write the LiRA statistics as
-$\mu_{\textrm{in}}=\operatorname{\mathbb{E}}_{D_{m}}[s_{y,m}^{(\textrm{in})}],\;\mu_{\textrm{out}}=\operatorname{\mathbb{E}}_{D_{m}}[s_{y,m}^{(\textrm{out})}]=\langle x,m_{y}\rangle$ (A7)
-$\Rightarrow$ $\mu_{\textrm{in}}-\mu_{\textrm{out}}=\frac{1}{S}\langle x,x-\mathbb{E}_{D_{m}}[x_{S,m}]\rangle=\frac{1}{S}\langle x,x-m_{y}\rangle,$ (A8)
-where $m_{y}$ is the true center of the $y$th cluster. For $\sigma$ we have
-$\sigma^{2}$ $=\operatorname{Var}(s_{y,m}^{(\textrm{out})})$ (A9)
-$=\frac{1}{S^{2}}\sum_{i=1}^{S}\operatorname{Var}(\langle x,x_{i,m}\rangle)$ (A10)
-$=\frac{1}{S}\operatorname{Var}(\langle x,x_{m}^{(j)}\rangle)$ (A11)
-$=\frac{1}{S}\operatorname{Var}(x_{m}^{(j)})||x||^{2}=\frac{1}{S}s^{2}||x||^{2},$ (A12)
-where $x_{m}^{(j)}$ denotes the $j$th dimension of the vector $x_{m}$ and $s$ is the std of each dimension of $x_{m}$. Now we can write the $\lambda_{\textsc{fpr}}$ as
-$\lambda_{\textsc{fpr}}=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(-2t_{x}+2\langle x,m_{y}\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A13)
-In the fpr case, we have that $x\notin\mathcal{D}_{\textrm{target}}$, therefore for the $t_{x}$ we have
-$t_{x}=\langle x,r_{y}^{\textrm{(target)}}\rangle\sim\langle x,m_{y}\rangle+\frac{s}{\sqrt{S}}\langle x,z\rangle,$ (A14)
-where $z\sim N(0,I_{d})$. Substituting this into $\lambda_{\textsc{fpr}}$ we have
-$\lambda_{\textsc{fpr}}=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A15)
-We see that $\lambda_{\textsc{fpr}}$ evolves as $O(1/\sqrt{S})$.
-Next, we will focus on the $\lambda_{\textsc{tpr}}$. The expression remains otherwise the same as for $\lambda_{\textsc{fpr}}$, but since $x\in\mathcal{D}_{\textrm{target}}$ the $t_{x}$ changes to
-$t_{x}\sim\langle x,m_{y}\rangle+\frac{1}{S}\langle x,x\rangle+s\frac{1}{\sqrt{S}}\sqrt{1-\frac{1}{S}}\langle x,z\rangle.$ (A16)
-Hence, for large enough $S$ we have
-$\lambda_{\textsc{tpr}}\approx-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle-\frac{2}{S}||x||^{2}+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A17)
+F.3 Empirical results for RMIA
+Figures A.6, A.7 and A.8 report additional results for RMIA Zarifzadeh et al. (2024).
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]

### v4

@@
 Abstract
-We analyse the relationship between privacy vulnerability and dataset properties, such as examples per class and number of classes, when applying two state-of-the-art membership inference attacks (MIAs) to fine-tuned neural networks.
-We derive per-example MIA vulnerability in terms of score distributions and statistics computed from shadow models.
-We introduce a simplified model of membership inference and prove that in this model, the logarithm of the difference of true and false positive rates depends linearly on the logarithm of the number of examples per class.
-We complement the theoretical analysis with empirical analysis by systematically testing the practical privacy vulnerability of fine-tuning large image classification models and obtain the previously derived power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate.
-Finally, we fit a parametric model of the previously derived form to predict true positive rate based on dataset properties and observe good fit for MIA vulnerability on unseen fine-tuning scenarios.
+Membership inference attacks (MIAs) are used to test practical privacy of machine learning models.
+MIAs complement formal guarantees from differential privacy (DP) under a more realistic adversary model.
+We analyse MIA vulnerability of fine-tuned neural networks both empirically and theoretically, the latter using a simplified model of fine-tuning.
+We show that the vulnerability of non-DP models when measured as the attacker advantage at fixed false positive rate reduces according to a simple power law as the number of examples per class increases, even for the most vulnerable points, but the dataset size needed for adequate protection of the most vulnerable points is very large.
 1 Introduction
-Machine learning models are prone to memorising their training data, which makes them vulnerable to privacy attacks such as membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and reconstruction attacks (e.g.
-Balle et al., 2022; Nasr et al., 2023).
-Differential privacy (DP; Dwork et al., 2006) provides protection against these attacks, but strong formal protection often comes at the cost of significant loss of model utility.
-Finding the correct balance between making models resistant to attacks while maintaining a high utility is important for many applications.
-In health, for example, many European countries and soon also the EU within the European Health Data Space have requirements that models trained on health data that are made publicly available must be anonymous, i.e.
-they must not contain information that can be linked to an identifiable individual.
-On the other hand, loss of utility of the model due to privacy constraints may compromise the health benefits that might be gained from it.
-In this paper, our aim is to theoretically understand and apply two state-of-the-art MIAs systematically, to help understand practical privacy risks when fine-tuning deep-learning-based classifiers without DP protections.
-We focus on transfer learning using fine-tuning because this is increasingly used for all practical applications of deep learning.
+Membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and differential privacy (DP; Dwork et al., 2006) provide complementary means of deriving lower and upper bounds for the privacy loss of a machine learning algorithm.
+Yet, the two operate under slightly different threat models.
+DP implicitly assumes a very powerful adversary with access to all training data except the target point and provides guarantees against every target point.
+MIAs assume an often more realistic adversary model with access to just the data distribution, and the unknown training data becoming latent variables introducing stochasticity into the attack, but the practical evaluation is statistical and cannot provide universal guarantees.
+In this paper, we seek to explore the area between DP and typical MIA evaluations.
+Inspired by an empirical finding that average MIA vulnerability of neural network fine-tuning strongly reduces as the number of samples in the target class increases (see Figure 1), we develop theory of optimal MIA against a simple model of neural network fine-tuning and reproduce the decrease in vulnerability.
+Furthermore, the theoretical model predicts that the vulnerability of all individual samples should reduce as the number of samples increases, which we are able to verify empirically.
+Figure 1: We observe a power-law relation between MIA vulnerability and $S$ (shots, i.e., examples per class) when attacking a ViT-B Head fine-tuned using LiRA.
+The solid line is median and the error bars the min/max bounds for the Clopper-Pearson CIs over six seeds.
+To achieve our goal, we theoretically analyze and systematically apply two state-of-the-art black-box MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), to help understand practical privacy risks when fine-tuning deep-learning-based classifiers without DP protections.
+For the theoretical model that we analyse, LiRA is the optimal attack by the Neyman–Pearson lemma (Neyman and Pearson, 1933).
+Under the black-box threat model, in which the adversary does not have access to the model parameters, LiRA and RMIA have been shown to empirically outperform other attacks, especially when the number of shadow models is sufficiently large.
+We focus on transfer learning using fine-tuning because this is increasingly used for all practical applications of deep learning and especially important when labeled examples are limited, which is often the case in privacy-sensitive applications.
 Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep transfer learning models to MIA.
-Since the earliest MIA methods, there has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017).
-Similarly, there is evidence that models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023).
-Feldman & Zhang (2020) showed that neural networks trained from scratch can memorise a large fraction of their training data, while the memorisation is greatly reduced for fine-tuning.
-In this paper, we theoretically study the relationship between examples per class, which we denote as shots ($S$), and MIA vulnerability (true positive rate tpr at fixed false positive rate fpr) for a simplified model of fine-tuning and derive a power-law relationship in the form
+In particular, we theoretically study the relationship between the number of examples per class, which we denote as shots ($S$), and MIA vulnerability (true positive rate tpr at fixed false positive rate fpr) for a simplified model of fine-tuning, and derive a power-law relationship (Figure 1) in the form
 $\log(\textsc{tpr}-\textsc{fpr})=-\beta_{S}\log(S)-\beta_{0}.$ (1)
-We complement the theoretical analysis with extensive experiments over many datasets with varying sizes in the transfer learning setting for image classification tasks and observe the same power law.
-Previously, Tobaben et al.
-(2023) and Yu et al.
-(2023) have presented limited studies of how dataset properties affect the MIA vulnerability.
-Tobaben et al.
-(2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots (i.e.
-the number of examples available per class).
+We complement the theoretical analysis with extensive experiments over many datasets with varying sizes, in the transfer learning setting for image classification tasks, and observe the same power-law.
+Based on extrapolation from our results, the number of examples per each class that are needed for adequate protection of the most vulnerable samples appears very high.
+Related work There has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017), models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023), and classes with less examples tend to be more vulnerable (Chang and Shokri, 2021; Kulynych et al., 2022; Tonni et al., 2020).
+A larger generalisation error, which is related to dataset size, has also been shown to be sufficient for MIA success (Song and Mittal, 2021), though not necessary (Yeom et al., 2018).
+Similarly, minority subgroups tend to be more affected by DP (Suriyakumar et al., ; Bagdasaryan et al., 2019).
+Feldman and Zhang (2020) showed that neural networks trained from scratch can memorise a large fraction of their training data, while the memorisation is greatly reduced for fine-tuning.
+Additionally, Tobaben et al.
+(2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots.
 Yu et al.
-(2023) studied how the MIA vulnerability of a class in an image classifier trained from scratch depends on the average individual privacy parameters of that class.
-Our work significantly expands on both of these by a more systematic study and modelling of dataset properties.
-List of contributions We analyze the MIA vulnerability of deep transfer learning using two state-of-the-art MIAs LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024).
-Both MIAs are based on the assumption that the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack.
-We first analyze both MIAs theoretically by introducing a simplified model of membership inference (Section 3).
-We support our theoretical findings by an extensive empirical study on the MIA vulnerability of deep learning models.
-We focus on a transfer learning setting for image classification tasks, where a large pre-trained neural network is fine-tuned on a sensitive dataset, and the MIA is performed on the final classifier.
-We find that the MIA vulnerability has a strong correlation with certain dataset properties (Section 4).
+(2023) studied the relationship between the MIA vulnerability and individual privacy parameters for different classes.
+Recently, worst-case MIA vulnerability has gained more attention (Guépin et al., 2024; Meeus et al., 2024; Azize and Basu, 2024).
+Nonetheless, the prior works do not consider the rate of change in the vulnerability evaluated at a low fpr, as dataset properties change.
+Our work significantly expands on these works by explicitly identifying a quantitative relationship between dataset properties and MIA vulnerability, i.e., the power-law in Equation 1.
+List of contributions We analyze the MIA vulnerability of deep transfer learning using two state-of-the-art score-based MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), which are a strong realistic threat model.
+We first analytically derive the power-law relationship in Equation 1 for both MIAs by introducing a simplified model of the optimal membership inference (Section 3).
+We support our theoretical findings by an extensive empirical study on the MIA vulnerability of deep learning models by focusing on a transfer-learning setting for image classification task, where a large pre-trained neural network is fine-tuned on a sensitive dataset.
   1. 1.
-    Per-example vulnerability: We derive per-example LiRA and RMIA vulnerability (tpr at fixed fpr) in terms of MIA attack score distributions and statistics computed from shadow models (see Section 3.3).
+    Power-law in a simplified model of the optimal MIA: We formulate a simplified model of MIA to quantitatively relate dataset properties and MIA vulnerability.
+In this model LiRA is the optimal attack.
+For this model, we prove a power-law relationship for both average and worst-case between the LiRA as well as RMIA vulnerability and the number of examples per class (See Section 3.4).
   2. 2.
-    Power-law in simplified model of MI: We formulate a simplified model of membership inference to quantitatively relate dataset properties and MIA vulnerability.
-For this model, we prove a power-law relationship between the per-example LiRA and RMIA vulnerability and the number of examples per class.
-We then extend the the power-law relationship to the average-case LiRA and RMIA vulnerability (See Section 3.4).
+    MIA experiments on the average case vulnerability: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power-law between the number of examples per class and the vulnerability to MIA (see Figures 1 and 4.2).
+We fit a regression model which follows the functional form of the theoretically derived power-law.
+We show both a very good fit on the training data as well as a good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Section 4.3).
   3. 3.
-    Few-shot MIA: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power law between the number of examples per class and the vulnerability to MIA (see Figure 1).
-  4. 4.
-    Regression model: We utilise our empirical observations to fit a regression model to predict MIA vulnerability ($\log(\textsc{tpr}-\textsc{fpr})$ at fixed low fpr) based on examples per class ($\log S$) and number of classes ($\log C$), which follows the functional form of the theoretically derived power-law.
-We show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Figure 4).
+    MIA experiments on the worst-case vulnerability: We extend the experiments to worst-case individual sample vulnerabilities and observe a similar decrease in vulnerability for quantiles of vulnerable data points and a slower decrease for the maximum individual vulnerability (Section 4.4).
+By extrapolation we find that an adequate protection of the most vulnerable samples would require an extremely large dataset (Section 4.5).
 2 Background
@@
 Membership inference attacks (MIAs) aim to infer whether a particular sample was part of the training set of the targeted model (Shokri et al., 2017).
-Thus, they can be used to determine lower bounds on the privacy leakage of models to complement the theoretical upper bounds obtained through differential privacy.
+Thus, they can be used to determine lower bounds on the privacy leakage of models to complement the theoretical upper bounds obtained through DP.
 Likelihood Ratio attack (LiRA; Carlini et al., 2022) While many different MIAs have been proposed (Hu et al., 2022), in this work we consider the Likelihood Ratio Attack (LiRA).
@@
 Based on this information, the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack but for which the attacker knows the training dataset.
-LiRA exploits the observation that the loss function value used to train a model is often lower for examples that were part of the training set compared to those that were not.
-For a target sample $(x,y_{x})$, LiRA trains the shadow models: 1 with $(x,y_{x})$ as a part of the training set ($(x,y_{x})\in\mathcal{D}$) and 2 without $x$ in the training set ($(x,y_{x})\notin\mathcal{D}$) .
-After training the shadow models, $(x,y_{x})$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $(x,y_{x})\in\mathcal{D}$ shadow models, and one for the $(x,y_{x})\notin\mathcal{D}$.
-Finally, the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $(x,y_{x})\in\mathcal{D}$ or $(x,y_{x})\notin\mathcal{D}$.
+LiRA exploits the observation that the loss function value used to train a model is often lower for the examples that were part of the training set compared to those that were not.
+For a target tuple $(\bm{x},y)$, where $y$ is a label, LiRA trains the shadow models: 1 with $(\bm{x},y)$ as a part of the training set ($(\bm{x},y)\in\mathcal{D}$) and 2 without $\bm{x}$ in the training set ($(\bm{x},y)\notin\mathcal{D}$) .
+After training the shadow models, $(\bm{x},y)$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $(\bm{x},y)\in\mathcal{D}$ shadow models, and one for the $(\bm{x},y)\notin\mathcal{D}$.
+Finally, the attacker computes the loss for the point $\bm{x}$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $(\bm{x},y)\in\mathcal{D}$ or $(\bm{x},y)\notin\mathcal{D}$.
+When the true distributions of the shadow models are Gaussians, LiRA is the optimal attack provided by the Neyman–Pearson lemma (Neyman and Pearson, 1933).
 We use an optimization by Carlini et al.
@@
 At attack time each model will be attacked once using all other models as shadow models.
-Robust Membership Inference Attack (RMIA; Zarifzadeh et al., 2024) Recently Zarifzadeh et al.
-(2024) proposed a new MIA algorithm called RMIA, which aims to improve performance when the number of shadow models is limited.
-Similar to LiRA, RMIA is based on shadow model training and computing the attack statistics based on a likelihood ratio.
-The main difference to LiRA is that RMIA does not compute the likelihood ratio based on aggregated IN/OUT statistics, but instead compares the target data point against random samples $(z,y_{z})$ from the target data distribution.
-After computing the likelihood ratios over multiple $(z,y_{z})$ values, the MIA score is estimated as a proportion of the ratios exceeding a preset bound.
-This approach makes RMIA a more effective attack when the number of shadow models is low.
+Robust Membership Inference Attack (RMIA; Zarifzadeh et al., 2024) RMIA is a new MIA algorithm, which aims to improve performance when the number of shadow models is limited.
+We show both theoretically and empirically that the power-law also holds for RMIA.
 Measuring MIA vulnerability Using the chosen MIA score of our attack, we can build a binary classifier to predict whether a sample belongs to the training data or not.
@@
 Identifying even a small number of examples with high confidence is considered harmful (Carlini et al., 2022) and thus we focus on the regions of small fpr.
+Measuring the uncertainty for tpr The tpr values from the LiRA-based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples.
+Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimators.
+Therefore, when we report the tpr values for a single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals (Clopper and Pearson, 1934).
+Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
+$\begin{split}&B(\alpha/2;\textsc{tp},\textsc{p}-\textsc{tp}+1)<\textsc{tpr}<B(1-\alpha/2;\textsc{tp}+1,\textsc{p}-\textsc{tp}),\end{split}$ (2)
+where $B(q;a,b)$ is the $q$th-quantile of $\text{Beta}(a,b)$ distribution.
 3 Theoretical analysis
@@
 Aerni et al., 2024; Leemann et al., 2024).
-Therefore, we start with analysing per-example vulnerabilities for LiRA and RMIA.
+Therefore, we start with analysing per-example MIA vulnerabilities.
 In order to quantitatively relate dataset properties to these vulnerabilities, a simplified model is formulated.
@@
 Finally, the per-example power-law is analytically extended to average-case MIA vulnerability, for which we provide empirical evidence in Section 4.
+We primarily focus on the analysis of LiRA, since it is the optimal attack in our simplified model.
+In Appendix B we show that similar theoretical results also hold for RMIA.
 3.1 Preliminaries
@@
 (2022).
-Denoting the logit of a target model $\mathcal{M}$ applied on a target data point $(x,y_{x})$ as $\ell(\mathcal{M}(x),y_{x})$, the LiRA computes the MIA score as the likelihood ratio
-$\mathrm{LR}(x)=\frac{p(\ell(\mathcal{M}(x),y_{x})\mid\mathbb{Q}_{\text{in}}(x,y_{x}))}{p(\ell(\mathcal{M}(x),y_{x})\mid\mathbb{Q}_{\text{out}}(x,y_{x}))},$ (2)
-where the $\mathbb{Q}_{\text{in}/\text{out}}$ denote the hypotheses that $(x,y_{x})$ was or was not in the training set of $\mathcal{M}$.
+Denoting the logit of a target model $\mathcal{M}$ applied on a target data point $(\bm{x},y)$ as $\ell(\mathcal{M}(\bm{x}),y)$, LiRA computes the MIA score as the likelihood ratio
+$\mathrm{LR}(\bm{x})=\frac{p(\ell(\mathcal{M}(\bm{x}),y)\mid\mathbb{Q}_{\text{in}}(\bm{x},y))}{p(\ell(\mathcal{M}(\bm{x}),y)\mid\mathbb{Q}_{\text{out}}(\bm{x},y))},$ (3)
+where the $\mathbb{Q}_{\text{in}/\text{out}}$ denote the hypotheses that $(\bm{x},y)$ was or was not in the training set of $\mathcal{M}$.
 Carlini et al.
 (2022) approximate the IN/OUT hypotheses as normal distributions.
-Denoting $t_{x}=\ell(\mathcal{M}(x),y_{x})$, the score becomes
-$\mathrm{LR}(x)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}^{2}(x))}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}^{2}(x))},$ (3)
-where the $\hat{\mu}_{\text{in/out}}(x)$ and $\hat{\sigma}_{\text{in/out}}(x)$ are the means and standard deviations for the IN/OUT shadow model losses for $(x,y_{x})$.
-Larger values of $\mathrm{LR}(x)$ suggest that $(x,y_{x})$ is more likely in the training set and vice versa.
-Now, to build a classifier from this score, the LiRA tests if $\mathrm{LR}(x)>\beta$ for some threshold $\beta$.
-Next, let us restate how RMIA (Zarifzadeh et al., 2024) builds the MIA score.
-RMIA augments the likelihood-ratio with a sample from the target data distribution to calibrate how likely you would obtain the target model if $(x,y_{x})$ is replaced with another sample $(z,y_{z})$.
-Denoting the target model parameters with $\theta$, RMIA computes
-$\mathrm{LR}(x,z)=\frac{p(\theta\mid x,y_{x})}{p(\theta\mid z,y_{z})},$ (4)
-and the corresponding MIA score is given as
-$\text{Score}_{\text{RMIA}}(x)=\Pr_{(z,y_{z})\sim\mathbb{D}}(\mathrm{LR}(x,z)>\gamma),$ (5)
-where $\mathbb{D}$ denotes the training data distribution.
-Similar to LiRA, the classifier is built by checking if $\text{Score}_{\text{RMIA}}(x)>\beta$.
-In the following, we will use the direct computation of likelihood-ratio as described in Equation 11 of Zarifzadeh et al.
-(2024) which approximates $\mathrm{LR}(x,z)$ using normal distributions.
-3.2 Computing the tpr for LiRA and RMIA
-Using the LiRA formulation of Equation 3, the tpr for the target point $(x,y_{x})$ for LiRA is defined as
-$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}(x)^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})}\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\right),$ (6)
-where $\beta$ is a threshold that defines a rejection region of the likelihood ratio test, $\hat{\mu}_{\mathrm{in}}(x),\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{in}}(x)$ and $\hat{\sigma}_{\mathrm{out}}(x)$ are LiRA statistics computed from shadow models, and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Appendix A for derivation).
-For theoretical analysis of RMIA, we focus on the direct approach that is an approximation of the efficient Bayesian approach, as Zarifzadeh et al.
-(2024) empirically demonstrates that these approaches exhibit similar performances.
-Let $\hat{\mu}_{a,b}$ and $\hat{\sigma}_{a,b}$ denote, respectively, the mean and standard deviation of $t_{b}$ estimated from shadow models, where $a$ denotes which of $(x,y_{x})$ and $(z,y_{z})$ is in the training set.
-By Equation 11 in (Zarifzadeh et al., 2024), the per-example performance for RMIA is given as
-$\textsc{tpr}_{\mathrm{RMIA}}(x)=\\ \Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\Pr_{(z,y_{z})\sim{\mathbb{D}}}(\mathrm{LR}(x,z)\geq\gamma)\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\notin\mathcal{D}_{\mathrm{target}}\right)$ (7)
-$\mathrm{LR}(x,z)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{x,x},\hat{\sigma}_{x,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{x,z},\hat{\sigma}_{x,z}^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{z,x},\hat{\sigma}_{z,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{z,z},\hat{\sigma}_{z,z}^{2})},$ (8)
-where $t_{z}$ is the score on $z$ similar to $t_{x}$ and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Appendix A for derivation).
-We define the average-case tprs for LiRA and RMIA by taking the expectation over the data distribution:
-$\overline{\textsc{tpr}}_{\mathrm{LiRA}}=\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}[\textsc{tpr}_{\mathrm{LiRA}}(x)]$ (9)
-$\overline{\textsc{tpr}}_{\mathrm{RMIA}}=\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}[\textsc{tpr}_{\mathrm{LiRA}}(x)]$ (10)
+Denoting $t_{\bm{x}}=\ell(\mathcal{M}(\bm{x}),y)$, the score becomes
+$\mathrm{LR}(\bm{x})=\frac{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\mathrm{in}}(\bm{x}),\hat{\sigma}_{\mathrm{in}}(\bm{x})^{2})}{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\mathrm{out}}(\bm{x}),\hat{\sigma}_{\mathrm{out}}(\bm{x})^{2})},$ (4)
+where the $\hat{\mu}_{\text{in/out}}(\bm{x})$ and $\hat{\sigma}_{\text{in/out}}(\bm{x})$ are the means and standard deviations for the IN/OUT shadow model losses for $(\bm{x},y)$.
+Larger values of $\mathrm{LR}(\bm{x})$ suggest that $(\bm{x},y)$ is more likely in the training set and vice versa.
+Now, to build a classifier from this score, the LiRA tests if $\mathrm{LR}(\bm{x})>\tau$ for some threshold $\tau$.
+Note that the attacker only has a finite set of shadow models to estimate the IN/OUT parameters.
+Therefore, the MIA scores become random variables over the true population level IN/OUT distributions.
+3.2 Computing the tpr for LiRA
+Using the LiRA formulation of Equation 4, the tpr for the target point $(\bm{x},y)$ for LiRA is defined as
+$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\mathrm{LR}(\bm{x})\geq\tau\mid(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}\right),$ (5)
+where $\tau$ is a threshold that defines a rejection region of the likelihood ratio test, and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Section A.1 for derivation).
+We define the average-case tpr for LiRA by taking the expectation over the data distribution:
+$\overline{\textsc{tpr}}_{\mathrm{LiRA}}=\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}[\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})]$ (6)
 3.3 Per-example MIA vulnerability
-Although LiRA models $t_{x}$ by a normal distribution, we consider a more general case where the true distribution of $t_{x}$ is of the location-scale family. That is,
-$t_{x}=\begin{cases}\mu_{\mathrm{in}}(x)+\sigma_{\mathrm{in}}(x)t&\text{ if }(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\\ \mu_{\mathrm{out}}(x)+\sigma_{\mathrm{out}}(x)t&\text{ if }(x,y_{x})\notin\mathcal{D}_{\mathrm{target}},\end{cases}$ (11)
-where $t$ has the standard location and unit scale, and $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and $\sigma_{\mathrm{in}}(x),\sigma_{\mathrm{out}}(x)$ are the locations and scales of IN/OUT distributions of $t_{x}$.
+Although LiRA models $t_{\bm{x}}$ by a normal distribution, we consider a more general case where the true distribution of $t_{\bm{x}}$ is of the location-scale family. That is,
+$t_{\bm{x}}=\begin{cases}\mu_{\mathrm{in}}(\bm{x})+\sigma_{\mathrm{in}}(\bm{x})Z&\text{ if }(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}\\ \mu_{\mathrm{out}}(\bm{x})+\sigma_{\mathrm{out}}(\bm{x})Z&\text{ if }(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}},\end{cases}$ (7)
+where $Z$ has the standard location and unit scale, and $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and $\sigma_{\mathrm{in}}(\bm{x}),\sigma_{\mathrm{out}}(\bm{x})$ are the locations and scales of IN/OUT distributions of $t_{\bm{x}}$.
 We assume that the target and shadow datasets have a sufficient number of examples.
-This allows us to also assume that $\hat{\sigma}(x)=\hat{\sigma}_{\mathrm{in}}(x)=\hat{\sigma}_{\mathrm{out}}(x)$ and $\sigma(x)=\sigma_{\mathrm{in}}(x)=\sigma_{\mathrm{out}}(x)$, where $\hat{\sigma}(x)$ is the standard deviation of $t_{x}$ estimated from shadow models and $\sigma(x)$ is the true scale parameter of $t_{x}$.
-(See Appendix B for the validity of these assumptions).
-The following result reduces the LiRA vulnerability to the location and scale parameters of $t_{x}$.
+This allows us to also assume that $\hat{\sigma}(\bm{x})=\hat{\sigma}_{\mathrm{in}}(\bm{x})=\hat{\sigma}_{\mathrm{out}}(\bm{x})$ and $\sigma(\bm{x})=\sigma_{\mathrm{in}}(\bm{x})=\sigma_{\mathrm{out}}(\bm{x})$, where $\hat{\sigma}(\bm{x})$ is the standard deviation of $t_{\bm{x}}$ estimated from shadow models and $\sigma(\bm{x})$ is the true scale parameter of $t_{\bm{x}}$.
+(See Section A.2 for the validity of these assumptions).
+The following result reduces the LiRA vulnerability to the location and scale parameters of $t_{\bm{x}}$.
 Lemma 1 (Per-example LiRA vulnerability).
-Suppose that the true distribution of $t_{x}$ is of location-scale family with locations $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and scale $\sigma(x)$, and that LiRA models $t_{x}$ by $\mathcal{N}(\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}(x))$ and $\mathcal{N}(\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}(x))$.
+Suppose that the true distribution of $t_{\bm{x}}$ is of location-scale family with locations $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and scale $\sigma(\bm{x})$, and that LiRA models $t_{\bm{x}}$ by $\mathcal{N}(\hat{\mu}_{\mathrm{in}}(\bm{x}),\hat{\sigma}(\bm{x})^{2})$ and $\mathcal{N}(\hat{\mu}_{\mathrm{out}}(\bm{x}),\hat{\sigma}(\bm{x})^{2})$.
 Assume that an attacker has access to the underlying distribution ${\mathbb{D}}$.
 Then for a large enough number of examples per class and infinitely many shadow models, the LiRA vulnerability of a fixed target example is
-$\textsc{tpr}_{\mathrm{LiRA}}(x)=\begin{cases}1-F_{t}\left(F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)>\hat{\mu}_{\mathrm{out}}(x)\\ F_{t}\left(F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)<\hat{\mu}_{\mathrm{out}}(x),\end{cases}$ (12)
-where $F_{t}$ is the cdf of $t$ with the standard location and unit scale.
+$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})=\begin{cases}1-F_{Z}\left(F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(\bm{x})>\hat{\mu}_{\mathrm{out}}(\bm{x})\\ F_{Z}\left(F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(\bm{x})<\hat{\mu}_{\mathrm{out}}(\bm{x}),\end{cases}$ (8)
+where $F_{Z}$ is the cdf of $Z$ with the standard location and unit scale, assuming that the inverse of $F_{Z}$ exists.
 Proof.
-See Section C.1. ∎
+See Section A.3. ∎
 Here we assume that an attacker trains shadow models with the true underlying distribution.
 However, in real-world settings the precise underlying distribution may not be available for an attacker.
-We relax this assumption in Appendix B so that the attacker only needs an approximated underlying distribution for the optimal LiRA as in Lemma 1.
-Next we focus on the per-example RMIA performance. As in the case of LiRA, we assume that $t_{x}$ and $t_{z}$ follow distributions of the location-scale family. We have
-$t_{x}$ =$=\begin{cases}\mu_{x,x}+\sigma_{x,x}t&\text{ if }(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{z,x}+\sigma_{z,x}t&\text{ if }(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\in\mathcal{D}_{\mathrm{target}}\end{cases}$ (13)
-$t_{z}$ =$=\begin{cases}\mu_{x,z}+\sigma_{x,z}t&\text{ if }(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{z,z}+\sigma_{z,z}t&\text{ if }(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\wedge(z,y_{z})\in\mathcal{D}_{\mathrm{target}}.\end{cases}$ (14)
-It is important to note that $\mu_{a,b}$ and $\sigma_{a,b}$ denote, respectively, a location and a scale, while previously defined $\hat{\mu}_{a,b}$ and $\hat{\sigma}_{a,b}$ are, respectively, a mean and a standard deviation.
-As for the analysis of LiRA, we assume that the target and shadow sets have a sufficient number of examples per class, and that $\sigma_{x}=\sigma_{x,x}=\sigma_{z,x}$, $\sigma_{z}=\sigma_{x,z}=\sigma_{z,z}$, $\hat{\sigma}_{x}=\hat{\sigma}_{x,x}=\hat{\sigma}_{z,x}$ and $\hat{\sigma}_{z}=\hat{\sigma}_{x,z}=\hat{\sigma}_{z,z}$, where $\sigma_{x}$ and $\sigma_{z}$ are, respectively, the true scales of $t_{x}$ and $t_{z}$, and $\hat{\sigma}_{x}$ and $\hat{\sigma}_{z}$ are, respectively, standard deviations of $t_{x}$ and $t_{z}$ estimated from shadow models (see Appendix B for the validity of these assumptions).
-Lemma 2 (Per-example RMIA vulnerability).
-Suppose that the true distributions of $t_{x}$ and $t_{z}$ are of location-scale family with locations $\mu_{x,x},\mu_{z,x},\mu_{x,z},\mu_{z,z}$ and scales $\sigma_{x},\sigma_{z}$, and that RMIA models $t_{x}$ and $t_{z}$ by normal distributions with parameters computed from shadow models.
-For a large enough number of examples per class and infinitely many shadow models, the RMIA vulnerability of a fixed target example is bounded by
-$\textsc{tpr}_{\mathrm{RMIA}}(x)\leq\begin{cases}1-F_{t}\left(F_{t}^{-1}(1-\alpha)-\frac{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[q]}{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]}\right)&\text{ if }\ \mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]>0\\ F_{t}\left(F_{t}^{-1}(\alpha)-\frac{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[q]}{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]}\right)&\text{ if }\ \mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}[A]<0,\end{cases}$ (15)
-for some constant $\alpha\geq\textsc{fpr}_{\mathrm{RMIA}}(x)$, where
-$q$ $=\frac{(\mu_{x,x}-\mu_{z,x})(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})}{\hat{\sigma}_{x}^{2}}-\frac{(\mu_{x,z}-\mu_{z,z})(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}{\hat{\sigma}_{z}^{2}}$ (16)
-$A$ $=\frac{\sigma_{x}}{\hat{\sigma}_{x}^{2}}(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\sigma_{z}}{\hat{\sigma}_{z}^{2}}(\hat{\mu}_{x,z}-\hat{\mu}_{z,z}).$ (17)
-Proof.
-See Section C.2. ∎
-Note that here we must assume that the attacker has access to the underlying distribution for the optimal RMIA as the Equations 16 and 17 depend on the parameters computed from shadow models.
-3.4 A simplified model of membership inference
+We relax this assumption in Section A.4 so that the attacker only needs an approximated underlying distribution for the optimal LiRA as in Lemma 1.
+3.4 A simplified model of the optimal membership inference
 Now we construct a simplified model of membership inference that streamlines the data generation and shadow model training.
-We sample vectors on a high dimensional unit sphere and classify them based on inner product with estimated class mean.
+We sample vectors on a high-dimensional unit sphere and classify them based on inner product with estimated class mean.
 This model is easier to analyse theoretically than real-world deep learning examples.
@@
   1. 1.
-    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in\{1,\dots,C\}:m_{i}\perp m_{j}\lor i=j$).
+    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in\{1,\dots,C\},i\neq j:m_{i}\perp m_{j}$).
   2. 2.
-    We sample $2S$ vectors $x_{c}$ for each class.
-We assume that they are Gaussian distributed around the the true class mean $x_{c}\sim\mathcal{N}(m_{c},s^{2}I)$ where the $s^{2}$ is the in-class variance.
+    We sample $2S$ vectors $\bm{x}_{c}$ for each class.
+We assume that they are Gaussian distributed around the the true class mean $\bm{x}_{c}\sim\mathcal{N}(m_{c},\Sigma)$ where the $\Sigma$ is the in-class covariance.
   3. 3.
-    For each ”target model” we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $r_{c}$.
+    For each ”target model” we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $\hat{m}_{c}$.
   4. 4.
-    The computed mean is used to classify sample $x$ by computing the inner product $\langle x,r_{c}\rangle$ as a metric of similarity.
+    The computed mean is used to classify sample $\bm{x}$ by computing the inner product $\langle\bm{x},\hat{m}_{c}\rangle$ as a metric of similarity.
 The attacker has to infer which vectors have been used for training the classifier.
 Instead of utilising the logits (like in many image classification tasks), the attacker can use the inner products of a point with the cluster means.
+Since the inner product score follows a normal distribution, LiRA with infinitely many shadow models is the optimal attack by the Neyman–Pearson lemma (Neyman and Pearson, 1933), which states that the likelihood ratio test is the most powerful test for a given FPR.
 This simplified model resembles a linear (Head) classifier often used in transfer learning when adapting to a new dataset.
@@
 Theorem 2 (Per-example LiRA power-law).
-Fix a target example $(x,y_{x})$. For the simplified model with arbitrary $C$ and infinitely many shadow models, the per-example LiRA vulnerability is given as
-$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))+\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}s||x||}\right),$ (18)
-where $m_{x}$ is the true mean of class $y_{x}$. In addition, for large $S$ we have
-$\log(\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}+\log\frac{|\langle x,x-m_{x}\rangle|}{||x||s\sqrt{2\pi}}.$ (19)
+Fix a target example $(\bm{x},y)$. For the simplified model with arbitrary $C$ and infinitely many shadow models, the per-example LiRA vulnerability is given as
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S))$ (9)
+where $m_{\bm{x}}$ is the true mean of class $y$ and $\xi(S)=O(1/\sqrt{S})$. For large $S$ we have
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (10)
 Proof.
-See Section C.3. ∎
+See Section A.5. ∎
 An immediate upper bound is obtained from Theorem 2 by the Cauchy-Schwarz inequality:
-$\log(\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x))\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}+\log\frac{||x-m_{x}||}{s\sqrt{2\pi}}.$ (20)
-This implies that if $||x-m_{x}||$ is bounded, then the worst-case vulnerability is also bounded.
-Hence we can have the LiRA vulnerability arbitrarily close to zero in this non-DP setting by simply increasing $S$.
-Similarly, employing Lemma 2 and the simplified model, we obtain the following upper bound for RMIA performance.
-Theorem 4 (Per-example RMIA power-law).
-Fix a target example $(x,y_{x})$. For the simplified model with infinitely many shadow models, the per-example RMIA vulnerability is given as
-$\textsc{tpr}_{\mathrm{RMIA}}(x)\leq 1-F_{|t|}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\psi(x,C)}{\sqrt{S}s}\right),$ (21)
-where $\alpha\geq\textsc{fpr}_{\mathrm{RMIA}}(x)$ and
-$\psi(x,C)=\\ \frac{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[2||x-z||^{2}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[(||x-m_{x}||^{2}+||z-m_{z}||^{2})\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[2||x-z||\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{(z,y_{z})\sim{\mathbb{D}}}\left[(||x-m_{x}||+||z-m_{z}||)\mid y_{z}\not=y_{x}\right]}.$ (22)
-In addition, for large $S$ we have
-$\log(\textsc{tpr}_{\mathrm{RMIA}}(x)-\textsc{fpr}_{\mathrm{RMIA}}(x))\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\alpha)^{2}+\log\frac{\psi(x,C)}{\sqrt{\pi/2}}.$ (23)
-Proof.
-See Section C.4. ∎
-As for the LiRA power-law, bounding $||x-z||,||x-m_{x}||$ and $||z-m_{z}||$ will provide a worst-case upper bound for which the power-law holds.
-Now the following corollaries extend the power-law to the average-case MIA vulnerabilities.
-We will also empirically validate these results in Section 4.
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{||\bm{x}-m_{\bm{x}}||}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}\\ +\log(1+\xi(S)).$ (11)
+This implies that if $||\bm{x}-m_{\bm{x}}||$ is bounded, then the worst-case vulnerability is also bounded.
+Hence we can significantly reduce the MIA vulnerability of all examples in this non-DP setting by simply increasing the number of examples per class.
+Now the following corollary extends the power-law to the average-case MIA vulnerability. We will also empirically validate this result in Section 4.
 Corollary 3 (Average-case LiRA power-law).
 For the simplified model with arbitrary $C$, sufficiently large $S$ and infinitely many shadow models, we have
-$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\langle x,x-m_{x}\rangle}{\sqrt{2\pi}||x||s}\right]\right).$ (24)
+$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}\right]\right).$ (12)
 Proof.
-See Section C.5. ∎
-Corollary 6 (Average-case RMIA power-law).
-For the simplified model with sufficiently large $S$ and infinitely many shadow models, we have
-$\log(\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}})\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\alpha)^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\psi(x,C)}{\sqrt{2\pi}}\right]\right).$ (25)
-Proof.
-See Section C.6. ∎
+See Section A.6. ∎
 4 Empirical evaluation of MIA vulnerability and dataset properties
-In this section, we investigate how different properties of datasets affect the MIA vulnerability.
+In this section, we investigate how different properties of datasets (shots $S$ and number of classes $C$) affect the MIA vulnerability.
 Based on our observations, we propose a method to predict the vulnerability to MIA using these properties.
-4.1 Experimental setup
 We focus on a image classification setting where we fine-tune pre-trained models on sensitive downstream datasets and assess the MIA vulnerability using LiRA and RMIA with $M=256$ shadow/reference models.
 We base our experiments on a subset of the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a test classification accuracy $>80\%$ (see Table A2).
+4.1 Experimental setup
+Figure 2: Small effect of number of classes $C$ (classes) on MIA vulnerability when attacking a fine-tuned ViT-B Head.
+The solid line is median and the error bars the min/max bounds for the Clopper-Pearson CIs over 12 seeds ($S=32$).
 We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k (Russakovsky et al., 2015).
-The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section F.1.
-We optimise the hyperparameters (batch size, learning rate and number of epochs) using the library Optuna (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section E.2).
+The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section D.1.
+We optimise the hyperparameters (batch size, learning rate and number of epochs) using the Optuna library (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section C.2).
 We provide the the code for reproducing the experiments in the supplementary material.
-Measuring the uncertainty for tpr The tpr values from the LiRA-based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples.
-Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimators.
-Therefore, when we report the tpr values for a single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals (Clopper & Pearson, 1934).
-Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
-$\begin{split}&B(\alpha/2;\textsc{tp},\textsc{p}-\textsc{tp}+1)<\textsc{tpr}\\ &\textsc{tpr}<B(1-\alpha/2;\textsc{tp}+1,\textsc{p}-\textsc{tp}),\end{split}$ (26)
-where $B(q;a,b)$ is the $q$th-quantile of $\text{Beta}(a,b)$ distribution.
 4.2 Experimental results
@@
 We make the following observations:
-  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in the Appendix (Figures A.2, A4 and A5).
-  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 1(b) with further evidence in Figures A.3, A6 and A7 in the Appendix.
+  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1.
+We provide tabular data and experiments using ResNet-50 in the Appendix (Figures A.1, A3 and A4).
+  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 2 with tabular data and experiments using ResNet-50 in the Appendix (Figures A.2, A5 and A6).
 However, the trend w.r.t.
 $C$ is not as clear as with $S$.
-(a) as function of $S$ (shots)
-(b) as function of $C$ (classes) at $S=32$
-[…]
 RMIA In Figure 3 we compare the vulnerability of the models to LiRA and RMIA as a function of the number of $S$ (shots) at $\textsc{fpr}=0.1$.
 We observe the power-law for both attacks, but the RMIA is more unstable than LiRA (especially for lower fpr).
-More results for RMIA are in Figures A.6, A.7 and A.8 in the Appendix.
+More results for RMIA are in Figures A.5, A.6 and A.7 in the Appendix.
 […]
@@
 The trends seen in Figure 1 suggest the same power law relationship that we derived for the simplified model of membership inference in Section 3.
-We fit a linear regression model to predict $\log(\textsc{tpr}-\textsc{fpr})$ for each $\textsc{fpr}=10^{-k},k=1,\dots,5$ separately using the $\log C$ and $\log S$ as covariates with statsmodels (Seabold & Perktold, 2010).
+We fit a linear regression model to predict $\log(\textsc{tpr}-\textsc{fpr})$ for each $\textsc{fpr}=10^{-k},k=1,\dots,5$ separately using the $\log C$ and $\log S$ as covariates with statsmodels (Seabold and Perktold, 2010).
 The general form of the model can be found in Equation 13, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
@@
 For larger fpr, the coefficient $\beta_{S}$ is around $-0.5$, as our theoretical analysis predicts.
-However, the coefficient value decreases for small fpr.
-This is perhaps because the power-law in Equation (24) only holds for large $S$, and for small fpr Equation (24) significantly underestimates the vulnerability in small-$S$ regime (see Appendix D).
 Figure 4: Coefficient values for different fpr when fitting a regression model based on Equation 13 fitted on data from ViT-B (Head) with LiRA (Table A3).
 The error bars display the 95% confidence intervals based on Student’s t-distribution.
+Theoretical values in the simplified model is shown by pink dotted lines ($\beta_{S}=0.5$ and $\beta_{C}=0$).
 Prediction quality on other MIA target models
@@
   • R-50 (FiLM): Figure 5(b) shows that the prediction quality is good for R-50 (FiLM) models.
-These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section E.1).
+These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section C.1).
 Tobaben et al.
@@
 We are not using this optimization.
+(a) The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100).
+The linear model is robust to changing the feature extractor from ViT-B to R-50.
+(b) Regression model is robust to changing the fine-tuning method from Head to FiLM, but from scratch training seems to be more vulnerable than predicted.
+1 left: fine-tuned with FiLM (see Table A7) 2 right: trained from scratch.
+Data is from Carlini et al.
+(2022) .
+Figure 5: Performance of the regression model based on Equation 13 fitted on data from Table A3.
+4.4 Individual MIA vulnerability
+In order to assess the per-sample MIA vulnerability, we run the experiment with $257$ models with each of them once acting as a target and as a shadow model otherwise, compute the tpr at fpr for every sample separately.
+In Figure 6, we display the individual vulnerability as a function of $S$ (shots) for Patch Camelyon.
+The plot shows the maximal vulnerability of all samples and different quantiles over six seeds.
+The solid line is the median and the errorbars display the min and max over seeds.
+The quantiles are more robust to extreme outliers and show decreasing trends already at much lower $S$ than the maximum vulnerability.
+When fitting the model in Equation 13 we observe that the coefficients $\beta_{S}$ that model the relationship between vulnerability and examples per class for the quantiles are $-0.5603$, $-0.5688$ and $-0.4796$ which is close to the theoretical value of $-0.5$ derived in the theoretical analysis in Section 3.4.
+However, the maximum vulnerability decreases with a lower slope of $-0.2695$ which is considerably smaller.
+With larger $S$ the slope of the max vulnerability increases, e.g., with $S\geq 32768$ the coefficient is $-0.3478$ suggesting that higher $S$ is required for the most vulnerable points.
+Figure 6: Individual vulnerability for ViT-B (Head) when fine-tuning on Patch Camelyon.
+We observe a similar power-law relationship for individuals when looking at the quantiles but the max decreases slower.
+Table 1: The minimum $S$ with $C$=$2$ predicted by the models in Sections 4.3 and 4.4 to empirically match the DP bounds ($\delta$=$10^{-5}$) calculated through (Kairouz et al., 2015) in terms of tpr at fpr.
+$\epsilon$ min $S$ in average case min $S$ in worst-case fpr=$0.1$ fpr=$0.01$ fpr=$0.001$ fpr=$0.1$ $0.25$ 5 40054005\,4005 400 69 0006900069\,00069 000 320 000320000320\,000320 000 $5.5\times 10^{9}$ $0.50$ 1 10011001\,1001 100 16 0001600016\,00016 000 88 0008800088\,00088 000 $2.6\times 10^{8}$ $0.75$ $360$ 5 90059005\,9005 900 38 0003800038\,00038 000 $3.5\times 10^{7}$ $1.00$ $160$ 2 70027002\,7002 700 19 0001900019\,00019 000 $7.0\times 10^{6}$
+4.5 Comparison between empirical models and universal DP bounds
+While the practical evaluation through MIAs is statistical and does not provide universal formal guarantees like DP, the power-law can aid understanding about how the practical vulnerability to MIA behaves in a more realistic threat model when the examples per class increase.
 […]
-(b) Evaluating prediction performance for target models 1 left: fine-tuned with FiLM (data partially from Tobaben et al.
-(2023), see Table A8) 2 right: trained from scratch (data from Carlini et al.
-(2022)) .
-The linear model is robust to changing the fine-tuning method from Head to FiLM, but from scratch training seems to be more vulnerable than predicted.
-Note that the results from Carlini et al.
-(2022) use data augmentation while we do not.
-Figure 4: Performance of the regression model based on Equation 27 fitted on data from Table A4.
 5 Discussion
-Under the GDPR and similar legal regimes, machine learning (ML) models trained on personal data that memorise the data are personal data and need to be carefully protected.
-Our work analyses in which cases trained models would most likely be personal data and in which cases they might not be.
-This will help in evaluating the risk of different kinds of models, favouring less risky models when possible and paying extra attention to more risky cases.
-As the best means of protecting privacy, differential privacy, reduces the utility of models, it is important to understand when it is necessary.
-Our results highlight that models are most vulnerable to MIA when the number of examples per class is low.
-This reinforces the well-known unfortunate dilemma: differential privacy seems most important in the regime where it likely reduces the model utility the most.
-One major reason for MIA vulnerability can be memorisation of the training data.
-Feldman & Zhang (2020) experimentally test memorisation in neural network training, and find that according to their definition, a large fraction of the training data are memorised when training from scratch, while only few are when fine-tuning.
-This is aligned with our results that indicate from scratch training to be more vulnerable than fine-tuning.
-Our model singles out the number of examples per class $S$ as the key predictor of model vulnerability.
-While our results are based on balanced datasets, it seems likely that minority classes and outliers would be more vulnerable as predicted by our model given their smaller $S$.
-Studying the vulnerability in unbalanced data and in the presence of outliers is an important area of future work.
+Trying to bridge empirical MIA vulnerability and formal DP guarantees is not an easy task because of different threat models and different nature of bounds (statistical vs.
+universal).
+While we were able to show both theoretically (Section 3) and empirically (Section 4) that having more examples per class provides protection against MIA in fine-tuned neural networks, the numbers required for significant protection (Section 4.5) limit the practical utility of this observation.
+Using a level of formal DP bounds that provide meaningful protection as a yardstick, at least tens of thousands of examples per class are needed for every class even at $\textsc{fpr}=0.001$.
+The $\epsilon$ values used here are formal upper bounds and must not be compared to actual DP deep learning with comparable privacy budgets, as the latter would be far less vulnerable.
+At this number of samples per class and a good pre-trained model, the impact of DP training is often negligible.
+This stresses the importance of formal guarantees like DP for privacy protection.
+Broader Impact
+Our work systematically studies factors influencing privacy risk of trained ML models.
+This has significant positive impact on users training ML models on personal data by allowing them to understand and limit the risks.
 Limitations
-Our work measures the vulnerability through running two particular MIAs.
-There is no guarantee that models that are not vulnerable to this attack might not be vulnerable to some stronger attacks.
-Formal bounds on attack susceptibility would require something like differential privacy.
-Both our theoretical and empirical analysis focus on deep transfer learning using fine-tuning.
+We mostly consider LiRA in our paper which is optimal for our simiplified model of membership inference (Section 3.4) but for the transfer-learning experiments (Section 4) there might be stronger attacks in the future.
+Furthermore, our results assume well-behaved underlying distributions.
+Formal bounds on MIA vulnerability would require something like DP.
+In addition, both our theoretical and empirical analysis focus on deep transfer learning using fine-tuning.
 Models trained from scratch are likely to be more vulnerable.
+We will publish the code in an open repository after acceptance of the paper.
 Acknowledgments
@@
 Bronskill for helpful discussions regarding few-shot learning.
-Appendix A Formulating LiRA and RMIA for Section 3
-Let $\mathcal{M}$ be our target model and $\ell(\mathcal{M}(x),y_{x})$ be the loss of the model on a target example $(x,y_{x})$.
-The goal of MIA is to determine whether $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$.
+Supplementary Material
+Appendix A Details of Section 3
+A.1 Formulating LiRA
+Let $\mathcal{M}$ be our target model and $\ell(\mathcal{M}(\bm{x}),y)$ be the loss of the model on a target example $(\bm{x},y)$.
+The goal of MIA is to determine whether $(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}$.
 This can be formulated as a hypothesis test:
-$H_{0}:(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$ (A1)
-$H_{1}:(x,y_{x})\in\mathcal{D}_{\mathrm{target}}.$ (A2)
-A.1 LiRA
+$H_{0}:(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}}$ (A1)
+$H_{1}:(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}.$ (A2)
 Following (Carlini et al., 2022), we formulate the Likelihood Ratio Attack (LiRA).
@@
 To model the IN/OUT loss distributions with few shadow models, LiRA employs a parametric modelling.
-Particularly, LiRA models $t_{x}$ by a normal distribution.
+Particularly, LiRA models $t_{\bm{x}}$ by a normal distribution.
 That is, the hypothesis test formulated above can be rewritten as
-$H_{0}:t_{x}\sim\mathcal{N}(\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})$ (A3)
-$H_{1}:t_{x}\sim\mathcal{N}(\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}(x)^{2}).$ (A4)
+$H_{0}^{\prime}:t_{\bm{x}}\sim\mathcal{N}(\hat{\mu}_{\mathrm{out}},\hat{\sigma}_{\mathrm{out}}^{2})$ (A3)
+$H_{1}^{\prime}:t_{\bm{x}}\sim\mathcal{N}(\hat{\mu}_{\mathrm{in}},\hat{\sigma}_{\mathrm{in}}^{2}).$ (A4)
 The likelihood ratio is now
-$\mathrm{LR}(x)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x)^{2},\hat{\sigma}_{\mathrm{in}}(x)^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})}.$ (A5)
-LiRA rejects $H_{0}$ if and only if
-$\mathrm{LR}(x)\geq\beta,$ (A6)
-concluding that $H_{1}$ is true, i.e., identifying the membership of $(x,y_{x})$. Thus, the true positive rate of this hypothesis test given as
-$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}_{\mathrm{in}}(x)^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}_{\mathrm{out}}(x)^{2})}\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\right),$ (A7)
+$\mathrm{LR}(\bm{x})=\frac{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\mathrm{in}},\hat{\sigma}_{\mathrm{in}}^{2})}{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\mathrm{out}},\hat{\sigma}_{\mathrm{out}}^{2})}.$ (A5)
+LiRA rejects $H_{0}^{\prime}$ if and only if
+$\mathrm{LR}(\bm{x})\geq\tau,$ (A6)
+concluding that $H_{1}^{\prime}$ is true, i.e., identifying the membership of $(\bm{x},y)$. Thus, the true positive rate of this hypothesis test given as
+$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\mathrm{LR}(\bm{x})\geq\tau\mid(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}\right),$ (A7)
 where $\phi^{M}$ denotes the randomness in the shadow set sampling and shadow model training.
-A.2 RMIA
-By modelling $t_{z}$ by a normal distribution, Zarifzadeh et al. (2024) approximate the pairwise likelihood ratio as
-$\mathrm{LR}(x,z)=\frac{p(\theta\mid x,y_{x})}{p(\theta\mid z,y_{z})}\approx\frac{\mathcal{N}(t_{x};\hat{\mu}_{x,x},\hat{\sigma}_{x,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{x,z},\hat{\sigma}_{x,z}^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{z,x},\hat{\sigma}_{z,x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{z,z},\hat{\sigma}_{z,z}^{2})},$ (A8)
-where $\hat{\mu}_{a,b}$ and $\hat{\sigma}_{a,b}$ are, respectively, the mean and standard deviation of $t_{b}$ estimated from shadow models when the training set contains $a$ but not $b$.
-Then RMIA exploits the probability of rejecting the pairwise likelihood ratio test over $(z,y_{z})\sim{\mathbb{D}}$:
-$\mathrm{Score}_{\mathrm{RMIA}}(x)=\Pr_{(z,y_{z})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,z)\geq\gamma\right).$ (A9)
-Thus, RMIA rejects $H_{0}$ if and only if
-$\Pr_{(z,y_{z})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta,$ (A10)
-identifying the membership of $x$. Hence the true positive rate of RMIA is given as
-$\textsc{tpr}_{\mathrm{RMIA}}(x)=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\Pr_{(z,y_{z})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\right),$ (A11)
-where $\phi^{M}$ denotes the randomness in the shadow set sampling and shadow model training.
-Appendix B On the assumptions in Section 3
-B.1 The assumption of shared variances
-In Section 3 we assumed that for LiRA $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}=\hat{\sigma}_{\mathrm{out}}$, and that for RMIA $\sigma_{x}=\sigma_{x,x}=\sigma_{z,x}$, $\sigma_{z}=\sigma_{x.z}=\sigma_{z,z}$, $\hat{\sigma}_{x}=\hat{\sigma}_{x,x}=\hat{\sigma}_{z,x}$ and $\hat{\sigma}_{z}=\hat{\sigma}_{x.z}=\hat{\sigma}_{z,z}$.
-Utilising the simplified model formulated in Section 3.4, we show that for large enough number $S$ of examples per class these assumptions are reasonable.
-From the proof of Theorem 3 (see Section C.3), the LiRA scores are
-$s_{y_{x}}^{\mathrm{(in)}}$ $=\langle x,\frac{1}{S}\left(\sum_{i=1}^{S-1}x_{y_{x},i}+x\right)\rangle=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle+\langle x,\frac{1}{S}(x-x_{y_{x},S})\rangle$ (A12)
-$s_{y_{x}}^{\mathrm{(out)}}$ $=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle,$ (A13)
-respectively, when $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$ and when $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$.
-Then, since for the simplified model scores follow Gaussian distributions, we have
-$\sigma_{\mathrm{in}}^{2}$ $=\hat{\sigma}_{\mathrm{in}}^{2}=\operatorname{Var}(s_{y_{x}}^{\mathrm{(in)}})=\frac{1}{S}\left(1-\frac{1}{S}\right)||x||^{2}s^{2}$ (A14)
-$\sigma_{\mathrm{out}}^{2}$ $=\hat{\sigma}_{\mathrm{out}}^{2}=\operatorname{Var}(s_{y_{x}}^{\mathrm{(out)}})=\frac{1}{S}||x||^{2}s^{2}.$ (A15)
-Thus, the differences $\sigma_{\mathrm{in}}-\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}-\hat{\sigma}_{\mathrm{out}}$ are negligible for large $S$. Similarly, we have
-$\sigma_{x,x}^{2}$ $=\hat{\sigma}_{x,x}^{2}=\operatorname{Var}(s_{y_{x}}^{(x)}(x))=\frac{1}{S}\left(1-\frac{1}{S}\right)||x||^{2}s^{2}$ (A16)
-$\sigma_{z,x}^{2}$ =$\hat{\sigma}_{z,x}^{2}=\operatorname{Var}(s_{y_{x}}^{(z)}(x))=\begin{cases}\frac{1}{S}\left(1-\frac{1}{S}\right)||x||^{2}s^{2}\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}||x||^{2}s^{2}\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A17)
-$\sigma_{x,z}^{2}$ =$\hat{\sigma}_{x,z}^{2}=\operatorname{Var}(s_{y_{z}}^{(x)}(z))=\begin{cases}\frac{1}{S}\left(1-\frac{1}{S}\right)||z||^{2}s^{2}\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}||z||^{2}s^{2}\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A18)
-$\sigma_{z,z}^{2}$ $=\hat{\sigma}_{z,z}^{2}=\operatorname{Var}(s_{y_{z}}^{(z)}(z))=\frac{1}{S}\left(1-\frac{1}{S}\right)||z||^{2}s^{2}.$ (A19)
-Therefore, the differences $\sigma_{x,x}-\sigma_{z,x}$, $\sigma_{x.z}-\sigma_{z,z}$, $\hat{\sigma}_{x,x}-\hat{\sigma}_{z,x}$ and $\hat{\sigma}_{x.z}-\hat{\sigma}_{z,z}$ are negligible for large enough $S$.
-Hence as long as the simplified model approximates classification tasks to which Lemmas 1 and 2 are applied, these assumptions are reasonably justified.
-B.2 Relaxing the assumption of Lemma 1
-In Lemma 1 we assume that an attacker has access to the true underlying distribution.
-However, this is not necessarily the case in real-world settings.
-Noting that the Equation 12 mainly relies on the true location parameters $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and scale parameter $\sigma(x)$, we may relax this assumption.
-Notice that if we completely drop this assumption so that an attacker trains shadow models with an arbitrary underlying distribution, then we may not be able to choose a desired $\textsc{fpr}_{\mathrm{LiRA}}(x)$.
-From Equation A36 in the proof of Lemma 1 we have
-$\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}=\begin{cases}F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}\end{cases}$ (A20)
-$\log\beta=\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}},\end{cases}$ (A21)
-where we abuse notations by denoting $\mu_{\mathrm{in}}$ to refer to $\mu_{\mathrm{in}}(x)$ and similarly for other parameters.
-Since we need to choose a rejection region of the likelihood ratio test such that $\beta\geq 1$, we have
-$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A22)
-Therefore, the sufficient condition about attacker’s knowledge on the underlying distribution for Lemma 1 to hold is
-$\begin{cases}\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\leq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A23)
-We summarise this discussion in the following:
-Lemma 7 (Lemma 1 with a relaxed assumption).
-Suppose that the true distribution of $t_{x}$ is of location-scale family with locations $\mu_{\mathrm{in}}(x),\mu_{\mathrm{out}}(x)$ and scale $\sigma(x)$, and that LiRA models $t_{x}$ by $\mathcal{N}(\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}(x))$ and $\mathcal{N}(\hat{\mu}_{\mathrm{out}}(x),\hat{\sigma}(x))$.
-Assume that an attacker estimates parameters $\hat{\mu}_{\mathrm{in}}(x),\hat{\sigma}(x),\hat{\mu}_{\mathrm{out}}(x)$ and $\hat{\sigma}_{\mathrm{out}}(x)$ with an approximated underlying distribution such that
-$\begin{cases}\sigma F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}(x)+\hat{\mu}_{\mathrm{out}}(x)}{2}+\mu_{\mathrm{out}}(x)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}(x)>\hat{\mu}_{\mathrm{out}}(x)\\ \sigma F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\hat{\mu}_{\mathrm{in}}(x)+\hat{\mu}_{\mathrm{out}}(x)}{2}+\mu_{\mathrm{out}}(x)\leq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}(x)<\hat{\mu}_{\mathrm{out}}(x).\end{cases}$ (A24)
-Then the LiRA vulnerability of a fixed target example is
-$\textsc{tpr}_{\mathrm{LiRA}}(x)=\begin{cases}1-F_{t}\left(F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)>\hat{\mu}_{\mathrm{out}}(x)\\ F_{t}\left(F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}(x)-\mu_{\mathrm{out}}(x)}{\sigma(x)}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}(x)<\hat{\mu}_{\mathrm{out}}(x),\end{cases}$ (A25)
-where $F_{t}$ is the cdf of $t$ with the standard location and unit scale.
-Appendix C Missing proofs of Section 3
-C.1 Proof of Lemma 1
+A.2 On the assumption of shared scale
+In Section 3 we assumed that for LiRA $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}=\hat{\sigma}_{\mathrm{out}}$.
+Using the simplified model formulated in Section 3.4, we show that for large enough number $S$ of examples per class these assumptions are reasonable.
+Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),...,(\bm{x}_{j,S},j)\}_{j=1}^{C}$. Then the IN/OUT LiRA scores are given as
+$s_{y}^{\mathrm{(in)}}$ $=\langle\bm{x},\frac{1}{S}\left(\sum_{i=1}^{S-1}\bm{x}_{y,i}+\bm{x}\right)\rangle=\langle\bm{x},\frac{1}{S}\sum_{i=1}^{S}\bm{x}_{y,i}\rangle+\langle\bm{x},\frac{1}{S}(\bm{x}-\bm{x}_{y,S})\rangle$ (A8)
+$s_{y}^{\mathrm{(out)}}$ $=\langle\bm{x},\frac{1}{S}\sum_{i=1}^{S}\bm{x}_{y,i}\rangle.$ (A9)
+Since for the simplified model scores follow Gaussian distributions, $\sigma_{\mathrm{in}}=\hat{\sigma}_{\mathrm{in}}$ and $\sigma_{\mathrm{out}}=\hat{\sigma}_{\mathrm{out}}$. It follows that
+$\sigma_{\mathrm{in}}^{2}$ $=\hat{\sigma}_{\mathrm{in}}^{2}=\operatorname{Var}(s_{y}^{\mathrm{(in)}})=\frac{1}{S}\operatorname{Var}(\langle\bm{x},\bm{x}_{y,i}\rangle)-\frac{1}{S^{2}}\operatorname{Var}(\langle\bm{x},\bm{x}_{y,i}\rangle)=\frac{1}{S}\left(1-\frac{1}{S}\right)\bm{x}^{T}\Sigma\bm{x}$ (A10)
+$\sigma_{\mathrm{out}}^{2}$ $=\hat{\sigma}_{\mathrm{out}}^{2}=\operatorname{Var}(s_{y}^{\mathrm{(out)}})=\frac{1}{S}\operatorname{Var}(\langle\bm{x},\bm{x}_{y,i}\rangle)=\frac{1}{S}\bm{x}^{T}\Sigma\bm{x}.$ (A11)
+Thus, the differences $\sigma_{\mathrm{in}}-\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}-\hat{\sigma}_{\mathrm{out}}$ are negligible for large $S$.
+A.3 Proof of Lemma 1
 See 1
 Proof.
-We abuse notations by denoting $\mu_{\mathrm{in}}$ to refer to $\mu_{\mathrm{in}}(x)$ and similarly for other statistics. We have
-$\log\frac{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{in}},\hat{\sigma})}{\mathcal{N}(t_{x};\hat{\mu}_{\mathrm{out}},\hat{\sigma})}\geq\log\beta$ (A26)
-$-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{\mathrm{in}}}{\hat{\sigma}}\right)^{2}+\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}}\right)^{2}\geq\log\beta$ (A27)
-$\frac{1}{2\hat{\sigma}^{2}}(2t_{x}\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{in}}^{2}-2t_{x}\hat{\mu}_{\mathrm{out}}+\hat{\mu}_{\mathrm{out}}^{2})\geq\log\beta$ (A28)
-$\frac{1}{2\hat{\sigma}^{2}}(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})(2t_{x}-\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})\geq\log\beta$ (A29)
-$\begin{cases}t_{x}\geq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ t_{x}\leq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A30)
-When $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$, $t_{x}\sim\mathcal{N}(\mu_{\mathrm{out}},\sigma)$.
-Therefore, if $\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}$, in the limit of infinitely many shadow models
-$\textsc{fpr}_{\mathrm{LiRA}}(x)$ $=\Pr_{t}\left(\mu_{\mathrm{out}}+\sigma t\geq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A31)
-$=\Pr_{t}\left(t\geq\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}\right)$ (A32)
-$=1-F_{t}\left(\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}\right),$ (A33)
+We abuse notations by denoting $\mu_{\mathrm{in}}$ to refer to $\mu_{\mathrm{in}}(\bm{x})$ and similarly for other statistics. We have
+$\log\frac{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\mathrm{in}},\hat{\sigma}^{2})}{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\mathrm{out}},\hat{\sigma}^{2})}\geq\log\tau$ (A12)
+$-\frac{1}{2}\left(\frac{t_{\bm{x}}-\hat{\mu}_{\mathrm{in}}}{\hat{\sigma}}\right)^{2}+\frac{1}{2}\left(\frac{t_{\bm{x}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}}\right)^{2}\geq\log\tau$ (A13)
+$\frac{1}{2\hat{\sigma}^{2}}(2t_{\bm{x}}\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{in}}^{2}-2t_{\bm{x}}\hat{\mu}_{\mathrm{out}}+\hat{\mu}_{\mathrm{out}}^{2})\geq\log\tau$ (A14)
+$\frac{1}{2\hat{\sigma}^{2}}(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})(2t_{\bm{x}}-\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})\geq\log\tau$ (A15)
+$\begin{cases}t_{\bm{x}}\geq\frac{\hat{\sigma}^{2}\log\tau}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ t_{\bm{x}}\leq\frac{\hat{\sigma}^{2}\log\tau}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A16)
+Then if $\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}$, in the limit of infinitely many shadow models
+$\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Pr_{Z}\left(\mu_{\mathrm{out}}+\sigma Z\geq\frac{\hat{\sigma}^{2}\log\tau}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A17)
+$=\Pr_{Z}\left(Z\geq\frac{\hat{\sigma}^{2}\log\tau}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}\right)$ (A18)
+$=1-F_{Z}\left(\frac{\hat{\sigma}^{2}\log\tau}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}\right),$ (A19)
 and if $\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}$, similarly,
-$\textsc{fpr}_{\mathrm{LiRA}}(x)$ $=\Pr_{t}\left(\mu_{\mathrm{out}}+\sigma t\leq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A34)
-$=F_{t}\left(\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}\right).$ (A35)
+$\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Pr_{Z}\left(\mu_{\mathrm{out}}+\sigma Z\leq\frac{\hat{\sigma}^{2}\log\tau}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A20)
+$=F_{Z}\left(\frac{\hat{\sigma}^{2}\log\tau}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}\right).$ (A21)
 Thus
-$\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}=\begin{cases}F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A36)
+$\frac{\hat{\sigma}^{2}\log\tau}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}=\begin{cases}F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A22)
 It follows that if $\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}$,
-$\textsc{tpr}_{\mathrm{LiRA}}(x)$ $=\Pr_{t}\left(\mu_{\mathrm{in}}+\sigma t\geq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A37)
-$=\Pr_{t}\left(t\geq\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{in}}}{\sigma}\right)$ (A38)
-$=1-F_{t}\left(F_{t}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right).$ (A39)
+$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Pr_{Z}\left(\mu_{\mathrm{in}}+\sigma Z\geq\frac{\hat{\sigma}^{2}\log\tau}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A23)
+$=\Pr_{Z}\left(Z\geq\frac{\hat{\sigma}^{2}\log\tau}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{in}}}{\sigma}\right)$ (A24)
+$=1-F_{Z}\left(F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right).$ (A25)
 If $\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}$, then
-$\textsc{tpr}_{\mathrm{LiRA}}(x)$ $=\Pr_{t}\left(\mu_{\mathrm{in}}+\sigma t\leq\frac{\hat{\sigma}^{2}\log\beta}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A40)
-$=\Pr_{t}\left(t\leq\frac{\hat{\sigma}^{2}\log\beta}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{in}}}{\sigma}\right)$ (A41)
-$=F_{t}\left(F_{t}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right).$ (A42)
-C.2 Proof of Lemma 2
+$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Pr_{Z}\left(\mu_{\mathrm{in}}+\sigma Z\leq\frac{\hat{\sigma}^{2}\log\tau}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}\right)$ (A26)
+$=\Pr_{Z}\left(Z\leq\frac{\hat{\sigma}^{2}\log\tau}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{in}}}{\sigma}\right)$ (A27)
+$=F_{Z}\left(F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right).$ (A28)
+A.4 Relaxing the assumption of Lemma 1
+In Lemma 1 we assume that an attacker has access to the true underlying distribution.
+However, in real-world settings the precise underlying distribution may not be available for an attacker.
+In the following, noting that the Equation 8 mainly relies on the true location parameters $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and scale parameter $\sigma(\bm{x})$, we relax this assumption of distribution availability so that the attacker only needs an approximated underlying distribution for the optimal LiRA that achieves the performance in Lemma 1.
+First, notice that if we completely drop this assumption so that an attacker trains shadow models with an arbitrary underlying distribution, then we may not be able to choose a desired $\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$.
+From Equation A22 we have
+$\frac{\hat{\sigma}^{2}\log\tau}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}=\begin{cases}F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}\end{cases}$ (A29)
+$\log\tau=\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A30)
+Since it does not make sense to choose a rejection region of the likelihood ratio test such that $\tau<1$, we assume $\tau\geq 1$. Then we have
+$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A31)
+Therefore, a sufficient condition about attacker’s knowledge on the underlying distribution for Lemma 1 to hold is
+$\begin{cases}\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\leq 0&\quad\text{if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A32)
+Rearranging the terms yields
+$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\leq\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\geq\sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A33)
+For example, if the estimated mean $\hat{\mu}_{\mathrm{out}}(<\hat{\mu}_{\mathrm{in}})$ is too large compared to the true parameter $\mu_{\mathrm{out}}$, the left hand side for the case $\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}$ becomes very large, thereby forcing us to choose sufficiently small $\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$.
+Similarly, if $\hat{\mu}_{\mathrm{out}}(>\hat{\mu}_{\mathrm{in}})$ is much smaller than $\mu_{\mathrm{out}}$, then the range of possible values of $\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$ will be limited.
+We summarise this discussion in the following:
+Lemma 4 (Lemma 1 with relaxed assumptions).
+Suppose that the true IN/OUT distributions of $t_{\bm{x}}$ are of a location-scale family with locations $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and a shared scale $\sigma(\bm{x})$ such that the distributions have finite first and second moments.
+Assume that LiRA models $t_{\bm{x}}$ by $\mathcal{N}(\hat{\mu}_{\mathrm{in}}(\bm{x}),\hat{\sigma}(\bm{x})^{2})$ and $\mathcal{N}(\hat{\mu}_{\mathrm{out}}(\bm{x}),\hat{\sigma}(\bm{x})^{2})$, and that in the limit of infinitely many shadow models, estimated parameters $\hat{\mu}_{\mathrm{in}}(\bm{x}),\hat{\mu}_{\mathrm{out}}(\bm{x})$ and $\hat{\sigma}(\bm{x})$ satisfy the following:
+$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\leq\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\geq\sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A34)
+Then in the limit of infinitely many shadow models, the LiRA vulnerability of a fixed target example $(\bm{x},y)$ is
+$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})=\begin{cases}1-F_{Z}\left(F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}\right)&\text{ if }\ \hat{\mu}_{\mathrm{in}}(\bm{x})>\hat{\mu}_{\mathrm{out}}(\bm{x})\\ F_{Z}\left(F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}\right)&\text{ if }\ \hat{\mu}_{\mathrm{in}}(\bm{x})<\hat{\mu}_{\mathrm{out}}(\bm{x}),\end{cases}$ (A35)
+where $F_{Z}$ is the cdf of $t$ with the standard location and unit scale, assuming that the inverse of $F_{Z}$ exists.
+A.5 Proof of Theorem 2
 See 2
 Proof.
-We abuse notations by denoting probabilities and expectations over sampling $\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|}$ and $(z,y_{z})\sim{\mathbb{D}}$ by subscripts $\mathcal{D}_{\mathrm{target}}$ and $z$.
-Recall that
-$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{\mathcal{D}_{\mathrm{target}}}\left(\Pr_{z}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta\mid(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\wedge(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\right)$ (A43)
-$\textsc{fpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{\mathcal{D}_{\mathrm{target}}}\left(\Pr_{z}\left(\mathrm{LR}(x,z)\geq\gamma\right)\geq\beta\mid(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}\wedge(x,y_{x})\in\mathcal{D}_{\mathrm{target}}\right),$ (A44)
-where
-$\mathrm{LR}(x,z)=\frac{\mathcal{N}(t_{x};\hat{\mu}_{x,x},\hat{\sigma}_{x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{x,z},\hat{\sigma}_{z}^{2})}{\mathcal{N}(t_{x};\hat{\mu}_{z,x},\hat{\sigma}_{x}^{2})\mathcal{N}(t_{z};\hat{\mu}_{z,z},\hat{\sigma}_{z}^{2})}.$ (A45)
-We have
-$\lambda:=$ $\log(\mathrm{LR}(x,z))$ (A46)
-$=$ $\log\left(\frac{\frac{1}{\sqrt{2\pi}\hat{\sigma}_{x}}\exp\left(-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{x,x}}{\hat{\sigma}_{x}}\right)^{2}\right)\frac{1}{\sqrt{2\pi}\hat{\sigma}_{z}}\exp\left(-\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{x,z}}{\hat{\sigma}_{z}}\right)^{2}\right)}{\frac{1}{\sqrt{2\pi}\hat{\sigma}_{x}}\exp\left(-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{z,x}}{\hat{\sigma}_{x}}\right)^{2}\right)\frac{1}{\sqrt{2\pi}\hat{\sigma}_{z}}\exp\left(-\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{z,z}}{\hat{\sigma}_{z}}\right)^{2}\right)}\right)$ (A47)
-$=$ $-\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{x,x}}{\hat{\sigma}_{x}}\right)^{2}+\frac{1}{2}\left(\frac{t_{x}-\hat{\mu}_{z,x}}{\hat{\sigma}_{x}}\right)^{2}-\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{x,z}}{\hat{\sigma}_{z}}\right)^{2}+\frac{1}{2}\left(\frac{t_{z}-\hat{\mu}_{z,z}}{\hat{\sigma}_{z}}\right)^{2}$ (A48)
-$=$ $\frac{1}{2\hat{\sigma}_{x}^{2}}(2t_{x}\hat{\mu}_{x,x}-\hat{\mu}_{x,x}^{2}-2t_{x}\hat{\mu}_{z,x}+\hat{\mu}_{z,x}^{2})+\frac{1}{2\hat{\sigma}_{z}^{2}}(2t_{z}\hat{\mu}_{x,z}-\hat{\mu}_{x,z}^{2}-2t_{z}\hat{\mu}_{z,z}+\hat{\mu}_{z,z}^{2})$ (A49)
-$=$ $\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2t_{x}-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2t_{z}-\hat{\mu}_{x,z}-\hat{\mu}_{z,z}).$ (A50)
-When $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$, $t_{x}=\mu_{x,x}+\sigma_{x}t$ and $t_{z}=\mu_{x,z}+\sigma_{z}t$, and thus $\lambda$ becomes
-$\lambda_{x}:=$ $\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2\mu_{x,x}+2\sigma_{x}t-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})$ (A51)
-$+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2\mu_{x,z}+2\sigma_{z}t-\hat{\mu}_{x,z}-\hat{\mu}_{z,z}).$ (A52)
-Similarly, when $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$, we have $t_{x}=\mu_{z,x}+\sigma_{x}t$ and $t_{z}=\mu_{z,z}+\sigma_{z}t$. Then $\lambda$ is now
-$\lambda_{z}:=$ $\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2\mu_{z,x}+2\sigma_{x}t-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})$ (A53)
-$+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2\mu_{z,z}+2\sigma_{z}t-\hat{\mu}_{x,z}-\hat{\mu}_{z,z})$ (A54)
-$=$ $\left(\underbrace{\frac{\sigma_{x}}{\hat{\sigma}_{x}^{2}}(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\sigma_{z}}{\hat{\sigma}_{z}^{2}}(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}_{A}\right)t$ (A55)
-$+\underbrace{\frac{\hat{\mu}_{x,x}-\hat{\mu}_{z,x}}{2\hat{\sigma}_{x}^{2}}(2\mu_{z,x}-\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\hat{\mu}_{x,z}-\hat{\mu}_{z,z}}{2\hat{\sigma}_{z}^{2}}(2\mu_{z,z}-\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}_{B}$ (A56)
-$=$ $At+B.$ (A57)
-Notice that $A$ and $B$ are functions of $z$ and independent of $t$.
-Also note that taking probability over $t$ corresponds to calculating probability over $\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|}$.
-Thus, Since we can set $\gamma>1$, using Markov’s inequality and the triangle inequality, we have in the limit of infinitely many shadow models
-$\textsc{fpr}_{\mathrm{RMIA}}(x)=$ $\Pr_{t}\left(\Pr_{z}(\lambda_{z}\geq\log(\gamma))\geq\beta\right)$ (A58)
-$\leq\Pr_{t}\left(\Pr_{z}(|\lambda_{z}|\geq\log(\gamma))\geq\beta\right)$ (A59)
-$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{z}|]}{\log(\gamma)}\geq\beta\right)$ (A60)
-$=\Pr_{t}\left(\mathbb{E}_{z}[|\lambda_{z}|]\geq\beta\log(\gamma)\right)$ (A61)
-$\leq\Pr_{t}\left(\mathbb{E}_{z}[|A|]|t|+\mathbb{E}_{z}[|B|]\geq\beta\log(\gamma)\right)$ (A62)
-$=\Pr_{t}\left(|t|\geq\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}\right)$ (A63)
-Therefore, we can upper-bound $\textsc{fpr}_{\mathrm{RMIA}}(x)\leq\alpha$ by setting
-$\alpha=1-F_{|t|}\left(\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}\right).$ (A64)
-That is,
-$\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}=F_{|t|}^{-1}(1-\alpha).$ (A65)
-Now let
-$q=\lambda_{x}-\lambda_{z}=\frac{(\mu_{x,x}-\mu_{z,x})(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})}{\hat{\sigma}_{x}^{2}}+\frac{(\mu_{x,z}-\mu_{z,z})(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}{\hat{\sigma}_{z}^{2}}.$ (A66)
-Note that $q$ is also independent of $t$, thereby $\mathbb{E}_{z}[|q|]$ being a constant. By the similar argument using Markov’s inequality and the triangle inequality, we have
-$\textsc{tpr}_{\mathrm{RMIA}}(x)=$ $\Pr_{t}\left(\Pr_{z}(\lambda_{x}\geq\log(\gamma))\geq\beta\right)$ (A67)
-$\leq\Pr_{t}\left(\Pr_{z}(|\lambda_{x}|\geq\log(\gamma))\geq\beta\right)$ (A68)
-$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{x}|]}{\log(\gamma)}\geq\beta\right)$ (A69)
-$=\Pr_{t}\left(\mathbb{E}_{z}[|\lambda_{x}|]\geq\beta\log(\gamma)\right)$ (A70)
-$\leq\Pr_{t}\left(\mathbb{E}_{z}[|A|]|t|+\mathbb{E}_{z}[|B|]+\mathbb{E}_{z}[|q|]\geq\beta\log(\gamma)\right)$ (A71)
-$=\Pr_{t}\left(|t|\geq\frac{\beta\log(\gamma)-\mathbb{E}_{z}[|B|]}{\mathbb{E}_{z}[|A|]}-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right)$ (A72)
-Hence we obtain
-$\textsc{tpr}_{\mathrm{RMIA}}(x)\leq 1-F_{|t|}^{-1}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right).$ (A73)
-C.3 Proof of Theorem 3
+Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),...,(\bm{x}_{j,S},j)\}_{j=1}^{C}$. Then the LiRA score of the target $(\bm{x},y)$ is
+$s_{y}^{\mathrm{(in)}}$ $=\langle\bm{x},\frac{1}{S}\left(\sum_{i=1}^{S-1}\bm{x}_{y,i}+\bm{x}\right)\rangle=\langle\bm{x},\frac{1}{S}\sum_{i=1}^{S}\bm{x}_{y,i}\rangle+\langle\bm{x},\frac{1}{S}(\bm{x}-\bm{x}_{y,S})\rangle$ (A36)
+$s_{y}^{\mathrm{(out)}}$ $=\langle\bm{x},\frac{1}{S}\sum_{i=1}^{S}\bm{x}_{y,i}\rangle,$ (A37)
+respectively, when $(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}$ and when $(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}}$. Thus we obtain
+$\mu_{\mathrm{in}}-\mu_{\mathrm{out}}$ $=\mathbb{E}[s_{y}^{\mathrm{(in)}}-s_{y}^{\mathrm{(out)}}]=\frac{1}{S}\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle$ (A38)
+$\sigma^{2}$ $=\operatorname{Var}(s_{y}^{\mathrm{(out)}})=\frac{1}{S}\operatorname{Var}(\langle\bm{x},\bm{x}_{y,i}\rangle)=\frac{1}{S}\bm{x}^{T}\Sigma\bm{x}.$ (A39)
+Noting that the LiRA score follows a normal distribution, by Lemma 1 we have
+$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right),$ (A40)
+where $\Phi$ is the cdf of the standard normal distribution. This completes the first half of the theorem.
+Now let $\phi(u)$ denote the pdf of the standard normal distribution, and let
+$r=\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}.$ (A41)
+Using Taylor expansion of $\Phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+r)$ around $r=0$, we have
+$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\sum_{k=0}^{\infty}\frac{\Phi^{(k)}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))}{k!}r^{k}$ (A42)
+$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+\sum_{k=1}^{\infty}\frac{\Phi^{(k)}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))}{k!}r^{k}$ (A43)
+$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+\sum_{k=1}^{\infty}\frac{\phi^{(k-1)}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))}{k!}r^{k}$ (A44)
+$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+\sum_{k=1}^{\infty}\frac{(-1)^{k-1}\mathrm{He}_{k-1}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))}{k!}r^{k}$ (A45)
+$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+r\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}r^{k}.$ (A46)
+It follows that
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))$ (A47)
+$=$ $\log r+\log\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+\log\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}r^{k}\right)$ (A48)
+$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A49)
+$+\log\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A50)
+$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A51)
+$+\log\left(1+\sum_{k=1}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A52)
+$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (A53)
+where we have $\xi(S)=O(1/\sqrt{S})$. For large enough $S$, ignoring the residual term, we approximate
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A54)
+A.6 Proof of Corollary 3
 See 3
 Proof.
-Let $\mathcal{D}_{\mathrm{target}}=\{(x_{j,1},j),...,(x_{j,S},j)\}_{j=1}^{C}$. Then the LiRA score of the target $(x,y_{x})$ is
-$s_{y_{x}}^{\mathrm{(in)}}$ $=\langle x,\frac{1}{S}\left(\sum_{i=1}^{S-1}x_{y_{x},i}+x\right)\rangle=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle+\langle x,\frac{1}{S}(x-x_{y_{x},S})\rangle$ (A74)
-$s_{y_{x}}^{\mathrm{(out)}}$ $=\langle x,\frac{1}{S}\sum_{i=1}^{S}x_{y_{x},i}\rangle,$ (A75)
-respectively, when $(x,y_{x})\in\mathcal{D}_{\mathrm{target}}$ and when $(x,y_{x})\notin\mathcal{D}_{\mathrm{target}}$. Thus, in the limit of infinitely many shadow models we obtain
-$\mu_{\mathrm{in}}-\mu_{\mathrm{out}}$ $=\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}=\mathbb{E}[s_{y_{x}}^{\mathrm{(in)}}-s_{y_{x}}^{\mathrm{(out)}}]=\frac{1}{S}\langle x,x-m_{x}\rangle$ (A76)
-$\sigma^{2}$ $=\operatorname{Var}(s_{y_{x}}^{\mathrm{(out)}})=\frac{1}{S}\operatorname{Var}(\langle x,x_{y_{x},i}\rangle)=\frac{1}{S}||x||^{2}s^{2}$ (A77)
-Noting that the LiRA score follows a normal distribution, by Lemma 1 we have
-$\textsc{tpr}_{\mathrm{LiRA}}(x)$ =$=\begin{cases}1-\Phi\left(\Phi^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right)&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}},\end{cases}$ (A78)
-=$=\begin{cases}1-\Phi\left(\Phi^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}\right)&\text{ if }\langle x,x-m_{x}\rangle>0\\ \Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))-\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}\right)&\text{ if }\langle x,x-m_{x}\rangle<0,\end{cases}$ (A79)
-By the symmetry of a normal distribution, this simplifies to
-$\textsc{tpr}_{\mathrm{LiRA}}(x)=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))+\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}s||x||}\right),$ (A80)
-where $\langle x,x-m_{x}\rangle\not=0$ by the assumption that $\hat{\mu}_{\mathrm{in}}\not=\hat{\mu}_{\mathrm{out}}$. This completes the first half of the theorem.
-Now we have
-$\textsc{tpr}_{\mathrm{LiRA}}(x)$ $=\Pr\left(\eta\leq\gamma_{0}+\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}s||x||}\right),$ (A81)
-$\textsc{fpr}_{\mathrm{LiRA}}(x)$ $=\Pr(\eta\leq\gamma_{0}),$ (A82)
-where $\gamma_{0}$ is a tunable constant and $\eta\sim\mathcal{N}(0,1)$. Thus for large enough $S$ we have
-$\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x)$ $\approx p_{\eta}(\gamma_{0})\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}||x||s}$ (A83)
-$=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}}\frac{|\langle x,x-m_{x}\rangle|}{\sqrt{S}||x||s},$ (A84)
-$\log(\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x))$ $\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))^{2}+\log\frac{|\langle x,x-m_{x}\rangle|}{||x||s\sqrt{2\pi}}.$ (A85)
-C.4 Proof of Theorem 4
-See 4
+By theorem 2 and the law of unconscious statistician, we have for large $S$
+$\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}}$ $=\int_{\mathscr{D}}p(\bm{x})(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\mathrm{d}\bm{x}$ (A55)
+$\approx\int_{\mathscr{D}}p(\bm{x})\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\mathrm{d}\bm{x}$ (A56)
+$=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\frac{1}{\sqrt{S}}\int_{\mathscr{D}}p(\bm{x})\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\mathrm{d}\bm{x}$ (A57)
+$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{2\pi}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right],$ (A58)
+where $p(\bm{x})$ is the density of ${\mathbb{D}}$ at $(\bm{x},y)$, and $\mathscr{D}$ is the data domain.
+Note that here we fixed $\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})=\overline{\textsc{fpr}}_{\mathrm{LiRA}}$ for all $\bm{x}$.
+Then we obtain
+$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{2\pi}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right]\right).$ (A59)
+Appendix B Theoretical analysis of RMIA
+Similar to LiRA, RMIA is based on shadow model training and computing the attack statistics based on a likelihood ratio.
+The main difference to LiRA is that RMIA does not compute the likelihood ratio based on aggregated IN/OUT statistics, but instead compares the target data point against random samples $(\bm{z},y_{\bm{z}})$ from the target data distribution.
+After computing the likelihood ratios over multiple $(\bm{z},y_{\bm{z}})$ values, the MIA score is estimated as a proportion of the ratios exceeding a preset bound.
+This approach makes RMIA a more effective attack when the number of shadow models is low.
+B.1 Formulating RMIA
+In the following, let us denote a target point by $(\bm{x},y_{\bm{x}})$ with $y_{\bm{x}}$ being the label of $\bm{x}$.
+Let us restate how RMIA (Zarifzadeh et al., 2024) builds the MIA score.
+RMIA augments the likelihood-ratio with a sample from the target data distribution to calibrate how likely you would obtain the target model if $(\bm{x},y_{\bm{x}})$ is replaced with another sample $(\bm{z},y_{\bm{z}})$.
+Denoting the target model parameters with $\theta$, RMIA computes the pairwise likelihood ratio
+$\mathrm{LR}(\bm{x},\bm{z})=\frac{p(\theta\mid\bm{x},y_{\bm{x}})}{p(\theta\mid\bm{z},y_{\bm{z}})},$ (A60)
+and the corresponding MIA score is given as
+$\text{Score}_{\text{RMIA}}(\bm{x})=\Pr_{(\bm{z},y_{\bm{z}})\sim\mathbb{D}}(\mathrm{LR}(\bm{x},\bm{z})>\gamma),$ (A61)
+where $\mathbb{D}$ denotes the training data distribution.
+Zarifzadeh et al.
+(2024) show two approaches to compute $\mathrm{LR}(\bm{x},\bm{z})$: the direct approach and the effecient Bayesian approach.
+In the following theoretical analysis, we focus on the direct approach that is an approximation of the efficient Bayesian approach, as Zarifzadeh et al.
+(2024) empirically demonstrates that these approaches exhibit similar performances.
+Let $\hat{\mu}_{\bm{a},\bm{b}}$ and $\hat{\sigma}_{\bm{a},\bm{b}}$ denote, respectively, the mean and standard deviation of $t_{b}=\ell(\mathcal{M}(b),y_{b})$ estimated from shadow models, where $a$ denotes which of $(\bm{x},y_{\bm{x}})$ and $(\bm{z},y_{\bm{z}})$ is in the training set.
+By Equation 11 in (Zarifzadeh et al., 2024), the pairwise likelihood ratio is given as
+$\mathrm{LR}(\bm{x},\bm{z})=\frac{p(\theta\mid\bm{x},y_{\bm{x}})}{p(\theta\mid\bm{z},y_{\bm{z}})}\approx\frac{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\bm{x},\bm{x}},\hat{\sigma}_{\bm{x},\bm{x}}^{2})\mathcal{N}(t_{\bm{z}};\hat{\mu}_{\bm{x},\bm{z}},\hat{\sigma}_{\bm{x},\bm{z}}^{2})}{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\bm{z},\bm{x}},\hat{\sigma}_{\bm{z},\bm{x}}^{2})\mathcal{N}(t_{\bm{z}};\hat{\mu}_{\bm{z},\bm{z}},\hat{\sigma}_{\bm{z},\bm{z}}^{2})},$ (A62)
+where $\hat{\mu}_{\bm{a},\bm{b}}$ and $\hat{\sigma}_{\bm{a},\bm{b}}$ are, respectively, the mean and standard deviation of $t_{b}$ estimated from shadow models when the training set contains $\bm{a}$ but not $\bm{b}$.
+Then RMIA exploits the probability of rejecting the pairwise likelihood ratio test over $(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}$:
+$\mathrm{Score}_{\mathrm{RMIA}}(\bm{x})=\Pr_{(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,\bm{z})\geq\gamma\right),$ (A63)
+where ${\mathbb{D}}$ is the underlying data distribution.
+Similar to LiRA, the classifier is built by checking if $\text{Score}_{\text{RMIA}}(\bm{x})>\tau$.
+In the following, we will use the direct computation of likelihood-ratio as described in Equation 11 of Zarifzadeh et al.
+(2024) which approximates $\mathrm{LR}(\bm{x},\bm{z})$ using normal distributions.
+Thus, RMIA rejects $H_{0}$ if and only if
+$\Pr_{(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}}\left(\mathrm{LR}(\bm{x},\bm{z})\geq\gamma\right)\geq\tau,$ (A64)
+identifying the membership of $\bm{x}$. Hence the true positive rate of per-example RMIA is given as
+$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})\\ =\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\Pr_{(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}}\left(\mathrm{LR}(\bm{x},\bm{z})\geq\gamma\right)\geq\tau\mid(\bm{x},y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}\wedge(\bm{x},y_{\bm{x}})\notin\mathcal{D}_{\mathrm{target}}\right),$ (A65)
+where $\phi^{M}$ denotes the randomness in the shadow set sampling and shadow model training.
+We define the average-case tpr for RMIA by taking the expectation over the data distribution:
+$\overline{\textsc{tpr}}_{\mathrm{RMIA}}=\mathbb{E}_{(\bm{x},y_{\bm{x}})\sim{\mathbb{D}}}[\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})].$ (A66)
+B.2 Per-example RMIA
+Next we focus on the per-example RMIA performance. As in the case of LiRA, we assume that $t_{\bm{x}}$ and $t_{\bm{z}}$ follow distributions of the location-scale family. We have
+$t_{\bm{x}}$ =$=\begin{cases}\mu_{\bm{x},\bm{x}}+\sigma_{\bm{x},\bm{x}}Z&\text{ if }(x,y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{\bm{z},\bm{x}}+\sigma_{\bm{z},\bm{x}}Z&\text{ if }(x,y_{\bm{x}})\notin\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\in\mathcal{D}_{\mathrm{target}}\end{cases}$ (A67)
+$t_{\bm{z}}$ =$=\begin{cases}\mu_{\bm{x},\bm{z}}+\sigma_{\bm{x},\bm{z}}Z&\text{ if }(x,y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{\bm{z},\bm{z}}+\sigma_{\bm{z},\bm{z}}Z&\text{ if }(x,y_{\bm{x}})\notin\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\in\mathcal{D}_{\mathrm{target}}.\end{cases}$ (A68)
+where $Z$ follows a distribution of location-scale family with the standard location and unit scale.
+It is important to note that $\mu_{\bm{a},\bm{b}}$ and $\sigma_{\bm{a},\bm{b}}$ denote, respectively, a location and a scale, while previously defined $\hat{\mu}_{\bm{a},\bm{b}}$ and $\hat{\sigma}_{\bm{a},\bm{b}}$ are, respectively, a mean and a standard deviation.
+As for the analysis of LiRA, we assume that the target and shadow sets have a sufficient number of examples per class, and that $\sigma_{\bm{x}}=\sigma_{\bm{x},\bm{x}}=\sigma_{\bm{z},\bm{x}}$, $\sigma_{\bm{z}}=\sigma_{\bm{x},\bm{z}}=\sigma_{\bm{z},\bm{z}}$, $\hat{\sigma}_{\bm{x}}=\hat{\sigma}_{\bm{x},\bm{x}}=\hat{\sigma}_{\bm{z},\bm{x}}$ and $\hat{\sigma}_{\bm{z}}=\hat{\sigma}_{\bm{x},\bm{z}}=\hat{\sigma}_{z,z}$, where $\sigma_{\bm{x}}$ and $\sigma_{\bm{z}}$ are, respectively, the true scales of $t_{\bm{x}}$ and $t_{\bm{z}}$, and $\hat{\sigma}_{\bm{x}}$ and $\hat{\sigma}_{\bm{z}}$ are, respectively, standard deviations of $t_{\bm{x}}$ and $t_{\bm{z}}$ estimated from shadow models.
+Similarly to the case of LiRA, these assumptions are justified using the simplified model.
+We have in the simplified model
+$\sigma_{\bm{x},\bm{x}}^{2}$ $=\hat{\sigma}_{\bm{x},\bm{x}}^{2}=\operatorname{Var}(s_{y_{\bm{x}}}^{(\bm{x})}(\bm{x}))=\frac{1}{S}\left(1-\frac{1}{S}\right)\bm{x}^{T}\Sigma\bm{x}$ (A69)
+$\sigma_{\bm{z},\bm{x}}^{2}$ =$\hat{\sigma}_{\bm{z},\bm{x}}^{2}=\operatorname{Var}(s_{y_{\bm{x}}}^{(\bm{z})}(\bm{x}))=\begin{cases}\frac{1}{S}\left(1-\frac{1}{S}\right)\bm{x}^{T}\Sigma\bm{x}\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \frac{1}{S}\bm{x}^{T}\Sigma\bm{x}\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A70)
+$\sigma_{\bm{x},\bm{z}}^{2}$ =$\hat{\sigma}_{\bm{x},\bm{z}}^{2}=\operatorname{Var}(s_{y_{\bm{z}}}^{(\bm{x})}(\bm{z}))=\begin{cases}\frac{1}{S}\left(1-\frac{1}{S}\right)\bm{z}^{T}\Sigma\bm{z}\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \frac{1}{S}\bm{z}^{T}\Sigma\bm{z}\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A71)
+$\sigma_{\bm{z},\bm{z}}^{2}$ $=\hat{\sigma}_{\bm{z},\bm{z}}^{2}=\operatorname{Var}(s_{y_{\bm{z}}}^{(\bm{z})}(\bm{z}))=\frac{1}{S}\left(1-\frac{1}{S}\right)\bm{z}^{T}\Sigma\bm{z}.$ (A72)
+Therefore, the differences $\sigma_{\bm{x},\bm{x}}-\sigma_{\bm{z},\bm{x}}$, $\sigma_{\bm{x},\bm{z}}-\sigma_{\bm{z},\bm{z}}$, $\hat{\sigma}_{\bm{x},\bm{x}}-\hat{\sigma}_{\bm{z},\bm{x}}$ and $\hat{\sigma}_{\bm{x},\bm{z}}-\hat{\sigma}_{\bm{z},\bm{z}}$ are negligible for large enough $S$.
+Now we derive the per-example RMIA vulnerability in terms of RMIA statistics computed from shadow models.
+Lemma 5 (Per-example RMIA vulnerability).
+Suppose that the true IN/OUT distributions of $t_{\bm{x}}$ and $t_{\bm{z}}$ are of location-scale family with locations $\mu_{\bm{x},\bm{x}},\mu_{\bm{z},\bm{x}},\mu_{\bm{x},\bm{z}},\mu_{\bm{z},\bm{z}}$ and scales $\sigma_{\bm{x}},\sigma_{\bm{z}}$.
+Assume that RMIA models $t_{\bm{x}}$ and $t_{\bm{z}}$ by normal distributions with parameters computed from shadow models, and that an attacker has access to the underlying distribution ${\mathbb{D}}$.
+Then with infinitely many shadow models, the RMIA vulnerability of a fixed target example $(\bm{x},y_{\bm{x}})$ satisfies
+$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})\leq 1-F_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}}[|q|]}{\mathbb{E}_{(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}}[|A|]}\right)$ (A73)
+for some constant $\alpha\geq\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$, where $F_{|Z|}$ is the cdf of $|Z|$ and
+$q$ $=\frac{(\mu_{\bm{x},\bm{x}}-\mu_{\bm{z},\bm{x}})(\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})}{\hat{\sigma}_{\bm{x}}^{2}}+\frac{(\mu_{\bm{x},\bm{z}}-\mu_{\bm{z},\bm{z}})(\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}})}{\hat{\sigma}_{\bm{z}}^{2}}$ (A74)
+$A$ $=\frac{\sigma_{\bm{x}}}{\hat{\sigma}_{\bm{x}}^{2}}(\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})+\frac{\sigma_{\bm{z}}}{\hat{\sigma}_{\bm{z}}^{2}}(\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}),$ (A75)
+assuming that the inverse of $F_{|Z|}$ exists.
 Proof.
-To apply Lemma 2, we will calculate $\mathbb{E}_{z}[q]$ and $\mathbb{E}_{z}[A]$.
-Let $s_{y_{x}}^{(x)}(x)$ (resp.
-$s_{y_{x}}^{(z)}(x)$) denote the score of the target $x$ for class $y_{x}$ when the dataset contains $(x,y_{x})$ but not $(z,y_{z})$ (resp.
-when the dataset contains $(z,y_{z})$ but not $(x,y_{x})$).
-Let $s_{y_{z}}^{(x)}(z)$ and $z_{y_{z}}^{(z)}(z)$ be corresponding scores of example $z$.
-Then we have
-$s_{y_{x}}^{(x)}(x)$ $=\frac{1}{S}\langle x,\sum_{i=1}^{S}x_{i}+x-x_{S}\rangle$ (A86)
-$s_{y_{x}}^{(z)}(x)$ =$\begin{cases}\frac{1}{S}\langle x,\sum_{i=1}^{S}x_{i}+z-x_{S}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}\langle x,\sum_{i=1}^{S}x_{i}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A87)
-$s_{y_{z}}^{(x)}(z)$ =$\begin{cases}\frac{1}{S}\langle z,\sum_{i=1}^{S}z_{i}+x-z_{S}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \frac{1}{S}\langle z,\sum_{i=1}^{S}z_{i}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A88)
-$s_{y_{z}}^{(z)}(z)$ $=\frac{1}{S}\langle z,\sum_{i=1}^{S}z_{i}+z-x_{S}\rangle$ (A89)
-where $x_{i}$ are samples with label $y_{x}$ and $z_{i}$ are samples with label $y_{z}$ when $y_{x}\not=y_{z}$. Thus we obtain
-$\mu_{x,x}$ $=\langle x,m_{x}\rangle+\frac{1}{S}\langle x,x-m_{x}\rangle$ (A90)
-$\mu_{z,x}$ =$\begin{cases}\langle x,m_{x}\rangle+\frac{1}{S}\langle x,z-m_{x}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \langle x,m_{x}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A91)
-$\mu_{x,z}$ =$\begin{cases}\langle z,m_{x}\rangle+\frac{1}{S}\langle z,x-m_{x}\rangle\quad&\text{ if }y_{x}=y_{z}\\ \langle z,m_{z}\rangle\quad&\text{ if }y_{x}\not=y_{z}\end{cases}$ (A92)
-$\mu_{z,z}$ $=\langle z,m_{x}\rangle+\frac{1}{S}\langle z,z-m_{x}\rangle\quad$ (A93)
-$\sigma_{x}$ $=\frac{1}{\sqrt{S}}s||x||$ (A94)
-$\sigma_{z}$ $=\frac{1}{\sqrt{S}}s||z||,$ (A95)
-where $m_{z}$ is the true class mean of $y_{z}$ (see Appendix B for derivation of $\sigma_{x}$ and $\sigma_{z}$).
+We abuse notations by denoting probabilities and expectations over sampling $\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{n}$ and $(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}$ by subscripts $\mathcal{D}_{\mathrm{target}}$ and $\bm{z}$.
+We have in the limit of infinitely many shadow models
+$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{\mathcal{D}_{\mathrm{target}}}\left(\Pr_{\bm{z}}\left(\mathrm{LR}(\bm{x},\bm{z})\geq\gamma\right)\geq\tau\mid(\bm{x},y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}\right)$ (A76)
+$\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{\mathcal{D}_{\mathrm{target}}}\left(\Pr_{\bm{z}}\left(\mathrm{LR}(\bm{x},\bm{z})\geq\gamma\right)\geq\tau\mid(\bm{x},y_{\bm{x}})\notin\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\in\mathcal{D}_{\mathrm{target}}\right),$ (A77)
+where
+$\mathrm{LR}(\bm{x},\bm{z})=\frac{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\bm{x},\bm{x}},\hat{\sigma}_{\bm{x}}^{2})\mathcal{N}(t_{\bm{z}};\hat{\mu}_{\bm{x},\bm{z}},\hat{\sigma}_{\bm{z}}^{2})}{\mathcal{N}(t_{\bm{x}};\hat{\mu}_{\bm{z},\bm{x}},\hat{\sigma}_{\bm{x}}^{2})\mathcal{N}(t_{\bm{z}};\hat{\mu}_{\bm{z},\bm{z}},\hat{\sigma}_{\bm{z}}^{2})}.$ (A78)
+Note that the probabilities are over target dataset sampling in the limit of infinitely many shadow models. We have
+$\lambda:=$ $\log(\mathrm{LR}(\bm{x},\bm{z}))$ (A79)
+$=$ $\log\left(\frac{\frac{1}{\sqrt{2\pi}\hat{\sigma}_{\bm{x}}}\exp\left(-\frac{1}{2}\left(\frac{t_{\bm{x}}-\hat{\mu}_{\bm{x},\bm{x}}}{\hat{\sigma}_{\bm{x}}}\right)^{2}\right)\frac{1}{\sqrt{2\pi}\hat{\sigma}_{\bm{z}}}\exp\left(-\frac{1}{2}\left(\frac{t_{\bm{z}}-\hat{\mu}_{\bm{x},\bm{z}}}{\hat{\sigma}_{\bm{z}}}\right)^{2}\right)}{\frac{1}{\sqrt{2\pi}\hat{\sigma}_{\bm{x}}}\exp\left(-\frac{1}{2}\left(\frac{t_{\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}}}{\hat{\sigma}_{\bm{x}}}\right)^{2}\right)\frac{1}{\sqrt{2\pi}\hat{\sigma}_{\bm{z}}}\exp\left(-\frac{1}{2}\left(\frac{t_{\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}}{\hat{\sigma}_{\bm{z}}}\right)^{2}\right)}\right)$ (A80)
+$=$ $-\frac{1}{2}\left(\frac{t_{\bm{x}}-\hat{\mu}_{\bm{x},\bm{x}}}{\hat{\sigma}_{\bm{x}}}\right)^{2}+\frac{1}{2}\left(\frac{t_{\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}}}{\hat{\sigma}_{\bm{x}}}\right)^{2}-\frac{1}{2}\left(\frac{t_{\bm{z}}-\hat{\mu}_{\bm{x},\bm{z}}}{\hat{\sigma}_{\bm{z}}}\right)^{2}+\frac{1}{2}\left(\frac{t_{\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}}{\hat{\sigma}_{\bm{z}}}\right)^{2}$ (A81)
+$=$ $\frac{1}{2\hat{\sigma}_{\bm{x}}^{2}}(2t_{\bm{x}}\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{x},\bm{x}}^{2}-2t_{\bm{x}}\hat{\mu}_{\bm{z},\bm{x}}+\hat{\mu}_{\bm{z},\bm{x}}^{2})+\frac{1}{2\hat{\sigma}_{\bm{z}}^{2}}(2t_{\bm{z}}\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{x},\bm{z}}^{2}-2t_{\bm{z}}\hat{\mu}_{\bm{z},\bm{z}}+\hat{\mu}_{\bm{z},\bm{z}}^{2})$ (A82)
+$=$ $\frac{\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}}}{2\hat{\sigma}_{\bm{x}}^{2}}(2t_{\bm{x}}-\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})+\frac{\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}}{2\hat{\sigma}_{\bm{z}}^{2}}(2t_{\bm{z}}-\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}).$ (A83)
+When $(\bm{x},y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}$ and $(z,y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}$, $t_{\bm{x}}=\mu_{\bm{x},\bm{x}}+\sigma_{\bm{x}}Z$ and $t_{\bm{z}}=\mu_{\bm{x},\bm{z}}+\sigma_{\bm{z}}Z$.
+Thus, $\lambda$ becomes
+$\lambda_{\bm{x}}:=$ $\frac{\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}}}{2\hat{\sigma}_{\bm{x}}^{2}}(2\mu_{\bm{x},\bm{x}}+2\sigma_{\bm{x}}Z-\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})$ (A84)
+$+\frac{\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}}{2\hat{\sigma}_{\bm{z}}^{2}}(2\mu_{\bm{x},\bm{z}}+2\sigma_{\bm{z}}Z-\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}).$ (A85)
+Similarly, when $(x,y_{\bm{x}})\notin\mathcal{D}_{\mathrm{target}}$ and $(z,y_{\bm{z}})\in\mathcal{D}_{\mathrm{target}}$, $t_{\bm{x}}=\mu_{\bm{z},\bm{x}}+\sigma_{\bm{x}}Z$ and $t_{\bm{z}}=\mu_{\bm{z},\bm{x}}+\sigma_{\bm{z}}Z$.
+Then $\lambda$ becomes
+$\lambda_{\bm{z}}:=$ $\frac{\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}}}{2\hat{\sigma}_{\bm{x}}^{2}}(2\mu_{\bm{z},\bm{x}}+2\sigma_{\bm{x}}Z-\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})$ (A86)
+$+\frac{\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}}{2\hat{\sigma}_{\bm{z}}^{2}}(2\mu_{\bm{z},\bm{z}}+2\sigma_{\bm{z}}Z-\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}})$ (A87)
+$=$ $\left(\underbrace{\frac{\sigma_{\bm{x}}}{\hat{\sigma}_{\bm{x}}^{2}}(\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})+\frac{\sigma_{\bm{z}}}{\hat{\sigma}_{\bm{z}}^{2}}(\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}})}_{A}\right)Z$ (A88)
+$+\underbrace{\frac{\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}}}{2\hat{\sigma}_{\bm{x}}^{2}}(2\mu_{\bm{z},\bm{x}}-\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})+\frac{\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}}{2\hat{\sigma}_{\bm{z}}^{2}}(2\mu_{\bm{z},\bm{z}}-\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}})}_{B}$ (A89)
+$=$ $AZ+B.$ (A90)
+Notice that $A$ and $B$ are functions of $\bm{z}$ and independent of $Z$.
+Also note that taking probability over $Z$ corresponds to calculating probability over $\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{n}$.
+Thus, since we can set $\gamma>1$, using Markov’s inequality and the triangle inequality, we have in the limit of infinitely many shadow models
+$\textsc{fpr}_{\mathrm{RMIA}}(x)=$ $\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{z}}\geq\log\gamma)\geq\tau\right)$ (A91)
+$\leq\Pr_{Z}\left(\Pr_{\bm{z}}(|\lambda_{\bm{z}}|\geq\log\gamma)\geq\tau\right)$ (A92)
+$\leq\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}\geq\tau\right)$ (A93)
+$=\Pr_{Z}\left(\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]\geq\tau\log\gamma\right)$ (A94)
+$\leq\Pr_{Z}\left(\mathbb{E}_{\bm{z}}[|A|]|Z|+\mathbb{E}_{\bm{z}}[|B|]\geq\tau\log\gamma\right)$ (A95)
+$=\Pr_{Z}\left(|Z|\geq\frac{\tau\log\gamma-\mathbb{E}_{\bm{z}}[|B|]}{\mathbb{E}_{\bm{z}}[|A|]}\right)$ (A96)
+Therefore, we can upper-bound $\textsc{fpr}_{\mathrm{RMIA}}(x)\leq\alpha$ by setting
+$\alpha=1-F_{|Z|}\left(\frac{\tau\log\gamma-\mathbb{E}_{\bm{z}}[|B|]}{\mathbb{E}_{\bm{z}}[|A|]}\right).$ (A97)
+That is,
+$\frac{\tau\log\gamma-\mathbb{E}_{\bm{z}}[|B|]}{\mathbb{E}_{\bm{z}}[|A|]}=F_{|Z|}^{-1}(1-\alpha).$ (A98)
+Now let
+$q=\lambda_{\bm{x}}-\lambda_{\bm{z}}=\frac{(\mu_{\bm{x},\bm{x}}-\mu_{\bm{z},\bm{x}})(\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})}{\hat{\sigma}_{\bm{x}}^{2}}+\frac{(\mu_{\bm{x},\bm{z}}-\mu_{\bm{z},\bm{z}})(\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}})}{\hat{\sigma}_{\bm{z}}^{2}}.$ (A99)
+Note that $q$ is also independent of $t$, thereby $\mathbb{E}_{\bm{z}}[|q|]$ being a constant. By the similar argument using Markov’s inequality and the triangle inequality, we have
+$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{x}}\geq\log\gamma)\geq\tau\right)$ (A100)
+$\leq\Pr_{Z}\left(\Pr_{\bm{z}}(|\lambda_{\bm{x}}|\geq\log\gamma)\geq\tau\right)$ (A101)
+$\leq\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{x}}|]}{\log\gamma}\geq\tau\right)$ (A102)
+$=\Pr_{Z}\left(\mathbb{E}_{\bm{z}}[|\lambda_{\bm{x}}|]\geq\tau\log\gamma\right)$ (A103)
+$\leq\Pr_{Z}\left(\mathbb{E}_{\bm{z}}[|A|]|Z|+\mathbb{E}_{\bm{z}}[|B|]+\mathbb{E}_{\bm{z}}[|q|]\geq\tau\log\gamma\right)$ (A104)
+$=\Pr_{Z}\left(|Z|\geq\frac{\tau\log\gamma-\mathbb{E}_{\bm{z}}[|B|]}{\mathbb{E}_{\bm{z}}[|A|]}-\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\right).$ (A105)
+Hence we obtain
+$\textsc{tpr}_{\mathrm{RMIA}}(x)\leq 1-F_{|Z|}^{-1}\left(F_{|Z|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\right).$ (A106)
+Note that, unlike per-example LiRA, we must assume that the attacker has access to the underlying distribution for the optimal RMIA as the Equations A74 and A75 depend on the parameters computed from shadow models.
+B.3 RMIA power-law
+Employing Lemma 5 and the simplified model, we obtain the following upper bound for RMIA performance.
+Theorem 6 (Per-example RMIA power-law).
+Fix a target example $(\bm{x},y_{\bm{x}})$. For the simplified model with large $S$ and infinitely many shadow models, the per-example RMIA vulnerability is given as
+$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})\leq 1-F_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)-\frac{\Psi(\bm{x},C)}{\sqrt{S}}\right),$ (A107)
+where $\alpha\geq\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$, $F_{|Z|}$ is the cdf of the standard folded normal distribution, and
+$\Psi(\bm{x},C)=\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{||\bm{x}||^{2}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{||\bm{z}||^{2}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{||\bm{x}||^{2}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{||\bm{z}||^{2}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{||\bm{x}||}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{||\bm{z}||}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{||\bm{x}||}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{||\bm{z}||}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A108)
+In addition, we have
+$\log(\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}+\log\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}.$ (A109)
+Proof.
+To apply Lemma 5, we will calculate $\mathbb{E}_{\bm{z}}[|q|]$ and $\mathbb{E}_{\bm{z}}[|A|]$.
+Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),\dots,(\bm{x}_{j,S},j)\}_{j=1}^{C}$.
+Let $s_{y_{a}}^{(\bm{b})}(\bm{a})$ denote the score of $(\bm{a},y_{\bm{a}})$ when $\mathcal{D}_{\mathrm{target}}$ contains $(\bm{b},y_{\bm{b}})$ but not the other example.
+Using similar argument as in the proof of Theorem 2, we have
+$s_{y_{\bm{x}}}^{(\bm{x})}(\bm{x})$ $=\frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}+\bm{x}-\bm{x}_{y_{\bm{x}},S}\rangle$ (A110)
+$s_{y_{\bm{x}}}^{(\bm{z})}(\bm{x})$ =$\begin{cases}\frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}+\bm{z}-\bm{x}_{y_{\bm{x}},S}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A111)
+$s_{y_{\bm{z}}}^{(\bm{x})}(\bm{z})$ =$\begin{cases}\frac{1}{S}\langle\bm{z},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}+\bm{x}-\bm{x}_{y_{\bm{x}},S}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \frac{1}{S}\langle\bm{z},\sum_{i=1}^{S}\bm{x}_{y_{\bm{z}},i}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A112)
+$s_{y_{\bm{z}}}^{(\bm{z})}(\bm{z})$ $=\frac{1}{S}\langle\bm{z},\sum_{i=1}^{S}\bm{x}_{y_{\bm{z}},i}+\bm{z}-\bm{x}_{y_{\bm{z}},S}\rangle.$ (A113)
+Thus we obtain
+$\mu_{\bm{x},\bm{x}}$ $=\langle\bm{x},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle$ (A114)
+$\mu_{\bm{z},\bm{x}}$ =$\begin{cases}\langle\bm{x},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{z}-m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{x},m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A115)
+$\mu_{\bm{x},\bm{z}}$ =$\begin{cases}\langle\bm{z},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{x}-m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{z},m_{y_{\bm{z}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A116)
+$\mu_{\bm{z},\bm{z}}$ $=\langle\bm{z},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{z}-m_{y_{\bm{x}}}\rangle\quad$ (A117)
+$\sigma_{\bm{x}}$ $=\frac{1}{\sqrt{S}}\sqrt{\bm{x}^{T}\Sigma\bm{x}}$ (A118)
+$\sigma_{\bm{z}}$ $=\frac{1}{\sqrt{S}}\sqrt{\bm{z}^{T}\Sigma\bm{z}},$ (A119)
+where $m_{y_{\bm{z}}}$ is the true class mean of class $y_{\bm{z}}$.
 Now recall that
-$q=$ $\frac{(\mu_{x,x}-\mu_{z,x})(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})}{\hat{\sigma}_{x}^{2}}+\frac{(\mu_{x,z}-\mu_{z,z})(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})}{\hat{\sigma}_{z}^{2}}$ (A96)
-$A=$ $\frac{\sigma_{x}}{\hat{\sigma}_{x}^{2}}(\hat{\mu}_{x,x}-\hat{\mu}_{z,x})+\frac{\sigma_{z}}{\hat{\sigma}_{z}^{2}}(\hat{\mu}_{x,z}-\hat{\mu}_{z,z})$ (A97)
-In the limit of infinitely many shadow models, estimated statisitcs converge to true parameters, thereby yielding
-$q=$ $\left(\frac{\mu_{x,x}-\mu_{z,x}}{\sigma_{x}}\right)^{2}+\left(\frac{\mu_{x,z}-\mu_{z,z}}{\sigma_{z}}\right)^{2}$ (A98)
-$A=$ $\frac{\mu_{x,x}-\mu_{z,x}}{\sigma_{x}}+\frac{\mu_{x,z}-\mu_{z,z}}{\sigma_{z}}.$ (A99)
+$q=$ $\frac{(\mu_{\bm{x},\bm{x}}-\mu_{\bm{z},\bm{x}})(\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})}{\hat{\sigma}_{\bm{x}}^{2}}+\frac{(\mu_{\bm{x},\bm{z}}-\mu_{\bm{z},\bm{z}})(\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}})}{\hat{\sigma}_{\bm{z}}^{2}}$ (A120)
+$A=$ $\frac{\sigma_{\bm{x}}}{\hat{\sigma}_{\bm{x}}^{2}}(\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})+\frac{\sigma_{\bm{z}}}{\hat{\sigma}_{\bm{z}}^{2}}(\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}).$ (A121)
+Since $t_{\bm{x}}$ and $t_{\bm{z}}$ follow normal distributions, the location and scale parameters of the true distributions correspond to the mean and standard deviations estimated from infinitely many shadow models, respectively.
+Thus, we have
+$q=$ $\left(\frac{\mu_{\bm{x},\bm{x}}-\mu_{\bm{z},\bm{x}}}{\sigma_{\bm{x}}}\right)^{2}+\left(\frac{\mu_{\bm{x},\bm{z}}-\mu_{\bm{z},\bm{z}}}{\sigma_{\bm{z}}}\right)^{2}$ (A122)
+$A=$ $\frac{\mu_{\bm{x},\bm{x}}-\mu_{\bm{z},\bm{x}}}{\sigma_{\bm{x}}}+\frac{\mu_{\bm{x},\bm{z}}-\mu_{\bm{z},\bm{z}}}{\sigma_{\bm{z}}}.$ (A123)
 Using the law of total expectation, we have
-$\mathbb{E}_{z}[|q|]=$ $\Pr_{z}(y_{z}=y_{x})\mathbb{E}_{z}[|q|\mid y_{z}=y_{x}]+\sum_{j=1,j\not=y_{x}}^{C}\Pr_{z}(y_{z}=j)\mathbb{E}_{z}[|q|\mid y_{z}=j]$ (A100)
-$=$ $\frac{1}{C}\mathbb{E}_{z}\left[\left.\left(\frac{\langle x,x-z\rangle}{\sqrt{S}s||x||}\right)^{2}+\left(\frac{\langle z,x-z\rangle}{\sqrt{S}s||z||}\right)^{2}\ \right|\ y_{z}=y_{x}\right]$ (A101)
-$+\frac{C-1}{C}\mathbb{E}_{z}\left[\left.\left(\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}\right)^{2}+\left(\frac{\langle z,z-m_{z}\rangle}{\sqrt{S}s||z||}\right)^{2}\ \right|\ y_{z}\not=y_{x}\right]$ (A102)
-$=$ $\frac{1}{CSs^{2}}\mathbb{E}_{z}\left[\left.\frac{\langle x,x-z\rangle^{2}}{||x||^{2}}+\frac{\langle z,x-z\rangle^{2}}{||z||^{2}}\ \right|\ y_{z}=y_{x}\right]$ (A103)
-$+\frac{C-1}{CSs^{2}}\mathbb{E}_{z}\left[\left.\frac{\langle x,x-m_{x}\rangle^{2}}{||x||^{2}}+\frac{\langle z,z-m_{z}\rangle^{2}}{||z||^{2}}\ \right|\ y_{z}\not=y_{x}\right],$ (A104)
+$\mathbb{E}_{\bm{z}}[|q|]=$ $\Pr_{\bm{z}}(y_{\bm{z}}=y_{\bm{x}})\mathbb{E}_{\bm{z}}[|q|\mid y_{\bm{z}}=y_{\bm{x}}]+\sum_{j=1,j\not=y_{\bm{x}}}^{C}\Pr_{\bm{z}}(y_{\bm{z}}=j)\mathbb{E}_{\bm{z}}[|q|\mid y_{\bm{z}}=j]$ (A124)
+$=$ $\frac{1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-z\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A125)
+$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{z}-m_{y_{\bm{z}}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A126)
+$=$ $\frac{1}{CS}\mathbb{E}_{\bm{z}}\left[\left.\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A127)
+$+\frac{C-1}{CS}\mathbb{E}_{\bm{z}}\left[\left.\frac{\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{y_{\bm{z}}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right],$ (A128)
 and
-$\mathbb{E}_{z}[|A|]=$ $\Pr_{z}(y_{z}=y_{x})\mathbb{E}_{z}[|A|\mid y_{z}=y_{x}]+\sum_{j=1,j\not=y_{x}}^{C}\Pr_{z}(y_{z}=j)\mathbb{E}_{z}[|A|\mid y_{z}=j]$ (A105)
-$=$ $\frac{1}{C}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-z\rangle}{\sqrt{S}s||x||}+\frac{\langle z,x-z\rangle}{\sqrt{S}s||z||}\right|\ \right|\ y_{z}=y_{x}\right]$ (A106)
-$+\frac{C-1}{C}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-m_{x}\rangle}{\sqrt{S}s||x||}+\frac{\langle z,z-m_{z}\rangle}{\sqrt{S}s||z||}\right|\ \right|\ y_{z}\not=y_{x}\right]$ (A107)
-$=$ $\frac{1}{C\sqrt{S}s}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-z\rangle}{||x||}+\frac{\langle z,x-z\rangle}{||z||}\right|\ \right|\ y_{z}=y_{x}\right]$ (A108)
-$+\frac{C-1}{C\sqrt{S}s}\mathbb{E}_{z}\left[\left.\left|\frac{\langle x,x-m_{x}\rangle}{||x||}+\frac{\langle z,z-m_{z}\rangle}{||z||}\right|\ \right|\ y_{z}\not=y_{x}\right].$ (A109)
-Hence we obtain by the Cauchy-Schwarz inequality
-$\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}=$ $\frac{1}{\sqrt{S}s}\cdot\frac{\mathbb{E}_{z}\left[\frac{\langle x,x-z\rangle^{2}}{||x||^{2}}+\frac{\langle z,x-z\rangle^{2}}{||z||^{2}}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[\frac{\langle x,x-m_{x}\rangle^{2}}{||x||^{2}}+\frac{\langle z,z-m_{z}\rangle^{2}}{||z||^{2}}\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{z}\left[\left|\frac{\langle x,x-z\rangle}{||x||}+\frac{\langle z,x-z\rangle}{||z||}\right|\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[\left|\frac{\langle x,x-m_{x}\rangle}{||x||}+\frac{\langle z,z-m_{z}\rangle}{||z||}\right|\mid y_{z}\not=y_{x}\right]}$ (A110)
-$\leq$ $\frac{1}{\sqrt{S}s}\cdot\frac{\mathbb{E}_{z}\left[2||x-z||^{2}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||^{2}+||z-m_{z}||^{2})\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{z}\left[2||x-z||\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||+||z-m_{z}||)\mid y_{z}\not=y_{x}\right]}.$ (A111)
+$\mathbb{E}_{\bm{z}}[|A|]=$ $\Pr_{\bm{z}}(y_{\bm{z}}=y_{\bm{x}})\mathbb{E}_{\bm{z}}[|A|\mid y_{\bm{z}}=y_{\bm{x}}]+\sum_{j=1,j\not=y_{\bm{x}}}^{C}\Pr_{\bm{z}}(y_{\bm{z}}=j)\mathbb{E}_{\bm{z}}[|A|\mid y_{\bm{z}}=j]$ (A129)
+$=$ $\frac{1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A130)
+$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A131)
+$=$ $\frac{1}{C\sqrt{S}}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A132)
+$+\frac{C-1}{C\sqrt{S}}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right].$ (A133)
+Hence we obtain
+$\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\\ =\frac{1}{\sqrt{S}}\cdot\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A134)
 Now Lemma 5 yields
-$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $\leq 1-F_{|t|}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right)$ (A112)
-$\leq 1-F_{|t|}\left(F_{|t|}^{-1}(1-\alpha)-\frac{\psi(x,C)}{\sqrt{S}s}\right)$ (A113)
-where $F_{|t|}$ is the cdf of the folded normal distribution and
-$\psi(x,C)=\frac{\mathbb{E}_{z}\left[2||x-z||^{2}\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||^{2}+||z-m_{z}||^{2})\mid y_{z}\not=y_{x}\right]}{\mathbb{E}_{z}\left[2||x-z||\mid y_{z}=y_{x}\right]+(C-1)\mathbb{E}_{z}\left[(||x-m_{x}||+||z-m_{z}||)\mid y_{z}\not=y_{x}\right]}.$ (A114)
+$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ $\leq 1-F_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\right)$ (A135)
+$=1-F_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)-\frac{\Psi(\bm{x},C)}{\sqrt{S}}\right),$ (A136)
+where $F_{|Z|}$ is the cdf of the folded normal distribution and
+$\Psi(\bm{x},C)=\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A137)
 This completes the first half of the theorem.
 Now that from Lemma 5 we have
-$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{t}\left(\Pr_{z}(\lambda_{z}+q\geq\log\gamma)\geq\beta\right)$ (A115)
-$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{z}|]+\mathbb{E}_{z}[|q|]}{\log(\gamma)}\geq\beta\right)$ (A116)
-$=\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{z}[|q|]}{\mathbb{E}_{z}[|A|]}\right)$ (A117)
-$\textsc{fpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{t}\left(\Pr_{z}(\lambda_{z}\geq\log\gamma)\geq\beta\right)$ (A118)
-$\leq\Pr_{t}\left(\frac{\mathbb{E}_{z}[|\lambda_{z}|]}{\log(\gamma)}\geq\beta\right)$ (A119)
-$=\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)\right).$ (A120)
-We claim that the bound for $\textsc{fpr}_{\mathrm{RMIA}}(x)$ (Equation A119) is tighter than that for $\textsc{tpr}_{\mathrm{RMIA}}(x)$ (Equation A116).
-Since this is not straightforward, we numerically show this in Table A1.
-Now, for large enough $S$ we obtain
-$\textsc{tpr}_{\mathrm{RMIA}}(x)-\textsc{fpr}_{\mathrm{RMIA}}(x)$ $\leq\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)-\left|\frac{\mathbb{E}_{z}[q]}{\mathbb{E}_{z}[A]}\right|\right)-\Pr_{t}(|t|\geq F_{|t|}^{-1}(1-\alpha))$ (A121)
-$\leq\Pr_{t}\left(|t|\geq F_{|t|}^{-1}(1-\alpha)+\frac{\psi(x,C)}{\sqrt{S}s}\right)-\Pr_{t}(|t|\leq F_{|t|}^{-1}(1-\alpha))$ (A122)
-$\approx p_{|t|}(F_{|t|}^{-1}(1-\alpha))\frac{\psi(x,C)}{\sqrt{S}}.$ (A123)
-Since $|t|$ follows the standard folded normal distribution with mean 00 and scale $1$, we have
-$p_{|t|}(F_{|t|}^{-1}(1-\alpha))=\frac{1}{\sqrt{\pi/2}}\exp\left(-\frac{1}{2}F_{|t|}^{-1}(1-\alpha)^{2}\right).$ (A124)
-Hence we obtain
-$\log(\textsc{tpr}_{\mathrm{RMIA}}(x)-\textsc{fpr}_{\mathrm{RMIA}}(x))\leq-\frac{1}{2}\log S-\frac{1}{2}F_{|t|}^{-1}(1-\alpha)^{2}+\log\frac{\psi(x,C)}{\sqrt{\pi/2}}.$ (A125)
-S FPR=0.2 FPR=0.1 FPR=0.01 FPR=0.001
-10 0.000 0.000 0.000 0.000
-100 0.000 0.000 0.000 0.000
-1000 0.000 0.000 0.041 0.068
-10000 0.329 0.353 0.365 0.368
-Table A1: The gap in the upper bound for $\textsc{tpr}_{\mathrm{RMIA}}(x)$ in Equation A116 minus the gap in the upper bound for $\textsc{fpr}_{\mathrm{RMIA}}(x)$ in Equation A119.
-5 random target examples are sampled from the simplified model distribution with $C=2,s=0.1,d=50$.
-Simulation is done with 100000 samples of test points $z$, 1000 samples of $t$ and $\gamma=1.0001$.
-Due to computational constraints, results are not stable especially for small $S$.
-C.5 Proof of Corollary 5
-See 5
+$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{z}}+q\geq\log\gamma)\geq\tau\right)$ (A138)
+$\leq\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]+\mathbb{E}_{\bm{z}}[|q|]}{\log\gamma}\geq\tau\right)$ (A139)
+$=\Pr_{Z}\left(|Z|\geq F_{|Z|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\right)$ (A140)
+$\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{z}}\geq\log\gamma)\geq\tau\right)$ (A141)
+$\leq\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}\geq\tau\right)$ (A142)
+$=\Pr_{Z}\left(|Z|\geq F_{|Z|}^{-1}(1-\alpha)\right).$ (A143)
+We claim that the bound for $\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation A142) is as tight as that for $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation A139) for sufficiently large $S$.
+Let us denote
+$\kappa_{\textsc{tpr}}(\gamma)$ $=\Pr_{\bm{z}}(\lambda_{\bm{z}}+q\geq\log\gamma)$ (A144)
+$\kappa_{\textsc{fpr}}(\gamma)$ $=\Pr_{\bm{z}}(\lambda_{\bm{z}}\geq\log\gamma).$ (A145)
+Since $q=O(1/S)$, for large $S$ we approximate
+$\kappa_{\textsc{tpr}}(\gamma)-\kappa_{\textsc{fpr}}(\gamma)\approx p_{\lambda_{\bm{z}}}(\log\gamma)\frac{c_{0}}{S},$ (A146)
+for some constant $c_{0}$.
+Since $\lambda_{\bm{z}}=O(1/\sqrt{S})$, the scaled random variable $\hat{\lambda}_{\bm{z}}=\sqrt{S}\lambda_{\bm{z}}$ is almost independent of $S$.
+Then by the change of variables formula, we have
+$p_{\lambda_{\bm{z}}}(\log\gamma)\frac{c_{0}}{S}=p_{\hat{\lambda}_{\bm{z}}}(\sqrt{S}\log\gamma)\frac{c_{0}}{\sqrt{S}}.$ (A147)
+Without loss of generality we may set $\log\gamma=1/\sqrt{S}$. Thus, we have
+$\kappa_{\textsc{tpr}}(1/\sqrt{S})-\kappa_{\textsc{fpr}}(1/\sqrt{S})\approx p_{\hat{\lambda}_{\bm{z}}}\left(\sqrt{S}\cdot\frac{1}{\sqrt{S}}\right)\frac{c_{0}}{\sqrt{S}}=p_{\hat{\lambda}_{\bm{z}}}(1)\frac{c_{0}}{\sqrt{S}}.$ (A148)
+This quantity scales as $O(1/\sqrt{S})$. Therefore,
+$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{t}(\kappa_{\textsc{tpr}}(1/\sqrt{S})\geq\tau)-\Pr_{t}(\kappa_{\textsc{fpr}}(1/\sqrt{S})\geq\tau)$ (A149)
+$\approx p_{\kappa_{\textsc{fpr}}(1/\sqrt{S})}(\tau)(\kappa_{\textsc{fpr}}(1/\sqrt{S})-\kappa_{\textsc{tpr}}(1/\sqrt{S}))$ (A150)
+$\approx p_{\kappa_{\textsc{fpr}}(1/\sqrt{S})}(\tau)p_{\hat{\lambda}_{\bm{z}}}(1)\frac{c_{0}}{\sqrt{S}}.$ (A151)
+Note that $\kappa_{\textsc{fpr}}(1/\sqrt{S})$ and, consequently, $\tau$ are independent of $S$ for large enough $S$ since
+$\kappa_{\textsc{fpr}}(1/\sqrt{S})=\Pr_{\bm{z}}\left(\lambda_{\bm{z}}\geq\frac{1}{\sqrt{S}}\right)=\Pr_{\bm{z}}\left(\frac{\hat{\lambda}_{\bm{z}}}{\sqrt{S}}\geq\frac{1}{\sqrt{S}}\right)=\Pr_{\bm{z}}(\hat{\lambda}_{\bm{z}}\geq 1).$ (A152)
+Hence $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})=O(1/\sqrt{S})$. On the other hand, from Equations A139 and A142 we have for sufficiently large $S$
+$\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}\geq\tau\right)-\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}+\frac{\mathbb{E}_{\bm{z}}[|q|]}{\log\gamma}\geq\tau\right)$ (A153)
+$=$ $\Pr_{Z}\left(\sqrt{S}\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]\geq\tau\right)-\Pr_{Z}\left(\sqrt{S}\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]+\frac{c_{1}}{\sqrt{S}}\geq\tau\right)$ (A154)
+$=$ $\Pr_{Z}\left(\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]\geq\tau\right)-\Pr_{Z}\left(\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]+\frac{c_{1}}{\sqrt{S}}\geq\tau\right)$ (A155)
+$\approx$ $p_{\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]}(\tau)\frac{c_{1}}{\sqrt{S}},$ (A156)
+where $c_{1}$ is some constant. Since for large $S$, $\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]$ and $\tau$ are independent of $S$, Equation A156 scales as $O(1/\sqrt{S})$.
+Now let
+$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]+\mathbb{E}_{\bm{z}}[|q|]}{\log\gamma}\geq\tau\right)-v_{\textsc{tpr}}$ (A157)
+$\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}\geq\tau\right)-v_{\textsc{fpr}}$ (A158)
+for some $v_{\textsc{tpr}},v_{\textsc{fpr}}\geq 0$ which evaluate the tightness of the bounds. Then we have
+$v_{\textsc{tpr}}-v_{\textsc{fpr}}$ (A159)
+$=\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]+\mathbb{E}_{\bm{z}}[|q|]}{\log\gamma}\geq\tau\right)-\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}\geq\tau\right)-(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))$ (A160)
+$=O(1/\sqrt{S}).$ (A161)
+Hence we conclude that for sufficiently large $S$, the bound for $\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$ is as tight as that for $\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$.
+Therefore, noting that $\mathbb{E}_{\bm{z}}[|q|]/\mathbb{E}_{\bm{z}}[|A|]=O(1/\sqrt{S})$, for large $S$ we obtain
+$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ (A162)
+$\approx\Pr_{Z}\left(|Z|\geq F_{|Z|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\right)-\Pr_{Z}\left(|Z|\geq F_{|Z|}^{-1}(1-\alpha)\right)$ (A163)
+$\approx p_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)\right)\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}$ (A164)
+$=p_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)\right)\frac{\Psi(\bm{x},C)}{\sqrt{S}}.$ (A165)
+Since $|Z|$ follows the standard folded normal distribution,
+$p_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)\right)=\frac{2}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}\right).$ (A166)
+It follows that
+$\log(\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}+\log\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}.$ (A167)
+As for the LiRA power-law, bounding $||\bm{x}-m_{\bm{x}}||$ and $||\bm{z}-m_{\bm{z}}||$ will provide a worst-case upper bound for which the power-law holds.
+Finally, the per-example RMIA power-law is also extended to the average-case:
+Corollary 7 (Average-case RMIA power-law).
+For the simplified model with sufficiently large $S$ and infinitely many shadow models, we have
+$\log(\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}})\approx-\frac{1}{2}\log S-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}+\log\left(\mathop{\mathbb{E}}_{(\bm{x},y_{\bm{x}})\sim{\mathbb{D}}}\left[\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}\right]\right),$ (A168)
+where $\alpha\geq\overline{\textsc{fpr}}_{\mathrm{RMIA}}$ and $F_{|Z|}$ is the cdf of the standard folded normal distribution.
 Proof.
 By Theorem 6 and the law of unconscious statistician, we have for large $S$
-[…]
-[…]
-[…]
-$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\langle x,x-m_{x}\rangle}{\sqrt{2\pi}||x||s}\right].$ (A129)
-Note that here we fixed $\textsc{fpr}_{\mathrm{LiRA}}(x)=\overline{\textsc{fpr}}_{\mathrm{LiRA}}$ for all $x$. Then we obtain
-$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr})^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\langle x,x-m_{x}\rangle}{\sqrt{2\pi}||x||s}\right]\right).$ (A130)
-C.6 Proof of Corollary 6
-See 6
-Proof.
-By theorem 4 and the law of unconscious statistician, we have for large $S$
-[…]
-[…]
-[…]
-$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\alpha)^{2}}\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\psi(x,C)}{\sqrt{2\pi}}\right].$ (A134)
-Hence we obtain
-$\log(\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}})\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\alpha)^{2}+\log\left(\mathbb{E}_{(x,y_{x})\sim{\mathbb{D}}}\left[\frac{\psi(x,C)}{\sqrt{2\pi}}\right]\right).$ (A135)
-Appendix D LiRA vulnerability for Small FPR
-In Section 3.4 we proved for the simplified model that for LiRA $\log(\textsc{tpr}-\textsc{fpr})\approx-\frac{1}{2}\log S$ ignoring additive constants when fpr is fixed (Corollary 5).
-In Section 4.3 we observed that the coefficient $\beta_{S}$ for $\log S$ is around $-0.5$ for larger fpr, aligning with the theoretical value.
-However, for smaller fpr the coefficient is smaller than $-0.5$.
-To understand this phenomenon, it is important to note that the power-law (Theorem 3 and Corollary 5) only holds for sufficiently large $S$.
-Thus, it can be explained that the difference of coefficient values $\beta_{S}$ for small and large fpr comes from the small-$S$ regime as follows.
-In the proof of Corollary 5 and Theorem 3 the only approximation that could introduce some bias is Equation A84. That is,
-$\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x)$ $=\Pr\left(\eta\leq\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))+r\right)-\Pr\left(\eta\leq\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))\right)$ (A136)
-$\approx p_{\eta}\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(x))\right)r,$ (A137)
-where $\eta\sim\mathcal{N}(0,1)$, $r$ is a shift that scales $O(1/\sqrt{S})$ and $p_{\eta}$ is the pdf of $\eta$.
-Figure A.1 numerically illustrates this approximation.
-It can be observed that for small $S$ the approximation does not hold, underestimating the true vulnerability.
-Particularly, this effect is remarkably larger for small $\textsc{fpr}_{\mathrm{LiRA}}(x)$.
-In other words, the simplified model overestimates the coefficient value $\beta_{S}$ for small $\textsc{fpr}_{\mathrm{LiRA}}(x)$ by underestimating the true vulnerability in the small-$S$ regime.
-(a) $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.1$
-(b) $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.00001$
-Figure A.1: True vs.
-approximated $\textsc{tpr}_{\mathrm{LiRA}}(x)-\textsc{fpr}_{\mathrm{LiRA}}(x)$ for the simplified model with in-class standard deviation $s=0.1$ and dimension $d=100$.
-One target example $(x,y_{x})$ is randomly sampled from the underlying distribution.
-Numbers in the plots are percentages of the approximated vulnerabilities over the true vulnerabilities.
-Figures illustrate the approximation for different $S$ (a) when $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.1$ and (b) when $\textsc{fpr}_{\mathrm{LiRA}}(x)=0.00001$.
-Appendix E Training details
-E.1 Parameterization
+$\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}}$ $=\int_{\mathscr{D}}p(\bm{x})(\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x}))\mathrm{d}\bm{x}$ (A169)
+$\approx\int_{\mathscr{D}}p(\bm{x})\frac{1}{\sqrt{\pi/2}}e^{-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}}\frac{\Psi(\bm{x},C)}{\sqrt{S}}\mathrm{d}\bm{x}$ (A170)
+$=\frac{1}{\sqrt{\pi/2}}e^{-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}}\int_{\mathscr{D}}p(\bm{x})\frac{\Psi(\bm{x},C)}{\sqrt{S}}\mathrm{d}\bm{x}.$ (A171)
+$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}}\mathop{\mathbb{E}}_{(\bm{x},y_{\bm{x}})\sim{\mathbb{D}}}\left[\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}\right],$ (A172)
+where $p(\bm{x})$ is the density of ${\mathbb{D}}$ at $(\bm{x},y_{\bm{x}})$, and $\mathscr{D}$ is the data domain. Hence we obtain
+$\log(\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}})\approx-\frac{1}{2}\log S-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}+\log\left(\mathop{\mathbb{E}}_{(\bm{x},y_{\bm{x}})\sim{\mathbb{D}}}\left[\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}\right]\right).$ (A173)
+Appendix C Training details
+C.1 Parameterization
 We utilise pre-trained feature extractors BiT-M-R50x1 (R-50) (Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B) (Dosovitskiy et al., 2021) with 85.8M parameters, both pretrained on the ImageNet-21K dataset (Russakovsky et al., 2015).
@@
 While a diverse set of adapters has been proposed, we utilise FiLM as it has been shown to be competitive in prior work (Shysheya et al., 2023; Tobaben et al., 2023).
-E.1.1 Licenses and access
+C.1.1 Licenses and access
 The licenses and means to access the model checkpoints can be found below.
-E.2 Hyperparameter tuning
+C.2 Hyperparameter tuning
 […]
@@
 learning rate 1e-7 1e-2
-E.3 Datasets
+C.3 Datasets
 Table A2 shows the datasets used in the paper.
@@
 Pets (Parkhi et al., 2012) 37 8 32 82.3% 90.7%
-Resics45 (Cheng et al., 2017) 45 32 256 83.5% 91.6%
+Resisc45 (Cheng et al., 2017) 45 32 256 83.5% 91.6%
 CIFAR100 (Krizhevsky, 2009) 100 16 128 82.2% 87.6%
-E.3.1 Licenses and access
+C.3.1 Licenses and access
 The licenses and means to access the datasets can be found below. We downloaded all datasets from TensorFlow datasets https://www.tensorflow.org/datasets but Resics45 which required manual download.
-E.4 Compute resources
+C.4 Compute resources
 All experiments but the R-50 (FiLM) experiments are run on CPU with 8 cores and 16 GB of host memory.
@@
 We estimate that in total we spend around 7 days of V100 and some dozens of weeks of CPU core time but more exact measurements are hard to make.
-Appendix F Additional results
+Appendix D Additional results
 In this section, we provide tabular results for our experiments and additional figures that did not fit into the main paper.
-F.1 Additional results for Section 4
+D.1 Additional results for Section 4
 This Section contains additional results for Section 4.
-F.1.1 Vulnerability as a function of shots
-This section displays additional results to Figure 1(a) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A4 and A5.
+D.1.1 Vulnerability as a function of shots
+This section displays additional results to Figure 1 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.1, A3 and A4.
 […]
@@
 dataset classes ($C$) shots ($S$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 2 256 0.272 0.076 0.022 512 0.195 0.045 0.011 1024 0.201 0.048 0.011 2048 0.178 0.041 0.010 4096 0.163 0.033 0.008 8192 0.143 0.026 0.006 16384 0.124 0.019 0.004 32768 0.118 0.016 0.003 65536 0.106 0.012 0.002 CIFAR10 (Krizhevsky, 2009) 10 8 0.911 0.574 0.324 16 0.844 0.526 0.312 32 0.617 0.334 0.183 64 0.444 0.208 0.106 128 0.334 0.159 0.084 256 0.313 0.154 0.086 512 0.251 0.103 0.051 1024 0.214 0.082 0.038 EuroSAT (Helber et al., 2019) 10 8 0.846 0.517 0.275 16 0.699 0.408 0.250 32 0.490 0.236 0.121 64 0.410 0.198 0.105 128 0.332 0.151 0.075 256 0.269 0.111 0.056 512 0.208 0.077 0.036 Pets (Parkhi et al., 2012) 37 8 0.937 0.631 0.366 16 0.745 0.427 0.227 32 0.588 0.321 0.173 Resics45 (Cheng et al., 2017) 45 32 0.671 0.405 0.235 64 0.534 0.289 0.155 128 0.445 0.231 0.121 256 0.367 0.177 0.088 CIFAR100 (Krizhevsky, 2009) 100 16 0.897 0.638 0.429 32 0.763 0.549 0.384 64 0.634 0.414 0.269
-F.1.2 Vulnerability as a function of the number of classes
-This section displays additional results to Figure 1(b) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.3, A6 and A7.
+D.1.2 Vulnerability as a function of the number of classes
+This section displays additional results to Figure 2 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A5 and A6.
 […]
@@
 dataset shots ($S$) classes ($C$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 32 2 0.452 0.151 0.041 CIFAR10 (Krizhevsky, 2009) 32 2 0.404 0.146 0.060 4 0.560 0.266 0.123 8 0.591 0.318 0.187 EuroSAT (Helber et al., 2019) 32 2 0.309 0.111 0.050 4 0.356 0.144 0.064 8 0.480 0.233 0.123 Pets (Parkhi et al., 2012) 32 2 0.249 0.068 0.029 4 0.326 0.115 0.056 8 0.419 0.173 0.075 16 0.493 0.245 0.127 32 0.559 0.294 0.166 Resics45 (Cheng et al., 2017) 32 2 0.310 0.103 0.059 4 0.415 0.170 0.083 8 0.510 0.236 0.119 16 0.585 0.311 0.174 32 0.644 0.382 0.218 CIFAR100 (Krizhevsky, 2009) 32 2 0.356 0.132 0.054 4 0.423 0.176 0.087 8 0.545 0.288 0.163 16 0.580 0.338 0.196 32 0.648 0.402 0.244 64 0.711 0.476 0.320
-F.1.3 Data for FiLM and from scratch training
+D.1.3 Data for FiLM and from scratch training
 Table A7: MIA vulnerability data used in Figure 5(b).
@@
 (2022) 0.700 0.400 0.276 (Krizhevsky, 2009)
-F.1.4 Predicting dataset vulnerability as function of $S$ and $C$
+D.1.4 Predicting dataset vulnerability as function of $S$ and $C$
 This section provides additional results for the model based on Equation 13
-Table A9: Results for fitting Equation A138 with statsmodels Seabold & Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$.
+Table A8: Results for fitting Equation 13 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$.
 We utilize an ordinary least squares.
@@
 $[0.025$ coeff.
-$0.975]$ $\beta_{S}$ (for $S$) 0.1 0.952 0.907 -0.506 0.011 -44.936 $0$0.529 -0.484 0.01 0.946 0.854 -0.555 0.014 -39.788 $0$0.582 -0.527 0.001 0.930 0.790 -0.627 0.019 -32.722 $0$0.664 -0.589 0.0001 0.852 0.618 -0.741 0.035 -21.467 $0$0.809 -0.673 0.00001 0.837 0.404 -0.836 0.045 -18.690 $0$0.924 -0.748 $\beta_{C}$ (for $C$) 0.1 0.952 0.907 0.090 0.021 4.231 0.000 0.048 0.131 0.01 0.946 0.854 0.182 0.026 6.960 0.000 0.131 0.234 0.001 0.930 0.790 0.300 0.036 8.335 0.000 0.229 0.371 0.0001 0.852 0.618 0.363 0.065 5.616 0.000 0.236 0.491 0.00001 0.837 0.404 0.569 0.085 6.655 0.000 0.400 0.737 $\beta_{0}$ (intercept) 0.1 0.952 0.907 0.314 0.045 6.953 0.000 0.225 0.402 0.01 0.946 0.854 0.083 0.056 1.491 0.137 -0.027 0.193 0.001 0.930 0.790 -0.173 0.077 -2.261 0.025 -0.324 -0.022 0.0001 0.852 0.618 -0.303 0.138 -2.202 0.029 -0.575 -0.032 0.00001 0.837 0.404 -0.615 0.180 -3.414 0.001 -0.970 -0.260
+$0.975]$ $\beta_{S}$ (for $S$) 0.1 0.952 0.907 -0.506 0.011 -44.936 $0$0.529 -0.484 0.01 0.946 0.854 -0.555 0.014 -39.788 $0$0.582 -0.527 0.001 0.930 0.790 -0.627 0.019 -32.722 $0$0.664 -0.589 $\beta_{C}$ (for $C$) 0.1 0.952 0.907 0.090 0.021 4.231 0.000 0.048 0.131 0.01 0.946 0.854 0.182 0.026 6.960 0.000 0.131 0.234 0.001 0.930 0.790 0.300 0.036 8.335 0.000 0.229 0.371 $\beta_{0}$ (intercept) 0.1 0.952 0.907 0.314 0.045 6.953 0.000 0.225 0.402 0.01 0.946 0.854 0.083 0.056 1.491 0.137 -0.027 0.193 0.001 0.930 0.790 -0.173 0.077 -2.261 0.025 -0.324 -0.022
 Figure A.3 shows the performance for all considered fpr.
@@
 […]
-[…]
-[…]
-F.2 Simpler variant of the prediction model
+D.2 Simpler variant of the prediction model
 […]
@@
 Table A9 provides tabular results on the performance of the variant.
-Table A10: Results for fitting Equation A138 with statsmodels Seabold & Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$.
+Table A9: Results for fitting Equation A174 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$.
 We utilize an ordinary least squares.
@@
 $[0.025$ coeff.
-$0.975]$ $\beta_{S}$ (for $S$) 0.1 0.908 0.764 -0.248 0.008 -30.976 $0$0.264 -0.233 0.01 0.940 0.761 -0.416 0.011 -36.706 $0$0.438 -0.393 0.001 0.931 0.782 -0.553 0.017 -32.507 $0$0.586 -0.519 0.0001 0.865 0.628 -0.697 0.031 -22.274 $0$0.758 -0.635 0.00001 0.862 0.400 -0.802 0.040 -20.311 $0$0.880 -0.725 $\beta_{C}$ (for $C$) 0.1 0.908 0.764 0.060 0.015 3.955 0.000 0.030 0.089 0.01 0.940 0.761 0.169 0.021 7.941 0.000 0.127 0.211 0.001 0.931 0.782 0.297 0.032 9.303 0.000 0.234 0.360 0.0001 0.865 0.628 0.371 0.059 6.328 0.000 0.255 0.486 0.00001 0.862 0.400 0.580 0.076 7.679 0.000 0.431 0.729 $\beta_{0}$ (intercept) 0.1 0.908 0.764 0.029 0.032 0.913 0.362 -0.034 0.093 0.01 0.940 0.761 -0.118 0.045 -2.613 0.010 -0.208 -0.029 0.001 0.931 0.782 -0.295 0.068 -4.345 $0$0.429 -0.161 0.0001 0.865 0.628 -0.387 0.125 -3.104 0.002 -0.633 -0.141 0.00001 0.862 0.400 -0.683 0.159 -4.288 $0$0.996 -0.369
+$0.975]$ $\beta_{S}$ (for $S$) 0.1 0.908 0.764 -0.248 0.008 -30.976 $0$0.264 -0.233 0.01 0.940 0.761 -0.416 0.011 -36.706 $0$0.438 -0.393 0.001 0.931 0.782 -0.553 0.017 -32.507 $0$0.586 -0.519 $\beta_{C}$ (for $C$) 0.1 0.908 0.764 0.060 0.015 3.955 0.000 0.030 0.089 0.01 0.940 0.761 0.169 0.021 7.941 0.000 0.127 0.211 0.001 0.931 0.782 0.297 0.032 9.303 0.000 0.234 0.360 $\beta_{0}$ (intercept) 0.1 0.908 0.764 0.029 0.032 0.913 0.362 -0.034 0.093 0.01 0.940 0.761 -0.118 0.045 -2.613 0.010 -0.208 -0.029 0.001 0.931 0.782 -0.295 0.068 -4.345 $0$0.429 -0.161
 Figure A.4 plots the performance of the variant similar to Figure 5(a) in the main text.
@@
 (c) tpr at $\textsc{fpr}=0.001$
-(d) tpr at $\textsc{fpr}=0.0001$
-(e) tpr at $\textsc{fpr}=0.00001$
 Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation A174 fitted Table A3 (ViT-B).
@@
 Note that the triangles and dots for $C=10$ are for EuroSAT.
-F.3 Empirical results for RMIA
-Figures A.6, A.7 and A.8 report additional results for RMIA Zarifzadeh et al. (2024).
+D.3 Empirical results for RMIA
+Figures A.5, A.6 and A.7 report additional results for RMIA Zarifzadeh et al. (2024).
 […]
@@
 […]
+D.4 Tabular results for Section 4.4
+Table A10 displays the tabular results for Figure 6 in Section 4.4.
+Table A10: Tabular results for Figure 6 on when attacking a ViT-B (Head) fine-tuned on PatchCamelyon. We display the median over six seeds at $\textsc{fpr}=0.1$.
+$S$ Max tpr tpr of 0.999 Quantile tpr of 0.99 Quantile tpr of 0.95 Quantile
+16384 0.77 0.43 0.22 0.11
+23170 0.73 0.37 0.19 0.10
+32768 0.65 0.30 0.15 0.08
+49152 0.54 0.23 0.13 0.07
+65536 0.49 0.19 0.11 0.07
+Appendix E Details on Section 4.5
+In Section 4.5, we compare the results of our empirical models of vulnerability (Sections 4.3 and 4.4) to DP bounds. Below we explain how we make the connection between both.
+First, we compute the upper bound on the tpr for a given $(\epsilon,\delta)$-DP privacy budget at a given fpr using Equation A175 reformulated from Kairouz et al. (2015) below:
+Theorem 8 (Kairouz et al. (2015)).
+A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP if and only if for all adjacent $\mathcal{D}\sim\mathcal{D}^{\prime}$
+$\textsc{tpr}\leq\min\{e^{\epsilon}\textsc{fpr}+\delta,1-e^{-\epsilon}(1-\delta-\textsc{fpr})\}\ .$ (A175)
+For a given $(\epsilon,\delta)$ and fpr we then obtain a value for the tpr.
+Next, we use the linear model from Section 4.3 to solve for the minimum $S$ predicted to be required given $C=2$ classes in our example.
+The coefficients can be found in Table A8 for the average case and Section 4.4 for the worst-case.
+We solve the tpr from the linear model as
+$\log_{10}(\textsc{tpr}-\textsc{fpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (A176)
+$\Leftrightarrow$ $\textsc{tpr}=S^{\beta_{S}}C^{\beta_{C}}10^{\beta_{0}}+\textsc{fpr}.$ (A177)
+Now, we find the minimum $S$ that the tpr from Equation A177 upper bounds the tpr of Equation A175 as
+$S^{\beta_{S}}C^{\beta_{C}}10^{\beta_{0}}+\textsc{fpr}=\min\{e^{\epsilon}\textsc{fpr}+\delta,1-e^{-\epsilon}(1-\delta-\textsc{fpr})\}$ (A178)
+$\Rightarrow$ $S=\left(\frac{\min\{e^{\epsilon}\textsc{fpr}+\delta,1-e^{-\epsilon}(1-\delta-\textsc{fpr})\}-\textsc{fpr}}{C^{\beta_{C}}10^{\beta_{0}}}\right)^{1/{\beta_{S}}}$ (A179)

### v5

@@
 We analyse MIA vulnerability of fine-tuned neural networks both empirically and theoretically, the latter using a simplified model of fine-tuning.
-We show that the vulnerability of non-DP models when measured as the attacker advantage at fixed false positive rate reduces according to a simple power law as the number of examples per class increases, even for the most vulnerable points, but the dataset size needed for adequate protection of the most vulnerable points is very large.
+We show that the vulnerability of non-DP models when measured as the attacker advantage at a fixed false positive rate reduces according to a simple power law as the number of examples per class increases.
+A similar power-law applies even for the most vulnerable points, but the dataset size needed for adequate protection of the most vulnerable points is very large.
 1 Introduction
+Figure 1: We observe a power-law relation between MIA vulnerability and examples per class (denoted as $S$ or shots) when attacking a fine-tuned ViT-B Head using LiRA.
+Each colored line denotes a different fine-tuning dataset where $C$ specifies the number of classes.
+The solid line is median and the error bars the min/max bounds for the Clopper-Pearson CIs over six seeds.
 Membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and differential privacy (DP; Dwork et al., 2006) provide complementary means of deriving lower and upper bounds for the privacy loss of a machine learning algorithm.
@@
 DP implicitly assumes a very powerful adversary with access to all training data except the target point and provides guarantees against every target point.
-MIAs assume an often more realistic adversary model with access to just the data distribution, and the unknown training data becoming latent variables introducing stochasticity into the attack, but the practical evaluation is statistical and cannot provide universal guarantees.
-In this paper, we seek to explore the area between DP and typical MIA evaluations.
-Inspired by an empirical finding that average MIA vulnerability of neural network fine-tuning strongly reduces as the number of samples in the target class increases (see Figure 1), we develop theory of optimal MIA against a simple model of neural network fine-tuning and reproduce the decrease in vulnerability.
+MIAs assume an often more realistic adversary model with access to just the data distribution and the unknown training data becoming latent variables that introduce stochasticity into the attack.
+However, the practical evaluation is statistical and cannot provide universal guarantees.
+In this paper, we seek to explore MIA vulnerability to extrapolate this gap.
+Inspired by an empirical finding that average MIA vulnerability of neural network fine-tuning strongly reduces as the number of samples in the target class increases (see Figure˜1), we develop theory of optimal MIA against a simple model of neural network fine-tuning and reproduce the decrease in vulnerability.
 Furthermore, the theoretical model predicts that the vulnerability of all individual samples should reduce as the number of samples increases, which we are able to verify empirically.
-Figure 1: We observe a power-law relation between MIA vulnerability and $S$ (shots, i.e., examples per class) when attacking a ViT-B Head fine-tuned using LiRA.
-The solid line is median and the error bars the min/max bounds for the Clopper-Pearson CIs over six seeds.
 To achieve our goal, we theoretically analyze and systematically apply two state-of-the-art black-box MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), to help understand practical privacy risks when fine-tuning deep-learning-based classifiers without DP protections.
@@
 Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep transfer learning models to MIA.
-In particular, we theoretically study the relationship between the number of examples per class, which we denote as shots ($S$), and MIA vulnerability (true positive rate tpr at fixed false positive rate fpr) for a simplified model of fine-tuning, and derive a power-law relationship (Figure 1) in the form
+In particular, we theoretically study the relationship between the number of examples per class, which we denote as shots ($S$), and MIA vulnerability (true positive rate tpr at fixed false positive rate fpr) for a simplified model of fine-tuning, and derive a power-law relationship (Figure˜1) in the form
 $\log(\textsc{tpr}-\textsc{fpr})=-\beta_{S}\log(S)-\beta_{0}.$ (1)
@@
 Based on extrapolation from our results, the number of examples per each class that are needed for adequate protection of the most vulnerable samples appears very high.
-Related work There has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017), models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023), and classes with less examples tend to be more vulnerable (Chang and Shokri, 2021; Kulynych et al., 2022; Tonni et al., 2020).
+Related work There has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017), models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2025), and classes with less examples tend to be more vulnerable (Chang and Shokri, 2021; Kulynych et al., 2022; Tonni et al., 2020).
 A larger generalisation error, which is related to dataset size, has also been shown to be sufficient for MIA success (Song and Mittal, 2021), though not necessary (Yeom et al., 2018).
-Similarly, minority subgroups tend to be more affected by DP (Suriyakumar et al., ; Bagdasaryan et al., 2019).
-Feldman and Zhang (2020) showed that neural networks trained from scratch can memorise a large fraction of their training data, while the memorisation is greatly reduced for fine-tuning.
+Similarly, minority subgroups tend to be more affected by DP (Suriyakumar et al., 2021; Bagdasaryan et al., 2019).
+Feldman and Zhang (2020) showed that neural networks trained from scratch are required to memorize a significant fraction of their training data to obtain high utility, while the memorisation is greatly reduced for fine-tuning.
 Additionally, Tobaben et al.
@@
 Nonetheless, the prior works do not consider the rate of change in the vulnerability evaluated at a low fpr, as dataset properties change.
-Our work significantly expands on these works by explicitly identifying a quantitative relationship between dataset properties and MIA vulnerability, i.e., the power-law in Equation 1.
+Our work significantly expands on these works by a) explicitly identifying a quantitative relationship between dataset properties and MIA vulnerability, i.e., the power-law in Equation˜1, and b) focusing on the worst-case vulnerability, both evaluated at a low FPR.
+This in turn allows us to extrapolate MIA vulnerability to DP guarantee.
 List of contributions We analyze the MIA vulnerability of deep transfer learning using two state-of-the-art score-based MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), which are a strong realistic threat model.
-We first analytically derive the power-law relationship in Equation 1 for both MIAs by introducing a simplified model of the optimal membership inference (Section 3).
+We first analytically derive the power-law relationship in Equation˜1 for both MIAs by introducing a simplified model of the optimal membership inference (Section˜3).
 We support our theoretical findings by an extensive empirical study on the MIA vulnerability of deep learning models by focusing on a transfer-learning setting for image classification task, where a large pre-trained neural network is fine-tuned on a sensitive dataset.
@@
 In this model LiRA is the optimal attack.
-For this model, we prove a power-law relationship for both average and worst-case between the LiRA as well as RMIA vulnerability and the number of examples per class (See Section 3.4).
+For this model, we prove a power-law relationship for both average and worst-case between the LiRA as well as RMIA vulnerability and the number of examples per class (See Section˜3.4).
   2. 2.
-    MIA experiments on the average case vulnerability: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power-law between the number of examples per class and the vulnerability to MIA (see Figures 1 and 4.2).
+    MIA experiments on the average case vulnerability: We conduct a comprehensive study of MIA vulnerability (tpr at a fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power-law between the number of examples per class and the vulnerability to MIA (see Figures˜1 and 4.2).
 We fit a regression model which follows the functional form of the theoretically derived power-law.
-We show both a very good fit on the training data as well as a good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Section 4.3).
+We show both a very good fit on the training data as well as a good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Section˜4.3).
   3. 3.
-    MIA experiments on the worst-case vulnerability: We extend the experiments to worst-case individual sample vulnerabilities and observe a similar decrease in vulnerability for quantiles of vulnerable data points and a slower decrease for the maximum individual vulnerability (Section 4.4).
-By extrapolation we find that an adequate protection of the most vulnerable samples would require an extremely large dataset (Section 4.5).
+    MIA experiments on the worst-case vulnerability: We extend the experiments to worst-case individual sample vulnerabilities and observe a similar decrease in vulnerability for quantiles of vulnerable data points and a slower decrease for the maximum individual vulnerability (Section˜4.4).
+By extrapolation we find that an adequate protection of the most vulnerable samples would require an extremely large dataset (Section˜4.5).
 2 Background
-Notation for the properties of the training dataset $\mathcal{\mathcal{D}}$: 1 $C$for the number of classes 2 $S$for shots (examples per class) 3 $\mathcal{|\mathcal{D}|}$for training dataset size ($\mathcal{|\mathcal{D}|}=CS$).
+Notation for the properties of the training dataset $\mathcal{\mathcal{D}}$: 1 CCfor the number of classes 2 SSfor shots (examples per class) 3 $\mathcal{|\mathcal{D}|}$for training dataset size ($\mathcal{|\mathcal{D}|}=CS$).
 We denote the number of MIA shadow models with $M$.
@@
 Within this model, we prove a power-law between the per-example vulnerability and the number $S$ of examples per class.
-Finally, the per-example power-law is analytically extended to average-case MIA vulnerability, for which we provide empirical evidence in Section 4.
-We primarily focus on the analysis of LiRA, since it is the optimal attack in our simplified model.
-In Appendix B we show that similar theoretical results also hold for RMIA.
+Finally, the per-example power-law is analytically extended to average-case MIA vulnerability, for which we provide empirical evidence in Section˜4.
+We primarily focus on the analysis of (online) LiRA, since it is the optimal attack in our simplified model.
+We show that similar theoretical results also hold for RMIA in Appendix˜B and offline LiRA Section˜A.4.
 3.1 Preliminaries
-First, let us restate the MIA score from LiRA as defined by Carlini et al.
+First, let us restate the MIA score from (online) LiRA as defined by Carlini et al.
 (2022).
@@
 3.2 Computing the tpr for LiRA
-Using the LiRA formulation of Equation 4, the tpr for the target point $(\bm{x},y)$ for LiRA is defined as
+Using the LiRA formulation of Equation˜4, the tpr for the target point $(\bm{x},y)$ for LiRA is defined as
 $\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\mathrm{LR}(\bm{x})\geq\tau\mid(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}\right),$ (5)
-where $\tau$ is a threshold that defines a rejection region of the likelihood ratio test, and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Section A.1 for derivation).
+where $\tau$ is a threshold that defines a rejection region of the likelihood ratio test, and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Section˜A.1 for derivation).
 We define the average-case tpr for LiRA by taking the expectation over the data distribution:
@@
 This allows us to also assume that $\hat{\sigma}(\bm{x})=\hat{\sigma}_{\mathrm{in}}(\bm{x})=\hat{\sigma}_{\mathrm{out}}(\bm{x})$ and $\sigma(\bm{x})=\sigma_{\mathrm{in}}(\bm{x})=\sigma_{\mathrm{out}}(\bm{x})$, where $\hat{\sigma}(\bm{x})$ is the standard deviation of $t_{\bm{x}}$ estimated from shadow models and $\sigma(\bm{x})$ is the true scale parameter of $t_{\bm{x}}$.
-(See Section A.2 for the validity of these assumptions).
+(See Section˜A.2 for the validity of these assumptions).
 The following result reduces the LiRA vulnerability to the location and scale parameters of $t_{\bm{x}}$.
@@
 Proof.
-See Section A.3. ∎
+See Section˜A.3. ∎
 Here we assume that an attacker trains shadow models with the true underlying distribution.
 However, in real-world settings the precise underlying distribution may not be available for an attacker.
-We relax this assumption in Section A.4 so that the attacker only needs an approximated underlying distribution for the optimal LiRA as in Lemma 1.
+We relax this assumption in Section˜A.5 so that the attacker only needs an approximated underlying distribution for the optimal LiRA as in Lemma˜1.
 3.4 A simplified model of the optimal membership inference
@@
 This simplified model resembles a linear (Head) classifier often used in transfer learning when adapting to a new dataset.
-We also focus on the linear (Head) classifier in our empirical evaluation in Section 4.
+We also focus on the linear (Head) classifier in our empirical evaluation in Section˜4.
 In the linear classifier, we find a matrix $W$ and biases $b$, to optimize the cross-entropy between the labels and logits $Wv+b$, where $v$ denotes the feature space representation of the data.
@@
 Fix a target example $(\bm{x},y)$. For the simplified model with arbitrary $C$ and infinitely many shadow models, the per-example LiRA vulnerability is given as
-$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S))$ (9)
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (9)
 where $m_{\bm{x}}$ is the true mean of class $y$ and $\xi(S)=O(1/\sqrt{S})$. For large $S$ we have
@@
 Proof.
-See Section A.5. ∎
+See Section˜A.6. ∎
 An immediate upper bound is obtained from Theorem 2 by the Cauchy-Schwarz inequality:
-$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{||\bm{x}-m_{\bm{x}}||}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}\\ +\log(1+\xi(S)).$ (11)
-This implies that if $||\bm{x}-m_{\bm{x}}||$ is bounded, then the worst-case vulnerability is also bounded.
+[…]
+This implies that if $\bm{x}-m_{\bm{x}}||$ is bounded, then the worst-case vulnerability is also bounded.
 Hence we can significantly reduce the MIA vulnerability of all examples in this non-DP setting by simply increasing the number of examples per class.
-Now the following corollary extends the power-law to the average-case MIA vulnerability. We will also empirically validate this result in Section 4.
+Remark 3.
+By Lemma˜1 and the proof of Theorem˜2, a necessary condition for the power-law is that $(\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x}))/\sigma(\bf{x})$ converges to zero at rate $O(1/S^{\alpha})$ with $\alpha>0$.
+In our simplified model, this holds with $\alpha=1/2$.
+However, $\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x})\rightarrow 0$ would not always be the case for larger neural networks trained from scratch.
+Therefore, we do not expect similar power-laws in more general training algorithms.
+Now the following corollary extends the power-law to the average-case MIA vulnerability. We will also empirically validate this result in Section˜4.
 Corollary 4 (Average-case LiRA power-law).
@@
 Proof.
-See Section A.6. ∎
+See Section˜A.7. ∎
 4 Empirical evaluation of MIA vulnerability and dataset properties
@@
 Based on our observations, we propose a method to predict the vulnerability to MIA using these properties.
+4.1 Experimental setup
 We focus on a image classification setting where we fine-tune pre-trained models on sensitive downstream datasets and assess the MIA vulnerability using LiRA and RMIA with $M=256$ shadow/reference models.
-We base our experiments on a subset of the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a test classification accuracy $>80\%$ (see Table A2).
-4.1 Experimental setup
+We base our experiments on a subset of the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a test classification accuracy $80\%$ (see Table˜A2).
+We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k (Russakovsky et al., 2015).
+The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section˜D.1.
+We optimise the hyperparameters (batch size, learning rate and number of epochs) using the Optuna library (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section˜C.2).
+We provide the the code for reproducing the experiments in an open repository [^1] https://github.com/DPBayes/impact-dataset-properties-MI-vulnerability-deep-TL.
+4.2 Experimental results
 Figure 2: Small effect of number of classes $C$ (classes) on MIA vulnerability when attacking a fine-tuned ViT-B Head.
 The solid line is median and the error bars the min/max bounds for the Clopper-Pearson CIs over 12 seeds ($S=32$).
-We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k (Russakovsky et al., 2015).
-The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section D.1.
-We optimise the hyperparameters (batch size, learning rate and number of epochs) using the Optuna library (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section C.2).
-We provide the the code for reproducing the experiments in the supplementary material.
-4.2 Experimental results
-Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 2) using LiRA.
+Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section˜2) using LiRA.
 We make the following observations:
-  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1.
-We provide tabular data and experiments using ResNet-50 in the Appendix (Figures A.1, A3 and A4).
-  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 2 with tabular data and experiments using ResNet-50 in the Appendix (Figures A.2, A5 and A6).
-However, the trend w.r.t.
-$C$ is not as clear as with $S$.
-RMIA In Figure 3 we compare the vulnerability of the models to LiRA and RMIA as a function of the number of $S$ (shots) at $\textsc{fpr}=0.1$.
+[…]
+[…]
+RMIA In Figure˜3 we compare the vulnerability of the models to LiRA and RMIA as a function of the number of $S$ (shots) at $\textsc{fpr}=0.1$.
 We observe the power-law for both attacks, but the RMIA is more unstable than LiRA (especially for lower fpr).
-More results for RMIA are in Figures A.5, A.6 and A.7 in the Appendix.
-[…]
+More results for RMIA are in Figures˜A.5, A.6 and A.7 in the Appendix.
+Figure 3: LiRA and RMIA vulnerability ($(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.1$) as a function of shots ($S$) when attacking a ViT-B Head fine-tuned on different datasets.
+For better visibility, we split the datasets into two panels.
+We observe the power-law for both attacks, but the RMIA is more unstable than LiRA.
+The lines display the median over six seeds.
 4.3 Model to predict dataset vulnerability
-The trends seen in Figure 1 suggest the same power law relationship that we derived for the simplified model of membership inference in Section 3.
+The trends seen in Figure˜1 suggest the same power law relationship that we derived for the simplified model of membership inference in Section˜3.
 We fit a linear regression model to predict $\log(\textsc{tpr}-\textsc{fpr})$ for each $\textsc{fpr}=10^{-k},k=1,\dots,5$ separately using the $\log C$ and $\log S$ as covariates with statsmodels (Seabold and Perktold, 2010).
-The general form of the model can be found in Equation 13, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
+The general form of the model can be found in Equation˜13, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
 $\log_{10}(\textsc{tpr}-\textsc{fpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (13)
-[…]
-We utilise MIA results of ViT-B (Head) (see Table A3) as the training data.
+In Section˜D.2, we propose a variation of the regression model that predicts $\log_{10}(\textsc{tpr})$ instead of $\log_{10}(\textsc{tpr}-\textsc{fpr})$ but this alternative model performs worse on our empirical data and predicts $\textsc{tpr}<\textsc{fpr}$ in the tail when $S$ is very large.
+We utilise MIA results of ViT-B (Head) (see Table˜A3) as the training data.
 Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well.
-We provide further evidence for other fpr in Figure A.3 and Table A8 in the Appendix.
-Figure 4 shows the parameters of the prediction model fitted to the training data.
-For larger fpr, the coefficient $\beta_{S}$ is around $-0.5$, as our theoretical analysis predicts.
-Figure 4: Coefficient values for different fpr when fitting a regression model based on Equation 13 fitted on data from ViT-B (Head) with LiRA (Table A3).
+We provide further evidence for other fpr in Figure˜A.3 and Table˜A8 in the Appendix.
+Figure˜4 shows the parameters of the prediction model fitted to the training data.
+For larger fpr, the coefficient $\beta_{S}$ is around 0.5-0.5, as our theoretical analysis predicts.
+Figure 4: Coefficient values for different fpr when fitting a regression model based on Equation˜13 fitted on data from ViT-B (Head) with LiRA (Table˜A3).
 The error bars display the 95% confidence intervals based on Student’s t-distribution.
@@
 We analyse how the regression model trained on the ViT-B (Head) data generalizes to other target models. The main points are:
-  • R-50 (Head): Figure 5(a) shows that the regression model is robust to a change of the feature extractor, as it is able to predict the tpr for R-50 (Head) (test $R^{2}=0.790$).
-  • R-50 (FiLM): Figure 5(b) shows that the prediction quality is good for R-50 (FiLM) models.
-These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section C.1).
+  • R-50 (Head): Figure˜5(a) shows that the regression model is robust to a change of the feature extractor, as it is able to predict the tpr for R-50 (Head) (test $R^{2}=0.790$).
+  • R-50 (FiLM): Figure˜5(b) shows that the prediction quality is good for R-50 (FiLM) models.
+These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section˜C.1).
 Tobaben et al.
@@
 (2023) with own FiLM training runs.
-Refer to Table A7 in the Appendix.
+Refer to Table˜A7 in the Appendix.
   • From-Scratch-Training: Carlini et al.
@@
 To the best of our knowledge these are the only published LiRA results on image classification models.
-Figure 5(b) displays that our prediction model underestimates the vulnerability of the from-scratch trained target models.
-We have identified two potential explanations for this 1 In from-scratch-training all weights of the model need to be trained from the sensitive data and thus potentially from-scratch-training could be more vulnerable than fine-tuning.
+Figure˜5(b) displays that our prediction model underestimates the vulnerability of the from-scratch trained target models.
+We have identified two potential explanations for this: 1 In from-scratch-training all weights of the model need to be trained from the sensitive data and thus potentially from-scratch-training could be more vulnerable than fine-tuning.
 2 The strongest attack in Carlini et al.
@@
 We are not using this optimization.
-(a) The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100).
+Additionally, as noted in Remark˜3, our theoretical analysis suggests that the power-law would not always hold for general from-scratch training.
+(a) The dots show the median tpr for the train set (ViT-B; Table˜A3) and the test set (R-50; Table˜A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100).
 The linear model is robust to changing the feature extractor from ViT-B to R-50.
 (b) Regression model is robust to changing the fine-tuning method from Head to FiLM, but from scratch training seems to be more vulnerable than predicted.
-1 left: fine-tuned with FiLM (see Table A7) 2 right: trained from scratch.
+1 left: fine-tuned with FiLM (see Table˜A7) 2 right: trained from scratch.
 Data is from Carlini et al.
 (2022) .
-Figure 5: Performance of the regression model based on Equation 13 fitted on data from Table A3.
+Figure 5: Performance of the regression model based on Equation˜13 fitted on data from Table˜A3.
 4.4 Individual MIA vulnerability
 In order to assess the per-sample MIA vulnerability, we run the experiment with $257$ models with each of them once acting as a target and as a shadow model otherwise, compute the tpr at fpr for every sample separately.
-In Figure 6, we display the individual vulnerability as a function of $S$ (shots) for Patch Camelyon.
+In Figure˜6, we display the individual vulnerability as a function of $S$ (shots) for Patch Camelyon.
 The plot shows the maximal vulnerability of all samples and different quantiles over six seeds.
@@
 The quantiles are more robust to extreme outliers and show decreasing trends already at much lower $S$ than the maximum vulnerability.
-When fitting the model in Equation 13 we observe that the coefficients $\beta_{S}$ that model the relationship between vulnerability and examples per class for the quantiles are $-0.5603$, $-0.5688$ and $-0.4796$ which is close to the theoretical value of $-0.5$ derived in the theoretical analysis in Section 3.4.
-However, the maximum vulnerability decreases with a lower slope of $-0.2695$ which is considerably smaller.
-With larger $S$ the slope of the max vulnerability increases, e.g., with $S\geq 32768$ the coefficient is $-0.3478$ suggesting that higher $S$ is required for the most vulnerable points.
+When fitting the model in Equation˜13 we observe that the coefficients $\beta_{S}$ that model the relationship between vulnerability and examples per class for the quantiles are 0.5603-0.5603, 0.5688-0.5688 and 0.4796-0.4796 which is close to the theoretical value of 0.5-0.5 derived in the theoretical analysis in Section˜3.4.
+However, the maximum vulnerability decreases with a lower slope of 0.2695-0.2695 which is considerably smaller.
+With larger $S$ the slope of the max vulnerability increases, e.g., with $S\geq 32768$ the coefficient is 0.3478-0.3478 suggesting that higher $S$ is required for the most vulnerable points.
 Figure 6: Individual vulnerability for ViT-B (Head) when fine-tuning on Patch Camelyon.
 We observe a similar power-law relationship for individuals when looking at the quantiles but the max decreases slower.
-Table 1: The minimum $S$ with $C$=$2$ predicted by the models in Sections 4.3 and 4.4 to empirically match the DP bounds ($\delta$=$10^{-5}$) calculated through (Kairouz et al., 2015) in terms of tpr at fpr.
-$\epsilon$ min $S$ in average case min $S$ in worst-case fpr=$0.1$ fpr=$0.01$ fpr=$0.001$ fpr=$0.1$ $0.25$ 5 40054005\,4005 400 69 0006900069\,00069 000 320 000320000320\,000320 000 $5.5\times 10^{9}$ $0.50$ 1 10011001\,1001 100 16 0001600016\,00016 000 88 0008800088\,00088 000 $2.6\times 10^{8}$ $0.75$ $360$ 5 90059005\,9005 900 38 0003800038\,00038 000 $3.5\times 10^{7}$ $1.00$ $160$ 2 70027002\,7002 700 19 0001900019\,00019 000 $7.0\times 10^{6}$
+Table 1: The minimum $S$ with CC=22 predicted by the models in Sections˜4.3 and 4.4 to empirically match the DP bounds ($\delta$$10^{-5}$) calculated through (Kairouz et al., 2015) in terms of tpr at fpr.
+$\epsilon$ min $S$ in average case min $S$ in worst-case fpr=$0.1$ fpr=$0.01$ fpr=$0.001$ fpr=$0.1$ $0.25$ 5 4005\,400 69 00069\,000 320 000320\,000 $5.5\times 10^{9}$ $0.50$ 1 1001\,100 16 00016\,000 88 00088\,000 $2.6\times 10^{8}$ $0.75$ $360$ 5 9005\,900 38 00038\,000 $3.5\times 10^{7}$ $1.00$ $160$ 2 7002\,700 19 00019\,000 $7.0\times 10^{6}$
 4.5 Comparison between empirical models and universal DP bounds
 While the practical evaluation through MIAs is statistical and does not provide universal formal guarantees like DP, the power-law can aid understanding about how the practical vulnerability to MIA behaves in a more realistic threat model when the examples per class increase.
-[…]
+Using the translation between the $(\textsc{tpr},\textsc{fpr})$ to $(\epsilon,\delta)$-DP guarantees proposed by Kairouz et al.
+(2015), we compute the minimum $S$ predicted by our empirical models such that the predicted tpr matches the theoretical bound at target DP level.
+(See Appendix˜E for a more detailed description.) Table˜1 shows the resulting lower bounds of $S$ for various DP levels and values of fpr.
+While our empirical observations do not provide any formal guarantees, the comparison serves as an illustration to better understand the different requirements of the average and individual vulnerability.
+We can see that both for the average and the worst case, obtaining a low fpr would require a large amount of samples per class in order to match the tpr for meaningfully strong DP bounds.
 5 Discussion
@@
 universal).
-While we were able to show both theoretically (Section 3) and empirically (Section 4) that having more examples per class provides protection against MIA in fine-tuned neural networks, the numbers required for significant protection (Section 4.5) limit the practical utility of this observation.
+While we were able to show both theoretically (Section˜3) and empirically (Section˜4) that having more examples per class provides protection against MIA in fine-tuned neural networks, the numbers required for significant protection (Section˜4.5) limit the practical utility of this observation.
 Using a level of formal DP bounds that provide meaningful protection as a yardstick, at least tens of thousands of examples per class are needed for every class even at $\textsc{fpr}=0.001$.
@@
 This stresses the importance of formal guarantees like DP for privacy protection.
+Our formal analysis focuses on a setting that can be linked to fine-tuning.
+As shown in Figure˜5(b), from-scratch training likely has higher vulnerability.
+Formally analyzing more models is an interesting area for future research.
+Since our MIA evaluation assumes that only a target point is known to the adversary and the rest of the dataset is random, this stochasticity likely introduces some protection.
+Hence the power-law may be invalidated under stronger MIA settings where the adversary has access to other points in the private dataset (Bai et al., 2025).
+Our experiments show that there is a difference between classes in terms of vulnerability and an interesting direction for future work is to understand the properties of classes that influence this vulnerability, e.g., variability within or between classes or their separability.
+Another direction for future work is understanding the impact of pre-training data on the vulnerability of fine-tuning data.
 Broader Impact
@@
 Limitations
-We mostly consider LiRA in our paper which is optimal for our simiplified model of membership inference (Section 3.4) but for the transfer-learning experiments (Section 4) there might be stronger attacks in the future.
-Furthermore, our results assume well-behaved underlying distributions.
+We mostly consider LiRA in our paper which is optimal for our simplified model of membership inference (Section˜3.4) but for the transfer-learning experiments (Section˜4) there might be stronger attacks in the future.
+Furthermore, our simplified model assumes well-behaved underlying distributions, meaning that the data is normally distributed around the class centres.
+We leave the analysis of other data distributions (e.g., heavy-tailed distributions) to future work..
 Formal bounds on MIA vulnerability would require something like DP.
@@
 Models trained from scratch are likely to be more vulnerable.
-We will publish the code in an open repository after acceptance of the paper.
+We provide the code in an open repository [^2] https://github.com/DPBayes/impact-dataset-properties-MI-vulnerability-deep-TL. All used pre-trained models and datasets are publicly available.
 Acknowledgments
-This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617).
+This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247), the European Union (Project 101070617) as well as JSPS KAKENHI Grant Number 25KJ1515.
 Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission.
 Neither the European Union nor the granting authority can be held responsible for them.
-The authors wish to thank the CSC – IT Center for Science, Finland for supporting this project with computational and data storage resources.
+he authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.
 We thank Mikko A.
@@
 Supplementary Material
-Appendix A Details of Section 3
+Contents
+Appendix A Details of Section˜3
 A.1 Formulating LiRA
@@
 A.2 On the assumption of shared scale
-In Section 3 we assumed that for LiRA $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}=\hat{\sigma}_{\mathrm{out}}$.
-Using the simplified model formulated in Section 3.4, we show that for large enough number $S$ of examples per class these assumptions are reasonable.
+In Section˜3 we assumed that for LiRA $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}=\hat{\sigma}_{\mathrm{out}}$.
+Using the simplified model formulated in Section˜3.4, we show that for large enough number $S$ of examples per class these assumptions are reasonable.
 Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),...,(\bm{x}_{j,S},j)\}_{j=1}^{C}$. Then the IN/OUT LiRA scores are given as
@@
 Thus, the differences $\sigma_{\mathrm{in}}-\sigma_{\mathrm{out}}$ and $\hat{\sigma}_{\mathrm{in}}-\hat{\sigma}_{\mathrm{out}}$ are negligible for large $S$.
-A.3 Proof of Lemma 1
+A.3 Proof of Lemma˜1
 See 1
@@
 $\frac{1}{2\hat{\sigma}^{2}}(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})(2t_{\bm{x}}-\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})\geq\log\tau$ (A15)
-$\begin{cases}t_{\bm{x}}\geq\frac{\hat{\sigma}^{2}\log\tau}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ t_{\bm{x}}\leq\frac{\hat{\sigma}^{2}\log\tau}{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A16)
+[…]
 Then if $\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}$, in the limit of infinitely many shadow models
@@
 $=F_{Z}\left(F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\mu_{\mathrm{in}}-\mu_{\mathrm{out}}}{\sigma}\right).$ (A28)
-A.4 Relaxing the assumption of Lemma 1
-In Lemma 1 we assume that an attacker has access to the true underlying distribution.
+A.4 Offline LiRA
+Carlini et al. (2022) proposes offline LiRA that only trains OUT shadow models for computational efficiency. Instead of likelihood ratio, the score for offline LiRA is given as
+$\Lambda(\bm{x})=\Pr_{Z}(\hat{\mu}_{\mathrm{out}}(\bm{x})+\hat{\sigma}_{\mathrm{out}}(\bm{x})Z\leq t_{\bm{x}}),$ (A29)
+where $Z$ is a location-scale distribution with standard location and unit scale. Then tpr of offline LiRA becomes
+$\textsc{tpr}_{\textrm{off}\mathrm{LiRA}}(\bm{x})=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|},\phi^{M}}\left(\Lambda(\bm{x})\geq\tau\mid(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}\right),$ (A30)
+where $\gamma$ is a tunable parameter.
+fpr is also given similarly.
+Below we show a result for offline LiRA similar to Lemma˜1, but under a slightly more strict assumption that an attacker accurately estimates the IN/OUT score distributions.
+In the following, as for online LiRA, we assume $\sigma_{\mathrm{in}}=\sigma_{\mathrm{out}}$.
+Lemma 5 (Per-example offline LiRA vulnerability).
+Suppose that $t_{\bm{x}}$ follows the normal distribution with means $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and standard deviation $\sigma(\bm{x})$.
+Assume that an attacker has access to the underlying distribution $\mathbb{D}$}.
+Then for a large enough number of examples per class and infinitely many shadow models, the offline LiRA vulnerability of a fixed target example is
+$\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})=\Phi\left(\Phi^{-1}\left(\textsc{fpr}_{\mathrm{offLiRA}}\right)+\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}\right)$ (A31)
+where $\Phi$ is the standard normal cdf.
+Proof.
+For infinitely many shadow models, we have
+$\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x})=\Pr_{\mathcal{D}_{\mathrm{target}}\sim{\mathbb{D}}^{|\mathcal{D}|}}\left(\Lambda(\bm{x})\geq\tau\mid(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}}\right),$ (A32)
+and the score is now
+$\Lambda(\bm{x})=\Pr_{\eta}(\mu_{\mathrm{out}}(\bm{x})+\sigma(\bm{x})\eta\leq t_{\bm{x}}),$ (A33)
+where $\eta\sim\mathcal{N}(0,1)$. When $(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}}$, $t_{\bm{x}}=\mu_{\mathrm{out}}(\bm{x})+\sigma(\bm{x})Z$. Thus we have
+$\textsc{fpr}_{\mathrm{offLiRA}}$ $=\Pr_{Z}\left(\Lambda(\bm{x})\geq\gamma\right)$ (A34)
+$=\Pr_{Z}\left(\Pr_{\eta}(\mu_{\mathrm{out}}(\bm{x})+\sigma(\bm{x})\eta\leq\mu_{\mathrm{out}}(\bm{x})+\sigma(\bm{x})Z)\geq\gamma\right)$ (A35)
+$=\Pr_{Z}\left(\Pr_{\eta}(\eta\leq Z)\geq\gamma\right)$ (A36)
+$=\Pr_{Z}(\Phi(Z)\geq\gamma)$ (A37)
+$=1-\gamma.$ (A38)
+On the other hand, when $(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}$, $t_{\bm{x}}=\mu_{\mathrm{in}}(\bm{x})+\sigma(\bm{x})Z$. Thus we obtain
+$\textsc{tpr}_{\mathrm{offLiRA}}$ $=\Pr_{Z}\left(\Pr_{\eta}(\mu_{\mathrm{out}}(\bm{x})+\sigma(\bm{x})\eta\leq\mu_{\mathrm{in}}(\bm{x})+\sigma(\bm{x})Z)\geq\gamma\right)$ (A39)
+$=\Pr_{Z}\left(\Pr_{\eta}\left(\eta\leq\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}+Z\right)\geq 1-\textsc{fpr}_{\mathrm{offLiRA}}\right)$ (A40)
+$=\Pr_{Z}\left(\Pr_{\eta}\left(\eta\leq-\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}-Z\right)\leq\textsc{fpr}_{\mathrm{offLiRA}}\right)$ (A41)
+$=\Pr_{Z}\left(-\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}-Z\leq\Phi^{-1}(\textsc{fpr}_{\mathrm{offLiRA}})\right)$ (A42)
+$=\Phi\left(\Phi^{-1}\left(\textsc{fpr}_{\mathrm{offLiRA}}\right)+\frac{\mu_{\mathrm{in}}(\bm{x})-\mu_{\mathrm{out}}(\bm{x})}{\sigma(\bm{x})}\right).$ (A43)
+Consequently, the power-law also holds for offline LiRA in the simplified model:
+Corollary 6 (Per-example offline LiRA power-law).
+Fix a target example $(\bm{x},y)$. For the simplified model with arbitrary $C$ and infinitely many shadow models, the per-example offline LiRA vulnerability is given as
+$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S))$ (A44)
+where $m_{\bm{x}}$ is the true mean of class $y$ and $\xi(S)=O(1/\sqrt{S})$. For large $S$ we have
+$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A45)
+A.5 Relaxing the assumption of Lemma˜1
+In Lemma˜1 we assume that an attacker has access to the true underlying distribution.
 However, in real-world settings the precise underlying distribution may not be available for an attacker.
-In the following, noting that the Equation 8 mainly relies on the true location parameters $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and scale parameter $\sigma(\bm{x})$, we relax this assumption of distribution availability so that the attacker only needs an approximated underlying distribution for the optimal LiRA that achieves the performance in Lemma 1.
+In the following, noting that the Equation˜8 mainly relies on the true location parameters $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and scale parameter $\sigma(\bm{x})$, we relax this assumption of distribution availability so that the attacker only needs an approximated underlying distribution for the optimal LiRA that achieves the performance in Lemma˜1.
 First, notice that if we completely drop this assumption so that an attacker trains shadow models with an arbitrary underlying distribution, then we may not be able to choose a desired $\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$.
-From Equation A22 we have
+From Equation˜A22 we have
 $\frac{\hat{\sigma}^{2}\log\tau}{\sigma(\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}})}+\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2\sigma}-\frac{\mu_{\mathrm{out}}}{\sigma}=\begin{cases}F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\text{ if }\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\text{ if }\hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}\end{cases}$ (A46)
@@
 Since it does not make sense to choose a rejection region of the likelihood ratio test such that $\tau<1$, we assume $\tau\geq 1$. Then we have
-$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}-\hat{\mu}_{\mathrm{out}}}{\hat{\sigma}^{2}}\left(\sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\right)\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A31)
-Therefore, a sufficient condition about attacker’s knowledge on the underlying distribution for Lemma 1 to hold is
+[…]
+Therefore, a sufficient condition about attacker’s knowledge on the underlying distribution for Lemma˜1 to hold is
 $\begin{cases}\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\geq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))-\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}+\mu_{\mathrm{out}}\leq 0&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A49)
 Rearranging the terms yields
-$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\leq\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\geq\sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A33)
+[…]
 For example, if the estimated mean $\hat{\mu}_{\mathrm{out}}(<\hat{\mu}_{\mathrm{in}})$ is too large compared to the true parameter $\mu_{\mathrm{out}}$, the left hand side for the case $\hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}$ becomes very large, thereby forcing us to choose sufficiently small $\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})$.
@@
 We summarise this discussion in the following:
-Lemma 4 (Lemma 1 with relaxed assumptions).
+Lemma 7 (Lemma˜1 with relaxed assumptions).
 Suppose that the true IN/OUT distributions of $t_{\bm{x}}$ are of a location-scale family with locations $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and a shared scale $\sigma(\bm{x})$ such that the distributions have finite first and second moments.
 Assume that LiRA models $t_{\bm{x}}$ by $\mathcal{N}(\hat{\mu}_{\mathrm{in}}(\bm{x}),\hat{\sigma}(\bm{x})^{2})$ and $\mathcal{N}(\hat{\mu}_{\mathrm{out}}(\bm{x}),\hat{\sigma}(\bm{x})^{2})$, and that in the limit of infinitely many shadow models, estimated parameters $\hat{\mu}_{\mathrm{in}}(\bm{x}),\hat{\mu}_{\mathrm{out}}(\bm{x})$ and $\hat{\sigma}(\bm{x})$ satisfy the following:
-$\begin{cases}\frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\leq\sigma F_{Z}^{-1}(1-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}>\hat{\mu}_{\mathrm{out}}\\ \frac{\hat{\mu}_{\mathrm{in}}+\hat{\mu}_{\mathrm{out}}}{2}-\mu_{\mathrm{out}}\geq\sigma F_{Z}^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))&\quad\text{ if }\ \hat{\mu}_{\mathrm{in}}<\hat{\mu}_{\mathrm{out}}.\end{cases}$ (A34)
+[…]
 Then in the limit of infinitely many shadow models, the LiRA vulnerability of a fixed target example $(\bm{x},y)$ is
@@
 where $F_{Z}$ is the cdf of $t$ with the standard location and unit scale, assuming that the inverse of $F_{Z}$ exists.
-A.5 Proof of Theorem 2
+A.6 Proof of Theorem˜2
 See 2
@@
 $\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A71)
-A.6 Proof of Corollary 3
-See 3
+A.7 Proof of Corollary˜4
+See 4
 Proof.
@@
 (2024) empirically demonstrates that these approaches exhibit similar performances.
-Let $\hat{\mu}_{\bm{a},\bm{b}}$ and $\hat{\sigma}_{\bm{a},\bm{b}}$ denote, respectively, the mean and standard deviation of $t_{b}=\ell(\mathcal{M}(b),y_{b})$ estimated from shadow models, where $a$ denotes which of $(\bm{x},y_{\bm{x}})$ and $(\bm{z},y_{\bm{z}})$ is in the training set.
+Let $\hat{\mu}_{\bm{a},\bm{b}}$ and $\hat{\sigma}_{\bm{a},\bm{b}}$ denote, respectively, the mean and standard deviation of $t_{\bm{b}}=\ell(\mathcal{M}(\bm{b}),y_{\bm{b}})$ estimated from shadow models, where $\bm{a}$ denotes which of $(\bm{x},y_{\bm{x}})$ and $(\bm{z},y_{\bm{z}})$ is in the training set.
 By Equation 11 in (Zarifzadeh et al., 2024), the pairwise likelihood ratio is given as
@@
 Then RMIA exploits the probability of rejecting the pairwise likelihood ratio test over $(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}$:
-$\mathrm{Score}_{\mathrm{RMIA}}(\bm{x})=\Pr_{(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,\bm{z})\geq\gamma\right),$ (A63)
+$\mathrm{Score}_{\mathrm{RMIA}}(\bm{x})=\Pr_{(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}}\left(\mathrm{LR}(\bm{x},\bm{z})\geq\gamma\right),$ (A80)
 where $\mathbb{D}$} is the underlying data distribution.
@@
 Next we focus on the per-example RMIA performance. As in the case of LiRA, we assume that $t_{\bm{x}}$ and $t_{\bm{z}}$ follow distributions of the location-scale family. We have
-$t_{\bm{x}}$ =$=\begin{cases}\mu_{\bm{x},\bm{x}}+\sigma_{\bm{x},\bm{x}}Z&\text{ if }(x,y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{\bm{z},\bm{x}}+\sigma_{\bm{z},\bm{x}}Z&\text{ if }(x,y_{\bm{x}})\notin\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\in\mathcal{D}_{\mathrm{target}}\end{cases}$ (A67)
-$t_{\bm{z}}$ =$=\begin{cases}\mu_{\bm{x},\bm{z}}+\sigma_{\bm{x},\bm{z}}Z&\text{ if }(x,y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{\bm{z},\bm{z}}+\sigma_{\bm{z},\bm{z}}Z&\text{ if }(x,y_{\bm{x}})\notin\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\in\mathcal{D}_{\mathrm{target}}.\end{cases}$ (A68)
+[…]
+[…]
 where $Z$ follows a distribution of location-scale family with the standard location and unit scale.
@@
 $=$ $\frac{\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}}}{2\hat{\sigma}_{\bm{x}}^{2}}(2t_{\bm{x}}-\hat{\mu}_{\bm{x},\bm{x}}-\hat{\mu}_{\bm{z},\bm{x}})+\frac{\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}}{2\hat{\sigma}_{\bm{z}}^{2}}(2t_{\bm{z}}-\hat{\mu}_{\bm{x},\bm{z}}-\hat{\mu}_{\bm{z},\bm{z}}).$ (A100)
-When $(\bm{x},y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}$ and $(z,y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}$, $t_{\bm{x}}=\mu_{\bm{x},\bm{x}}+\sigma_{\bm{x}}Z$ and $t_{\bm{z}}=\mu_{\bm{x},\bm{z}}+\sigma_{\bm{z}}Z$.
+When $(\bm{x},y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}$ and $(\bm{z},y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}$, $t_{\bm{x}}=\mu_{\bm{x},\bm{x}}+\sigma_{\bm{x}}Z$ and $t_{\bm{z}}=\mu_{\bm{x},\bm{z}}+\sigma_{\bm{z}}Z$.
 Thus, $\lambda$ becomes
@@
 Note that $q$ is also independent of $t$, thereby $\mathbb{E}_{\bm{z}}[|q|]$ being a constant. By the similar argument using Markov’s inequality and the triangle inequality, we have
-$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{x}}\geq\log\gamma)\geq\tau\right)$ (A100)
+$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{x}}\geq\log\gamma)\geq\tau\right)$ (A117)
 $\leq\Pr_{Z}\left(\Pr_{\bm{z}}(|\lambda_{\bm{x}}|\geq\log\gamma)\geq\tau\right)$ (A118)
@@
 $\textsc{tpr}_{\mathrm{RMIA}}(x)\leq 1-F_{|Z|}^{-1}\left(F_{|Z|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\right).$ (A123)
-Note that, unlike per-example LiRA, we must assume that the attacker has access to the underlying distribution for the optimal RMIA as the Equations A74 and A75 depend on the parameters computed from shadow models.
+Note that, unlike per-example LiRA, we must assume that the attacker has access to the underlying distribution for the optimal RMIA as the Equations˜A91 and A92 depend on the parameters computed from shadow models.
 B.3 RMIA power-law
@@
 where $\alpha\geq\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$, $F_{|Z|}$ is the cdf of the standard folded normal distribution, and
-$\Psi(\bm{x},C)=\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{||\bm{x}||^{2}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{||\bm{z}||^{2}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{||\bm{x}||^{2}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{||\bm{z}||^{2}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{||\bm{x}||}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{||\bm{z}||}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{||\bm{x}||}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{||\bm{z}||}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A108)
+[…]
 In addition, we have
@@
 Proof.
-To apply Lemma 5, we will calculate $\mathbb{E}_{\bm{z}}[|q|]$ and $\mathbb{E}_{\bm{z}}[|A|]$.
+To apply Lemma˜8, we will calculate $\mathbb{E}_{\bm{z}}[|q|]$ and $\mathbb{E}_{\bm{z}}[|A|]$.
 Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),\dots,(\bm{x}_{j,S},j)\}_{j=1}^{C}$.
 Let $s_{y_{a}}^{(\bm{b})}(\bm{a})$ denote the score of $(\bm{a},y_{\bm{a}})$ when $\mathcal{D}_{\mathrm{target}}$ contains $(\bm{b},y_{\bm{b}})$ but not the other example.
-Using similar argument as in the proof of Theorem 2, we have
+Using similar argument as in the proof of Theorem˜2, we have
 $s_{y_{\bm{x}}}^{(\bm{x})}(\bm{x})$ $=\frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}+\bm{x}-\bm{x}_{y_{\bm{x}},S}\rangle$ (A127)
-$s_{y_{\bm{x}}}^{(\bm{z})}(\bm{x})$ =$\begin{cases}\frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}+\bm{z}-\bm{x}_{y_{\bm{x}},S}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A111)
+[…]
 $s_{y_{\bm{z}}}^{(\bm{x})}(\bm{z})$ $=\begin{cases}\frac{1}{S}\langle\bm{z},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}+\bm{x}-\bm{x}_{y_{\bm{x}},S}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \frac{1}{S}\langle\bm{z},\sum_{i=1}^{S}\bm{x}_{y_{\bm{z}},i}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A129)
@@
 $\mu_{\bm{x},\bm{x}}$ $=\langle\bm{x},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle$ (A131)
-$\mu_{\bm{z},\bm{x}}$ =$\begin{cases}\langle\bm{x},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{z}-m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{x},m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A115)
+[…]
 $\mu_{\bm{x},\bm{z}}$ $=\begin{cases}\langle\bm{z},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{x}-m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{z},m_{y_{\bm{z}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A133)
@@
 $\mathbb{E}_{\bm{z}}[|q|]=$ $\Pr_{\bm{z}}(y_{\bm{z}}=y_{\bm{x}})\mathbb{E}_{\bm{z}}[|q|\mid y_{\bm{z}}=y_{\bm{x}}]+\sum_{j=1,j\not=y_{\bm{x}}}^{C}\Pr_{\bm{z}}(y_{\bm{z}}=j)\mathbb{E}_{\bm{z}}[|q|\mid y_{\bm{z}}=j]$ (A141)
-$=$ $\frac{1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-z\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A125)
+$=$ $\frac{1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A142)
 $+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{z}-m_{y_{\bm{z}}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A143)
@@
 $\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\\ =\frac{1}{\sqrt{S}}\cdot\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A151)
-Now Lemma 5 yields
+Now Lemma˜8 yields
 $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ $\leq 1-F_{|Z|}\left(F_{|Z|}^{-1}(1-\alpha)-\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\right)$ (A152)
@@
 This completes the first half of the theorem.
-Now that from Lemma 5 we have
+Now that from Lemma˜8 we have
 $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{z}}+q\geq\log\gamma)\geq\tau\right)$ (A155)
@@
 $=\Pr_{Z}\left(|Z|\geq F_{|Z|}^{-1}(1-\alpha)\right).$ (A160)
-We claim that the bound for $\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation A142) is as tight as that for $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation A139) for sufficiently large $S$.
+We claim that the bound for $\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation˜A159) is as tight as that for $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation˜A156) for sufficiently large $S$.
 Let us denote
@@
 $\kappa_{\textsc{fpr}}(1/\sqrt{S})=\Pr_{\bm{z}}\left(\lambda_{\bm{z}}\geq\frac{1}{\sqrt{S}}\right)=\Pr_{\bm{z}}\left(\frac{\hat{\lambda}_{\bm{z}}}{\sqrt{S}}\geq\frac{1}{\sqrt{S}}\right)=\Pr_{\bm{z}}(\hat{\lambda}_{\bm{z}}\geq 1).$ (A169)
-Hence $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})=O(1/\sqrt{S})$. On the other hand, from Equations A139 and A142 we have for sufficiently large $S$
+Hence $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})=O(1/\sqrt{S})$. On the other hand, from Equations˜A156 and A159 we have for sufficiently large $S$
 $\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}\geq\tau\right)-\Pr_{Z}\left(\frac{\mathbb{E}_{\bm{z}}[|\lambda_{\bm{z}}|]}{\log\gamma}+\frac{\mathbb{E}_{\bm{z}}[|q|]}{\log\gamma}\geq\tau\right)$ (A170)
@@
 $\approx$ $p_{\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]}(\tau)\frac{c_{1}}{\sqrt{S}},$ (A173)
-where $c_{1}$ is some constant. Since for large $S$, $\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]$ and $\tau$ are independent of $S$, Equation A156 scales as $O(1/\sqrt{S})$.
+where $c_{1}$ is some constant. Since for large $S$, $\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]$ and $\tau$ are independent of $S$, Equation˜A173 scales as $O(1/\sqrt{S})$.
 Now let
@@
 $\log(\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}+\log\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}.$ (A184)
-As for the LiRA power-law, bounding $||\bm{x}-m_{\bm{x}}||$ and $||\bm{z}-m_{\bm{z}}||$ will provide a worst-case upper bound for which the power-law holds.
+As for the LiRA power-law, bounding $\bm{x}-m_{\bm{x}}||$ and $\bm{z}-m_{\bm{z}}||$ will provide a worst-case upper bound for which the power-law holds.
 Finally, the per-example RMIA power-law is also extended to the average-case:
@@
 Proof.
-By Theorem 6 and the law of unconscious statistician, we have for large $S$
+By Theorem˜9 and the law of unconscious statistician, we have for large $S$
 $\overline{\textsc{tpr}}_{\mathrm{RMIA}}-\overline{\textsc{fpr}}_{\mathrm{RMIA}}$ $=\int_{\mathscr{D}}p(\bm{x})(\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x}))\mathrm{d}\bm{x}$ (A186)
@@
 C.3 Datasets
-Table A2 shows the datasets used in the paper.
-We base our experiments on a subset of the the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by a practitioner.
+Table˜A2 shows the datasets used in the paper.
+We base our experiments on a subset of the the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a classification accuracy $80\%$ and thus would considered to be used by a practitioner.
 Additionally, we add CIFAR10 which is not part of the original VTAB benchmark.
@@
 dataset (max.) min. max. test accuracy test accuracy
-$C$ $S$ $S$ (min. $S$) (max. $S$)
+[…]
 Patch Camelyon (Veeling et al., 2018) 2 256 65536 82.8% 85.6%
@@
 In this section, we provide tabular results for our experiments and additional figures that did not fit into the main paper.
-D.1 Additional results for Section 4
-This Section contains additional results for Section 4.
+D.1 Additional results for Section˜4
+This Section contains additional results for Section˜4.
 D.1.1 Vulnerability as a function of shots
-This section displays additional results to Figure 1 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.1, A3 and A4.
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+This section displays additional results to Figure˜1 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures˜A.1, A3 and A4.
+(a) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.1$
+(b) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.1$
+(c) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.01$
+(d) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.01$
+(e) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.001$
+(f) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.001$
 Figure A.1: MIA vulnerability as a function of shots (examples per class) when attacking a pre-trained ViT-B and R-50 Head trained without DP on different downstream datasets.
@@
 D.1.2 Vulnerability as a function of the number of classes
-This section displays additional results to Figure 2 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A5 and A6.
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+This section displays additional results to Figure˜2 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures˜A.2, A5 and A6.
+(a) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.1$
+(b) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.1$
+(c) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.01$
+(d) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.01$
+(e) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.001$
+(f) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.001$
 Figure A.2: MIA vulnerability as a function of $C$ (classes) when attacking a ViT-B and R-50 Head fine-tuned without DP on different datasets where the classes are randomly sub-sampled and $S=32$.
@@
 D.1.3 Data for FiLM and from scratch training
-Table A7: MIA vulnerability data used in Figure 5(b).
+Table A7: MIA vulnerability data used in Figure˜5(b).
 Note that the data from Carlini et al.
@@
 D.1.4 Predicting dataset vulnerability as function of $S$ and $C$
-This section provides additional results for the model based on Equation 13
-Table A8: Results for fitting Equation 13 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$.
+This section provides additional results for the model based on Equation˜13
+Table A8: Results for fitting Equation˜13 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$.
 We utilize an ordinary least squares.
 The test $R^{2}$ assesses the fit to the data of R-50 Head.
-coeff.
-fpr $R^{2}$ test $R^{2}$ coeff.
-value std.
-error $t$ $p>|z|$ coeff.
-$[0.025$ coeff.
-$0.975]$ $\beta_{S}$ (for $S$) 0.1 0.952 0.907 -0.506 0.011 -44.936 $0$0.529 -0.484 0.01 0.946 0.854 -0.555 0.014 -39.788 $0$0.582 -0.527 0.001 0.930 0.790 -0.627 0.019 -32.722 $0$0.664 -0.589 $\beta_{C}$ (for $C$) 0.1 0.952 0.907 0.090 0.021 4.231 0.000 0.048 0.131 0.01 0.946 0.854 0.182 0.026 6.960 0.000 0.131 0.234 0.001 0.930 0.790 0.300 0.036 8.335 0.000 0.229 0.371 $\beta_{0}$ (intercept) 0.1 0.952 0.907 0.314 0.045 6.953 0.000 0.225 0.402 0.01 0.946 0.854 0.083 0.056 1.491 0.137 -0.027 0.193 0.001 0.930 0.790 -0.173 0.077 -2.261 0.025 -0.324 -0.022
-Figure A.3 shows the performance for all considered fpr.
 […]
-[…]
-[…]
-[…]
+Figure˜A.3 shows the performance for all considered fpr.
+(a) ($(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.1$
+(b) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.01$
+(c) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.001$
+Figure A.3: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation˜13 fitted Table˜A3 (ViT-B).
+The triangles show the median $\textsc{tpr}-\textsc{fpr}$ for the train set (ViT-B; Table˜A3) and circle the test set (R-50; Table˜A4) over six seeds.
+Note that the triangles and dots for $C=10$ are for EuroSAT.
 D.2 Simpler variant of the prediction model
-[…]
-The general form this variant can be found in Equation A174, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
+The prediction model in the main text (Equation˜13) avoids predicting $\textsc{tpr}<\textsc{fpr}$ in the tail when $S$ is very large.
+In this section, we analyse a variation of the regression model that is simpler and predicts $\log_{10}(\textsc{tpr})$ instead of $\log_{10}(\textsc{tpr}-\textsc{fpr})$.
+This variation fits worse to the empirical data and will predict $\textsc{tpr}<\textsc{fpr}$ for high $S$.
+The general form this variant can be found in Equation˜A191, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
 $\log_{10}(\textsc{tpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (A191)
-Table A9 provides tabular results on the performance of the variant.
-Table A9: Results for fitting Equation A174 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$.
+Table˜A9 provides tabular results on the performance of the variant.
+Table A9: Results for fitting Equation˜A191 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$.
 We utilize an ordinary least squares.
 The test $R^{2}$ assesses the fit to the data of R-50 Head.
-coeff.
-fpr $R^{2}$ test $R^{2}$ coeff.
-value std.
-error $t$ $p>|z|$ coeff.
-$[0.025$ coeff.
-$0.975]$ $\beta_{S}$ (for $S$) 0.1 0.908 0.764 -0.248 0.008 -30.976 $0$0.264 -0.233 0.01 0.940 0.761 -0.416 0.011 -36.706 $0$0.438 -0.393 0.001 0.931 0.782 -0.553 0.017 -32.507 $0$0.586 -0.519 $\beta_{C}$ (for $C$) 0.1 0.908 0.764 0.060 0.015 3.955 0.000 0.030 0.089 0.01 0.940 0.761 0.169 0.021 7.941 0.000 0.127 0.211 0.001 0.931 0.782 0.297 0.032 9.303 0.000 0.234 0.360 $\beta_{0}$ (intercept) 0.1 0.908 0.764 0.029 0.032 0.913 0.362 -0.034 0.093 0.01 0.940 0.761 -0.118 0.045 -2.613 0.010 -0.208 -0.029 0.001 0.931 0.782 -0.295 0.068 -4.345 $0$0.429 -0.161
-Figure A.4 plots the performance of the variant similar to Figure 5(a) in the main text.
+[…]
+Figure˜A.4 plots the performance of the variant similar to Figure˜5(a) in the main text.
 (a) tpr at $\textsc{fpr}=0.1$
@@
 (c) tpr at $\textsc{fpr}=0.001$
-Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation A174 fitted Table A3 (ViT-B).
-The triangles show the median tpr for the train set (ViT-B; Table A3) and circle the test set (R-50; Table A4) over six seeds.
+Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation˜A191 fitted Table˜A3 (ViT-B).
+The triangles show the median tpr for the train set (ViT-B; Table˜A3) and circle the test set (R-50; Table˜A4) over six seeds.
 Note that the triangles and dots for $C=10$ are for EuroSAT.
 D.3 Empirical results for RMIA
-Figures A.5, A.6 and A.7 report additional results for RMIA Zarifzadeh et al. (2024).
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Figures˜A.5, A.6 and A.7 report additional results for RMIA Zarifzadeh et al. (2024).
+(a) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.1$
+(b) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.01$
+(c) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.001$
+Figure A.5: RMIA (Zarifzadeh et al., 2024) vulnerability ($\textsc{tpr}-\textsc{fpr}$ at fixed fpr) as a function of $S$ (shots) when attacking a ViT-B Head fine-tuned without DP on different datasets.
+We observe at power-law relationship but especially at low fpr the relationship is not as clear as with LiRA (compare to Figure˜A.1).
+The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over six seeds.
+(a) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.1$
+(b) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.01$
+(c) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.001$
+Figure A.6: Predicted MIA vulnerability ($(\textsc{tpr}-\textsc{fpr})$ at fpr) based on LiRA vulnerability data as a function of $S$ (shots) in comparison to observed RMIA (Zarifzadeh et al., 2024) vulnerability on the same settings.
+The triangles show the highest tpr when attacking (ViT-B Head) with RMIA over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100).
+Especially at $\textsc{fpr}=0.1$ the relationship behaves very similar for both MIAs, but RMIA shows more noisy behavior at lower fpr.
+(a) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.1$
+(b) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.01$
+(c) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.001$
+Figure A.7: LiRA and RMIA vulnerability ($(\textsc{tpr}-\textsc{fpr})$) as a function of shots ($S$) when attacking a ViT-B Head fine-tuned without DP on different datasets.
+For better visibility, we split the datasets into two panels.
+We observe the power-law for both attacks, but the RMIA is more unstable than LiRA.
+The lines display the median over six seeds.
 D.4 Tabular results for Section 4.4
-Table A10 displays the tabular results for Figure 6 in Section 4.4.
-Table A10: Tabular results for Figure 6 on when attacking a ViT-B (Head) fine-tuned on PatchCamelyon. We display the median over six seeds at $\textsc{fpr}=0.1$.
+Table˜A10 displays the tabular results for Figure˜6 in Section˜4.4.
+Table A10: Tabular results for Figure˜6 on when attacking a ViT-B (Head) fine-tuned on PatchCamelyon. We display the median over six seeds at $\textsc{fpr}=0.1$.
 $S$ Max tpr tpr of 0.999 Quantile tpr of 0.99 Quantile tpr of 0.95 Quantile
@@
 Appendix E Details on Section 4.5
-In Section 4.5, we compare the results of our empirical models of vulnerability (Sections 4.3 and 4.4) to DP bounds. Below we explain how we make the connection between both.
-First, we compute the upper bound on the tpr for a given $(\epsilon,\delta)$-DP privacy budget at a given fpr using Equation A175 reformulated from Kairouz et al. (2015) below:
+In Section˜4.5, we compare the results of our empirical models of vulnerability (Sections˜4.3 and 4.4) to DP bounds. Below we explain how we make the connection between both.
+First, we compute the upper bound on the tpr for a given $(\epsilon,\delta)$-DP privacy budget at a given fpr using Equation˜A192 reformulated from Kairouz et al. (2015) below:
 Theorem 11 (Kairouz et al. (2015)).
@@
 For a given $(\epsilon,\delta)$ and fpr we then obtain a value for the tpr.
-Next, we use the linear model from Section 4.3 to solve for the minimum $S$ predicted to be required given $C=2$ classes in our example.
-The coefficients can be found in Table A8 for the average case and Section 4.4 for the worst-case.
+Next, we use the linear model from Section˜4.3 to solve for the minimum $S$ predicted to be required given $C=2$ classes in our example.
+The coefficients can be found in Table˜A8 for the average case and Section˜4.4 for the worst-case.
 We solve the tpr from the linear model as
@@
 $\Leftrightarrow$ $\textsc{tpr}=S^{\beta_{S}}C^{\beta_{C}}10^{\beta_{0}}+\textsc{fpr}.$ (A194)
-Now, we find the minimum $S$ that the tpr from Equation A177 upper bounds the tpr of Equation A175 as
+Now, we find the minimum $S$ that the tpr from Equation˜A194 upper bounds the tpr of Equation˜A192 as
 $S^{\beta_{S}}C^{\beta_{C}}10^{\beta_{0}}+\textsc{fpr}=\min\{e^{\epsilon}\textsc{fpr}+\delta,1-e^{-\epsilon}(1-\delta-\textsc{fpr})\}$ (A195)

### v6

@@
 MIAs complement formal guarantees from differential privacy (DP) under a more realistic adversary model.
-We analyse MIA vulnerability of fine-tuned neural networks both empirically and theoretically, the latter using a simplified model of fine-tuning.
+We analyze MIA vulnerability of fine-tuned neural networks both empirically and theoretically, the latter using a simplified model of fine-tuning.
 We show that the vulnerability of non-DP models when measured as the attacker advantage at a fixed false positive rate reduces according to a simple power law as the number of examples per class increases.
@@
 To achieve our goal, we theoretically analyze and systematically apply two state-of-the-art black-box MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), to help understand practical privacy risks when fine-tuning deep-learning-based classifiers without DP protections.
-For the theoretical model that we analyse, LiRA is the optimal attack by the Neyman–Pearson lemma (Neyman and Pearson, 1933).
+For the theoretical model that we analyze, LiRA is the optimal attack by the Neyman–Pearson lemma (Neyman and Pearson, 1933).
 Under the black-box threat model, in which the adversary does not have access to the model parameters, LiRA and RMIA have been shown to empirically outperform other attacks, especially when the number of shadow models is sufficiently large.
@@
 Related work There has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017), models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2025), and classes with less examples tend to be more vulnerable (Chang and Shokri, 2021; Kulynych et al., 2022; Tonni et al., 2020).
-A larger generalisation error, which is related to dataset size, has also been shown to be sufficient for MIA success (Song and Mittal, 2021), though not necessary (Yeom et al., 2018).
+A larger generalization error, which is related to dataset size, has also been shown to be sufficient for MIA success (Song and Mittal, 2021), though not necessary (Yeom et al., 2018).
 Similarly, minority subgroups tend to be more affected by DP (Suriyakumar et al., 2021; Bagdasaryan et al., 2019).
-Feldman and Zhang (2020) showed that neural networks trained from scratch are required to memorize a significant fraction of their training data to obtain high utility, while the memorisation is greatly reduced for fine-tuning.
+Feldman and Zhang (2020) showed that neural networks trained from scratch are required to memorize a significant fraction of their training data to obtain high utility, while the memorization is greatly reduced for fine-tuning.
 Additionally, Tobaben et al.
@@
 (2023) studied the relationship between the MIA vulnerability and individual privacy parameters for different classes.
-Recently, worst-case MIA vulnerability has gained more attention (Guépin et al., 2024; Meeus et al., 2024; Azize and Basu, 2024).
+Recently, worst-case MIA vulnerability has gained more attention (Guépin et al., 2024; Meeus et al., 2024; Azize and Basu, 2025).
 Nonetheless, the prior works do not consider the rate of change in the vulnerability evaluated at a low fpr, as dataset properties change.
@@
 We fit a regression model which follows the functional form of the theoretically derived power-law.
-We show both a very good fit on the training data as well as a good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Section˜4.3).
+We show both a very good fit on the training data as well as a good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterizations (see Section˜4.3).
   3. 3.
@@
 We sample vectors on a high-dimensional unit sphere and classify them based on inner product with estimated class mean.
-This model is easier to analyse theoretically than real-world deep learning examples.
+This model is easier to analyze theoretically than real-world deep learning examples.
 We generate the data and form the classifiers (which are our target models) as follows:
   1. 1.
-    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in\{1,\dots,C\},i\neq j:m_{i}\perp m_{j}$).
+    For each class, we first sample a true class mean $\bm{m}_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in\{1,\dots,C\},i\neq j:\bm{m}_{i}\perp\bm{m}_{j}$).
   2. 2.
     We sample $2S$ vectors $\bm{x}_{c}$ for each class.
-We assume that they are Gaussian distributed around the the true class mean $\bm{x}_{c}\sim\mathcal{N}(m_{c},\Sigma)$ where the $\Sigma$ is the in-class covariance.
+We assume that they are Gaussian distributed around the the true class mean $\bm{x}_{c}\sim\mathcal{N}(\bm{m}_{c},\Sigma)$ where the $\Sigma$ is the in-class covariance.
   3. 3.
-    For each ”target model” we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $\hat{m}_{c}$.
+    For each “target model” we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $\hat{\bm{m}}_{c}$.
   4. 4.
-    The computed mean is used to classify sample $\bm{x}$ by computing the inner product $\langle\bm{x},\hat{m}_{c}\rangle$ as a metric of similarity.
+    The computed mean is used to classify sample $\bm{x}$ by computing the inner product $\langle\bm{x},\hat{\bm{m}}_{c}\rangle$ as a metric of similarity.
 The attacker has to infer which vectors have been used for training the classifier.
@@
 Fix a target example $(\bm{x},y)$. For the simplified model with arbitrary $C$ and infinitely many shadow models, the per-example LiRA vulnerability is given as
-$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (9)
-where $m_{\bm{x}}$ is the true mean of class $y$ and $\xi(S)=O(1/\sqrt{S})$. For large $S$ we have
-$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (10)
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (9)
+where $\bm{m}_{\bm{x}}$ is the true mean of class $y$ and $\xi(S)=O(1/\sqrt{S})$. For large $S$ we have
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (10)
 Proof.
@@
 […]
-This implies that if $\bm{x}-m_{\bm{x}}||$ is bounded, then the worst-case vulnerability is also bounded.
+This implies that if $\bm{x}-\bm{m}_{\bm{x}}||$ is bounded, then the worst-case vulnerability is also bounded.
 Hence we can significantly reduce the MIA vulnerability of all examples in this non-DP setting by simply increasing the number of examples per class.
@@
 In our simplified model, this holds with $\alpha=1/2$.
-However, $\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x})\rightarrow 0$ would not always be the case for larger neural networks trained from scratch.
-Therefore, we do not expect similar power-laws in more general training algorithms.
+However, $\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x})\rightarrow 0$ might not always be the case for larger neural networks trained from scratch.
+Furthermore, even if $(\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x}))/\sigma(\bf{x})$ converges to zero, it is not clear what the convergence rate would be.
 Now the following corollary extends the power-law to the average-case MIA vulnerability. We will also empirically validate this result in Section˜4.
@@
 For the simplified model with arbitrary $C$, sufficiently large $S$ and infinitely many shadow models, we have
-$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}\right]\right).$ (12)
+$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}\right]\right).$ (12)
 Proof.
@@
 The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section˜D.1.
-We optimise the hyperparameters (batch size, learning rate and number of epochs) using the Optuna library (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section˜C.2).
+We optimize the hyperparameters (batch size, learning rate and number of epochs) using the Optuna library (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section˜C.2).
 We provide the the code for reproducing the experiments in an open repository [^1] https://github.com/DPBayes/impact-dataset-properties-MI-vulnerability-deep-TL.
@@
 Prediction quality on other MIA target models
-We analyse how the regression model trained on the ViT-B (Head) data generalizes to other target models. The main points are:
+We analyze how the regression model trained on the ViT-B (Head) data generalizes to other target models. The main points are:
   • R-50 (Head): Figure˜5(a) shows that the regression model is robust to a change of the feature extractor, as it is able to predict the tpr for R-50 (Head) (test $R^{2}=0.790$).
@@
 We are not using this optimization.
-Additionally, as noted in Remark˜3, our theoretical analysis suggests that the power-law would not always hold for general from-scratch training.
+Additionally, as noted in Remark˜3, our theoretical analysis is based on a simple model and the power-law might not occur at all for larger neural networks trained from scratch.
 (a) The dots show the median tpr for the train set (ViT-B; Table˜A3) and the test set (R-50; Table˜A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100).
@@
 Figure 5: Performance of the regression model based on Equation˜13 fitted on data from Table˜A3.
+Figure 6: Individual vulnerability for ViT-B (Head) when fine-tuning on Patch Camelyon.
+We observe a similar power-law relationship for individuals when looking at the quantiles but the max decreases slower.
+The Average Case shows the average LiRA vulnerability over the individuals also illustrated in the right panel of Figure˜3.
+Table 1: The minimum $S$ with CC=22 predicted by the models in Sections˜4.3 and 4.4 to empirically match the DP bounds ($\delta$$10^{-5}$) calculated through (Kairouz et al., 2015) in terms of tpr at fpr.
+$\epsilon$ min $S$ in average case min $S$ in worst-case fpr=$0.1$ fpr=$0.01$ fpr=$0.001$ fpr=$0.1$ $0.25$ 5 4005\,400 69 00069\,000 320 000320\,000 $5.5\times 10^{9}$ $0.50$ 1 1001\,100 16 00016\,000 88 00088\,000 $2.6\times 10^{8}$ $0.75$ $360$ 5 9005\,900 38 00038\,000 $3.5\times 10^{7}$ $1.00$ $160$ 2 7002\,700 19 00019\,000 $7.0\times 10^{6}$
+Figure 7: Illustration of the extrapolation for the average-case at $\textsc{fpr}=10^{-3}$ (left) and worst-case at $\textsc{fpr}=10^{-1}$ (right).
 4.4 Individual MIA vulnerability
 In order to assess the per-sample MIA vulnerability, we run the experiment with $257$ models with each of them once acting as a target and as a shadow model otherwise, compute the tpr at fpr for every sample separately.
-In Figure˜6, we display the individual vulnerability as a function of $S$ (shots) for Patch Camelyon.
+In Figure˜6, we display the individual vulnerability as a function of $S$ (shots) for Patch Camelyon and compare it with corresponding average case vulnerability from Figure˜3.
 The plot shows the maximal vulnerability of all samples and different quantiles over six seeds.
@@
 With larger $S$ the slope of the max vulnerability increases, e.g., with $S\geq 32768$ the coefficient is 0.3478-0.3478 suggesting that higher $S$ is required for the most vulnerable points.
-Figure 6: Individual vulnerability for ViT-B (Head) when fine-tuning on Patch Camelyon.
-We observe a similar power-law relationship for individuals when looking at the quantiles but the max decreases slower.
-Table 1: The minimum $S$ with CC=22 predicted by the models in Sections˜4.3 and 4.4 to empirically match the DP bounds ($\delta$$10^{-5}$) calculated through (Kairouz et al., 2015) in terms of tpr at fpr.
-$\epsilon$ min $S$ in average case min $S$ in worst-case fpr=$0.1$ fpr=$0.01$ fpr=$0.001$ fpr=$0.1$ $0.25$ 5 4005\,400 69 00069\,000 320 000320\,000 $5.5\times 10^{9}$ $0.50$ 1 1001\,100 16 00016\,000 88 00088\,000 $2.6\times 10^{8}$ $0.75$ $360$ 5 9005\,900 38 00038\,000 $3.5\times 10^{7}$ $1.00$ $160$ 2 7002\,700 19 00019\,000 $7.0\times 10^{6}$
 4.5 Comparison between empirical models and universal DP bounds
@@
 (2015), we compute the minimum $S$ predicted by our empirical models such that the predicted tpr matches the theoretical bound at target DP level.
+We illustrate this in Figure˜7.
 (See Appendix˜E for a more detailed description.) Table˜1 shows the resulting lower bounds of $S$ for various DP levels and values of fpr.
@@
 Limitations
-We mostly consider LiRA in our paper which is optimal for our simplified model of membership inference (Section˜3.4) but for the transfer-learning experiments (Section˜4) there might be stronger attacks in the future.
-Furthermore, our simplified model assumes well-behaved underlying distributions, meaning that the data is normally distributed around the class centres.
-We leave the analysis of other data distributions (e.g., heavy-tailed distributions) to future work..
+We mostly consider LiRA in our paper which is optimal for our simplified model of membership inference (Section˜3.4), but for the transfer-learning experiments (Section˜4) there might be stronger attacks in the future.
+Furthermore, our simplified model assumes well-behaved underlying distributions, meaning that the data is normally distributed around the class centers.
+We leave the analysis of other data distributions (e.g., heavy-tailed distributions) to future work.
 Formal bounds on MIA vulnerability would require something like DP.
@@
 Models trained from scratch are likely to be more vulnerable.
+Concurrently, a related paper (Hayes et al., 2025) has studied the relation between LiRA vulnerability and dataset size when attacking LLMs during the pre-training phase.
+Extending this analysis is promising but computationally very expensive.
 We provide the code in an open repository [^2] https://github.com/DPBayes/impact-dataset-properties-MI-vulnerability-deep-TL. All used pre-trained models and datasets are publicly available.
@@
 Fix a target example $(\bm{x},y)$. For the simplified model with arbitrary $C$ and infinitely many shadow models, the per-example offline LiRA vulnerability is given as
-$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S))$ (A44)
+$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S))$ (A44)
 where $m_{\bm{x}}$ is the true mean of class $y$ and $\xi(S)=O(1/\sqrt{S})$. For large $S$ we have
-$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A45)
+$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A45)
 A.5 Relaxing the assumption of Lemma˜1
@@
 respectively, when $(\bm{x},y)\in\mathcal{D}_{\mathrm{target}}$ and when $(\bm{x},y)\notin\mathcal{D}_{\mathrm{target}}$. Thus we obtain
-$\mu_{\mathrm{in}}-\mu_{\mathrm{out}}$ $=\mathbb{E}[s_{y}^{\mathrm{(in)}}-s_{y}^{\mathrm{(out)}}]=\frac{1}{S}\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle$ (A55)
+$\mu_{\mathrm{in}}-\mu_{\mathrm{out}}$ $=\mathbb{E}[s_{y}^{\mathrm{(in)}}-s_{y}^{\mathrm{(out)}}]=\frac{1}{S}\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle$ (A55)
 $\sigma^{2}$ $=\operatorname{Var}(s_{y}^{\mathrm{(out)}})=\frac{1}{S}\operatorname{Var}(\langle\bm{x},\bm{x}_{y,i}\rangle)=\frac{1}{S}\bm{x}^{T}\Sigma\bm{x}.$ (A56)
 Noting that the LiRA score follows a normal distribution, by Lemma 1 we have
-$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right),$ (A57)
+$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right),$ (A57)
 where $\Phi$ is the cdf of the standard normal distribution. This completes the first half of the theorem.
 Now let $\phi(u)$ denote the pdf of the standard normal distribution, and let
-$r=\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}.$ (A58)
+$r=\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}.$ (A58)
 Using Taylor expansion of $\Phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+r)$ around $r=0$, we have
@@
 $=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+\sum_{k=1}^{\infty}\frac{(-1)^{k-1}\mathrm{He}_{k-1}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))}{k!}r^{k}$ (A62)
-$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+r\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}r^{k}.$ (A63)
-It follows that
+$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+r\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}r^{k},$ (A63)
+where $\mathrm{He}$ denotes Hermite polynomials. It follows that
 $\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))$ (A64)
 $=$ $\log r+\log\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+\log\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}r^{k}\right)$ (A65)
-$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A66)
-$+\log\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A67)
-$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A68)
-$+\log\left(1+\sum_{k=1}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A69)
-$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (A70)
+$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A66)
+$+\log\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A67)
+$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A68)
+$+\log\left(1+\sum_{k=1}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A69)
+$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (A70)
 where we have $\xi(S)=O(1/\sqrt{S})$. For large enough $S$, ignoring the residual term, we approximate
-$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A71)
+$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A71)
 A.7 Proof of Corollary˜4
@@
 $\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}}$ $=\int_{\mathscr{D}}p(\bm{x})(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\mathrm{d}\bm{x}$ (A72)
-$\approx\int_{\mathscr{D}}p(\bm{x})\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\mathrm{d}\bm{x}$ (A73)
-$=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\frac{1}{\sqrt{S}}\int_{\mathscr{D}}p(\bm{x})\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\mathrm{d}\bm{x}$ (A74)
-$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{2\pi}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right],$ (A75)
+$\approx\int_{\mathscr{D}}p(\bm{x})\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\mathrm{d}\bm{x}$ (A73)
+$=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\frac{1}{\sqrt{S}}\int_{\mathscr{D}}p(\bm{x})\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\mathrm{d}\bm{x}$ (A74)
+$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{2\pi}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right],$ (A75)
 where $p(\bm{x})$ is the density of $\mathbb{D}$} at $(\bm{x},y)$, and $\mathscr{D}$ is the data domain.
@@
 Then we obtain
-$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{2\pi}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right]\right).$ (A76)
+$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{2\pi}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right]\right).$ (A76)
 Appendix B Theoretical analysis of RMIA
@@
 Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),\dots,(\bm{x}_{j,S},j)\}_{j=1}^{C}$.
-Let $s_{y_{a}}^{(\bm{b})}(\bm{a})$ denote the score of $(\bm{a},y_{\bm{a}})$ when $\mathcal{D}_{\mathrm{target}}$ contains $(\bm{b},y_{\bm{b}})$ but not the other example.
+Let $s_{y_{\bm{a}}}^{(\bm{b})}(\bm{a})$ denote the score of $(\bm{a},y_{\bm{a}})$ when $\mathcal{D}_{\mathrm{target}}$ contains $(\bm{b},y_{\bm{b}})$ but not the other example.
 Using similar argument as in the proof of Theorem˜2, we have
@@
 Thus we obtain
-$\mu_{\bm{x},\bm{x}}$ $=\langle\bm{x},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle$ (A131)
+$\mu_{\bm{x},\bm{x}}$ $=\langle\bm{x},\bm{m}_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{x}-\bm{m}_{y_{\bm{x}}}\rangle$ (A131)
 […]
-$\mu_{\bm{x},\bm{z}}$ $=\begin{cases}\langle\bm{z},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{x}-m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{z},m_{y_{\bm{z}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A133)
-$\mu_{\bm{z},\bm{z}}$ $=\langle\bm{z},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{z}-m_{y_{\bm{x}}}\rangle\quad$ (A134)
+$\mu_{\bm{x},\bm{z}}$ $=\begin{cases}\langle\bm{z},\bm{m}_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{x}-\bm{m}_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{z},\bm{m}_{y_{\bm{z}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A133)
+$\mu_{\bm{z},\bm{z}}$ $=\langle\bm{z},\bm{m}_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{z}-\bm{m}_{y_{\bm{x}}}\rangle\quad$ (A134)
 $\sigma_{\bm{x}}$ $=\frac{1}{\sqrt{S}}\sqrt{\bm{x}^{T}\Sigma\bm{x}}$ (A135)
 $\sigma_{\bm{z}}$ $=\frac{1}{\sqrt{S}}\sqrt{\bm{z}^{T}\Sigma\bm{z}},$ (A136)
-where $m_{y_{\bm{z}}}$ is the true class mean of class $y_{\bm{z}}$.
+where $\bm{m}_{y_{\bm{z}}}$ is the true class mean of class $y_{\bm{z}}$.
 Now recall that
@@
 $=$ $\frac{1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A142)
-$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{z}-m_{y_{\bm{z}}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A143)
+$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-\bm{m}_{y_{\bm{x}}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{z}-\bm{m}_{y_{\bm{z}}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A143)
 $=$ $\frac{1}{CS}\mathbb{E}_{\bm{z}}\left[\left.\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A144)
-$+\frac{C-1}{CS}\mathbb{E}_{\bm{z}}\left[\left.\frac{\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{y_{\bm{z}}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right],$ (A145)
+$+\frac{C-1}{CS}\mathbb{E}_{\bm{z}}\left[\left.\frac{\langle\bm{x},\bm{x}-\bm{m}_{y_{\bm{x}}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{y_{\bm{z}}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right],$ (A145)
 and
@@
 $=$ $\frac{1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A147)
-$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A148)
+$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A148)
 $=$ $\frac{1}{C\sqrt{S}}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A149)
-$+\frac{C-1}{C\sqrt{S}}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right].$ (A150)
+$+\frac{C-1}{C\sqrt{S}}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right].$ (A150)
 Hence we obtain
-$\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\\ =\frac{1}{\sqrt{S}}\cdot\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A151)
+$\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\\ =\frac{1}{\sqrt{S}}\cdot\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A151)
 Now Lemma˜8 yields
@@
 where $F_{|Z|}$ is the cdf of the folded normal distribution and
-$\Psi(\bm{x},C)=\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A154)
+$\Psi(\bm{x},C)=\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A154)
 This completes the first half of the theorem.
@@
 $\log(\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}F_{|Z|}^{-1}(1-\alpha)^{2}+\log\frac{\Psi(\bm{x},C)}{\sqrt{\pi/2}}.$ (A184)
-As for the LiRA power-law, bounding $\bm{x}-m_{\bm{x}}||$ and $\bm{z}-m_{\bm{z}}||$ will provide a worst-case upper bound for which the power-law holds.
+As for the LiRA power-law, bounding $\bm{x}-\bm{m}_{\bm{x}}||$ and $\bm{z}-\bm{m}_{\bm{z}}||$ will provide a worst-case upper bound for which the power-law holds.
 Finally, the per-example RMIA power-law is also extended to the average-case:
@@
 The prediction model in the main text (Equation˜13) avoids predicting $\textsc{tpr}<\textsc{fpr}$ in the tail when $S$ is very large.
-In this section, we analyse a variation of the regression model that is simpler and predicts $\log_{10}(\textsc{tpr})$ instead of $\log_{10}(\textsc{tpr}-\textsc{fpr})$.
+In this section, we analyze a variation of the regression model that is simpler and predicts $\log_{10}(\textsc{tpr})$ instead of $\log_{10}(\textsc{tpr}-\textsc{fpr})$.
 This variation fits worse to the empirical data and will predict $\textsc{tpr}<\textsc{fpr}$ for high $S$.
@@
 $\Rightarrow$ $S=\left(\frac{\min\{e^{\epsilon}\textsc{fpr}+\delta,1-e^{-\epsilon}(1-\delta-\textsc{fpr})\}-\textsc{fpr}}{C^{\beta_{C}}10^{\beta_{0}}}\right)^{1/{\beta_{S}}}$ (A196)
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The main claims that are summarized at the end of Section˜1 accuracly reflect the contributions in Sections˜3 and 4.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The limiations are discussed in Section˜5.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: The details for the proofs in Section˜3 are in Appendix˜A.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Experimental details are in Appendix˜C and the documented source code is in the supplementary material.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: The source code including instructions can be found in the supplementary material.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We provide details in Appendix˜C and the remaining details can be found in the documented source code.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: In Section˜4 we provide results over at least six seeds and plot the median as well as errorbars. Many errorbars are based on Clopper-Pearson intervals.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Details can be found in Section˜C.4.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conforms with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We discuss the broader impacts in Section˜5.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper does not pose such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We cite the original owners and mention the licenses in Sections˜C.3.1 and C.1.1.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: The documentation for the code is with the code in the supplementary material.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.